A Complete Bibliography of The Bell System Technical Journal, 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.08

Title word cross-reference

01 [CM77]. a_{n} [Ric73b]. B [Mes72]. μ [ACG79]. ±1 [Ric73b]. ∑ a_{n}/n [Ric73b].

-Law [ACG79].

0.9- [EKN76].

1 [Hey75, Van75b]. 1-Watt [MG75]. 1.4X [PS70]. 10-W [TG76]. 100-GHz [Sem77]. 11-GHz [HGLM77, Lin77a, Osb77]. 12 [Aca79]. 12-GHz [Rus78]. 12.4-GHz [EKN76]. 12/14 [Aca79]. 14 [Aca79]. 15.3-GHz [Pen70]. 18-GHz [Osb71, SBR78a]. 19- [ACH^{+}78, Cox78]. 19-GHz [ACG78, AC78]. 1972 [ALM74].

1978 [Ano78a]. 1A [AOR76, ABG^{+}77, BCKM70, BCE^{+}77, BDG^{+}77, BDN^{+}77, CFG^{+}77, DL76, DKP^{+}70, GAW^{+}76, GBN^{+}76, HW76, HMSV77, KLS76, Kle76, KU76, Sta77].

2 [Bro71a]. 2.6-Kilometer [Sem71]. 20- [Tur75]. 250-Element [TKK73]. 28-GHz [ACH^{+}78, Cox78]. 2C [CSW71].

3 [ALS71, BCS71, BJL71, BW71, CGL71, Die71, DSW71, Fen71, Giu71, HJ71, JP71, KS71, SW71a, SW71b]. 3-Phase [Seq74]. 3.5X [PS70]. 30-GHz [Sem79, Tur75]. 302A [JPS^{+}78]. 39A [FP74].

4 [ALM74, CT79]. 4-GHz [KS78]. 45-Mb [MSW78]. 45-Mb/s [MSW78]. 48 [Ano70l].

49 [Ano70k, Ano71k]. 4A [ALS71, CFMS78].

500-Element [Seq74]. 52 [NK73a]. 53 [Ano75k]. 56 [Ano78n]. 57 [CH79]. 58 [Ano79l].

6-GHz [TG76]. 6-Gigahertz [MG75]. 60-GHz [DDH75]. 60-mm [Doa78].

7-Meter [CWE^{+}78]. 70-MHz [CKS79].

8 [Rov78]. 850A [And71].

A. [Ano71k, Ano79l]. A/D [Net77]. ABC [Alb78]. Abelian [Mac70]. Absorption [Mar76d]. Abstracts [Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e]. Access [BBB^{+}79, CR78, GR78, Nea72, SW73, WP71]. Accuracy [Hil77, Hol73, KN72, Lev78, MS72]. Accurate [Sch78]. Achievements [ACTW77]. Acid [Ano71k, BW70a, BLV70, CFB70, Cus70, DS70, Gar70, HRH70, KFBL70, Lue70b, Mil70, OBW70, SSV70]. Acoustic [Ano78n, CS77b, Fla79, IF77]. Acoustical [GS70]. Acquisition [Cha70a]. Across [Kra71]. Active [Bau72, FL79, FDDM75, Mos70, Sze73, Tow75]. Adaptation [Ano79l, NMW79]. Adaptive [AS70, Boy76, Cro78b, CJF73, Cum75, FM73, Fal76b, Fal78, FM79, GW79, Goo71, Has75, Jay70, Jay73, Jay77, JC79, MS79a, MSG78, Mit74, Mit75, MG78, Mue75, NPM77, RT71, RB74, Sca79, Tho71b, Tho71c, Yeh70]. Adaptively [MS79b]. Added [BB78]. Adder [Wil72]. Addition [MPS77]. Additive [HY70, MF70]. Address [Wol72]. Addressing [BGK72, GP71]. ADF [AC70, AAS70, Ano70m, BGK70, CFP70, CLT70, EG70, KNS^{+}70, PSW70]. Adhesive [SP77]. Adhesives [Dah72]. Adjusted [Cho74]. Adjustment [Owe72, Raw70]. ADM [Aha78b]. ADM-Encoded [Aha78b].

Administering [CGHW74]. Administration [BZ74, BS78, Fre78, GLW74, GHHJ77, GHRS77, MOR^{+}77, Mor78b, Sna78]. Administrative [CFG^{+}77]. ADPCM [ACG79, RSR74]. Advanced [AFS79, CHPT79, DPZ79, EFW79, Fis79, FP79, Huf79, Mac79, Tsi79, Wal79, You79]. Advertisements [CD79]. After [Con72b]. AGC [FM79]. Aging [Wah77]. Aid [Fla79]. Aided [HN78, MS75a]. Aids [Fra78, O’N71]. Al [Dix76]. Alarm [Fan76]. Alerter [DMK78]. Algebra [Aha72a, Aha72b, Aha72d, Mac70]. Algorithm [CR73a, Els77, Fal76a, Mue75, NS76, OTW71a, RS75, RSR74]. Algorithmic [CCR76]. Algorithms [Kau77]. Alignment [HLL78, JZ70]. All-Glass [BS75a]. All-Pass [RC76]. Allocation [Mar78]. Alloy [PB75]. Almost [Eis72, Kog76]. Almost-Coherent [Eis72]. Almost-Periodic [Kog76]. Along [Coh71, MSW77]. Aluminum [CLB75]. AM [SSP71]. Ambient [Daw72]. Amplifier [Apr75, BW71, Fen71, Goe74b, KS78, MG75, TG76, WC73]. Amplifiers [Goe74a, Per73a, Wal77]. Amplitude [Bul71, RC76, Rob71]. Analog [BHO71, DT71, Rai73, Rao78]. Analyses [DL78]. Analysis [Aha75, AR79, AS78, BKS71, BS70a, Blu72, Cha72, CM75a, CM75b, Cro78a, Eck78, ES77, Fal76a, FM75, HS71, HHJ71, Hef73, Hol71, KN78, Kru76, Lak79, Las76, Lev78, LRF78, LF72, Liu74, Mar77b, MS75a, Mea75, Mit74, Moo70, Mor70, Peh70, RSA77, San70a, SR71, SRH75, Sch71, Sle78, SS71b, SK74, Tho71a, Wal77, Whi75, Yeh70, Zuc74]. Analytic [Log78, Ric73c]. Analytical [BB79, Dar70, Nag71, Par78b]. Analyzers [EKN76]. Angle [Las76, Maz77, Pra76]. Angle-Modulation [Pra76]. Angles [Hen73b]. Angular [CM78]. Anisotropic [Kau74, Mar75a, Set71]. Anniversary [Dan75]. Announcement [CSL79]. Anomalous [DO79]. Anomaly [Wan70]. Antenna [CWE^{+}78, Dra74, Lee71, Moo70, Ohm74, Sem77, Sem79, SBR78a, Tur75]. Antennas [Chu71a, Dra78, Gan76, IZ70, Tho71a, Zuc70a]. Anticausal [Wal77]. Aperture [JC79, Mar77a]. Apertures [And75a]. Apollo [Cap72]. Apparatus [FRSD72, Las76, WSS78]. Appearing [Ano76a]. Application [BHKJ74, CGL71, CFMS78, CS78, FR79, HO77, Jag78, JR72, KS71, Lap76, Liu74, Lue72, Mor78c, Mos70, Ols71, Osb77, PB74, RW79, Ric70, Rum79, SW71a, Smi79, SW71b, Tho71c]. Applications [Aha72a, Aha78a, Aha78c, Ben75a, HS78a, Koo78b, LM70, LMT78, MG75, Mus77, Per71b, Per71c, Per73a]. Applied [Bau72, Ano70l, Lee69]. Approach [Aca79, ACG79, BG72b, Den76, Goo71, HY70, Koo78a, LW74, Mos70, Mue73, RSA77, Rif75]. Approximate [Hal75]. Approximating [Net73]. Approximation [Hey75, Jag70, Jag78, MS72, Mes78, Rob71]. Approximations [Bud73, Dar70, FR79]. Arbitrary [Mar77a, RM72]. Architecture [FP79, Ols75]. Area [Aca78, Seq72]. Areas [Fre78]. Arithmetic [Maz78b, Mit77]. Arrangement [BDJ79, BPW79, BGH79]. Arrangements [DG79, FG71]. Array [AH79, DG79, Mil78, Pir79]. Arrays [Abe78, KT73, Yeh70]. Arsenide [KS78, MWW^{+}79]. Articles [Ano70m, Ano70n]. Aspects [HP72]. Assignment [CR71, CR72, Kar71, Sch71]. Assist [Pea72]. Assisted [CRAC79]. Associated [DSS78, SDS78]. Asymmetric [Aun73, Buc70, Mar71b, Mar73b, Mar77c, RW74, SDS78]. Asymptotic [Mor70]. Asynchronous [Cha74, Maz71]. Atlanta [BSS78, BM78, DW78, Jac78, KW78, MSW78, MHSS78, MP78, RC78, SRMC78, SCD78, SBR78b]. Atmosphere [Hen73b]. Atmospheric [Wri71]. Attempts [DM78]. Attenuation [BR74, Coh71, Hen73b, Lin73, Lin75, Lin77a, Man74, Mar71a, MCC73, MC73, Osb77, Rut70, Sem71, Wri71]. August [ALM74]. Authors [Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77r, Ano77s, Ano77t, Ano77u, Ano78o, Ano78p, Ano78q, Ano78r, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v]. Autocorrelation [Jay75a]. Automated [AHPR79, BDRP79, CSL79, EKN76, YBBW79]. Automatic [Alb78, Blu77, BPM79, BCM^{+}74, BCKV74, BWW74, CGHW74, Cha70a, Cha74, CCWW74, HHM^{+}74, MRR79, PMAB79, Ros76, RS79]. Autonomous [CLT70, Mor70]. Auxiliary [BBJM^{+}77, KS71]. Available [Che76]. Avalanche [MHSS78, Per71b, Per71c, Sch70]. Average [AS72]. Avoiding [Lun70]. Award [WP75]. Axial [CM77]. Axis [Mar70a]. Axisymmetric [MC74].

B.S.T.J. [Aha72b, ATS70, Ben71a, Ber79, BBL72, Bob74, BS70b, BG72b, Che76, CT79, Chu79a, CS71, Daw72, Fen70, FMOT74, Hen73b, HS72b, HL73, Hol73, KMM73, KS70, Kur70, Laa71, Mes72, ML79, MCC73, MC73, Pen70, Per71a, Ros72, SP77, SS74, Sem73, Sem79, Seq72, Sta74, TBCR72, WKS^{+}72, WHF77, Wil70a, WN72, Zuc70a]. Background [You79]. Balance [Gue75]. Balanced [Bel78, Chu78, HL77, HL78, Len77a, Mil72]. Balanced-Pair [Mil72]. Balancing [BN71]. Band [AI74, Ber70b, CS77a, Cro78a, Per71a, PK70, Ric73a, Wyn76, Cro77]. Band-Limited [Wyn76]. Bandlimited [Jag70]. Bandpass [Log77]. Bands [Tur75, CWF76]. Bandwidth [Ber70a, Hub73, LS71, Lun71, Pra70, Wyn78]. Bandwidths [RB70]. Bank [BG72a, BS72b, CT72, DMM72, FDDM75, HP72, MT72, VG72]. Banks [SR71, SRH75].

Barrier [Dra72b]. Base [CGHW74, Lee71, LMT78]. Base-Station [Lee71]. Baseband [Bro71c, Dut76, Per73b, RB70]. Based [Ben71a, CB72, Che76, Fan78, Gea79, MS79a, WS70]. Basic [CCR^{+}77, Fuk79, Gey76, Ken70]. Batch [Kuc73a]. Batteries [KFBL70]. Battery [Ano71k, BW70a, BLV70, CFB70, Cus70, DS70, Gar70, HRH70, KFBL70, Lue70b, Mil70, OBW70, SSV70]. Beacon [ACH^{+}78, ACG78, Cox78, DZ78, Mul78, Rus78]. Beam [Aca77, ALW79, DG79, Gan76, Glo70, Kai70, Mar70b, Ohm74, RY77, SRPR70, Sem77]. Beams [MM71a, May71]. Beamwidth [Pir79]. Beat [Hub71]. Bed [DPZ79]. Beginning [RSR74]. Behavior [Aha79, Bra71, Con72b, DM78, Ken70, Lin71, Lin73, Mil70, RS74, SR74]. Behaviors [LC70]. Bell [HN78, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77r, Ano77s, Ano77t, Ano77u, Ano78a, Ano78o, Ano78p, Ano78q, Ano78r, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v, Cox78, Dan75, Eck78, Kes71, KFBL70, OBW70, Pfe79, SSV70]. BELLPAC [HL79]. Bending [Arn74e, BY78, Mar71b]. Bends [Mar73d]. Benes [Hwa71]. Bent [Mar74b]. Between [And75a, Arn74c, BM71, CM75b, Chu71a, Dra77a, GH70, Jay74, Maz77, RKHD74, Som73]. Biagetti [Ano71k]. Bias [Alb78, Bra73, TKN74]. Bias-Circuit [Bra73]. Bias-Temperature-Stess [TKN74]. Biases [Des73]. Binary [Aha78b, Fer72, Gre74, Llo77, Mac70, Obe70]. Bipolar [Blu72, Gum70, GP70, MCKP72, MM75, Per72]. Bipolar-IGFET [MCKP72, MM75]. Biquad [FL79]. Bit [ACG79, Aha78a, Can74, Cro77, HS75b, Jay70, Jay74, Jay77, Mus77, PMN78, RJ72, TNMC79]. Bit-Rate [PMN78]. Blazed [Mar76a]. Blinders [Tho71a]. Block [Aha72c, Cha70b, Chu78, Gre74, Kro72, Pie72, SJS79]. Blocking [Ber79, CH77, CH78b, CH79, Jag75]. Blocks [FL79]. Blooming [Seq72]. Boards [Rai79]. Bonding [Cus70]. Bound [Chi70, Liu74, Maz75b, Wyn75a]. Boundary [Blu78, Lew76, Sch72]. Bounds [Pra71, YH71]. Breakdown [AS72]. Briefs [Aha72b, ATS70, Ben71a, Ber79, BBL72, Bob74, BS70b, BG72b, Che76, CT79, Chu79a, CS71, Daw72, Fen70, FMOT74, Hen73b, HS72b, HL73, Hol73, KMM73, KS70, Kur70, Laa71, Mes72, ML79, MCC73, MC73, Pen70, Per71a, Ros72, SP77, SS74, Sem73, Sem79, Seq72, Sta74, TBCR72, Wag70, WKS^{+}72, WHF77, Wil70a, WN72, Zuc70a]. Broadband [Dra77a]. Broadening [Arn75a]. BSTJ [Ano70k, Ano70l, Ano71k, Ano75k, Ano78n, Ano79l, CH79, NK73a, Ano72k, Wag70]. Bubble [Aha72a, Aha72d, BBL72, BBB^{+}79, BN73, FM71, TBCR72]. Bubbles [DeB72, Kur70, Ros72, WN72]. Buffer [Has72, MCKP72, MM78a]. Buffer-Driver [MCKP72]. Buffering [FGM78, Lim72]. Building [Eck78, FL79, SW71b]. Buildings [Pfe79]. Bump [Cho72]. Burial [Beb72, Con72a, Con72b, Dah72, DeC72, Kle72, Kwe72, Min72, Ven72]. Buried [Ken73, MSW74, WKS^{+}72]. Buried-Channel [Ken73, MSW74]. Burnout [Bra73]. Burst [BW70b, Cha73, GMS73, Jay75a, Peh70, SGM74]. Burst-Error [Jay75a]. Burst-Trapping [Peh70]. Bust [Ton70]. Bust-Trapping [Ton70]. Busy [Eis77]. Busyness [CB72]. Butt [CMG73].

C [Ano70l, AI74, JR78, RJLK78]. C-Band [AI74]. Cable [AOS70, AHK^{+}78, BB78, Bre70, BEST78, BY78, BCD^{+}78, BSS78, BHO71, CGS^{+}78, Cha70b, Cho72, CDD^{+}78, DeC72, DO79, EFG78, ES77, Fos71, FYMS78, HLL78, LS70, Mil75a, MAGS78, Sch78, Sti78, Wah77]. Cables [CR76, Len77a, Len77b, Len77c, Mil76a, SBGC75]. Calculating [Lin75]. Calculation [Hal75, Lap76, Len77b, Len77c, Lin77a, MM78a]. Calculations [Ben75b, GH70, Len77a, MC73]. Call [CCR^{+}77, DM78, FR73a, KN72]. Call-Carrying [FR73a]. Call-Congestion [KN72]. Calling [Bri71]. Calls [Ben70]. Camera [AEH^{+}70]. Cameras [PS70, Raw70]. Camp [Wol72]. Camp-On [Wol72]. Cancellation [Chu77, FM79, Mue79]. Cancellation/AGC [FM79]. Canceller [RT71, Tho71b]. Cancellers [Tho71c]. Canister [SBR78a]. Cannone [Ano71k]. Capacitance [Lee75]. Capacitances [Mil72]. Capacitor [FL79, Lak79]. Capacity [BW74, CR71, Ebe70, FR73a, Fri78, HN76, MF70, Nea72, Wal70, Wyn76]. Carbon [Mea75]. Carlo [BKS71, Ols71]. Carried [HH73]. Carrier [Ano74k, Cha70a, Fal76a, Fal76b, Ho71, HS72b, MNS78, Per72, Rao78]. Carriers [Hub73, MS73b, MS73a]. Carry [Nea72]. Carrying [FR73a, Rai76]. Cascade [CR73a, CR73b]. Cascaded [Hal70]. Case [BW74, Dic75, O’N71, Sle78]. Cases [DSS78, SDS78]. Cassegrain [IZ70, Zuc70a]. Casting [Ven72]. Caused [Mar71d, Mar73e, VP79]. Caustic [DSS78, SDS78]. Caustics [SDMS77]. CCD [Kra71, KWP72]. Cell [EFW79, KFBL70]. Cell-Site [EFW79]. Cells [KSBL74, TKN74]. Cellular [DPZ79, Mac79]. Centennial [Ano76h]. Center [BJO78, CK78, MB71, MP78]. Center-Clipping [MB71]. Centimeter [Chu74]. Central [BS71, Hag75, Ols75]. Centralized [TAB74]. CENTRAN [Dic75]. Ceramic [NS72]. Ceramics [MMF70]. Certain [Aha79]. Change [Aha71, But71, TKN74, Van75a]. Channal [BG72a]. Channel [Aca78, Aha78b, Bab72, BS72b, BKT78, BW74, CH78a, CT72, CMR78, CR71, CR72, CFMS78, Cro78a, Cro78b, CS78, DR78, DMM72, Ebe70, FM73, Fal78, FDDM75, GK76, Has72, HP72, KCM^{+}78, Ken73, KS71, Lit78, Lun71, MT72, MM75, MSW74, MW78b, Mot77, Nol75b, RM78, Sca79, Sna78, SOP73, VG72, WKS^{+}72, Wyn71, Wyn75b, Wyn76]. Channelization [HNW77]. Channels [Bab73, Cho72, Fer72, Hub73, Jay75a, MG78, MF70, SW73, Wal70]. Characteristics [Ame72, Avi71, BDJ79, BDRP79, Car71, Dix76, Dra77a, Man78, Mea75, Sch70, SOP73]. Characterization [Ben73, Bra70, Log71, Met79, OBW70, San79, Tho77b, Tho77a]. Charge [ATS70, Ame72, BS72a, BT73, BS70b, Gum70, GP70, JB71, KSBL74, Ken73, Len77c, LM76, MS72, MS73b, MS73a, MSW74, MS75a, Per72, Seq72, Seq74, SS71b, Tho73a, Tho73b, TKN74, Tho74a, TKK73, WKS^{+}72]. Charge-Control [Per72]. Charge-Coupled [Ame72, Ken73, MS73a, MSW74, MS75a, Seq74]. Charge-Coupled-Device [SS71b]. Charge-Storage [KSBL74, TKN74]. Charge-Transfer [BS72a, Tho74a]. Charged [WM70]. Chebyshev [Dar70]. Checkout [Yam72b]. Chemical [DW78]. Chicago [SRMC78]. Chip [CT79, Mat70]. Chips [Mil78, Sch78]. Chloride [HRH70]. Chords [IF72]. Circuit [Alb78, BKS71, Bra73, But71, CK71, CES74, DC71, Fra78, HY71a, HN78, Kar71, Ken70, Log71, Mat70, O’N71, Ols71, SWD71, Yam72b]. Circuits [Aha71, Aha72a, Aha72b, Aha72d, BBB^{+}79, Bra71, BCM^{+}74, CK71, Cha74, CDH70, Die71, GS73b, Lak79, Maz78a, Ric70, Sch74, SWD71, Wad78a, Wad78b, Wag71]. Circular [Avi71, Car71, CD73, Moo70, Sem73, Sem74]. Circulation [Owe72]. Clad [Glo73]. Cladding [Mar71a]. Class [Aha71, CH78b, CH79, HHJ71, LM79, Net73, OTW71a, OTW71b, Per71c]. Clear [Hen73b]. Climatological [Che76]. Clipping [Bra72, MB71]. Clock [Aha78c]. Clocks [Mor71]. Clos’ [Ben75c]. Closing [Ben71a]. Clustering [RW79]. Clusters [CFHM71]. CMOS [Wad78b]. Co [PB75]. Coated [Mat70]. Coatings [Beb72, CM75a, Pre76]. Coaxial [Ano74k, Cho72, GRW71]. Coaxial-Carrier [Ano74k]. Code [Slo76]. Codec [Aha78c, LM70, MS79a]. Codecs [Bob73, DC78, Kan70, Can74]. Coded [Bos72, GMN76, Gre74, RSF71, RSR74]. Coder [CHM73, Cro78a, HS75b, LP71]. Coders [Cro77, Cro78b, GSC^{+}79, Has72, Jay75b, Laa70, Sca79, TNMC79, Wit79]. Codes [Aha72c, Chi70, GMS74, Mac70, MSG72, Ton70]. Coding [Aca78, AS70, Bob74, CPS71, CWF76, CS77a, CJF73, DMM72, DC79, GW74, Has75, Jay77, JC79, Lim72, Lim73, LPW74, MNP77, MBN79, NMB76, NPM77, NMB77, NR79, NS79, Neu71, PL71, Pea72, RB74, SW73, SJS79, WHF77]. Coefficient [GGO79]. Coefficients [Mar73b]. Coherent [Eis72, RB74]. Coin [AHPR79, BDRP79, CSL79, YBBW79]. Combinatorial [Mor78c, Sch71]. Combined [SSP71]. Combining [LPW74, Mue79, Nea71a]. Come [Hef73]. COMET [CH74]. Commissioning [HLL78]. Common [BKT78, CMR78, CFMS78, CS78, DR78, KCM^{+}78, Lit78, MW78b, RM78, Sna78]. Communication [Aca77, BBJM^{+}77, BDS78, Cha70a, CH72, Chu71b, Chu73, Cro77, Doa78, ES77, Fal76b, FGS75, Hay74, Hen73a, MS76, Mil76a, ML79, MS75b, Per71c, Per73a, Per73b, Per73c, Per73d, Per74, RSCG77, Rai73]. Communications [Cha71, CR71, CGRS77, Gol73, GH70, Maz79b, SRMC78]. Companded [Aha78c]. Companding [Kan70]. Compandors [Kan70]. Companion [Tow75]. Comparative [Nol75a]. Comparator [Obe70]. Comparison [BT73, CBK79, DC78, Dra77a, Maz78a, Per76, TNMC79]. Comparisons [RKHD74, Wil71]. Compatibility [BLV70]. Compensated [NR79, NS79]. Component [Kar71]. Components [FYMS78, GOS74, GS71b, Len77c, WMLT70]. Compression [MNW78]. Compromise [BFG73]. Computation [FM71, Hil71, IZ70, MP71, San70b, Zuc70a]. Computations [HO70]. Computed [Set70]. Computer [Ame72, AMS70, FRSD72, GRW70, Hay74, HY71a, HN78, Kno77, Kuo75, LS71, Mea75, O’N71, RSF71, WP71, Yam72b]. Computer-Aided [HN78]. Computer-Controlled [HY71a]. Computer-Generated [FRSD72]. Computer-Sided [Mea75]. Computers [Cok72, Con75]. Computing [CHK^{+}74, Rai73]. COMSTAR [ACH^{+}78, ACG78, CWE^{+}78, Cox78, DZ78, Mul78]. Concatenation [RSF71]. Concentrated [Lew76]. Concentrating [Mor78c]. Concentration [GLW74, OP74]. Concentration-Fluctuation [OP74]. Concentrators [Chu79b]. Concept [ATS70, Mac79]. Concerning [San78a]. Conditional [Pea72]. Conditions [Dra72a, Sca79]. Conductive [Mat70]. Conductor [Rai76]. Conductors [Mor79]. Confirmation [HY71a]. Confocal [And75a]. Congestion [KN72]. Conjecture [Ben76, Ber79]. Connect [Fre78]. Connect-Through [Fre78]. Connecting [Ben70, Ben75a, Ben75b, Hwa76, HL77, Kes71, SBGC75]. Connection [BKKM71, DT71, FH71]. Connections [CHS76, CGRS77, Cro78a, DMNT75, GSC^{+}79, RSCG77, Sul71]. Connector [CR76, Mil75a, Sch78]. Connectors [RC78]. Consecutive [Jay74]. Considerations [Aca77, Bau72, BS72a, BPW79, BHO71, Gey76, HW78, JP71, MSW74, Pra71, SW71a, SW71b, Tho71c]. Consisting [Dra74]. Constant [BM71, Bur71, Kuc73a, Pir79]. Constellation [FGW73]. Constituents [CDF^{+}78]. Constrained [Chi70]. Constraints [RC76]. Constriction [Rai76]. Construction [HL77, HL78]. Contact [AOR76, GBN^{+}76, PK70]. Containing [CHM73, FM75, San70b]. Content [MM78a]. Continuous [Bra72, Stu76b]. Contract [WP75]. Contrast [Nag71]. Contributors [Ano70a, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano71a, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75m, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i]. Contributors [Ano78b, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k]. Control [Alb78, BCKM70, BDN^{+}77, CFG^{+}77, DL76, DCRZ70, DKP^{+}70, FP79, Gaw75, Gop71, GRW70, Gum70, GP70, Hag75, MP78, Ols75, OTW71a, Per72, Smi79, Sto76, Van75a, Wal79, WSS78]. Controllability [CH74, RM72]. Controlled [AMS70, HY71a, NS72, PB74, Wag70]. Convergence [Maz79a]. Converging [MS75b, Mue75]. Conversational [Bra71].

Conversion [BS72b, CG70, Dra77b, Sch71]. Converter [Ric73d]. Converters [Net77]. Converting [NP78]. Convexity [Mes72]. Coordinate [AMS70]. Copper [MM79a]. Cord [IF77]. Cords [FIS75]. Core [MG73, MP75, Mil76b]. Correcting [Aha72c, Slo76]. Correction [BW70b, Peh70, SGM79]. Correlated [GM77, MM78a, Nea71a, SW73]. Correlation [Jay74, Ano70l, Lee69]. Corrugated [Dra74, Dra77a, Dra77b]. Cost [Gib78, LN72, WP75]. Cost-Plus-Award-Fee [WP75]. Costs [DL78, Koo78a]. Count [Des76]. Counterexample [Ber79]. Coupled [ATS70, Ame72, Ano78n, BS70b, CS77b, IS73, Ken73, Mar74a, Mar72a, Mar72b, Mar73a, Mar75a, MS72, MS73b, MS73a, MSW74, MS75a, MM72b, Per75, Seq72, Seq74, SS71b, TKK73, WKS^{+}72]. Coupled-Mode [Mar75a]. Coupled-Power [Per75]. Coupler [Oga77, RS76]. Coupler-Pair [RS76]. Couplers [KS70, RS78]. Coupling [Arn74c, Arn74d, Arn74e, Arn75b, Coh72, CMG73, GH70, Mar71c, Mar73b, Mar74e, MP75, MM72b, MAGS78, RW74]. Couplings [LS70]. Cover [DS70]. Coverage [Aca78, Wad78a]. Covers [HRH70]. Crawford [ACH^{+}78, ACG78, CWE^{+}78, Rus78]. Criteria [MS79c, Mit77]. Criterion [Jay75a, Rei72]. Cross [Ano78n, Chu74, Gan76, Mar74c, MSW77]. Cross-Polarization [Chu74]. Cross-Sectional [MSW77]. Crossbar [CFMS78, Gue75]. Crossings [Bra72, Log77]. Crosslinked [Beb72]. Crosspoints [Ben71a]. Crosstalk [Arn75b, BM78, BHO71, CM75b, Fos71, Glo70, Gol73, IS73, Lap76, Len72, Mar71d, Mil75c, Par78a, Sen70]. Crystal [Gea79]. CTS [Rus78]. Cubic [WS70]. Current [BBB^{+}79, Dix76, PB74, Rai76, Wil71].

Current-Access [BBB^{+}79]. Current-Carrying [Rai76]. Curvature [CM77]. Curved [Mar70a]. Customer [BJO78, Bre71, DM78, Man78, Par78a]. Cycles [Aha79, LM79]. Cyclic [Aha72c, Coo70, Hal75, Kue79, MS75b]. Cyclic-Type [Hal75]. Cylinders [MSW77]. Cylindrical [Dra74, TBCR72, Thi71, TBTG71].

D [Ano71k, Net77]. D2 [BG72a, BS72b, CT72, DMM72, HP72, MT72, VG72]. D3 [FDDM75]. Data [AH79, AFG79, Ano75n, AFS79, Avi71, BKKM71, BH71, BKL75, BBCC75, BN73, BPS75, BFR75, Bud73, CGHW74, CH72, Che76, CBK79, Con75, CLT70, Cro75, DJ75, Dic75, DM75, DG75, Dut76, Fal78, FG78, FM79, FR75, FH71, Fre75, Gaw75, GS73a, GHM73, GH75, GHRS77, Hag75, HS75a, HS71, HS72a, Ho71, HS72b, HJ75, ILR75, Kau77, Kro72, Kuo75, Lee73, Lim72, LW74, MMM75, MS76, Mes75, MS75b, MW75, Nic75, Ols75, Osb77, Phi75, Pie72, Rif79, Rif75, Rum79, SZ75, SSP71, SK75, Van75a, Van75b, Wil71, WP75, Wyn78]. Data-Processing [Ano75n, BFR75, Con75, Cro75, DJ75, Dic75, DM75, DG75, Fre75, Gaw75, Hag75, HS75a, HJ75, Kuo75, Mes75, MW75, Nic75, Ols75, Phi75, Rif75, Van75a, Van75b, WP75]. Data/Trunk [GHRS77]. dc [Chi70, Gre74, WC73, Wil70b]. dc-Constrained [Chi70]. DC-to-2.3-GHz [WC73]. Debugging [Phi75]. Deception [GMS74]. Decision [FF73, Mes73a, Mue79, Sal74, Sal73]. Decision-Feedback [Mes73a]. Decoders [Aha72c]. Decomposition [Lue70a]. Deformation [Sch72]. Degenerate [Mar71c, RW74]. Delay [Ano75k, Bra71, Gue75, HS72a, Lee75, Mil74, Sch72]. Delays [Col73, Hal75]. Delta [Boy76, Cum75, Ger72, Goo71, Gre73, Jay70, JR71, Jay74, LM70, Laa70, Sle72b, GG73]. Delta-Modulation [Laa70]. Demodulator [Gla71b]. Demountable [RC78]. Density [Bos72, Nea71b, Slo76]. Dependence [AC78, CAK76, Hol71, Lin76a]. Dependent [Per75]. Depletion [MM75, MS72]. Depletion-Load [MM75]. Depolarization [AC78, Coh71, Cox75, Lin77b, Sem74]. Deposition [DW78, JMM78]. Derivation [Mar72a, Maz75b, Yam72a]. Derivations [Per75]. Derivative [Dix76]. Descent [Els77]. Description [BDJ79, BDRP79, CSW74, DR78, Gol73, KH74, Slo76]. Design [Aha72c, AHK^{+}78, Ano71k, Apr75, BKS71, BC70, BBW78, BB79, BCE^{+}77, BR74, BCDT77, BFG73, BS75b, But71, BKY70, CFB70, CD79, CK71, CHJZ70, Cro77, DC71, Dic75, DG75, Fra78, GGL^{+}77, HY71a, HHJ71, HW78, HRC73, HGLM77, HN78, ILR75, Kar71, KH74, KL73, Kle76, KU76, KFBL70, KLRS70, Las76, LN72, LF72, Log71, MSW74, MAGS78, Mos70, MNP77, NS76, O’N71, Ols71, Per73c, Per73d, Pir79, PS70, Rab72, RC76, Raw70, Rut70, Sal74, SR71, SWD71, SBR78a, Sti78, Sto76, Tho71a, Tow75, VG72, Whi75, WA74, Wyn78, Yam72b]. Designed [MG78]. Designing [CCR76, CB72]. Designs [ALW79, But71, Chu78]. Detect [GMS74]. Detecting [CRSB77]. Detection [Den76, Eis72, FGS75, Hub71, Kau77, MS76, PB74, Per71b, RB74, SG77, Wyn75a]. Detector [Mes73b, SBR78b]. Detectors [KG76, Per71c]. Determination [Fuk79, GS70, Kau74]. Determining [Aha79, Mit77, RS75]. Deterministic [MM72a]. Developed [Sne71]. Development [AOR76, BCKV74, BHKJ74, CMR78, EFG78, Ele77, Fre75, JL78, SVW79, Tho77b, Tho77a, Tsi79, WP75]. Developmental [Huf79]. Developments [SH79]. Device [AEH^{+}70, ATS70, Ame72, AMS70, Ble70, BP70, Bro70, CHJZ70, DCRZ70, Fuk79, GRW70, Her70, HP70, JZ70, Ker70, KLRS70, MS73b, MS73a, MSW74, MS75a, Poo70, PS70, Raw70, SRPR70, Seq74, SS71b, Thi71, WKS^{+}72]. Devices [BT73, BBL72, BS70b, DPR77, FYMS78, JB71, Ken73, Kno77, MS72, MMF70, Seq72, Tho73a, Tho73b, Tho74a, TKK73, Wah77]. Devises [Gea79]. Diagnosis [OTW71b]. Diagnostic [HS75a, Pek78]. Dial [CR78, Gue75]. Dial-Tone [Gue75]. Dialing [DM78, Spa76]. Dialogue [PR73]. Diameters [Mil76b]. Dictionary [Rif75]. Dielectric [Ano70k, CD71, Car71, CM73, CD73, KSBL74, LM76, Mar70a, Mar70d, Mar70c, Mar71c, Mar72d, Mar72e, Mar72g, Mar76a, Mar77c, Per71d, RW74, TKN74]. Dielectric-Lined [CM73, CD73]. Dielectrics [Sal75]. Dietrich [Ano70k]. Difference [OP74, Per74]. Differences [CFHM71]. Differential [CB72, Can74, CHM73, CJF73, Hel79, Jay75b, MSG78, MCC73, MC73, San79, Sha70]. Differentially [GHM73]. Differentiators [SR74]. Differing [Wal70]. Diffusion [Hey75, SS71b]. Digit [SR75]. Digit-Recognition [SR75]. Digital [ABLR77, BKL75, BBCC75, Bos72, Bri78, Bro71a, BS75b, BPS75, CCR76, CR73a, CR73b, Cha70a, Cha71, Cha72, Cha74, CT72, CRS75, CWF76, DC78, Fal76b, FG78, FR75, Fla79, FGS75, Gla71b, Gla71a, Goo71, GK76, GRW71, HRC73, Hil71, HY71b, ILR75, Jac70, Kan70, KG76, LM79, Lee73, Lin77b, LR79, MMM75, MNS78, Maz78b, MM71b, Mit77, MG78, Net77, O’N71, Per73a, Per73b, Per73c, Per73d, Per74, Pra70, Pra71, PR74, Pra76, Rab72, RS74, RKHD74, RRS70, Rif75, RP75, RB74, SZ75, SR71, SR74, SRH75, SRB79, Shn70, SK75, Tho73a, Wad78a, Wil72, YH71]. Digitally [GW79]. Dimensional [Blu72, CPS71, CR72, Fal76b, FGW73, Mor79]. Dimensionality [Len72]. Dimensions [Gol74]. Diode [Sch70, Sch74]. Diodes [Bra70, San70b, Sch70, WMLT70]. Dioxide [CLB75]. Diplexed [KS71]. Diplexers [AP75]. Diplexing [CGL75]. Direct [Aha75, DM78, FGS75, FHH^{+}79, MS76, Spa76]. Direct-Distance-Dialing [DM78]. Directional [RS78]. Directory [RS79]. Disconnected [Wag71]. Discrete [GG70, GM77, JB71, Mor78c, RB76, Sle72a, Sle78, Stu76a, RV70]. Discrete-Charge-Transfer [JB71]. Discrete-Time [GG70, GM77, Mor78c, RB76]. Discrimination [Dra78]. Disk [AAS70]. Dispersion [CDF^{+}78, Hen73a, JB71, Mar77a, Mar74e, Per71d, SOP73]. Displacement [IF77]. Display [MMF70]. Displays [Kno77]. Distance [BW70b, Doa78, DM78, Spa76]. Distinguishing [Stu76a, Stu76b]. Distortion [Ano75k, DO79, Lue70a, Mil74, Rai73, Ric73a, Ric73d, Sch72]. Distortions [MP75]. Distributed [Bel78, Rei72]. Distributing [HMS72, Sch71]. Distribution [BKL75, Bur71, GM73, Ham70, Ken73, Lin76a, Mar72e, Ric73b, Sti78]. Distributions [Che76, Lin75, Lin76b, Lin78]. Distributor [CLT70]. Disturbances [ALM74]. Dither [JR72, RJ72]. Dithered [NMB77]. Diversity [Bab73, Bar70, DZ78, Jay75a, Lee71, Osb71, Vig70, Vig75, VP79, Wil70a]. Division [AD78, LW74, Maz71]. Document [KLO78]. Domain [Aha72b, Aha72c, Gea79, Gra70a, Mue73, Per77, SN79, TBCR72, Zuc74]. Domains [Thi71, TBTG71]. Dominant [Mar70b, Mar70d]. Dopants [KSBL74]. Doped [Aha75, KWP72, PSMO75]. Double [CBK79, CLB75, Dix76]. Double-Heterostructure [Dix76]. Double-Insulator [CLB75]. Downconverter [SS74]. DPCM [Jay77, MS79a, MS79b, SGM79, SJS79]. DPCM-Encoded [SJS79]. Drawdown [SDMS77]. Drawing [Gey76, Gra72, MP78, Smi79]. Driver [MCKP72]. Dual [Cha72, KSBL74, Mot77, Sem71, TKN74]. Dual-Dielectric [KSBL74, TKN74]. Dual-Polarized [Mot77]. Due [Bis71a, CM78, Doa78, GH70, LR79, Sch72, Set71, WM70]. Duplex [FM79, MB71]. Duration [Hub71, Vig71]. During [DM78, HO70, Lin77b, Mot77, SOP73]. Dynamic [BS72a, BN73, CR71, CR72, Cro78a, FIS75, Gey76, Jac70].

Each [Whi75]. EADAS [Kau77]. EADAS/ICUR [Kau77]. Early [BGK70, CGWN77]. Earth [Moo70, Ohm74, Pen70, Rus78, Wri71]. Earthquake [LF72]. Echo [Bra71, CHS76, DMNT75, FM79, MB71, Mue79, RT71, Spa76, Tho71b, Tho71c]. Echo-Free [Bra71]. Echoes [Doa78]. Eclipses [Lun70]. Economic [Koo78b, Sti78]. Edge [CB72, LW70]. Effect [And75a, AD79, Beb72, CM75a, CF79, CMG73, Dah72, Daw72, DeC72, Gue75, Ho71, Kle72, KS78, Kwe72, Lee75, Min72, Ros72, Sem73, Ven72, Wag71]. Effective [Kuo75]. Effects [AS72, Bra70, Bra71, CHS76, Chu79a, Cox75, DO79, Hil77, Hol71, Lev78, MRR79, Mil75c, Nol75b, Per72, PB73, RJ72, RSSF78, Sha75, Wil72]. Efficiency [Chi70, CMG73, Sch70]. Efficient [AD78, KL70, Ric73c, Sam75]. Eight [Con72b]. EL2 [BB79]. Elastic [BY78, MSW77]. Electret [BB79]. Electric [Car71, MP78]. Electrical [GHWY77, GS73b, Kar71, Kwe72, Mat70]. Electrochemical [BLV70, OBW70]. Electrochemically [Wag70]. Electrode [LW70]. Electrodeposits [AD79]. Electrolyte [Lew76]. Electromagnetic [MC74, Mor79]. Electron [Bro70, Has78, SRPR70]. Electron-Sensitive [Bro70]. Electronic [Aun73, DL76, Gea79, HL79, SS78, Sto76]. Electronics [And78, ACH^{+}78, AB78, HS78a, HW78, MNS78, Mor78b, Rao78, SS78]. Element [Ano79l, Kau74, NMW79, Seq74, TKK73]. Elementary [Kur70, Tho71a]. Elevation [Hen73b]. Elimination [Bra73, Gra72]. Elliptic [Sch72]. Embedding [WC73]. Embodiment [Aha72c]. Emphasis [Aun73]. Emphasizing [Lue72]. Empirical [Lin77a]. Employing [Cha72, FF73]. Encoded [ACG79, Aha78b, MS79b, SJS79]. Encoders [Aha72c, GG73]. Encoding [Aha78a, Buc70, CS77a, Lim73, MM71b, Nol75a, Nol75b]. Encryption [KJ77, SJ76]. End [Bis71a, BM78, CM78, GSBC73, RSR74]. End-to-End [BM78]. Endpoints [RS75]. Ends [CGBS75]. Energy [CM75a, TBTG71]. Engineered [HN76]. Engineering [AOS70, AGW77, Cap72, CH74, Eis77, MP78, Osb77, Vig75]. Ensemble [Wal70]. Entity [BB78]. Entrance [CKS79]. Entropy [Has75]. Enumeration [OTW71b, WS70]. Environment [LF72, Pek78, RT71, Rov78, Tho71b, Yeh70]. Environmental [KU76, Sha75]. Epitaxial [BBL72, MHSS78]. Equal [Ric73b]. Equalization [BBJST77, Cha70a, Cho72, Cho74, FF73, Fal76a, Fal76b, Fal78, GHM73, Hen73a, HO70, Ho71, Mes73a, MS75b, Mue79, Per73b, Sal74, Sal73, Shn70]. Equalizer [AHK^{+}78, BFG73, BKY70, Cha71, Maz75c, Maz79a]. Equalizers [Cha70b, Cho72, Cho74, GW79, Mue75, O’N71]. Equalizing [Per76]. Equation [Blu78]. Equations [Hel79, Mar72a, MM72b, Mor79, Per75, San79, Wil70b]. Equipment [BCD^{+}77, BCD^{+}78, BDS78, Fis79, MT72, MS79c, Pfe79, SW71b]. Equivalent [Hol73, Lak79, Nea72]. Erlang [Jag74, Mes72]. Errata [Ano78n, Ano79l]. Erratum [Ano70k, Ano70l, Ano71k, Ano75k, CH79, NK73a]. Error [Aha72c, Bal76, Bri78, Cho74, FF73, Gaw75, Hil71, HY70, HY71b, Jay75a, MS79b, Peh70, Pra71, SR74, Slo76, Tho73a, Wyn75a, YH71]. Error-Correcting [Slo76]. Errors [And75a, Bri78, Can74, Cox75, Cro78a, FHH^{+}79, NK73a, NK73b, Nol75b, SG77, SGM79]. ESS [ABLR77, AC70, AAS70, Ano70m, AOR76, BGK70, BCD^{+}77, BHP71, CFP70, CCR^{+}77, CGWN77, CLT70, CS78, DL76, EG70, GAW^{+}76, GBN^{+}76, GHHJ77, GHRS77, HW76, HJJ^{+}77, KLS76, KNS^{+}70, Kle76, KU76, LO77, MOR^{+}77, MRY77, Pek78, PSW70, RT77, Spe77]. Estimate [Mes78]. Estimated [Che71]. Estimating [Che76, Nea72]. Estimation [AR79, FM73, Mor78a, RB76, SN79, Tho77b, Tho77a]. Estimators [Den76]. Etched [Mil78]. Evaluating [HY70]. Evaluation [ACTW77, Bal76, Bri71, BM78, CHM^{+}77, CRSB77, DVW79, GMN76, GBN^{+}76, HLL78, JZ70, Koo78b, LRF78, MOR^{+}77, MSG78, RJ72, RSA77, Ric73c, Ric75, Ros76, Sca79, Tow75]. Evaluations [CB72]. Eventual [Sha70]. Evolution [Pfe79]. Exact [Mar76a, MM72a]. Excess [Wyn78]. Exchange [PL71]. Excitation [Mar70b, MM71a, Mar75b]. Excited [CM78, LC70]. Exerciser [DM75]. Existing [Ben70]. Expand [Pra70]. Expansion [Fri78]. Experience [BGK70, CGWN77]. Experiment [ACH^{+}78, ACG78, BR74, CWE^{+}78, Cox78, DDH75, DZ78, Dra77a, Fre75, Hol75, MSW78, Mul78, Sei71, Wil70a, BSS78, BM78, DW78, Jac78, KW78, MHSS78, MP78, RC78, SRMC78, SCD78, SBR78b]. Experimental [ATS70, CM73, CKK^{+}71, Cok72, Hay74, Kro72, Len77a, Mus77, Seq74, SBR78a, SH74]. Experiments [Cox75, Cox78, LS71, NPM77, WSS78]. Exposure [Beb72, Con72b, Dah72, DeC72, Kle72, Kwe72, Min72, Ven72]. extended [Ano70l, Lee69]. Extendible [Dic75].

Extension [Aha72b]. Extensions [Kai71, MNW78]. Extracting [Dut76]. Extreme [Lin76b, Lin78]. Extreme-Value [Lin76b].

F [Ano79l, CH79, PK70]. F. [CH79].

Fabrication [BS72a, BCDT77, MP78]. Face [GHL72]. Facilitate [Rif79]. Facilities [CFP70, LTV70, RSS79]. Facility [AS78, BS76]. Facsimile [Ano79l, MWW^{+}79, MNW78, MBN79, NMB76, NMW79]. Factors [YBBW79]. Fade [SOP73]. Faded [Bab73]. Fades [Vig70, Vig71]. Fading [Bab72, Bul71, Lin71, Lin77b, LR79, Mot77, Rum79, Rut71, VP79, Yeh70]. Failure [IL78]. Family [FL79, Hal75]. Far [GS75, IZ70, Zuc70a]. Far-Field [GS75]. Fast [Cha71, CH72, MM79a, MCKP72, Mue75, RRS70, Rob71]. Fast-Converging [Mue75]. Faster [Maz75a]. Faster-Than-Nyquist [Maz75a]. Fault [LTV70, OTW71b, Wad78a, Wad78b]. Faulty [Kau77]. Faulty-Trunk [Kau77]. FDM [Lun71]. FDM-FM [Lun71]. Fe [PB75]. Fe/Co/V [PB75]. Features [Pfe79]. Feder [Ano71k]. Fee [WP75]. Feed [Dra77a, Dra77b, Gol77, HM74, Sei71]. Feed-Forward [Sei71]. Feedback [Apr75, Bau72, Ben71b, Ebe70, FF73, Mes73a, Mue79, Rei72, Sal74, Sal73, Sze73, Wal77, WR73]. Feeder [Fri78, Mar78]. Ferrite [Owe72]. Ferroelectric [MMF70]. Ferroeletric [Gea79]. FETs [IL78]. Fiber [Arn74c, Arn74d, Arn74e, Arn74b, Arn75b, Bis71b, BS75a, BPM79, BSS78, BM78, CM75a, CRAC79, CGBS75, CMG73, Daw72, DMK78, DW78, FGS75, GSBC73, Glo75, Glo76, Hen73a, Jac78, KMM73, KW78, MSW78, Mar70b, Mar73c, Mar73d, Mar73e, Mar74f, MP75, Mar76d, Mar76e, Mar77b, MHSS78, Mil75a, Mil76a, Mil76b, Mil78, MM78b, ML79, Mus77, MP78, Oga77, Per73b, Per73c, Per73d, Per74, Per76, Per77, PMAB79, RC78, SKCM78, SP77, Sch78, SRMC78, SCD78, SBGC75, SBR78b, Smi79, Sta74, SH74, Wil77]. Fiber-Optic [FGS75, Mil78]. Fiber-Optic-Cable [Mil75a]. Fiberguide [Bal76]. Fibers [Arn74a, Arn75a, AD76, Bis71a, CM75b, CM78, Coh71, Coh72, CAK76, CDF^{+}78, Coo77, Dab75, DS73, DW78, ES77, Gar75, Gey76, Glo72, Glo73, GM73, KA74, MG73, Mar77a, Mar73a, Mar74c, Mar74d, Mar75b, Mar76b, Mar76c, Mar76d, Mar76e, Met79, Mil75b, OP74, PSMO75, Pre76, RKHD74, RSSF78, Sch72, Sha75, SDMS77, Som73]. Fidelity [Llo77]. Field [ACTW77, BGK70, CHM^{+}77, GS75, IZ70, KS78, LW70, LM76, Lit78, MS75a, OBW70, PB74, Pir79, Zuc70a]. Field-Aided [MS75a]. Fiftieth [Dan75]. Fifty [Ano72k]. Figures [Yam72a]. Filaments [PB74]. File [HHM^{+}74]. Film [KS70, Ram74]. Films [BBL72, May71, WN72]. Filter [Kog76, Mit77, SR71, SRH75]. Filtering [GGO79, Mue73, Per74]. Filters [CCR76, CR73a, CR73b, Fla79, FDDM75, GG70, GK76, Hal70, HRC73, Jac70, LM79, Maz78b, Net77, Rab72, Sal74, Shn70, Sne71, Sze73, Tow75, Wil72, Zuc74]. Final [FHH^{+}79, HLL78]. Finding [Slo72]. Fine [MMF70]. Fine-Grain [MMF70]. Finite [CR73a, CR73b, HH73, HRC73, Kau74, Rab72, Whi75]. Finite-Element [Kau74]. Finline [HL73]. FIR [CCR76, RS74, RKHD74, SR74, SRH75]. FIR/IIR [CCR76]. Fires [Eck78]. First [Hef73, Pen70]. First-Come [Hef73]. First-Served [Hef73]. Five [Che76]. Five-Minute [Che76]. Fixed [Jay77, NS76]. Fixed-Point [NS76]. Flat [Lew70]. Float [Mil70]. Flow [Glo72]. Fluctuation [OP74]. Fluctuations [Mar72b, Tho74b]. Fluid [Gey76]. Fluid-Dynamic [Gey76]. FM [ALS71, Lun71, Rai73, Ric73a, VP79]. Force [TBTG71]. Forced [SGM79, Wil72]. Forcing [Mes73a]. Forecast [FHH^{+}79].

Foreword [Ano74k, Bai71, Cro75, Gar70]. Form [DG79, Tow75]. Formant [RSF71]. Formant-Coded [RSF71]. Formed [Yam72a]. Formula [Mes72]. Formulas [GMS73, SGM74]. Formulation [Kan70]. Forney [Maz75b]. Forward [Ble70, Bre70, MPT78, Sei71]. Foundation [Par78b]. Four [TNMC79]. Fourier [RV70, Maz77, Sle78]. Fourth [Maz78a]. Fourth-Power [Maz78a]. Frame [CFHM71, CHM73, Gue75, HMS72, Has75, LPW74, Sch71]. Frame-to-Frame [CFHM71, CHM73, Has75, LPW74]. Frames [BB78]. Free [Bra71, Has78, Mit77]. Frequencies [MCC73, MC73, RV70, Set70, Set71]. Frequency [Aha75, Bab72, Bar71, Bar70, BNO74, Ben71b, CAK76, HS78a, Hol75, Hub71, Hub73, KS71, Lue72, MSW77, Par78a, Par78b, Ric73d, RP75, Rut71, Sch70, Sem71, Sne71, VP79]. Frequency-Diplexed [KS71]. Frequency-Diversity [VP79]. Frequency-Emphasizing [Lue72]. Frequency-Modulation [RP75]. Frequency-Response [CAK76]. Fresnel [Chu71a]. Frontiers [AI74]. Full [FM79, MB71]. Full-Duplex [FM79, MB71]. Function [JB71, Lue70a, Nea71b, Net73, Rut71, SKCM78, Ano70l, Lee69, Jag74]. Functional [San79, Sha70]. Functions [CT72, Jag70, Len72, Ric73c, San78a, Sle78]. Fundamental [BT73]. Fused [Dab75, KA74]. Fused-Silica [Dab75].

G [Ano71k, Hey75]. Ga [Dix76]. GaAlAs [CBK79, SCD78]. GaAs [CBK79, Coh72, Fuk79, IL78, TG76]. Gain [Dra72a, Koo78b, Lap76, Sch73]. Gain-Induced [Sch73]. Gallium [KS78, MWW^{+}79]. Gallium-Arsenide [MWW^{+}79]. Gaps [Kra71]. Garnet [BBL72, WN72]. Gas [Kai70]. Gaseous [Wan70]. Gate [Col73]. Gauge [Lin76a]. Gauging [FHT77]. Gaussian [BW74, Ebe70, FGW73, Gol74, Gre73, HY70, HY71b, HS72b, Mar70b, MF70, Nea71b, Wyn71, Wyn76]. General [Aha71, Lue72, Met79, Mos70, Per71c, TBTG71, WR73]. Generalization [CH78a, Gol74]. Generalized [Ano75k, Mil74, Yam72a]. Generalizers [Chu79b]. Generated [FRSD72, Lim72, Maz78a]. Generating [RRS70]. Generation [AI74, Cha74, Sch70]. Generator [BJL71, CHJZ70, DCRZ70, GK76, GRW70, JZ70, KLRS70, KG76, Poo70, SRPR70]. Generator-Detectors [KG76]. Generators [Bar71]. Geomagnetic [ALM74]. Geometric [Maz75b, Mes73a, Mes73b]. Geometrical [Pre76]. Geometry [Mor79]. Geostationary [MP71]. Germanium [PSMO75, Sch70]. Germanium-Doped [PSMO75]. GHz [Aca79, ACH^{+}78, ACG78, AC78, Bar72, Cox78, DDH75, EKN76, Hen73b, HGLM77, KS78, Lin77a, Osb71, Osb77, Pen70, Rus78, SS71a, Sem73, Sem74, Sem77, Sem79, SBR78a, TG76, Tur75, Vig71, WC73, Wri71]. GI [Hey75]. GI/G/1 [Hey75]. Gigahertz [MG75]. Gilbert [Cha73]. Glass [BS75a, Coh71, Daw72, DSS78, PB75, SDS78]. Glass-to-Metal [PB75]. Glossary [Ano75l, Ano77q]. Gold [AD79, MM79a]. Grade [Kle72, Kwe72]. Graded [CM78, Coo77, MG73, OP74]. Graded-Core [MG73]. Graded-Index [Coo77, OP74]. Gradient [Aha78a, Fal76a]. Grain [MMF70]. Granularity [JR71]. Graphs [Ber79, Chu75, CH78b, CH78a, KL70, CH79]. Grating [Mar77c]. Gratings [Mar76a]. Green [JB71]. Grid [Ano71k, AP75, CFB70, Chu77]. Grids [And75b]. Ground [Liu70]. Group [Ben75a, Ben75b, Ben75c, HH73, HN76, HL77, Mac70, Nea72, Wah77]. Group-Balanced [HL77].

Group-Theoretic [Ben75c]. Groups

[FM75, FHH^{+}79, RM72]. Growth [HMS72]. Guide [Mar70a, SKCM78]. Guided [HMG70]. Guiding [Sch72].

Hadamard [Ber70a, MNP77, NPM77]. Hadamard-Transform [MNP77]. Half [DeB72]. Half-Bubbles [DeB72]. Handling [CCR^{+}77]. Hard [BBL72, Ros72, WN72]. Hardware [BGH79, CSL79, EFW79, KCM^{+}78]. Harmonic [AI74, Bra70]. Haul [Bro71c, Che71, CGL71, KS71, MG75, SW71a, SW71b]. Heat [Aun73]. Heating [Aun73]. Heavy [Avi71, Hey75]. Helicoidal [Arn74b]. Heterodyne [Per71a]. Heterostructure [CBK79, Dix76]. Heuristic [KL70, KL73]. Hexagonal [WS70]. Hexagonal-Based [WS70]. Hidden [Gra72]. Hierarchical [GS73a]. High [AS72, BKKM71, CR71, Dra72a, Goe74b, Mit77, MM79b, MSW77, Sch70, Slo76, Whi71]. High-Capacity [CR71]. High-Density [Slo76]. High-Frequency [MSW77]. High-Impedance [Goe74b]. High-Order [Mit77]. High-Speed [BKKM71]. Higher [Gol74, Mar72c, Mar72d, Maz78b]. Higher-Order [Mar72c, Mar72d]. Hilbert [RS74]. Hill [ACH^{+}78, ACG78, CWE^{+}78, Rus78]. HILO [Laa71]. History [DR78, DC71, Dic75]. Hits [FG78]. Hold [Ben76]. Holding [Bur71, FM75]. Holograms [Ber70b]. Holographic [KS70]. Homogeneous [Moo70]. Hop [Osb71, Sha72]. Horn [Dra74, Sem79, Tho71a]. Horn-Reflector [Sem79]. Hot [Per72]. Hour [Eis77]. Human [GHL72, GS70, MRM78, YBBW79]. Human-Face [GHL72]. Hybrid [Aca78, AFG79, CCR76, Mos70, SS71a]. Hyperbolic [LC70].

ICUR [Kau77]. Ideals [Mac70].

Identification [AR79, BGM^{+}71, GHL72, Owe72]. Identifying [Goo74]. Identities [MSG72]. Idle [Aha78b, Cro78b]. Idle-Channel [Cro78b]. if [Mit77, Fen71]. IGFET [MCKP72, MM75]. II [Arn74d, Chu73, KLRS70, Len77c, Mes73b, OTW71b, Per73d, SDS78, SGM74, Stu76b, Tho77a]. III [Arn74e, DCRZ70]. IIR [CCR76, RKHD74]. Image [MS79a, May71, MMF70, Per71a]. Images [Lim72]. Imaging [DG79, Nag71, Seq72]. Imbalance [Mil75c]. Impact [Lin77b]. Impairments [Ols71]. Imparted [ES77]. IMPATT [Bra70, Bra73, MG75, Sch74, TG76]. Impedance [Goe74b, Man74]. Imperfect [Mar73b]. Implantation [WN72]. Implementation [BKT78, BGH79, CRS75, Lit78, Tho78]. Implementations [MG78]. Implemented [GW79]. Implications [Thi71]. Importance [Bal76]. Improved [Cro78b, Dra74, Kai70, Laa71, YH71]. Improving [AS78]. Impulse [CR73a, CR73b, Coo77, Glo73, GM73, HRC73, Mar73c, Mar73d, Mar74c, Mar74d, Rab72]. Inaccuracy [GGO79]. Inband [Lue70a]. Incidence [Sal75]. Incident [AC78]. Incoherent [CM78, Mar75b]. Incomplete [BT73, JB71]. Incorporating [BFG73, Lue70b]. Increase [Aca79]. Independence [Maz79a]. Independent [Sal75, SR75]. Index [Ano70m, Ano70n, AD76, BPM79, CM78, Coo77, GM73, Mar70a, Mar73c, Mar73d, Mar74c, Mar75b, Mar76b, Mar76c, OP74, Per74, PMAB79]. Induced [Bra73, Chu74, Des73, MCC73, MC73, Sch73, Sem71]. Inductive [Abe78]. Inefficiency [TKK73]. Inequality [HO77]. Infinite [Maz75c, Ric75]. Influence [Glo75, HGLM77]. Influences [PB75]. Information [BP70, CW73, GS73a, HR73, Jag70, Log77, PR73, Whi71]. Infrared [Daw72]. Inherent [BN71]. Injection [Alb78, CR76, Coh72, Eis72, Lee75]. Injection-Locked [Eis72]. Injection-Molded [CR76]. Innovations [Ben76, Kai71]. Input [Ben71b, Goe74a, Goe74b, Gop71, HH73, Kno77, KN72, Kuc73a, NK73b, Ric73d, NK73a]. Input-Output [Gop71]. Inputs [GM77, Gre73, Hef73, Hol73, Kuc72, MM78a]. Inserted [Man74]. Installation [AGW77, CDD^{+}78, HO70, RBH77]. Installations [BM78]. Insulator [CLB75]. Integral [Blu78, GP70, Mor79, Per72]. Integrals [Ric73c, Ric75]. Integrands [Ric75]. Integrated [BKS71, Die71, GS74a, MM71a, Mat70, Mos70, SS71a, Wad78a, Wad78b, Yam72b]. Integration [CGWN77, Cum75, Hel79, HMSV77, Laa70, Lin76a]. Intelligible [Par78a]. Intentional [Mar74e]. Interaction [GHL72, Jac70]. Interactions [FM71, Gra70a]. Interactive [CD79, CW73]. Intercept [BCM^{+}74, BCKV74, BWW74, CGHW74, CCWW74, HHM^{+}74]. Interconnected [San78b]. Interconnection [BZ74, Cok72]. Interesting [LM79]. Interface [BJO78, CR78, GR78]. Interfaces [BCD^{+}77]. Interfacial [KSBL74]. Interference [Bel78, Bul75, Fen70, Gol73, HY70, HY71b, MP71, Mes73a, Mes73b, Par78b, Wyn71, Wyn75a, YH71]. Interframe [Bob74, Has72, HS75b, LP71]. Interim [ACG78]. Intermodulation [DO79]. Interoffice [BKT78, CMR78, CFMS78, CS78, DR78, KCM^{+}78, Lit78, MW78b, RM78, Sna78]. Interpolative [Net77]. Interpretation [Per71a]. Interrupted [Kuc73b]. Intersection [Yam72a]. Intersymbol [Fen70, HY70, HY71b, Mes73a, Mes73b, Wyn71, Wyn75a, YH71]. Interval [WHF77]. Intraframe [Bob73, LPW74, MS79a]. Introduction [BFR75, DC71, EFG78, Poo70, War77, You79]. Introductory [VG72]. Invariance [BFG73, San79]. Invariant [RT71]. Inversion [Jag78]. Investigation [Seq74]. Investigations [Nag71]. Ion [WN72]. Ions [Wan70]. Isolated [RS75, RW79, TKK73]. Isomers [TBCR72]. Issue [Ano70a, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano71a, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78b]. Issue [Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k]. IV [Arn75b, JZ70].

Jacket [BY78, Mar74b]. Jar [DS70]. Jar-Cover [DS70]. Jars [HRH70]. Jitter [Dut72, Dut76, FGW73, HS72b, Maz78a]. Joining [Bis71b, CRAC79]. Joint [Cha70a]. Jointly [Fal76b]. Joints [CMG73, Som73]. Journals [Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e]. Jumbogroup [BNO74, Mau74]. Jump [Dra72a]. Jumpers [HMS72]. Junction [Lee75]. Junctions [Owe72].

Kailath [Ben76]. kb [CS77a, Cro78a]. kb/s [CS77a, Cro78a]. Key [Das71]. Keyed [Eis72]. Kilometer [Sem71]. Kruithof [Kru79].

L4 [ALM74]. L5 [ACKL74, Ano74k, BNO74, BZ74, CHK^{+}74, DJNP74, FP74, GOS74, GS74b, HM74, KH74, Mau74, TAB74, WA74]. Laboratories [Ano77f, Ano77r, Ano77s, Ano77t, Ano77u, Ano78a, Ano78o, Ano78p, Ano78q, Ano78r, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v, Cox78, Dan75]. Laboratory [LR79, Sul71]. Laminated [Mil76a]. Laminates [Kle72, Kwe72]. LAMP [BHKJ74, CH74, CSW74, Cha74, CES74]. Language [Dic75, JL78, RJLK78]. Laplace [Blu78, Jag78]. Large [BB78, But71, CR72, RB70, Tho71a, WP75]. Large-Scale [CR72]. Laser [Bur70, CKS79, Has78, Hub71, MM71a, May71, MWW^{+}79, SCD78]. Lasers [Alb78, Coh72, Dix76, Lee75, Mar76d]. Last [LMY76]. Last-Trunk-Usage [LMY76]. Lattices [WS70]. Launcher [GS75]. Laureate [Ano77f]. Laureates [Ano78a]. Law [ACG79, MPS77]. Laws [Kan70]. Layer [Lew76, MS72, Ros72]. Layers [MM79a]. Lead [Ano71k, BW70a, BLV70, CFB70, Cus70, DS70, Gar70, HRH70, KFBL70, Lue70b, Mil70, OBW70, SSV70]. Lead-Acid [Ano71k, BW70a, BLV70, CFB70, Cus70, DS70, Gar70, HRH70, KFBL70, Lue70b, Mil70, OBW70, SSV70]. LED [AH79, Lee75]. LEDs [CBK79]. Lee [Ano70l]. Lemma [Mor78c]. Length [AS72, CKK^{+}71, RSSF78, Rut71]. Lens [Ano75k, Kai70, Mil74]. Lens-Like [Ano75k, Mil74]. Lenses [Her70]. Level [MNW78, NMB77, PMN78]. Levels [RV70]. Librarians [Nic75]. Light [Coh71, Dix76, Las76, MS76]. Light-Current-Voltage [Dix76]. Lightguide [BSS78, BM78, ML79]. Lightguides [CF79]. Lightwave [MSW78, SRMC78, SCD78]. Like [Ano75k, Mil74]. Likelihood [FM73, Mes73b]. Limit [Aha79, LM79]. Limitation [Ric73a]. Limitations [Tho73b]. Limited [Wyn76]. Limiting [Can74, LC70]. Limits [Ric75]. Line [ACKL74, Bab72, Bar70, BBJST77, FG71, Gil75, Gra72, Gra70b, GS74b, HM74, KLS76, LM76, Man74, MM72a, Mil72, MM72b, O’N71, Rut71, SOP73]. Line-of-Sight [Bab72, Bar70, Gil75, Gra70b, Rut71, SOP73]. Line-Power [HM74]. Line-Protection [GS74b]. Linear [Ber79, CH78b, CH79, Gop71, Gre73, Has75, Liu71, Mos70, MF70, Rab72, RT71, RM72, RB70, Sam75, SWD71, Sem74, Seq74, Sha72, SG77]. Linear-Prediction [Sam75]. Linear-Predictive [Has75]. Linearity [Per73b]. Linearized [JB71]. Lined [CD71, Car71, CM73, CD73]. Lines [Bri78, CRSB77, IS73, Ros76, Sta74]. Link [CGRS77, Mus77, RSCG77]. Links [AH79, CKS79, CBK79, GS73a]. Lipschitz [Sha70]. List [Ano75m]. Listener [Bri71]. Listeners [MRM78]. Listings [RS79]. Lithium [Wil77]. Load [BN71, Des75, Des76, GLW74, Gue75, Hil77, MM75, Mes78, Wil71]. Load-Balancing [BN71]. Loaded [Owe72, Ram74]. Loading [MPS77]. Local [BKL75]. Locating [RSR74]. Location [LTV70, RC78]. Locked [Cha72, Dut76, Eis72, Obe70]. Logic [CES74, Hag75, Ols75, Wad78b]. Logic-Circuit [CES74]. Long [AS72, Che71, Doa78, Lin77a]. Long-Distance [Doa78]. Long-Haul [Che71]. Long-Term [Lin77a]. Longitudinal [ES77, IF77, Mil75c]. Look [BG72b]. Loop [And78, AS78, AB78, BB78, BGK72, BG72b, Cok72, DL78, Fre78, Fri78, Gib78, GP71, HS78a, HW78, Koo78b, Koo78a, Lap76, Lon78, MNS78, Man78, Mar78, Moo70, Mor78b, Obe70, Rao78, SS78, Sti78]. Loops [Bro71c, Cha72, Dut76, Par78a]. Loose [Mil75b]. Loss [Bis71a, CD71, Car71, CM73, CHS76, CGBS75, CM78, DS73, Gar75, GSBC73, Glo75, Glo76, Jag74, KA74, Kra71, KN72, Mar72c, Mar77b, Met79, MM78b, NK73a, NK73b, Sne71, Som73, Spa76, Sul71, Wil71]. Loss-Noise-Echo [Spa76]. Losses [Ano70k, Arn74e, FMOT74, Lue70a, Mar70d, Mar70c, Mar71b, Mar72d, Mar72e, Mar73d, Mar73e, Mar74c, Mar76b, Mar76c, Mar76d, Mar76e]. Lossy [CM75a, IS73, Mar74b]. Loudness [Sul71]. Low [Aha78a, Aha78c, CD71, CM73, CGBS75, CS71, Cro77, DS73, FH71, FMOT74, GS74a, GSBC73, HS75b, Hen73b, HRC73, Hub73, IL78, KA74, KS78, Mus77, Neu71, Par78b, RJ72, Sne71, Som73, TNMC79]. Low- [Hub73]. Low-Bit-Rate [Aha78a, Cro77, HS75b, Mus77, TNMC79]. Low-Clock-Rate [Aha78c]. Low-Frequency [Par78b]. Low-Loss [CD71, CM73, CGBS75, DS73, GSBC73, KA74, Sne71, Som73]. Low-Noise [CS71, GS74a, IL78, KS78]. Low-Pass [HRC73]. Low-Speed [FH71]. LPC [GSC^{+}79, SJ76]. Lumped [Ken70]. Lumped-Circuit [Ken70].

M [GG73]. M/PCM [GG73]. MAC [CT79, Rov78]. MAC-4 [CT79]. MAC-8 [Rov78]. Machine [AMS70, GHL72]. Machinery [WSS78]. MacWilliams [MSG72]. Made [KA74]. Magnetic [Aha72c, BBL72, BBB^{+}79, Car71, FM71, GOS74, Gra70a, HN78, PB74, PB75, PSW70, Ros72, Slo76, TBCR72, Thi71, TBTG71, WN72]. Main [BB78, Fen71, HMS72, Sch71]. Maintenance [AC70, BS78, BBJM^{+}77, BDG^{+}77, CH74, CDD^{+}78, DPS71, FG71, FR75, GH77, GHRS77, HS75a, MRY77, Mor78b, Rif75, Sna78]. Major [Apr75]. Making [HP70]. Man [GHL72].

Man-Machine [GHL72]. Management [CW73, DJ75, GS73a, GHHJ77, HR73, MW75, PR73]. Manifolds [Net73]. Manipulations [SJ76]. Manually [Cho74]. Manufacture [AHK^{+}78, BSS78, Cha70b]. Manufacturing [BG72a, Ele77, Gar72]. Map [Goo74]. Marcuse [Ano70k]. Margin [Aca79]. Markov [Gib78]. Markovian [Des75]. Mask [BP70, HP70, Yam72b]. Mask-Making [HP70]. Masking [Sen70]. Mass [IF72]. MASTER [GS73a]. Mastergroup [GRW71]. Material [Arn74b, BW70a, CDF^{+}78, Con72b, KA74, Mar74f, PSMO75, SH74]. Materials [Bro70, Con72a, Ker70]. Mathematical [Ber70a, FM71, Gra70a, Mit74, Mit75]. Matrix [PR74]. Maxentropic [Sle72a]. Maximum [Chu71a, FM73, Mes73b]. Mb/s [MSW78]. Mean [Cho74, FF73, GW79, Mue75, Sal73]. Mean-Square [Mue75, Sal73]. Mean-Square-Error [FF73]. Mean-Squared [GW79]. Mean-Squared-Error [Cho74]. Means [Kuo75]. Measured [Coh71, Laa70]. Measurement [CM78, Cox75, Des73, Doa78, FHH^{+}79, GHWY77, Hil77, Hub71, NK73a, NK73b, RV70, RC78, Rus78]. Measurements [Bis71a, CHK^{+}74, CAK76, Des75, Dix76, GS70, Has75, KN72, LMY76, MM79a, RS78, Rif79, Sem71, Sem74, Sem77, TKK73, Wri71]. Measures [Bra72, Stu76a, Stu76b]. Measuring [AD76, AMS70, Sul71]. Mechanical [FHT77, GGL^{+}77, KLRS70, PS70]. Mechanisms [IL78]. Media [Mil74, Ano75k]. Medium [ACTW77, CGL71, Ele77, GGL^{+}77, GH77, KS71, SW71a, SW71b]. Medium-Haul [CGL71, KS71, SW71a, SW71b]. Melt [DSS78, SDS78]. Memories [BN73]. Memory [ABG^{+}77, BW74, FM73, Jay70, Jay73].

MERT [LB78, NP78]. MESFET [Fuk79]. Message [ADM78, AAS70, BS76, KNS^{+}70]. Message-Switching [BS76]. Metal [CS71, PB75]. Metal-Semiconductor-Metal [CS71]. Metallic [WM70]. Metallurgy [PB73]. Meter [CWE^{+}78, Fen70]. Method [Che76, Hal70, Hal75, Hol73, Hub71, IZ70, Kau74, Kru79, Lin75, Liu74, Nea72, RRS70]. Methodology [AS78]. Methods [AR79, CB72, CRSB77, PR74]. MHz [CKS79]. Michailov [Rei72]. Microbending [Gar75, Mar76b]. Microcomputer [CT79, WSS78]. Micromachining [May71]. Microphone [NS72, Pir79]. Microprocessor [Fan78, Lyc78]. Microprocessor-Based [Fan78]. Microprogram [Sto76]. Microscopic [Tho74b]. Microstrip [Sne71]. Microwave [ALS71, Bab72, Bab73, BCS71, Bar70, BJL71, BW71, Bul71, Bul75, Chu79a, CS71, Die71, Dra74, Dra77a, DSW71, Fen71, Giu71, HJ71, JP71, Lin75, Lin77a, Lin77b, MCC73, MC73, Ohm74, Rut71, Sei71, Sem71, Set70, Sha72, SOP73, TG76, VP79]. Microwaves [CGL75]. Mid [Cro78b]. Mid-Range [Cro78b]. Mid-Range/Mid-Tread [Cro78b]. Mid-Tread [Cro78b]. Miller [Ano75k]. Millimeter [ABH77, AGW77, ACTW77, BBJH^{+}77, BBJM^{+}77, BCDT77, BBJST77, CM77, CHM^{+}77, Chu74, CWE^{+}78, DPR77, Ele77, FHT77, GHWY77, GS74a, Gol77, GGL^{+}77, GH77, HNW77, HMG70, LOOS77, RBH77, Tho77a, War77]. Millimeter-Wave [GS74a]. Minicomputer [LC78]. Minimax [RS74, SR74]. Minimize [Wyn78]. Minimizing [CR73a, Lue70a]. Minimum [Coo77, FF73]. Minority [MS73b, MS73a]. Minute [Che76]. Mirrors [Zuc70b]. Misalignment [CM78]. Misalignments [CMG73]. Mixed [Kuc72]. Mixers [Dra72a, Dra72b]. Mixing [Mar76c, Per74]. mm [Doa78, CKK^{+}71]. MM-Wave [CKK^{+}71]. Mobile [AFS79, CHPT79, CR71, CR72, DPZ79, EFW79, Fis79, FP79, Huf79, Jay75b, Lee71, Mac79, MS73b, MS73a, Tsi79, Wal79, You79, Ano70l, Lee69]. Mobility [Wan70]. Modal [Mar77a, Owe72]. Mode [Arn74d, Cha72, CMG73, Dra77b, Mar70b, Mar70d, Mar72f, Mar73a, Mar74e, Mar75a, MP75, Mar76b, Mar76c, Mar77b, Per74, RSSF78, Sch70, Som73]. Mode-Mixing [Per74]. Model [AFG79, BB78, Ben71a, Bud73, Bur70, Cha73, FIS75, FR73b, FHH^{+}79, FR79, Fri78, GS73b, Gra70a, GP70, Hey75, IF72, Lim73, LN72, LW74, MS79a, MP71, Mea75, MM78b, Per72, Rum79, Sem79]. Modeling [Ame72, AS78, BB79, BB78, DL78, Fre78, Fri78, Gib78, HMS72, Koo78b, Koo78a, Log71, Lon78, Mar78, Sti78, Wad78b]. Models [Ben71a, CHS76, FM71, Gib78, Hay74]. Moderate [Hub73]. Moderate-Bandwidth [Hub73]. Modern [GRW71]. Modes [Arn74c, CD71, Car71, CM73, CD73, Doa78, Mar71a, Mar71c, Mar72b, RW74, Sch73]. Modification [WHF77]. Modified [DW78]. Modular [HL79]. Modulated [GHM73, Gla71a, WR73]. Modulation [Aha78c, Cum75, Ger72, Goo71, Hol75, HMG70, Jay70, JR71, LM70, Laa70, Log78, Pra70, PR74, Pra76, RP75, RB74, San78a, SKCM78, Sle72b, Whi71]. Modulations [Jay74]. Modulator [Boy76, CKK^{+}71, RB70, Sha72]. Modulators [Giu71, Gre73]. Molded [CR76, Min72, SBGC75]. Molded-Plastic [SBGC75]. Moment [Den76]. Monitor [Fan76]. Monolithic [Gea79]. Monte [BKS71, Ols71]. Moon [Cap72]. MOS [CLB75, Wad78b]. Motion [Ber70a, Gea79, NR79, NS79, SN79]. Motion-Compensated [NR79, NS79]. Motions [Liu70]. Mounting [Mat70]. Mounts [Ano79l]. Moving [Lim72]. MSM [CS71]. Multibeam [AD78, Tur75]. Multicomponent [Lew76]. Multiconnection [Hwa79]. Multidimensional [Aha72a, Aha72b, Aha72d]. Multidisturber [Lap76]. Multihour [Els77]. Multilayer [BBL72, Sal75]. Multilevel [Gla71a, HY71b, HMG70]. Multiloop [Apr75]. Multimode [Arn74a, Arn75a, CM75a, CM75b, CAK76, Glo72, Glo73, MG73, Mar72c, Mar72e, Mar72g, Mar74e, Mar76b, Mar76d, Met79, Oga77, SKCM78]. Multipath [Bar72, Bul71, Lin77b, Mot77, VP79]. Multiple [Aca77, Aha72b, BW70b, Dab75, Glo70, Gop71, Ohm74, RB76, Rut71, Sem77, SW73, Sze73, Wol72]. Multiple-Address [Wol72]. Multiple-Beam [Glo70, Ohm74, Sem77]. Multiple-Burst [BW70b]. Multiple-Feedback [Sze73]. Multiple-Path [Rut71]. Multiple-Spot-Beam [Aca77]. Multiplex [Mau74]. Multiplexer [AFG79, GLW74]. Multiplexers [BBCC75]. Multiplexing [DMM72, HS72a]. Multiplication [Per71b]. Multiplier [SS71a]. Multiprocessor [Phi75]. Multiqueue [Kue79, Whi75]. Multireflector [Dra78]. Multiserver [Kuc73a, Whi75]. Multistage [Hwa76]. Multivariate [BW74]. Mutually [Mor71].

N [Ano79l]. Named [Ano77f, Ano78a]. Names [RS79]. Narrow [Ber70b]. Narrow-Band [Ber70b]. Narrowband [Bab72, Bab73, Bur70, CGRS77, RSCG77]. Narrowband-to-Wideband [CGRS77]. Nationwide [Lin77a]. Natural [PR73]. Nature [Aha79]. Near [LM76, Pir79]. Nearly [RW74]. Necessary [Lun71]. Negative [Man74]. Negative-Impedance [Man74]. Neighbors [WS70]. Netravali [Ano79l]. Network [ABLR77, ADM78, Avi71, BDK78, BKKM71, BBW78, BS78, Ben75c, Ben78, BJO78, BPS75, BDS78, CR78, CK78, DL76, Dor71, DT71, DM78, EKN76, FW78, FH71, FR73b, FR73a, FMSZ78, GR78, GBN^{+}76, GW74, GHHJ77, Hay74, HS78b, Koo78a, MMM75, MW78a, MW78b, Pie72, Rod78, SZ75, Slo72, Spa76]. Networks [AOR76, Bau72, Ben70, Ben73, Ben75a, Ben75b, BS75b, BDS78, CH77, Chu78, Chu79b, DL76, DSW71, Eis77, Els77, GOS74, GAW^{+}76, GBN^{+}76, HW76, Hol71, Hwa71, Hwa76, HL77, HO77, HL78, Hwa79, Kar71, KLS76, Kle76, KU76, KT73, Kru76, Laa71, Lak79, Lue72, Mos70, OTW71a, OTW71b, RM72, San70a, San70b, Sti78, Tho71c, Wil70b]. Nickel [AD79, MM79a]. Niobate [Wil77]. No [ABLR77, AHPR79, AC70, AAS70, Ano70m, Ano70n, Ano79l, AOR76, BCKM70, BC70, BGK70, BDJ79, BDRP79, BCD^{+}77, BHP71, BPW79, BGH79, CFP70, CCR^{+}77, CSL79, CDH70, CGWN77, CLT70, CS78, DL76, DVW79, DKP^{+}70, EG70, GAW^{+}76, GBN^{+}76, GHHJ77, Gue75, GHRS77, HW76, HST70, HJJ^{+}77, JJ70, KLS76, KPS70, KNS^{+}70, Kle76, KU76, LO77, MOR^{+}77, MS73b, MS73a, MRY77, NMW79, Pek78, PSW70, RSS79, RT77, Spe77, SH79, SVW79, YBBW79]. Nobel [Ano78a, Ano77f]. Noise [Aha78b, Bur70, CHS76, CR73a, CR73b, Cha73, CS71, CHM73, FGW73, GS74a, Gol74, GG73, Gre73, HY70, HY71b, HS72b, IL78, Jac70, KS78, Laa70, MPS77, MF70, Nol75b, Per71a, Ric70, Sen70, Spa76, SK74, Tho74b, Tho74a]. Noisy [MG78, RT71, Sca79]. Nonadaptive [Has75]. Nonblocking [Ben73, Chu79b, Hwa79]. Nondiversity [Bab73]. Nonequalizing [Per76]. Nonexhaustive [Kue79]. Nonideal [Tho71b]. Nonlinear [Bau72, CK71, KN78, MSG72, Mor70, Nea71b, San70a, San70b, SS71b, Tho71b, Tho71c, Wil70b]. Nonlinearities [Fal78]. Nonorthogonal [Arn70]. Nonrandom [Nea71a]. Nonstationary [Jag75, MS79a].

Nonuniform [Kog76]. Normal [CD73].

Notch [Zuc74]. Note

[Net73, San78a, Wyn76]. Notes [Mul78]. Novel [AD76, CRS75]. Nuclear [Sha75]. Number [Vig70, Vig71]. Numbers [Ham70, RRS70]. Numerical [Blu77, Hel79, Mar77a, Ric75]. Numerics [JPS^{+}78]. Nyquist [Maz75a].

O [Ano71k]. Objective

[Bra72, GSC^{+}79, MSG78]. Objectives [BWW74, EG70, JJ70, RT77, Sta77, You79]. Oblique [Sal75]. Observability [CH74, RM72]. Observation [LR79]. Observations [Bri78, Gol77, Maz79b, RB76]. Obtaining [RS79]. Occupancy [Pra76]. Occurrence [SDMS77]. Ocean [Cha70b, LS70]. Off [HR73, Per74]. Offered [Mes78]. Office [BS71, CHPT79, CGWN77]. Offset [Chu77, CM78, Dra78, GS75, Glo76, Mil76b, Sem77]. Offsets [Bis71a]. One [Can74, Eis77, Jay70, Jay73, Ric73d]. One-Bit [Can74, Jay70]. One-Word [Jay73]. Open [IZ70, Zuc70a]. Operating [Fre78, Hag75, Koo78a, LMT78, LB78, NP78, RT71]. Operation [BGK70, FR73b, Mar72c, SBR78a]. Operational [BDJ79, BDRP79, CCWW74, KPS70, Rod78, Tho73b]. Operations [AS78, CK78, DL78, Gib78]. Operator [RSS79]. Optic [FGS75, Hen73a, Mil75a, Mil78, Per73b, Per73c, Per73d, Per74, RC78]. Optical [AH79, Arn70, Arn74a, Arn75a, AP75, AD76, Bis71a, Bis71b, BS75a, BM78, CM75a, CM75b, CR76, CRAC79, CBK79, CGBS75, CGL75, Coh72, CAK76, CMG73, CHJZ70, DMK78, DS73, DW78, ES77, FMOT74, Gar75, Gey76, Glo72, Glo73, GSBC73, Glo75, Glo76, Goe74a, Goe74b, Hol75, Hub73, JMM78, Kai70, KMM73, KN78, Las76, Mar77a, Mar73a, Mar73c, MM73, Mar74b, Mar74d, Mar75a, MP75, Mar76d, Mar76e, MS76, Met79, Mil75b, Mil76a, Mus77, Per71a, Per71b, Per71c, Per73a, Per74, Per76, Per77, PSMO75, Pre76, RS76, RS78, Raw70, San78a, SKCM78, SP77, Sch72, Sch78, Sha75, SBGC75, SBR78b, Smi79, Som73, Sta74, SH74, Whi71, Zuc70b]. Optical-Fiber [BS75a, CGBS75, Glo75, Mus77, Per77, SKCM78, Sch78, SBGC75]. Optical-Frequency [Hol75, Hub73]. Optical-Waveguide [San78a]. Optically [Kno77, ML79]. Optics [Arn74c, Arn74d, Arn74e, Arn75b, MM71a]. Optimal [Cho72, Chu75, CH78a, Fer72, Fre78, Gar72, Net73, Per74, Whi75]. Optimization [BB79, Cha70a, Hal70, KL73, LN72, SBR78a, Smi79]. Optimum [FG78, FGS75, FR73a, HRC73, Ho71, Kar71, Maz75c, Mue73, NS76, Net77, OP74, Sal73, Shn70]. Order [Coo70, DO79, Lue72, Mar72c, Mar72d, Maz78b, Mit77]. Ordering [Ano79l, MNW78, MBN79, NMB76, NMW79]. Organization [BWW74, CCR^{+}77, EG70, JJ70, KNS^{+}70, RT77, Sta77]. Orthogonality [Chu71b, Chu73]. Oscillating [Ric75]. Oscillations [Bra73, Mit77, Sch74]. Oscillator [BKS71, CS71, Eis72, FP74, Ken70]. Oscillators [Bar71, Bra73]. Other [Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ben71a]. Outage [ALM74, Che71, LR79, Osb71]. Outages [Lun70, Mot77]. Output [Gop71, Hub71]. Overall [BDJ79, BDRP79, KH74]. Overflow [Bur71, Kuc73b, Mit77, Nea72, Wil72]. Overload [FR73b, Gre73, JR71]. Overmoded [CD73]. Overview [AE70, And78, BFR75, Con75, Cox78, DJ75, FW78, HW76, HP70, Jac78, Lon78, RM78, SK75, SH79]. Oxide [CLB75]. Oxide/Silicon [CLB75]. p [MM75, Fen70]. p-Channel [MM75]. P. [Ano78n]. P/AR [Fen70]. Package [AH79, SBR78b]. Packaged [HR73]. Packaging [Glo75, HL79]. Packet [WHF77]. Packing [Gil79]. Pair [Bel78, Koo78b, Len77a, Mil72, RS76]. Paired [BHO71]. Pairs [BB78, Gil79, Rob71]. PAM [GS71a, Wyn78]. Papers [Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77r, Ano77s, Ano77t, Ano77u, Ano78o, Ano78p, Ano78q, Ano78r, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v]. Parabolic [GM73, Mar70a, Mar73c, Mar73d, Mar74c, Mar75b, Mar76b, Mar76c, Mil76b, MM78b]. Parabolic-Index [Mar75b, Mar76b, Mar76c]. Parabolic-Profile [Mil76b, MM78b]. Paraboloid [Chu77]. Parallel [Mar71c, RS78, RW74]. Parameter [Len77a, RS78, RB76]. Parameters [Fuk79, Len77b, SJ76, Wal70]. Parametric [Sch74]. Part [BPM79, CGRS77, NR79, PMAB79, RSCG77, Arn74c, Arn74d, Arn74e, Arn75b, CHJZ70, DCRZ70, DSS78, JZ70, KLRS70, Len77b, Len77c, Mes73a, Mes73b, OTW71a, OTW71b, SDS78, Stu76a, Stu76b]. Partial [Des73, Fer72, Mue75, Nea72, SGM79]. Partial-Access [Nea72]. Partial-Response [Mue75]. Partitioning [KL70]. Pass [HRC73, RC76]. Passband [Fal76a, GHM73]. Passive [FYMS78, GM71]. Paste [BW70a]. Path [CKK^{+}71, Osb71, Osb77, Rut70, Rut71, Sem71, Wil70a, Wri71]. Path-Diversity [Wil70a]. Paths [BR74, Gil75, Gra70b, Lin75, Slo72]. Pattern [Bur70, CHJZ70, DCRZ70, Dra78, GRW70, JZ70, KLRS70, Poo70, SRPR70]. Patterns [DSS78, MM79b, RS78, SDS78].

PBX [And71]. PCM [CJF73, Goe74a, GG73, GMN76, HHJ71, MSG78, MG78, RJ72, SG77]. Peak [Bra72, GLW74, RC76]. Peak-Load [GLW74]. Peaked [Hef73]. Peakedness [HH73]. Penalty [Mar72c]. Per-Channel [MT72]. Perceptual [MSG78, RJ72]. Perfect [Dra78, Mor79]. Performance [AE70, BKKM71, BS72a, BEST78, BFG73, BSS78, CHM^{+}77, Cro78a, Cro78b, Dra72b, DT71, DMNT75, DM78, Dut76, FH71, GSC^{+}79, HS72a, Hay74, HS72b, JZ70, Kes71, Las76, Lin77b, Mes75, Mes73b, Mor71, Nol75b, Osb71, Par78a, Pra70, RT71, Ton70, TNMC79]. Periodic [Kog76, Rif79]. Periodically [Ric70]. Peripheral [BCM^{+}74, CDH70, HJJ^{+}77, Sna78]. Permanent [Dab75]. Perspective [Gra72]. Perturbation [MC73]. Phase [And75a, BS75b, Bul71, CRS75, Dut76, Eis72, Fal76a, FG78, FGW73, GHM73, Gla71b, Gla71a, Ho71, HS72b, IZ70, KWP72, Maz75c, MSW74, MCC73, MC73, Nag71, Obe70, PR74, RB70, RB74, SKCM78, Seq74, Sha72, Shn70, SOP73]. Phase-Locked [Dut76, Obe70]. Phase-Modulated [GHM73, Gla71a]. Phase-Shift-Keyed [Eis72]. Phased [RS76]. Phenomenon [Dra72a]. Phone [AFS79, CHPT79, DPZ79, EFW79, Fis79, FP79, Huf79, Mac79, Tsi79, Wal79, You79]. Photodiode [MHSS78]. Photolithographic [Her70]. Photolithography [AEH^{+}70, AMS70, Ble70, BP70, Bro70, CHJZ70, DCRZ70, GRW70, Her70, HP70, JZ70, Ker70, KLRS70, Poo70, PS70, Raw70, SRPR70]. Photon [Per77]. Photosensitive [Ker70]. Photovoltaic [KN78]. Physical [BC70, BBW78, BCE^{+}77, HW78, ILR75, Kle76, LF72, Man78, VG72, WA74]. Picosecond [Hub71]. Picture [LS71, Lim73]. Picturephone [BS70a, Bob73, Bob74, CHM73, Has72, And71, Bai71, BH71, Bre71, Bro71a, Bro71c, BHP71, Bro71b, BHO71, BS71, CSW71, Cra71, Das71, Dor71, DPS71, FG71, GS71b, GRW71, MM71b, Uri71, WP71]. Pictures [Ano79l, Ber70a, MS79b, MNP77, MNW78, NMB76, NPM77, NMB77, NMW79]. Pitch [Jay77]. Pitch-Adaptive [Jay77]. PL [Van75b]. PL/1 [Van75b]. Plan [ABLR77, AC70, AS78, Bro71b, DPS71, EFG78, Uri71]. Planar [Arn74a, KT73, MHSS78, Sch73]. Planarity [GS73b]. Plane [MC74]. Planning [BKT78, BPS75, SS78]. Plant [And78, AS78, AB78, BB78, DL78, Fos71, Fre78, Fri78, Gib78, HS78a, HW78, Koo78b, Koo78a, Lon78, Lue70b, MNS78, Man78, Mar78, Mor78b, Rao78, SS78, Sti78]. Plasma [JMM78]. Plastic [CR76, CRAC79, Kle72, Kwe72, Pre76, SBGC75]. Plastics [BLV70, DeC72, Min72]. Plate [Lew70, MS73b]. Plates [BW70a, Cus70]. Plus [WP75]. PM [SSP71]. PNIPN [PB74]. Point [Che76, Gar72, Mor78a, NS76]. Point-to-Point [Mor78a]. Poisson [Kuc72, Kuc73b]. Polarization [AC78, Chu74, CGL75, Chu77, Dra78, Gan76, RSSF78, Sal75, Sem73]. Polarization-Independent [Sal75]. Polarizations [Chu71b, Chu73, Sem74]. Polarized [Mot77]. Policies [Fre78]. Polled [GR78]. Polyethylene [Beb72]. Polygons [Yam72a]. Polynomial [Aha72a, Aha72b, Aha72d]. Polynomials [Kur70]. Polyvinyl [HRH70]. Port [Ric73d]. Port-Distortion [Ric73d]. Portability [JR78]. Position [AHPR79, BDJ79, BDRP79, BPW79, BGH79, CSL79, DVW79, Hol75, RSS79, SH79, SVW79, YBBW79]. Positive [Ano71k, BW70a, CFB70, Cus70]. Positivity [San79]. Post [Abe78, SSV70]. Potential [BM71, MS73b, MS73a, WM70]. Power [AS72, AI74, ADM78, BCS71, BS72b, CG70, CGS^{+}78, Chu71a, Coh72, Gla71a, Glo72, HM74, LO77, Mar72a, Mar72b, Mar72e, Maz78a, MPS77, Par78b, Per75, RP75, Sch70, Wah77]. Power-Frequency [Sch70]. Power-Telephone [Par78b]. Powered [DMK78, ML79]. Practical [HRC73, MSW78]. Precision [Bar71, CHK^{+}74, FP74, GW79]. Predicted [MS79b]. Prediction [CPS71, GG70, Jay77, LR79, Sam75]. Predictive [AS70, Has75, NPM77, Nol78]. Preface [Ano75n, Cro78c]. Preference [JR71]. Preferential [Sch71]. Prefilters [Bob73]. Preform [MP78]. Preforms [JMM78, PSMO75]. Preparation [CGBS75, GSBC73, JMM78, KLO78]. Prepared [DW78]. Presence [FG78, FGW73, HY70, HS72b]. Pressure [Dah72]. Pressure-Sensitive [Dah72]. Primary [CHJZ70, DCRZ70, JZ70, KLRS70, Poo70]. Principles [CFB70, Ano71k]. Printed [Rai76, Rai79]. Printer [MW78a, MWW^{+}79]. Prize [Ano78a]. Probabilities [Ber79, CH77, CH78b, CH79, Lap76, MM78a]. Probability [Hil71, HN76, HY70, HY71b, HO77, Stu76a, Stu76b, Wyn75a]. Probability-Engineered [HN76]. Problem [BGK72, Cap72, Gil79, GP71, Lew70, Wol72, Wyn71]. Problems [MM72a]. Procedure [CCR76, KL70, Peh70]. Procedures [DVW79]. Process [DW78, DM75, DG75, Ele77, Kuc73b, Mar78, Nea71b, Smi79]. Process-System [DM75]. Processed [MSG78]. Processes [Gar72, GG70, Gib78, GMS73, Mar72c, Sle72a, SGM74]. Processing [Ano75n, BFR75, Con75, Cro75, DJ75, Dic75, DM75, DG75, Fre75, Gaw75, Hag75, HS75a, HJ75, KNS^{+}70, Kuo75, LS71, MCM78, Mes75, MW75, Nic75, Ols75, Phi75, PB73, Rif75, Van75a, Van75b, WP75]. Processor [ABG^{+}77, BCE^{+}77, BDG^{+}77, BDN^{+}77, CFG^{+}77, HMSV77, LC78, Sta77, Sto76]. Produced [Ric73a]. Product [MP78].

Production [Nic75]. Profile [AD76,

Mar77a, Mar73c, Mil76b, MM78b, MM79b]. Profiling [BPM79, PMAB79]. Program [BCKM70, DKP^{+}70, KNS^{+}70, MOR^{+}77, Nic75, VG72, WP75, Yam72b]. Programmer [DHM78]. Programming [Nic75, Rab72, RJLK78, Van75b]. Programs [CCWW74, JR78, KPS70, Kur70, Rod78]. Project [MW75, SRMC78]. Projecting [GG70]. Projection [Kru79]. Prolate [Sle78]. Prologue [Spe77]. Proof [Ben75c, Mes72]. Propagation [Aha72b, ACH^{+}78, ACG78, Bar70, Bar72, Can74, Chu79a, Cox78, DDH75, Gra70b, Len77a, Len77b, Mar72g, Mar72f, MS79b, MSW77, Pen70, Rum79, Rus78, Sem71]. Propagation-Parameter [Len77a]. Proper [GS73b]. Properties [Beb72, Ber70a, Cha73, CDF^{+}78, Dah72, DeC72, Des76, Gol74, IF77, Jag74, Kle72, Kru79, Kwe72, Mat70, Min72, PB75, Rai79, Ven72]. Property [Mes72]. Proposed [Ohm74]. Protected [VP79]. Protection [BBJM^{+}77, Che71, CGL71, Fan78, GS74b, KU76, SJS79]. Proving [Ben75b]. Provisioning [FHH^{+}79]. Pseudo [CH72]. Pseudo-Random [CH72]. Pseudorandom [MPS77]. Publications [Ano76a]. Published [Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e]. Publishing [Ano76h]. Pull [Obe70]. Pull-In [Obe70]. Pulse [AS72, Arn74a, Arn75a, CDF^{+}78, Hol75, Mar72g, Mar72f, Mar74e, Mue73, O’N71, WR73]. Pulse-Position-Modulation [Hol75]. Pulse-Width-Modulated [WR73]. Pulses [Hub71, MS76]. Pump [SS74]. Pure [BY78, KA74]. Pyramidal [Sem79].

QAM [FF73, Fal78]. Quadrature

[Blu77, Rob71]. Quality [RSA77, WHF77]. Quantization [CJF73, JR72, Jay73, Jay77, Nol75a, Wil72].

Quantizer [Cro78b, Mit74, Mit75, NS76]. Quantizers [CB72, MNP77]. Quantizing [CHM73, GG73, Laa70, Nol78]. Quantum [Per73a]. Quasi [AP75, CM75b, CGL75]. Quasi-Optical [AP75, CGL75]. Quasi-Ray [CM75b]. Quasioptical [Gol77]. Queue [Hey75, Kuc73a, Whi75]. Queues [Col73, Coo70, GMS73, GM77, Hal75, Kuc72, Mor78c, SGM74]. Queuing [AFG79, Hef73].

R [Ano71k, Ano78n, CH79, NK73a]. Radiation [Ano70k, Mar70d, Mar70c, Mar72e, RS78, Sha75]. Radiator [HL73]. Radio [ALS71, Bab72, Bab73, BCS71, BJL71, BR74, BW71, Bro71a, Che71, Chu71b, Chu73, CR72, Die71, Dra74, DSW71, Fen71, Giu71, GH70, HJ71, HGLM77, JP71, KS78, Lee71, Lin77a, Lin77b, LR79, MP71, MG75, Osb71, Osb77, Rut70, Rut71, Set71, Sha72, SOP73, TG76, VP79, Ano70l, Lee69]. Radio-Relay [GH70, MP71, Set71]. Radioastronomical [Gol77]. Radiometer [Wil70a]. Rain [Aca79, BR74, Bul75, Che76, Chu74, Cox75, DDH75, GH70, HGLM77, Lin73, Lin75, Lin76a, Lin76b, Lin77a, Lin78, MCC73, MC73, Osb71, Osb77, Rut70, Sem71, Sem73, Sem74, Set70, Set71]. Rain-Gauge [Lin76a]. Rain-Induced [Chu74, MCC73, MC73]. Rain-Rate [Che76, Lin76a, Lin76b]. Rain-Scatter [Bul75]. Raindrops [MC74]. Random [CH72, Col73, Gil75, Hol73, Mar73d, MM72a, MM72b, RRS70, Sch71, Ano70l, Lee69]. Randomly [Ben71a, LC70, Ric73b]. Range [Cro78a, EKN76, Jac70, Obe70]. Range/Mid [Cro78b]. Rapid [BPM79, PMAB79]. Rapidly [MS75b]. Rate [ACG79, Aha71, Aha78a, Aha78c, Bal76, Cha72, Che76, Cro77, DC79, HS75b, Lin76a, Lin76b, Lin77a, Lin78, Llo77, Mus77, PMN78, RJ72, TNMC79]. Rate-Change [Aha71]. Rate-Locked [Cha72]. Rates [Bob73, Bri78, Tho73a, Whi71]. Ratio [MPS77]. Ray [CM75b]. Rayleigh [Mar74d]. RBCS [NP78]. RBCS/RCMAS [NP78]. RC [BKS71, FDDM75, Ric70]. RCMAS [NP78]. Real [DG75, MNP77, Phi75, Wit79]. Real-Time [DG75, MNP77, Phi75, Wit79]. Realizations [CR73a, CR73b]. Reallocation [BN73]. Rearrangeability [Ben75b, Ben75c, Hwa76]. Rearrangeable [Chu75, Hwa71, OTW71a, OTW71b]. Rearranged [Ben70]. Reassignment [PMN78]. Receiver [ALS71, DZ78, GS74a, GK76, HJ71, KN78, Per73c, Per73d, Sal74]. Receiver-Generator [GK76]. Receivers [Ben71b, Goe74a]. Receiving [ACH^{+}78, ACG78]. Reception [FG78, Fer72, Jay75a, Vig70]. Recognition [Lev78, LRF78, RW79, RS79, SR75]. Recording [May71]. Recovery [Fal76a, Fal76b, GS71a, GH75]. Rectangular [Abe78]. Recursive [GG70, LM79]. Reduce [ACG79, Col73, Man74]. Reduced [Mar76c]. Reducing [Ber70a]. Reduction [Ben78, HJ75, LS71, Mar74e, MS79b, PS70, PMN78, Raw70]. Redundancy [Neu71]. Reed [AOR76, PK70]. Reed-Contact [PK70]. Reflection [Dra77b]. Reflectivity [Zuc70b]. Reflectometer [Per77]. Reflector [Chu71a, DG79, Gan76, Sem79, Tho71a, Tur75]. Reflector-Type [Gan76]. Reflectors [Dra74]. Refractive [OP74]. Refractive-Index [OP74]. Regenerative [BBJH^{+}77]. Regenerator [MSW78]. Registers [BS72a]. Reinforced [Kle72, Kwe72]. Relating [FHH^{+}79]. Relation [Car71, Dra72a, Gum70, JB71]. Relationships [Ben71a]. Relative [SR74]. Relaxation [Aha78b]. Relay [Che71, GH70, HGLM77, MP71, Set71, Wag71]. Reliability [Aca77, Fan78, GH77, IL78, JPS^{+}78]. Remanent [AOR76]. Remendur [PB73]. Remote [BDJ79, BPW79, BGH79]. Remover [Fan76]. Remreed [AOR76, DL76, GAW^{+}76, GBN^{+}76, HW76, KLS76, Kle76, KU76]. Renewal [HH73, Hol73, KN72, Kuc72, NK73a, NK73b]. Repair [CRAC79]. Repeat [AEH^{+}70]. Repeater [AHK^{+}78, Bal76, BBJST77, BKY70, DPR77, Goe74b, LOOS77]. Repeatered [ACKL74, Bri78]. Repeaters [BBJH^{+}77, Cha70b, Per76]. Representation [Tho71c]. Representations [Liu70]. Reproducibility [DW78]. Required [GW79]. Requirement [Lee71]. Requirements [AE70, BEST78, Mes75, Sch71]. Research [Ele77]. Researches [Bel76]. Reserve [KFBL70]. Resins [Ven72]. Resist [MM79b]. Resistance [BM71]. Resolution [MM79b, PL71]. Resonances [Owe72]. Resonant [And75b, AP75]. Resonant-Grid [AP75]. Resonators [Arn70, CS77b, Mar76e, Zuc70b, Ano78n]. Resource [Aca79, Kuo75]. Response [Bel78, Boy76, CR73a, CR73b, CAK76, Coo77, Fer72, Glo73, GM73, HRC73, Kog76, Liu74, Mar73c, Mar73d, Mar74c, Mar74d, Mue75, NS72, Rab72, Ric70, San70b, Wil72]. Response-Bound [Liu74]. Responsible [DO79]. Restoring [Chu71b, Chu73]. Result [Pen70]. Resulting [OP74]. Results [KW78, Mit75, Per71b, SH74]. Retention [TKN74]. Retransmission [Yeh70]. Retrial [FR79]. Retrospect [Mes75]. Retrospective [Rit78]. Revolution [Kau74]. Ribbon [Mil76a, Sta74]. Ribbons [CRAC79, SP77]. Ring [GM73, HS71]. Ring-Shaped [GM73]. Ringer [HN78]. Rise [Lee75, Rai76]. Role [CHK^{+}74]. Rotation [Chu77]. Round [Mar70b, Mar70d, Mar73a]. Roundoff [CR73a, CR73b, Jac70]. Route [AOS70, AGW77]. Rubber [Beb72]. Rule [Ric73c]. Rules [HRC73]. Runs [NMW79, Ano79l]. Rural [GLW74].

S [Ano78n, CS77a, Cro78a, MSW78]. S.

[Ano75k, NK73a]. SAFEGUARD [Ano75n, BFR75, Con75, Cro75, DJ75, Dic75, DM75, DG75, Fre75, Gaw75, Hag75, HS75a, HJ75, Kuo75, Mes75, MW75, Nic75, Ols75, Phi75, Rif75, Van75a, Van75b, WP75]. Same [Wah77]. Sample [LP71]. Sampled [LW74, Mue73]. Sampled-Data [LW74]. Samples [BPM79, DSS78, PMAB79, SDS78]. Sampling [Bal76, Bob73, Des73]. Sandstorms [Chu79a]. Sandwich [SP77]. Satellite [AD78, Aca78, Aca79, AC78, GH70, Lun70, LC78, RY77, Rus78, Tur75]. Satellites [Aca77, ALW79, MP71, Ohm74]. Saturation [Maz78b, Mit77]. Scale [CR72, Sem79]. Scaled [SS71a]. Scaling [Sne71]. Scanner [CLT70, KLS76]. Scanning [AD78, ALW79, DG79, RY77]. Scanning-Beam [ALW79]. Scatter [Bul75]. Scattering [GH70, Las76, Mar71d, Mar72d, Mar73e, Mar74d, Mar76a, Mar76d, MC74, Mor79, OP74, Set71]. Scheme [Ber70a, Buc70, CS77a, Laa71, MBN79, WC73]. Schemes [Nol75a, Nol78]. Schmidt [Ano78n]. Schottky [Dra72b]. Sciences [Bel76]. Scientist [Ano77f]. Scientists [Ano78a]. Scrambler [Lee73]. Scramblers [GH75]. Scrambling [KJ77]. Sealed [AOR76]. Sealing [PB75]. Seals [DS70, SSV70]. Second [Bra70, Lue72, Ros72]. Second-Harmonic [Bra70]. Second-Order [Lue72]. Section [Mar74c]. Sectional [MSW77]. Segment [Kan70]. Seismic [LN72]. Selected [DC78]. Selection [FGW73, Gar72]. Selective [Bab72, LR79, Rum79, SG77]. Selectively [Bab73]. Self [And75b, Neu71]. Self-Resonant [And75b]. Self-Synchronizing [Neu71]. Semi [Gib78]. Semi-Markov [Gib78]. Semiconductor [BS70b, CS71, DPR77, FYMS78, Laa71, Lew70, LW70, Wah77, WM70]. Semideterminate [Aha79]. Semilattice [Ben73]. Sense [Dar70, Hwa79]. Sensitive [Bro70, Dah72]. Sensitivities [But71]. Separation [CM78, MS73b]. Separations [Bis71a]. Sequence [FM73]. Sequencer [ALW79]. Sequences [CH72, Jay74, Rif79]. Sequential [Gar72, Neu71]. Sequentially [Aha78c]. Serial [Lun70]. Series [Gol73, PK70]. Served [Coo70, Hef73]. Server [GM77]. Service [AHPR79, AFS79, BDJ79, BH71, BDRP79, Bob73, BJO78, BPW79, BGH79, CHPT79, CSL79, Col73, Cra71, DVW79, Des75, DPZ79, EFW79, Fis79, FP79, FR79, HS78b, Huf79, Kuc73a, Kue79, Mac79, Nea72, RSS79, SH79, SVW79, Tsi79, Wal79, You79, YBBW79]. Serving [Fre78]. Set [CSW71, Fis79, GS71b, Mea75, Rif79]. Sets [HY71a]. Several [Has72, MRM78, MRR79]. SG [AHK^{+}78, BEST78, BCD^{+}78, CGS^{+}78, CDD^{+}78, EFG78, FYMS78, HLL78, MAGS78]. Shape [GS70]. Shaped [GM73]. Shapes [MSW77]. Shaping [Mue73]. Shared [Aca79]. Sharing [Aca78, Bou78, CK78, Cro78c, DHM78, Fra78, Has72, JR78, JL78, KLO78, LMT78, LB78, Lyc78, LC78, MPT78, MCM78, NP78, Pek78, RT78, Rit78, RJLK78, Rov78, Tho78, WSS78]. Sheath [AGW77]. Sheathed [GGL^{+}77]. Shelf [HR73]. Shell [Bou78]. Shielded [Len77a, Mil72]. Shift [BS72a, Eis72, MCC73, MC73]. Shifted [Tow75]. Shifted-Companion-Form [Tow75]. Shifters [CRS75]. Shifting [KT73]. Shock [Liu74]. Shop [BP70]. Shore [LTV70]. Short [AR79, Bar71, BR74, Bro71c, FM75, MG75, Osb71, RSSF78, Sha72]. Short-Haul [Bro71c, MG75]. Short-Holding-Time [FM75]. Short-Hop [Sha72]. Short-Term [Bar71]. Short-Time [AR79]. Shot [Ric70]. Sided [Mea75]. Sight [Bab72, Bar70, Gil75, Gra70b, Rut71, SOP73]. Signal [ADM78, BZ74, BS70a, Ber70b, Blu72, Bos72, FGW73, Gre74, Jay74, KL73, Lin71, MM71b, MS79c, Nol75b, RP75, Wyn78]. Signal-to-Noise [Nol75b]. Signaling [BKT78, BPW79, CSL79, CMR78, CFMS78, CS78, DR78, KCM^{+}78, Lit78, Maz75a, MW78b, Mue75, RM78, Sna78]. Signals [AC78, AS70, Bri71, Bul71, BHO71, CHM73, Eis72, FG78, GHM73, Gla71a, Has75, JR72, Log77, Pea72, Pra76, RC76, Shn70, Tho73a]. Silence [RSA77]. Silica [Dab75, KA74, SDMS77]. Silicon [BS70a, CLB75, MHSS78, Mil78, Sch70, Sch78, Wag70]. Simple [Che76, CRSB77, Fri78, GW74, Per73a, Slo76, Som73]. Simplified [Oga77]. Simulation [LR79, SRB79, Wad78b]. Simulators [CES74]. Simultaneous [Fal76a, MM79a, Sem74]. Sine [Bud73]. Sine-Wave [Bud73]. Single [Aha72c, Ano79l, Arn74b, BKS71, CBK79, CT79, CMG73, Cum75, GM77, KA74, KG76, Laa70, Mar74f, Mar76b, Mar77b, NMW79, RSSF78, RC78, Som73, SH74]. Single-Chip [CT79]. Single-Error [Aha72c]. Single-Fiber [RC78]. Single-Integration [Cum75, Laa70]. Single-Material [Arn74b, KA74, Mar74f, SH74]. Single-Mode [CMG73, Mar76b, Mar77b, Som73]. Single-Substrate [BKS71]. Single-Tone [KG76]. Singularity [LW70]. Sinks [Arn74d]. Sinusoids [Gol74]. Site [EFW79]. Size [Jay75b]. Sizing [Els77]. Slab [Ano70k, Mar74a, Mar70c, Mar71b, Mar71d, Mar72e, Mar73b, Mar77c]. Slab-Coupled [Mar74a]. Slope [Gre73, JR71]. Slow [FGM78]. Small [And75a, Blu72, DG79, Las76]. Small-Signal [Blu72]. Smearing [Ber70a]. Smoothing [SG77]. Software [AHPR79, BDG^{+}77, BGH79, CCR^{+}77, CFG^{+}77, Con75, DG75, Fre75, MRY77, MW75, SVW79, Van75a, WP75]. Soil [Beb72, Con72a, Con72b, Dah72, DeC72, Kle72, Kwe72, Min72, Ven72]. Solidified [DSS78, SDS78]. Solids [Kau74]. Solution [KL73]. Solutions [Blu78, MM72a]. Some [Ber70a, Cox75, Des76, GS75, GH70, Hwa76, Jag74, Kai71, LS71, Maz71, Maz79b, MM72a, MNW78, Nag71, NPM77, Pra71, RKHD74, Tho71c, Wil71, Wil72]. Sound [DMK78]. Sounds [IF72]. Source [CM78, GW74, Lim73, Llo77, Wit79]. Sources [Mar75b, SW73]. Space [Bab73, Vig70, Whi75, Pen70, Rus78, Vig75, Wri71]. Space-Diversity [Vig70, Vig75]. Spacer [Sch78]. Spacing [Lee71, Lun71]. Spare [Sch71]. Spatial [CPS71, PL71]. Spatio [Bud73]. Spatio-Temporal [Bud73]. Speaker [MRM78, MRR79, RW79, Ros76, SR75]. Speaker-Independent [SR75]. Speaker-Trained [RW79]. Speaker-Verification [Ros76]. Special [Pfe79]. Specially [Aha75]. Speckle [Bur70]. Spectra [Gla71a, Liu71, PR74]. Spectral [Aca78, AR79, Bos72, Den76, Nea71b, Pra76]. Spectrum [Gre74, Jay77, Laa70, Maz79b, Rif79, RP75, Tho77b, Tho77a]. Speech [ACG79, Aha78a, Aha78b, Aha78c, ADM78, AS70, Bra72, CWF76, Cro77, CGRS77, CS77a, CJF73, DC78, DC79, FRSD72, FIS75, GMN76, JR71, JR72, Jay74, Jay75a, Jay77, JC79, KJ77, MSG78, MRM78, ML79, Nol75a, Nol75b, RSF71, RSA77, RSCG77, RSR74, SJ76, Sca79, SR71, SRH75, Sen70, SJS79, TNMC79, WHF77]. Speech-Encoding [Nol75b]. Speed [BKKM71, FH71]. Spelled [RS79]. Spherical [Tur75]. Spherical-Reflector [Tur75]. Spheroidal [Sle78]. Splice [CRAC79, Met79]. Splices [CGBS75, Dab75, DS73, GSBC73, Glo76, Mar77b, Mil75b, Mil76b, MM78b]. Splicing [CR76, Mil78, SBGC75]. Spoken [RS79]. Spot [Aca77, Aca78, RY77]. Spot-Beam [RY77]. Spread [Maz79b]. Spread-Spectrum [Maz79b]. Spreading [Arn74a, BM71]. Spurious [Doa78, MS79c, Sch74]. Square [FF73, Mar74c, Mue75, Sal73]. Squared [Cho74, GW79]. Squaring [Maz78a]. Stability [Bar71, Bau72, Ben71b, DeB72, Dra72b, Ger72, Maz78b, Rei72, San78b, Sha70, WR73]. Stable [Stu76a, Stu76b]. Stage [Hwa79]. Stalpeth [BY78]. Standards [Cra71]. Start [Cha71, CH72]. Start-Up [Cha71, CH72]. State [Bel78, BY78, Mar76e, MM78a]. States [Ben78, DMNT75]. Static [DeB72]. Station [CSW71, CSL79, GS71b, Lee71, LOOS77]. Stationary [IZ70, Sle72a]. Stations [BDK78, CFP70, Ohm74]. Statistical [BKS71, Bal76, BGM^{+}71, But71, CK71, Cha73, DC71, DL78, Eck78, Gol74, HY71a, Kar71, Lin71, Lin73, LF72, Log71, MCM78, O’N71, Ols71, RSA77, SWD71, SJS79, SK74]. Statistics [Fan76, Lin76b, Lin77a, Lin78, Per71b, Per71c]. Steady [Bel78, Mar76e, MM78a]. Steady-State [Bel78, Mar76e, MM78a]. Steep [MM79b]. Step [AEH^{+}70, Boy76, BN71, Jay75b, LMY76, Mar76b]. Step-and-Repeat [AEH^{+}70]. Step-by-Step [BN71, LMY76]. Step-Index [Mar76b]. Step-Size [Jay75b]. Steps [Ben71b]. Stess [TKN74]. Stimulated [Bri71]. Stochastic [FR79, Ger72, Hel79, Liu70]. Storage [KSBL74, MMF70, TKN74]. Store [AAS70, BCKM70]. Stored [Aha78a, BCKM70, DKP^{+}70]. Straightness [Glo75]. Streams [Nea71a]. Stress [BY78, ES77]. Stresses [Kau74]. Strip [Ram74, WM70]. Strip-Loaded [Ram74]. Stripline [SS74]. Structural [Kle72]. Structure [Cha71, CLB75, DG75, Mar73e, PB75, PSMO75, Wit79]. Structured [Nic75]. Structures [Bau72, Con72a, FM79, Kog76, LN72, Liu74, PB74, Sch73]. Studies [GS75, OTW71b, TKN74, YBBW79]. Study [Bab72, DM78, Gra70a, HS72a, Ken70, Mit75, Nol75a, O’N71, Ols71, Pen70, Spa76].

Studying [Fri78]. Stuffing [Buc70]. Styles [Rai79]. Sub [CWF76, Cro77, Cro78a]. Sub-Band [Cro78a, Cro77]. Sub-bands [CWF76]. Subharmonic [SS74]. Subjected [BY78]. Subjective [CB72, CHS76, DC78, GMN76, GSC^{+}79]. Submarine [AOS70, Bre70, Cha70b, Wah77]. Subsampling [Bob74, Pea72]. Subscriber [Fis79, Koo78b, Lap76]. Subscriber-Loop [Lap76]. Subspaces [Maz77]. Substrate [AD79, BKS71]. Substrates [MM79a, Mat70]. Subsystem [BPW79, CSL79, HS75a, HHM^{+}74, PSW70]. Sulfate [BW70a]. Sum [Gol74]. Sun [Lun70, Wri71]. Sun-Transit [Lun70]. Superconcentrators [Chu79b]. Superimposed [Kno77]. Supply [BNO74]. Support [BCKV74, Con75, Mar73e, Rov78, Sna78, Tsi79]. Suppression [Aha78b, BBL72, Gre74, Seq72, WN72]. Suppressor [MB71]. Surface [Ano78n, CS77b, DO79, KWP72, LW70, MSW77, WM70]. Surface-Acoustic-Wave [Ano78n, CS77b]. Surfaces [BM71, Gra72]. Surveillance [TAB74]. Survey [BKKM71, DT71, FH71]. Switch [Cro78b, Des76, PK70, RS76]. Switch-Count [Des76]. Switched [ABLR77, ADM78, BKKM71, DT71, FL79, FH71, FMSZ78, HS71, Lak79, Ric70]. Switches [GAW^{+}76, LW74, Maz71]. Switching [AOR76, AB78, BBJM^{+}77, BCD^{+}77, BS76, BGK72, BG72b, Bre71, BN71, BS71, BHKJ74, CHPT79, Che71, CH77, CGL71, DL76, Fan78, GAW^{+}76, GP71, GBN^{+}76, GS74b, HW76, HO77, KLS76, Kle76, KU76, Kro72, KT73, Kru76, LMY76, LM70, Laa71, OTW71a, OTW71b, Pie72, Sto76, Uri71, Wag71]. Switching-System [BHKJ74]. Symmetric [DSS78, Kur70]. Symmetrical [HL78]. Symmetries [Ben78]. Symmetry [Dra78, Goo74, RM72]. Synchronization [Cha72, SZ75]. Synchronized [Mor71]. Synchronizing [Neu71]. Synchronous [Ho71, MS75b]. Syntactic [Lev78]. Syntax [LRF78]. Synthesis [FIS75, IF72, Kan70, Lak79, Liu70, RSF71, Sze73, Zuc74]. System [AHPR79, AAS70, AOS70, AR79, ABH77, AE70, ALM74, AGW77, ACTW77, AHK^{+}78, ALS71, ACKL74, Ano74k, Ano75n, ACG78, AB78, BCS71, BBJH^{+}77, BGK70, BNO74, BZ74, BDJ79, BJL71, BKL75, BBCC75, BDRP79, BBJM^{+}77, Bou78, BCDT77, BS76, BW71, BCM^{+}74, Bre70, BEST78, BP70, Bro71a, BCD^{+}78, BPS75, BFR75, BBJST77, BPW79, BCKV74, BSS78, BM78, BDN^{+}77, BGH79, BHKJ74, BKY70, BWW74, CG70, CGS^{+}78, CGHW74, Cha72, CSW74, Cha70b, CHK^{+}74, CHM^{+}77, CSL79, CCWW74, CK78, Cok72, CGWN77, Con75, CDD^{+}78, Cro75, Cro78c, DJNP74, DR78, DJ75, DVW79, DW78, Dic75, Die71, Doa78, DHM78, DM75, DCRZ70, DG75, DSW71, DO79, DPR77, EFG78, Ele77, EG70, Fan78, Fen71, FR75, FP74, FHT77, FYMS78, Fra78, Fre75, FMSZ78, GOS74, Gaw75, GHWY77, GS73a, Giu71, Gol77]. System [GS74b, GGL^{+}77, GH77, Hag75, HS75a, HM74, HJ71, HNW77, HLL78, HL79, HW76, HS71, HR73, HST70, HP72, Her70, HO70, HY71b, HJ75, HHM^{+}74, HP70, Huf79, HJJ^{+}77, ILR75, Jac78, JJ70, JP71, JR78, JL78, KH74, KW78, KLO78, Koo78b, Kro72, Kuo75, LS70, LRF78, LOOS77, LW74, LMT78, LO77, Lun70, LB78, Lyc78, LC78, LTV70, MMM75, MSW78, Mau74, MRR79, MPT78, MCM78, MHSS78, Mes75, Mil70, MG78, Mor71, Mor70, MAGS78, MW75, MP78, NP78, Nic75, Ols71, Ols75, Pek78, Phi75, PR73, RY77, RSS79, Rif75, RT77, RT78, Rit78, RJLK78, Ros76, Rov78, RC78, RBH77, SZ75, SR75, SRB79, SRMC78, SW71a, Sha72, SCD78, SBR78b, Sna78, SK75, SH79, SVW79, Sto76, Sul71, TAB74, Tho78, Tho77a, Van75a]. System [Van75b, WMLT70, WR73, War77, Whi75, WP75, WA74, WSS78, YBBW79, And71, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Bai71, BH71, Bre71, Bro71a, Bro71c, BHP71, Bro71b, BHO71, BS71, CSW71, Cra71, Das71, Dor71, DPS71, Eck78, FG71, GS71b, GRW71, Kes71, KFBL70, MM71b, OBW70, Pfe79, SSV70, Uri71, WP71]. Systems [Aca78, Aca79, Aha79, ABG^{+}77, Aun73, Bre71, BN71, BCKV74, Bul75, Cap72, CW73, CH72, Che71, Chu71b, Chu73, CR71, CR72, CGRS77, Das71, Des75, DG75, Dra74, FF73, Fal76b, Fal78, Fos71, FGS75, FR79, GS71a, Gol73, Gop71, GRW70, GRW71, GH70, HHJ71, Hef73, Hen73a, Ho71, KN72, Kue79, LMY76, LC70, Log78, Lun71, MNS78, MP71, MRM78, MRR79, Mue73, Nea72, NK73a, NK73b, Nol75b, Osb71, Per73a, Per73b, Per73c, Per73d, Per74, Per76, Pra70, Pra71, RJ72, RSCG77, Rai73, Rao78, Rei72, Ric70, Ric73d, Rov78, Rut71, San78b, SS78, Sha70, SCD78, TG76, Tsi79, Van75b, VP79, YH71].

T [Lue72, Mos70]. T1 [Bri78]. T2 [O’N71]. Table [BG72b]. Takagi [CH78a]. Talker [BGM^{+}71, CHS76]. Tandem [CGRS77, Cro78a, GSC^{+}79, RSCG77]. Tangent [LC70]. Tap [GW79, Wyn75b]. Tap-Weight [GW79]. Tape [CGBS75, PSW70, Slo76]. Tapered [Ano70k, Mar70c]. Tapes [BS75a, Dah72, SBGC75]. Target [BS70a]. TD [Bro71a]. TD-2 [Bro71a]. TDMA [Aca79]. TE [CM77, Mar76a]. TE-Wave-Scattering [Mar76a]. Technique [AD76, Bis71b, CH74, Goo74, Jag78, MS75b, Pea72, PMN78, SBGC75]. Techniques [Aca78, Ano79l, Apr75, Bob74, BGM^{+}71, BS75b, Cus70, FHT77, HS72a, HMG70, MNW78, NMB76, NMB77, NMW79, Ols71, Rab72, RW79, Tho77b, Tho77a].

Technology [BCE^{+}77, KCM^{+}78, MM75]. Telecommunications [Ano76h, BKKM71, DT71, FH71]. Telephone [Ben71a, Bra71, Bri71, CSW71, CHS76, CHPT79, CRSB77, Das71, DO79, DMNT75, FRSD72, FR73b, FMSZ78, Has75, Hol71, Jag75, KN78, LM70, Liu74, Lue70b, Mea75, MS79c, Mor78a, Par78b, Pfe79, RSA77, Ros76, SRB79, SK74, Sul71, Wag71, Wal79]. Telephone-Quality [RSA77]. Telephones [BDK78, BBW78, BS78, BJO78, BDS78, CR78, FW78, FMSZ78, GR78, HS78b, MW78a, Rod78]. Telephony [Bel76, Jay75b, LP71]. Teletraffic [Wil71]. Teletypewriter [CFP70]. Television [Ber70b, CFHM71, CPS71, LPW74, NR79, PL71, Pea72]. Temperature [CF79, Daw72, Rai76, TKN74]. Temporal [Bob74, Bud73, PL71]. Ten [Wah77]. Tendem [Osb71]. TEon [Doa78]. Term [Bar71, Lin77a]. Terminal [BCD^{+}77, BCD^{+}78, CD79, Kur70, LTV70, Mau74]. Terminals [BDK78, BBW78, BS78, BJO78, BDS78, CR78, FW78, FGM78, FMSZ78, GR78, HS78b, MW78a, Rod78]. Terrain [Gil75]. Terrestrial [Bul75, GH70, HGLM77]. Test [ACTW77, BS76, Cha74, CHM^{+}77, DPZ79, Fis79, Gar72, HY71a, Len77a, MOR^{+}77, Wal79]. Testing [BGK70, BG72a, DM75, FR75, GS73b, HST70, HMSV77]. Tests [Beb72, Con72a, Con72b, Dah72, DeC72, HO70, Kle72, Kwe72, Min72, Ven72]. Tetrabasic [BW70a]. Text [MCM78]. TH [ALS71, BCS71, BJL71, BW71, CGL71, Die71, DSW71, Fen71, Giu71, HJ71, JP71, KS71, SW71a, SW71b]. TH-3 [ALS71, BCS71, BJL71, BW71, CGL71, Die71, DSW71, Fen71, Giu71, HJ71, JP71, KS71, SW71a, SW71b]. The-Off-The-Shelf [HR73]. Their [Dra72a, Owe72, RC78]. Theorem [Ben75c, CH78a, Kai71, SW73]. Theorems [San70a, San70b, Wil70b].

Theoretic [Ben75c]. Theoretical [Maz79b]. Theory [And75b, Arn74b, Ben75a, BR74, CR73b, Dra77a, FF73, Jag70, Len77b, Len77c, Log78, MG73, Mar73a, Mar74f, Mar75a, Mar76a, Maz71, Maz79a, Mes73a, Mes73b, NK73a, NK73b, Oga77, Smi79, Thi71, Tho74b, Tho74a, Wil71]. Thermal [Kai70, Kau74, Mat70]. Thick [Mar77c]. Thickness [MM79a]. Thin [Ker70, KS70, May71, Moo70]. Thinning [Wag70]. Third [DO79]. Third-Order [DO79]. Three [Blu72, Hwa79, Wil70a]. Three-Dimensional [Blu72]. Three-Radiometer [Wil70a]. Three-Stage [Hwa79]. Threshold [Bra72, Bud73]. TI [Mea75]. Tilt [Glo76]. Time [AD78, AR79, Bou78, Bur71, CK78, Cro78c, DHM78, DG75, Dut72, FM75, Fra78, GG70, GM77, Gra70b, Jay75a, JR78, JL78, KLO78, Lee75, Lin76a, Liu71, LW74, LMT78, LB78, Lyc78, LC78, Maz71, MPT78, MCM78, Mor78c, MNP77, Mue73, NP78, Pek78, Per71d, Per75, Per77, Phi75, RB76, RT78, Rit78, RJLK78, RT71, Rov78, RM72, Stu76a, Stu76b, Tho71b, Tho78, WHF77, Wit79, WSS78, Zuc74]. Time-Dependent [Per75]. Time-Diversity [Jay75a]. Time-Division [AD78, LW74, Maz71]. Time-Domain [Mue73, Per77]. Time-Interval [WHF77]. Time-Invariant [RT71]. Time-Sharing [Bou78, CK78, Cro78c, DHM78, Fra78, JR78, JL78, KLO78, LMT78, LB78, Lyc78, LC78, MPT78, MCM78, NP78, Pek78, RT78, Rit78, RJLK78, Rov78, Tho78, WSS78]. Time-Variable [RM72]. Time-Varying [Liu71]. Times [Coo70, Kuc73a]. Timing [Dut76, FG78, GS71a, GH75, Ho71, Maz75c]. Timing-Phase [FG78]. Tip [LM76]. TM [CM73]. Tolerance [BKS71]. Tolerances [Kar71]. Toll [AHPR79, BDRP79, BCD^{+}77, CSL79, CFMS78, DMNT75, Kes71, Kru76, YBBW79]. Tone [BKS71, Den76, GK76, Gue75, KG76, NMB76, RB76]. Tones [CRSB77, Maz78a, RV70]. Tools [CMR78, JL78, SVW79]. Touch [BKS71, GK76]. Touch-Tone [BKS71, GK76]. Tracker [Wri71]. Tract [FIS75, GS70]. Trade [Per74]. Trade-Off [Per74]. Traffic [AHPR79, Avi71, BDJ79, Ben70, Ben71a, BDRP79, BPW79, BGH79, CSL79, DVW79, Des73, Eis77, Els77, GLW74, GHHJ77, HS71, HH73, Hey75, HN76, Hil77, Jag75, Kau77, Mor78a, Nea71a, Nea72, NK73a, NK73b, RSS79, SH79, SVW79, Wol72, YBBW79]. Traffic-Measurement [NK73a, NK73b]. Trained [RW79]. Training [CH72, RSS79]. Transaction [BDK78, BBW78, BS78, BJO78, BDS78, CR78, FW78, FMSZ78, GR78, HS78b, MW78a, Rod78]. Transfer [Aun73, BS72a, BT73, JB71, Kra71, Lue70a, MS75a, San78a, SKCM78, SS71b, Tho73a, Tho73b, Tho74a, TKK73]. Transform [Jag78, MNP77, NPM77, NS79, SN79, RV70]. Transformation [Liu71]. Transformers [RS74]. Transient [San70b, Wag71]. Transistor [DO79, KS78, San70a, Wil70b]. Transistors [Blu72, DJNP74, GM71, Gum70, GP70, Per72, San70b, WMLT70]. Transit [Gra70b, Lun70]. Transit-Time [Gra70b]. Translation [TBTG71]. Transmission [ABH77, And75a, ACTW77, Ano74k, AFS79, BKKM71, Ber70b, Bob73, BCD^{+}77, Bra71, Bro71a, BCD^{+}78, Bro71c, Bro71b, BPW79, BHO71, CM77, CFP70, CM75a, Chu71a, CF79, Cok72, Daw72, Dra77b, DT71, Fal78, FM79, FH71, GOS74, GHWY77, GH75, GH77, GRW71, HS71, Hil71, HO70, HS72b, IS73, Kes71, KU76, LM70, Laa71, MSW78, Man78, Man74, MM73, MS79b, MRM78, MRR79, MM72a, Mea75, Mil72, Mil76b, O’N71, Ols71, Rai79, SSP71, Set70, SCD78, Sta74, SG77, SGM79, TAB74, VP79, WHF77, Wyn78, Lee69, Ano70l]. Transmission/Switching [BCD^{+}77]. Transmitted [Coh71]. Transmitter [ALS71, BB79, HJ71, Mea75, Sal74, SCD78].

Transmitters [CKS79]. Transmitting [CFHM71, Jay75b]. Transponders [AD78]. Transversal [GGO79]. Transverse [Arn74c, Arn74d, Arn74e, Arn75b, Mil76b]. TRAPATT [Sch70]. Trapezoidal [Ric73c]. Trapped [Arn74c]. Trapping [Peh70, Ton70]. Traveling [BW71]. Traveling-Wave [BW71]. Tread [Cro78b]. Treatment [Tho74b]. Trees [Mor78c]. Trends [Con72b]. Trials [OBW70]. Tripling [Aha75]. Tropospheric [Gra70b]. Truck [HN76]. Truncated [Mar70a]. Truncation [FM73]. Trunk [BDJ79, BPW79, BGH79, FG71, FM75, FHH^{+}79, GHRS77, HH73, Kau77, LMY76]. Trunking [Hol71]. Trunks [Bro71c, FM75, Kes71]. TSPS [Ano70n, BCKM70, BC70, CDH70, DKP^{+}70, HST70, JJ70, KPS70]. Tube [BW71, MM73, Mil75b, Per77]. Tuning [Bra73]. Tuning-Induced [Bra73]. Turn [Lee75]. Turn-on [Lee75]. Twin [Lue72, Mos70]. Twin-T [Lue72, Mos70]. Twisted [Gil79]. Twisting [ES77]. Two [BS75b, BM78, CRSB77, Chu71a, Chu71b, Chu73, CPS71, CR72, Dra74, Fal76b, FM79, FGW73, IF72, Jay77, KWP72, Mar71c, Mar72f, Maz77, MSW74, Mor79, MNW78, NMB76, NMB77, Per75, Ano70l, Lee69]. Two-Bit [Jay77]. Two-Dimensional [CPS71, CR72, Fal76b, FGW73, Mor79]. Two-Level [MNW78, NMB77]. Two-Mass [IF72]. Two-Mode [Mar72f]. Two-Phase [KWP72, MSW74]. Two-Tone [NMB76]. Two-Wire [FM79]. Type [Gan76, Hal75, TBCR72, WR73].

U.S.A. [DM78]. Ultralinear [DJNP74]. Unavailability [VP79]. Unbiased [AR79]. Unbounded [Wyn75a]. Uncertainty [Sle78]. Uncladded [Mar73e]. Unconditional [Bob74]. Undercut [TKK73]. Undercut-Isolated [TKK73]. Underplate [AD79]. Undersea [AHK^{+}78, BEST78, BCD^{+}78, CGS^{+}78, CDD^{+}78, DO79, EFG78, FYMS78, HLL78, MAGS78]. Unequal [Mil76b]. Unified [Kan70]. Uniform [Len77b, Len77c]. Uniformity [Pre76]. Uniformly [IS73, Man74]. Unit [ES77, Wal79]. United [DMNT75]. Units [Man74]. Universe [Lee73]. UNIX [Bou78, CK78, Cro78c, DHM78, Fra78, JR78, JL78, KLO78, LMT78, LB78, Lyc78, LC78, MPT78, MCM78, NP78, Pek78, RT78, Rit78, RJLK78, Rov78, Tho78, WSS78]. Unvoiced [RSA77]. Unwanted [Mar71a]. UPCO [Mes78]. Updating [SGM79]. Upon [And75a, MS79a]. Upper [Chi70, Maz75b, Wyn75a]. Usage [LMY76]. Use [Aha79, Col73, KFBL70, Lim73, Man74, Maz78b, Mea75, O’N71, Pir79, RV70, WHF77]. User [MMM75]. Using [AR79, Apr75, Bal76, Ber70b, Bob74, CH72, Cho72, Cho74, CPS71, DG79, Eis72, Fre78, FM71, Gib78, HY71a, KJ77, Kau77, LRF78, MMF70, MB71, Mit77, Mue73, NS76, Rab72, RSR74, Rus78, Sem74, WC73, Wri71]. Utility [CFG^{+}77]. Utilization [AD78, Hub73, Kuo75]. Utterance [RSR74]. Utterances [RS75].

V [Ano71k, Ano78n, PB75, Sle78]. Vacuum [CRAC79]. Vacuum-Assisted [CRAC79]. Value [Lin76b, Lin78]. Vapor [DW78]. Varactor [AI74]. Varactors [Aha75]. Variable [CS77a, DC79, RM72, WHF77, Zuc70b]. Variable-Band [CS77a]. Variable-Quality [WHF77]. Variables [Goo74, Ano70l, Lee69]. Variance [Col73, Des75, Des76, Mes78]. Variant [Tho71b]. Variation [Hil77]. Variations [Bul71, Gra70b, PB73]. Various [CDF^{+}78, MSW77, Nol75a, Rai79]. Varying [Liu71, Ric70]. Velocities [Aha72b]. Verification [ATS70, CM73, DVW79, MRM78, MRR79, Ros76]. Vertical [Bob74, Pea72]. Very [FMOT74, Las76]. Very-Small-Angle [Las76]. Via [MS76, AD78, Aca78, Rei72]. Vibration [IF77, Liu74]. Vibrations [Wil77]. Vibratory [Mor70]. Video [Ber70b, Bro71c, CSW71, Has75, LP71, MM71b]. Videotelephone [HS75b]. View [MMM75]. Viewer [Lim73]. Violation [Fan76]. Visible [Set70]. Visual [Bud73]. Vocal [FIS75, GS70, IF72, IF77]. Vocoder [GSC^{+}79, Sam75]. Voice [AFS79, Fla79, HS78a, Par78a]. Voice-Frequency [Par78a]. Voiceband [BKKM71, MS79c]. Voiced [IF72, RSA77]. Voiced-Unvoiced-Silence [RSA77]. Vol [Bel76]. Voltage [Dix76]. Volterra [Gol73, Ric73d, Tho71c]. vs [Llo77, LR79, Mil76b]. Vulcanized [Beb72].

W [Ano79l, TG76]. W. [Ano70l]. Waiting [Coo70, Dut72, Whi75]. Wall [Gea79]. Walsh [Ano79l]. Watt [MG75]. Wave [ABH77, Ano78n, BW71, Bud73, CWE^{+}78, CKK^{+}71, CS77b, GS74a, Mar76a, MC74, Ric73a, Sle78]. Waveform [GSC^{+}79, KJ77, TNMC79]. Waveforms [Dut76, JC79]. Waveforms-I [JC79]. Waveguide [Abe78, AS72, ABH77, AGW77, ACTW77, BBJH^{+}77, BBJM^{+}77, BCDT77, BBJST77, CD71, Car71, CM73, CD73, CM77, CHM^{+}77, Doa78, DPR77, Ele77, FHT77, GHWY77, GGL^{+}77, GH77, HNW77, JMM78, LOOS77, Mar71b, Mar72f, MM73, Mar74b, Mar77c, Ols71, Owe72, Ram74, RS78, RBH77, San78a, Tho77b, Tho77a, War77]. Waveguides [Arn70, FMOT74, Gan76, Glo70, Kai70, Mar74a, Mar70d, Mar70c, Mar71d, Mar71c, MM71a, Mar72d, Mar72e, Mar72g, Mar73b, Mar75a, Per71d, RW74, Ano70k]. Wavelength [CAK76]. Wavelengths [Chu74, Gol77]. Waves [DDH75, HMG70, MSW77]. Weakly [Sch72]. Wear [AD79]. Wedge [LM76]. Weight [GW79]. Well [Bel78].

Well-Balanced [Bel78]. Western [MP78]. Where [Cap72]. Which [GMS74, Hwa79, Mac70, Maz78b, Nea72]. Whole [BPM79, PMAB79]. Whole-Fiber [BPM79, PMAB79]. Wide [Hwa79]. Wideband [Apr75, BH71, Cho72, CGRS77, Pir79, RSCG77]. Wideband-to-Narrowband [RSCG77]. Width [WR73]. Windings [Wag71]. Wire [Beb72, Bel78, DeC72, FM79, Wyn75b]. Wire-Tap [Wyn75b]. Wiring [FRSD72, Rai79]. Without [Bar70]. Word [Buc70, Jay73, Lev78, LRF78, RW79]. Words [RSF71]. Work [Gib78]. Workbench [DHM78]. WT4 [ABH77, AGW77, ACTW77, BBJH^{+}77, BBJM^{+}77, BCDT77, BBJST77, CM77, CHM^{+}77, DPR77, Ele77, FHT77, GHWY77, GGL^{+}77, GH77, HNW77, LOOS77, RBH77, Tho77b, Tho77a, War77]. WT4/WT4A [ABH77]. WT4A [ABH77].

XII [Bel76]. XYTOLR [Yam72b].

Y-Junctions [Owe72]. Years

[Ano72k, Ano76h, Con72b, Wah77]. Yield [Kur70].

Zero [Kra71, Log77, MS73b, Mes73a]. Zero-Forcing [Mes73a]. Zone [Chu71a, Chu78, HL78, SDMS77]. Zone-Balanced [Chu78, HL78]. Zones [DSS78, SDS78].

References

Aitcheson:1970:NEAb

[AAS70] E. J. Aitcheson, C. F. Ault, and R. G. Spencer. No. 1 ESS ADF: Message store — A disk message system. The Bell System Technical Journal, 49(10): 2887–2913, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2887.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2887.pdf.

Avaneas:1978:LPE

[AB78] N. G. Avaneas and J. M.

Brown. Loop plant electronics: The loop switching system. The Bell System Technical Journal, 57(4):1157–1183, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-1157.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-1157.pdf.

Abele:1978:IPA

[Abe78] T. A. Abele. Inductive post arrays in rectangular waveguide. The Bell System Technical Journal, 57(3):577–594, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-577.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-577.pdf.

Ault:1977:PMS

[ABG^{+}77] C. F. Ault, J. H. Brewster, T. S. Greenwood, R. E. Haglund, W. A. Read, and M. W. Rolund. 1A processor: Memory systems. The Bell System Technical Journal, 56(2):181–205, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-181.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-181.pdf.

Alsberg:1977:WMW

[ABH77] D. A. Alsberg, J. C. Bankert, and P. T. Hutchison. WT4 millimeter waveguide system: The WT4/WT4A millimeter — wave transmission system. The Bell System Technical Journal, 56 (10):1829–1848, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-1829.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-1829.pdf.

Abate:1977:NES

[ABLR77] J. E. Abate, L. H. Brandenberg, J. C. Lawson, and W. L. Ross. No. 4 ESS: The switched digital network plan. The Bell System Technical Journal, 56(7):1297–1320, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1297.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1297.pdf.

Aitcheson:1970:NEAa

[AC70] E. J. Aitcheson and R. F.

Cook. No. 1 ESS ADF: Maintenance plan. The Bell System Technical Journal, 49(10): 2831–2856, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2831.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2831.pdf.

Arnold:1978:DDI

[AC78] H. W. Arnold and D. C. Cox.

Dependence of depolarization on incident polarization for 19-GHz satellite signals. The Bell System Technical Journal, 57(9):3267–3276, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3267.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3267.pdf.

Acampora:1977:RCM

[Aca77]  A. S. Acampora.  Reliabil-

ity considerations for multiplespot-beam communication satellites. The Bell System Technical Journal, 56(4):575–596, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-4-575.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-4-575.pdf.

Acampora:1978:SSH

[Aca78] A. S. Acampora. Spectral sharing in hybrid spot and area coverage satellite systems via channel coding techniques. The Bell System Technical Journal, 57(7):2613–2631, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-7-2613.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-7-2613.pdf.

Acampora:1979:SRT

[Aca79] A. S. Acampora. A shared resource TDMA approach to increase the rain margin of 12/14 GHz satellite systems. The Bell System Technical Journal, 58(9):2097–2111, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-9-2097.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-9-2097.pdf.

Arnold:1978:CEGb

[ACG78] H. W. Arnold, D. C. Cox, and D. A. Gray. COMSTAR experiment: The 19-GHz receiving system for an interim COMSTAR beacon propagation experiment at Crawford Hill. The Bell System Technical Journal, 57(5):1331–1339, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1331.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1331.pdf.

Adelmann:1979:AAR

[ACG79] H. W. Adelmann, Y. C. Ching, and B. Gotz. An ADPCM approach to reduce the bit rate of μ-law encoded speech. The Bell System Technical Journal, 58(7):1659–1671, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1659.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1659.pdf.

Arnold:1978:CEGa

[ACH^{+}78] H. W. Arnold, D. C. Cox, H. H.

Hoffman, R. H. Brandt, R. P. Leck, and M. F. Wazowicz. COMSTAR experiment: The 19and 28-GHz receiving electronics for the Crawford Hill COMSTAR beacon propagation experiment. The Bell System Technical Journal, 57(5):1289–1329, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1289.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1289.pdf.

Angell:1974:LSR

[ACKL74] H. E. Angell, Y. S. Cho, K. P. Kretsch, and M. M. Luniewicz. L5 system: Repeatered line. The Bell System Technical Journal, 53(10):1935–1985, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-1935.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-1935.pdf.

Anderson:1977:WMWb

[ACTW77] J. C. Anderson, J. W. Carlin, D. J. Thomson, and T. J. West. WT4 millimeter waveguide system: Field evaluation test — transmission medium achievements. The Bell System Technical Journal, 56(10): 2157–2178, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-2157.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-2157.pdf.

Arnaud:1976:NTM

[AD76] J. A. Arnaud and R. M. Derosier. Novel technique for measuring the index profile of optical fibers. The Bell System Technical Journal, 55(10):1489–1508, December 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol55/bstj55-10-1489.pdf;
http://www.alcatel-lucent.
com/bstj/vol55-1976/articles/
bstj55-10-1489.pdf.

Acampora:1978:EUS

[AD78] A. S. Acampora and B. R.

Davis. Efficient utilization of satellite transponders via time-division multibeam scanning. The Bell System Technical Journal, 57(8):2901–2914, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2901.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2901.pdf.

Antler:1979:WGE

[AD79] M. Antler and M. H. Drozdowicz. Wear of gold electrodeposits: Effect of substrate and of nickel underplate. The Bell System Technical Journal, 58(2):323–349, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-2-323.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-2-323.pdf.

Ahern:1978:SSP

[ADM78] W. C. Ahern, F. P. Duffy, and J. A. Maher. Speech signal power in the switched message network. The Bell System Technical Journal, 57(7):2695–2726, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2695.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2695.pdf.

Anderson:1970:SSO

[AE70] Cleo D. Anderson and Robert L. Easton. SF system: An overview: Requirements and performance. The Bell System Technical Journal, 49(5):605–630, May/June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-605.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-605.pdf.

Alles:1970:DPS

[AEH^{+}70] D. S. Alles, J. W. Elek, F. L. Howland, B. Nevis, R. J. Nielsen, W. A. Schlegel, J. G. Skinner, and C. E. Stout, Jr. Device photolithography: The step-and-repeat camera. The Bell System Technical Journal, 49(9):2145–2177, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2145.
pdf;   http://www.alcatel-

lucent.com/bstj/vol49-1970/
articles/bstj49-9-2145.pdf.

Anderson:1979:QMH

[AFG79] R. R. Anderson, G. J. Foschini, and B. Gopinath. A queuing model for a hybrid data multiplexer. The Bell System Technical Journal, 58 (2):279–300, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-2-279.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-2-279.pdf.

Arredondo:1979:AMP

[AFS79] G. A. Arredondo, J. C. Feggeler, and J. I. Smith. Advanced mobile phone service: Voice and data transmission. The Bell System Technical Journal, 58(1):97–122, January 1979. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol58/bstj58-1-
97.pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-1-97.pdf.

Anderson:1977:WMWa

[AGW77] J. C. Anderson, R. W. Gretter, and T. J. West. WT4 millimeter waveguide system: Route engineering and sheath installation. The Bell System Technical Journal, 56(10):1953–1982, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-1953.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-1953.pdf.

Albanese:1979:LAP

[AH79] A. Albanese and W. S. Holden. LED array package for optical data links. The Bell System Technical Journal, 58(3):713– 720, March 1979. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol58/bstj58-3-
                                        713.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-3-713.pdf.

Ahamed:1971:GCR

[Aha71] S. V. Ahamed. A general class of rate-change circuits. The Bell System Technical Journal, 50 (10):3177–3194, December 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3177.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3177.pdf.

Ahamed:1972:AMP

[Aha72a] S. V. Ahamed. Applications of multidimensional polynomial algebra to bubble circuits. The Bell System Technical Journal, 51(7):1559–1580, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1559.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1559.pdf.

Ahamed:1972:BJB

[Aha72b] S. V. Ahamed. B.S.T.J.

briefs: Extension of multidimensional polynomial algebra to domain circuits with multiple propagation velocities. The Bell System Technical Journal, 51(8):1919–1922, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1919.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1919.pdf.

Ahamed:1972:DEM

[Aha72c] S. V. Ahamed. The design and embodiment of magnetic domain encoders and single-error correcting decoders for cyclic block codes. The Bell System Technical Journal, 51(2):461–485, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-461.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-2-461.pdf.

Ahamed:1972:MPA

[Aha72d] S. V. Ahamed. Multidimensional polynomial algebra for bubble circuits. The Bell System Technical Journal, 51(7):1535–1558, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1535.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1535.pdf.

Ahamed:1975:ASD

[Aha75] S. V. Ahamed. Analysis of specially doped varactors for direct frequency tripling. The Bell System Technical Journal, 54(2):317–334, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-317.pdf.

Ahamed:1978:GEL

[Aha78a] S. V. Ahamed. Gradient encoding for low-bit-rate stored speech applications. The Bell System Technical Journal, 57(3):779–789, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-779.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-779.pdf.

Ahamed:1978:ICN

[Aha78b] S. V. Ahamed. Idle channel noise suppression by relaxation of binary ADM-encoded speech. The Bell System Technical Journal, 57(5):1699–1706, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1699.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1699.pdf.

Ahamed:1978:SCM

[Aha78c] S. V. Ahamed. Sequentially companded modulation for lowclock-rate speech codec applications. The Bell System Technical Journal, 57(3):765–778, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-765.pdf.

Ahamed:1979:NUL

[Aha79] S. V. Ahamed. The nature and use of limit cycles in determining the behavior of certain semideterminate systems. The Bell System Technical Journal, 58(8):1869–1884, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1869.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-8-1869.pdf.

Anderson:1978:SUC

[AHK^{+}78] C. D. Anderson, W. E. Hower, J. J. Kassig, V. M. Krygowski, R. L. Lynch, G. A. Reinold, and P. A. Yeisley. SG undersea cable system: Repeater and equalizer design and manufacture. The Bell System Technical Journal, 57(7):2355–2403, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-7-2355.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-7-2355.pdf.

Ahmari:1979:TSP

[AHPR79] R. Ahmari, J. C. Hsu, R. L. Potter, and S. C. Reed. Traffic service position system no. 1: Automated coin toll service: Software. The Bell System Technical Journal, 58(6):1251–1290, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1251.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1251.pdf.

Ahamed:1974:NFV

[AI74] S. V. Ahamed and J. C. Irvin. New frontiers of varactor harmonic power generation in the C-band. The Bell System Technical Journal, 53(9):1839–1843, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-9-1839.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-9-1839.pdf.

Albanese:1978:ABC

[Alb78]  A. Albanese.  An auto-

matic bias control (ABC) circuit for injection lasers. The Bell System Technical Journal, 57(5):1533–1544, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1533.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1533.pdf.

Anderson:1974:OLS

[ALM74] C. W. Anderson, L. J. Lanzerotti, and C. G. MacLennan. Outage of the L4 system and the geomagnetic disturbances of 4 August 1972. The Bell System Technical Journal, 53(9):1817–1837, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-9-1817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-9-1817.pdf.

Androski:1971:TMR

[ALS71] F. J. Androski, N. E. Lentz, and R. C. Salvage. TH-3 microwave radio system: 4A FM transmitter and receiver. The Bell System Technical Journal, 50(7):2249–2269, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2249.pdf.

Aranguren:1979:SDS

[ALW79] W. L. Aranguren, R. E. Langseth, and C. B. Woodworth. Sequencer designs for scanning-beam satellites. The Bell System Technical Journal, 58(9):1999–2011, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-9-1999.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-9-1999.pdf.

Amelio:1972:CMC

[Ame72]  G. F. Amelio.  Computer

modeling of charge-coupled device characteristics. The Bell System Technical Journal, 51(3):705–730, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-3-705.

pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-3-705.pdf.

Ashley:1970:DPC

[AMS70] F. R. Ashley, Miss E. B.

Murphy, and H. J. Savard, Jr. Device photolithography: A computer controlled coordinate measuring machine. The Bell System Technical Journal, 49(9):2193–2202, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2193.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2193.pdf.

Anderson:1971:PSP

[And71] H. P. Anderson. The Picturephone System: The 850A PBX. The Bell System Technical Journal, 50(2):585–604, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-585.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-585.pdf.

Anderson:1975:ESP

[And75a] I. Anderson. The effect of small phase errors upon transmission between confocal apertures. The Bell System Technical Journal, 54(4):783–795, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-4-783.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-4-783.pdf.

Anderson:1975:TSR

[And75b] I. Anderson. On the theory of self-resonant grids. The Bell System Technical Journal, 54(10): 1725–1731, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-10-1725.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol54-1975/articles/
                                        bstj54-10-1725.pdf.

Andrews:1978:LPE

[And78] F. T. Andrews, Jr. Loop plant electronics: Overview. The Bell System Technical Journal, 57(4):1025–1034, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-1025.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-1025.pdf.

Anonymous:1970:CIa

[Ano70a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(1):155–157, January 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol49/bstj49-1-155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-1-155.pdf.

Anonymous:1970:CIb

[Ano70b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(2):317–319, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-317.pdf.

Anonymous:1970:CIc

[Ano70c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(3):471–472, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-3-471.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-3-471.pdf.

Anonymous:1970:CId

[Ano70d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(4):585, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-4-585.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-4-585.pdf.

Anonymous:1970:CIe

[Ano70e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(5):911–918, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-5-911.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-5-911.pdf.

Anonymous:1970:CIf

[Ano70f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(6):1235–1238, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-1235.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1235.pdf.

Anonymous:1970:CIg

[Ano70g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(7): 1589–1594, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1589.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1589.pdf.

Anonymous:1970:CIh

[Ano70h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(8):1987–1990, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1987.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1987.pdf.

Anonymous:1970:CIi

[Ano70i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(9): 2405–2415, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2405.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2405.pdf.

Anonymous:1970:CIj

[Ano70j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 49(10):3009– 3019, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-3009.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-3009.pdf.

Anonymous:1970:EDM

[Ano70k] Anonymous. Erratum: Dietrich Marcuse, Radiation Losses of Tapered Dielectric Slab Waveguides, BSTJ 49(2) 273–290 (1970). The Bell System Technical Journal, 49(5):919, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-919.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-919.pdf.

See [Mar70c].

Anonymous:1970:EWC

[Ano70l] Anonymous. Erratum: W. C. Y. Lee, An extended correlation function of two random variables applied to mobile radio transmission, BSTJ 48(10) 3423–3440 (1969). The Bell System Technical Journal, 49(2):320, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-320.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-320.pdf.

See [Lee69].

Anonymous:1970:INE

[Ano70m] Anonymous. Index to no. 1 ESS ADF articles. The Bell System Technical Journal, 49(10): 3005–3007, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-3005.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-3005.pdf.

Anonymous:1970:ITN

[Ano70n] Anonymous. Index to TSPS no. 1 articles. The Bell System Technical Journal, 49(10): 2729–2731, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2729.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2729.pdf.

Anonymous:1971:CIa

[Ano71a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(1):209–211, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-209.pdf.

Anonymous:1971:CIb

[Ano71b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(2):701–709, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-701.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-701.pdf.

Anonymous:1971:CIc

[Ano71c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(3):1083–1088, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-1083.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-1083.pdf.

Anonymous:1971:CId

[Ano71d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(4):1495–1502, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1495.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1495.pdf.

Anonymous:1971:CIe

[Ano71e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(5):1691–1694, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1691.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1691.pdf.

Anonymous:1971:CIf

[Ano71f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(6):2079–2083, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-2079.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-2079.pdf.

Anonymous:1971:CIg

[Ano71g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(7): 2485–2493, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2485.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2485.pdf.

Anonymous:1971:CIh

[Ano71h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(8):2853–2856, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2853.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2853.pdf.

Anonymous:1971:CIi

[Ano71i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(9): 3071–3073, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-9-3071.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-9-3071.pdf.

Anonymous:1971:CIj

[Ano71j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 50(10):3271– 3273, December 1971. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3271.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3271.pdf.

Anonymous:1971:EGC

[Ano71k] Anonymous. Erratum: A. G. Cannone and D. O. Feder and R. V. Biagetti, Lead-Acid Battery: Positive Grid Design Principles, BSTJ 49(7) 1279–1303 (1970). The Bell System Technical Journal, 50(1):217, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-1-217.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-1-217.pdf.

See erratum [CFB70].

Anonymous:1972:CIa

[Ano72a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(1):317–320, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-1-317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-317.pdf.

Anonymous:1972:CIb

[Ano72b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(2):563–567, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-563.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-2-563.pdf.

Anonymous:1972:CIc

[Ano72c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(3):785–786, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-3-785.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-3-785.pdf.

Anonymous:1972:CId

[Ano72d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(4):947–949, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-947.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-4-947.pdf.

Anonymous:1972:CIe

[Ano72e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(5):1127, May/ June 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-5-1127.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-5-1127.pdf.

Anonymous:1972:CIf

[Ano72f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(6):1423–1426, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1423.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1423.pdf.

Anonymous:1972:CIg

[Ano72g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(7): 1631–1634, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-7-1631.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-7-1631.pdf.

Anonymous:1972:CIh

[Ano72h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(8):1913–1917, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1913.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1913.pdf.

Anonymous:1972:CIi

[Ano72i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(9): 2093–2094, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-9-2093.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-9-2093.pdf.

Anonymous:1972:CIj

[Ano72j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 51(10):2227– 2228, December 1972. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol51/bstj51-10-2227.pdf;
http://www.alcatel-lucent.
com/bstj/vol51-1972/articles/
bstj51-10-2227.pdf.

Anonymous:1972:FYB

[Ano72k] Anonymous. Fifty years of BSTJ. The Bell System Technical Journal, 51(6):1129–1132, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1129.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1129.pdf.

Anonymous:1973:CIa

[Ano73a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(1):143–146, January 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-1-143.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-1-143.pdf.

Anonymous:1973:CIb

[Ano73b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(2):263–264, February 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-2-263.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-2-263.pdf.

Anonymous:1973:CIc

[Ano73c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(3):449–451, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-449.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-3-449.pdf.

Anonymous:1973:CId

[Ano73d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(4):597–598, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-4-597.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-4-597.pdf.

Anonymous:1973:CIe

[Ano73e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(5):767–768, May/ June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-5-767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-5-767.pdf.

Anonymous:1973:CIf

[Ano73f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(6):1025–1028, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol52/bstj52-6-1025.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-6-1025.pdf.

Anonymous:1973:CIg

[Ano73g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(7): 1243–1247, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-7-1243.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-7-1243.pdf.

Anonymous:1973:CIh

[Ano73h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(8):1449–1451, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-8-1449.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-8-1449.pdf.

Anonymous:1973:CIi

[Ano73i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(9): 1669–1671, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-9-1669.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol52-1973/
articles/bstj52-9-1669.pdf.

Anonymous:1973:CIj

[Ano73j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 52(10):1903– 1906, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1903.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1903.pdf.

Anonymous:1974:CIa

[Ano74a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(1):171–176, January 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-1-171.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-1-171.pdf.

Anonymous:1974:CIb

[Ano74b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(2):391–394, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-391.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-391.pdf.

Anonymous:1974:CIc

[Ano74c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(3):553–555, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-3-553.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-3-553.pdf.

Anonymous:1974:CId

[Ano74d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(4):741–743, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-4-741.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-4-741.pdf.

Anonymous:1974:CIe

[Ano74e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(5):947–949, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-947.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-947.pdf.

Anonymous:1974:CIf

[Ano74f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(6):1175–1177, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-6-1175.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-6-1175.pdf.

Anonymous:1974:CIg

[Ano74g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(7): 1427–1429, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-7-1427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-7-1427.pdf.

Anonymous:1974:CIh

[Ano74h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(8):1677–1680, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1677.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1677.pdf.

Anonymous:1974:CIi

[Ano74i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(9): 1893–1896, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-9-1893.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-9-1893.pdf.

Anonymous:1974:CIj

[Ano74j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 53(10):2269– 2276, December 1974. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-2269.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-2269.pdf.

Anonymous:1974:LCC

[Ano74k] Anonymous. L5 coaxial-carrier transmission system, foreword. The Bell System Technical Journal, 53(10):1897–1899, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-1897.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-1897.pdf.

Anonymous:1975:CIa

[Ano75a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(1):203–206, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-1-203.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-203.pdf.

Anonymous:1975:CIb

[Ano75b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(2):467–470, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-467.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-467.pdf.

Anonymous:1975:CIc

[Ano75c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(3):663–666, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-3-663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-3-663.pdf.

Anonymous:1975:CId

[Ano75d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(4):807–809, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-4-807.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-4-807.pdf.

Anonymous:1975:CIe

[Ano75e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(5):965–969, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-5-965.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-5-965.pdf.

Anonymous:1975:CIf

[Ano75f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(6):1175–1178, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-6-1175.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-6-1175.pdf.

Anonymous:1975:CIg

[Ano75g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(7): 1353–1354, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-7-1353.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-7-1353.pdf.

Anonymous:1975:CIh

[Ano75h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(8):1503–1505, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-8-1503.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-8-1503.pdf.

Anonymous:1975:CIi

[Ano75i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(9): 1663–1664, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-9-1663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-9-1663.pdf.

Anonymous:1975:CIj

[Ano75j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 54(10):1795– 1796, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-1795.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-1795.pdf.

Anonymous:1975:EMD

[Ano75k] Anonymous. Erratum: S. E. Miller, Delay Distortion in Generalized Lens-Like Media, BSTJ 53(2) 177–193 (1974). The Bell System Technical Journal, 54(1):207, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-1-207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-207.pdf.

See [Mil74].

Anonymous:1975:G

[Ano75l] Anonymous. Glossary. The Bell System Technical Journal, 54 (10):S261, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S261.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S261.pdf.

Anonymous:1975:LC

[Ano75m] Anonymous. List of contributors. The Bell System Technical Journal, 54(10):S263– S269, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S263.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S263.pdf.

Anonymous:1975:SDP

[Ano75n] Anonymous. SAFEGUARD data-processing system, preface. The Bell System Technical Journal, 54(10):S1, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-10-S1.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-10-S1.pdf.

Anonymous:1976:ABS

[Ano76a] Anonymous. Abstracts of Bell System papers appearing in other publications. The Bell System Technical Journal, 55(4):499–501, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-4-499.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-4-499.pdf.

Anonymous:1976:APBa

[Ano76b] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 55(5):687–689, May/ June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-687.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-687.pdf.

Anonymous:1976:APBb

[Ano76c] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 55(6):827–830, July/ August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-6-827.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-6-827.pdf.

Anonymous:1976:APBc

[Ano76d] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 55(7):1007–1010, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-7-1007.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-7-1007.pdf.

Anonymous:1976:APBd

[Ano76e] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 55(8):1221–1224, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-8-1221.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol55-1976/
articles/bstj55-8-1221.pdf.

Anonymous:1976:APBe

[Ano76f] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 55(9):1441–1443, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-9-1441.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-9-1441.pdf.

Anonymous:1976:APBf

[Ano76g] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 55 (10):1577–1579, December 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol55/bstj55-10-1577.pdf;
http://www.alcatel-lucent.
com/bstj/vol55-1976/articles/
bstj55-10-1577.pdf.

Anonymous:1976:BSC

[Ano76h] Anonymous. Bell System centennial: 100 years of publishing on telecommunications. The Bell System Technical Journal, 55(3):273–276, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol55/bstj55-3-273.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-3-273.pdf.

Anonymous:1976:CIa

[Ano76i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(1):127–128, January 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-1-127.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-1-127.pdf.

Anonymous:1976:CIb

[Ano76j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(2):271–272, February 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-2-271.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-2-271.pdf.

Anonymous:1976:CIc

[Ano76k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(3):371–372, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-3-371.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-3-371.pdf.

Anonymous:1976:CId

[Ano76l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(4):497–498, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-4-497.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-4-497.pdf.

Anonymous:1976:CIe

[Ano76m] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(5):681–685, May/ June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-681.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-681.pdf.

Anonymous:1976:CIf

[Ano76n] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(6):825–826, July/ August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-6-825.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-6-825.pdf.

Anonymous:1976:CIg

[Ano76o] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(7):

1003–1005, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-1003.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-1003.pdf.

Anonymous:1976:CIh

[Ano76p] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(8):1219–1220, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1219.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1219.pdf.

Anonymous:1976:CIi

[Ano76q] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(9): 1437–1439, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-9-1437.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-9-1437.pdf.

Anonymous:1976:CIj

[Ano76r] Anonymous. Contributors to this issue. The Bell System Technical Journal, 55(10):1573– 1575, December 1976. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol55/bstj55-10-1573.pdf;
http://www.alcatel-lucent.
com/bstj/vol55-1976/articles/
bstj55-10-1573.pdf.

Anonymous:1977:APBa

[Ano77a] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 56(1):117–118, January 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-1-117.pdf.

Anonymous:1977:APBb

[Ano77b] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 56(2):325–327, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-325.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-325.pdf.

Anonymous:1977:APBc

[Ano77c] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 56(4):641–644, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-4-641.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-4-641.pdf.

Anonymous:1977:APBd

[Ano77d] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 56(5):831–833, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-5-831.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-5-831.pdf.

Anonymous:1977:APBe

[Ano77e] Anonymous. Abstracts of papers by Bell System authors published in other journals. The Bell System Technical Journal, 56(6):1009–1012, July/ August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-6-1009.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-6-1009.pdf.

Anonymous:1977:BLS

[Ano77f] Anonymous. Bell Laboratories scientist named Nobel Laureate. The Bell System Technical Journal, 56(9):i–iii, November 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol56/bstj56-9-
i.pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-i.pdf.

Anonymous:1977:CIa

[Ano77g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(1):115–116, January 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-1-115.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-1-115.pdf.

Anonymous:1977:CIb

[Ano77h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(2):317–324, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-317.pdf.

Anonymous:1977:CIc

[Ano77i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(3):483–485, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-3-483.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-3-483.pdf.

Anonymous:1977:CId

[Ano77j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(4):637–639, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-4-637.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-4-637.pdf.

Anonymous:1977:CIe

[Ano77k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(5):827–829, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-5-827.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-5-827.pdf.

Anonymous:1977:CIf

[Ano77l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(6):1007–1008, July/ August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-6-1007.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol56-1977/
articles/bstj56-6-1007.pdf.

Anonymous:1977:CIg

[Ano77m] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(7): 1321–1331, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1321.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1321.pdf.

Anonymous:1977:CIh

[Ano77n] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(8):1561–1564, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1561.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1561.pdf.

Anonymous:1977:CIi

[Ano77o] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(9): 1817–1820, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1817.pdf.

Anonymous:1977:CIj

[Ano77p] Anonymous. Contributors to this issue. The Bell System Technical Journal, 56(10):2197– 2207, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-2197.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-2197.pdf.

Anonymous:1977:G

[Ano77q] Anonymous. Glossary. The Bell System Technical Journal, 56(2):313–315, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-2-313.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-2-313.pdf.

Anonymous:1977:PBLa

[Ano77r] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 56(7):1333–1336, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-7-1333.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-7-1333.pdf.

Anonymous:1977:PBLb

[Ano77s] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 56(8):1565–1567, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1565.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1565.pdf.

Anonymous:1977:PBLc

[Ano77t] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 56(9):1821–1823, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1821.pdf.

Anonymous:1977:PBLd

[Ano77u] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 56 (10):2208, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2208.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2208.pdf.

Anonymous:1978:BLS

[Ano78a] Anonymous. Bell Laboratories scientists named 1978 Nobel Prize Laureates. The Bell System Technical Journal, 57(9):i–ii, November 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol57/bstj57-9-
                                        i.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-9-i.pdf.

Anonymous:1978:CIh

[Ano78b] Anonymous. Contributors of this issue. The Bell System Technical Journal, 57(7): 2565–2573, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-7-2565.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-7-2565.pdf.

Anonymous:1978:CIa

[Ano78c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(1):211–215, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-1-211.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-211.pdf.

Anonymous:1978:CIb

[Ano78d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(2):467–473, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-2-467.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-2-467.pdf.

Anonymous:1978:CIc

[Ano78e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(3):791–795, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-791.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-791.pdf.

Anonymous:1978:CId

[Ano78f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(4):1225–1230, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-1225.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-1225.pdf.

Anonymous:1978:CIe

[Ano78g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(5):1707–1715, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1707.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1707.pdf.

Anonymous:1978:CIf

[Ano78h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(6):1889–1895, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1889.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1889.pdf.

Anonymous:1978:CIg

[Ano78i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(6):2305–2312, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2305.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2305.pdf.

Anonymous:1978:CIi

[Ano78j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(7): 2765–2767, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2765.pdf.

Anonymous:1978:CIj

[Ano78k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(8):3091–3096, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-3091.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-3091.pdf.

Anonymous:1978:CIk

[Ano78l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(9): 3317–3320, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3317.pdf.

Anonymous:1978:CIl

[Ano78m] Anonymous. Contributors to this issue. The Bell System Technical Journal, 57(10):3531– 3537, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/

                                        Vol57/bstj57-10-3531.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol57-1978/articles/
                                        bstj57-10-3531.pdf.

Anonymous:1978:EPC

[Ano78n] Anonymous. Errata: P. S. Cross and R. V. Schmidt, Coupled Surface-Acoustic-Wave Resonators, BSTJ 56(8) 1447–1482 (1977). The Bell System Technical Journal, 57(1):217–218, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-1-217.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-217.pdf.

See [CS77b].

Anonymous:1978:PBLa

[Ano78o] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 57(1):219, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-1-219.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-219.pdf.

Anonymous:1978:PBLb

[Ano78p] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 57(2):475–477, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-2-475.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-2-475.pdf.

Anonymous:1978:PBLc

[Ano78q] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 57(9):3321–3323, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3321.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3321.pdf.

Anonymous:1978:PBLd

[Ano78r] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 57(10): 3539–3541, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3539.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3539.pdf.

Anonymous:1979:CIa

[Ano79a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(1):271–275, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol58/bstj58-1-271.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-271.pdf.

Anonymous:1979:CIb

[Ano79b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(2):539–543, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-539.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-539.pdf.

Anonymous:1979:CIc

[Ano79c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(3):799–803, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-3-799.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-3-799.pdf.

Anonymous:1979:CId

[Ano79d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(4):953–956, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-4-953.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-4-953.pdf.

Anonymous:1979:CIe

[Ano79e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(5):1101–1102, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-5-1101.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-5-1101.pdf.

Anonymous:1979:CIf

[Ano79f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(6):1359–1367, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1359.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1359.pdf.

Anonymous:1979:CIg

[Ano79g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(6):1541–1543, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1541.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1541.pdf.

Anonymous:1979:CIh

[Ano79h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(7):

1719–1723, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1719.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1719.pdf.

Anonymous:1979:CIi

[Ano79i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(8):1895–1898, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1895.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-8-1895.pdf.

Anonymous:1979:CIj

[Ano79j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(9): 2129–2133, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-9-2129.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-9-2129.pdf.

Anonymous:1979:CIk

[Ano79k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 58(10):2317– 2320, December 1979. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol58/bstj58-10-2317.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-10-2317.pdf.

Anonymous:1979:ENF

[Ano79l] Anonymous. Errata: A. N. Netravali and F. W. Mounts and K. A. Walsh, Adaptation of Ordering Techniques for Facsimile Pictures with No Single Element Runs, BSTJ 58(4) 857–865 (1979). The Bell System Technical Journal, 58(6):1549, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1549.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1549.pdf.

See [NMW79].

Anonymous:1979:PBLa

[Ano79m] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(1):277–278, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-1-277.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-1-277.pdf.

Anonymous:1979:PBLb

[Ano79n] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(2):545–547, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-545.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-545.pdf.

Anonymous:1979:PBLc

[Ano79o] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(3):805–807, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-3-805.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-3-805.pdf.

Anonymous:1979:PBLd

[Ano79p] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(4):957–958, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-4-957.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-4-957.pdf.

Anonymous:1979:PBLe

[Ano79q] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(5):1103–1105, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-5-1103.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-5-1103.pdf.

Anonymous:1979:PBLf

[Ano79r] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(6):1545–1548, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1545.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1545.pdf.

Anonymous:1979:PBLg

[Ano79s] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(7):1725–1733, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1725.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1725.pdf.

Anonymous:1979:PBLh

[Ano79t] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(8):1899–1908, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-8-1899.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-8-1899.pdf.

Anonymous:1979:PBLi

[Ano79u] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(9):2135–2141, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-9-2135.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-9-2135.pdf.

Anonymous:1979:PBLj

[Ano79v] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 58(10): 2321–2326, December 1979. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol58/bstj58-10-2321.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol58-1979/articles/
                                        bstj58-10-2321.pdf.

Archer:1976:RSN

[AOR76] W. E. Archer, K. M. Olsen, and P. W. Renaut. Remreed switching networks for no. 1 and no. 1A ESS: Development of a remanent reed sealed contact. The Bell System Technical Journal, 55(5):511–535, May/ June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-511.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-511.pdf.

Allen:1970:SSS

[AOS70] R. C. Allen, D. H. Ort, and L. M. Schindel. SF system: Submarine cable route engineering. The Bell System Technical Journal, 49(5):767–781, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-767.pdf.

Arnaud:1975:RGQ

[AP75] J. A. Arnaud and F. A.

Pelow. Resonant-grid quasioptical diplexers. The Bell System Technical Journal, 54 (2):263–283, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-263.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-263.pdf.

Aprille:1975:WAD

[Apr75] T. J. Aprille, Jr. Wideband amplifier design using major multiloop feedback techniques. The Bell System Technical Journal, 54(7):1253–1275, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-7-1253.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-7-1253.pdf.

Allen:1979:USE

[AR79] J. B. Allen and L. R. Rabiner. Unbiased spectral estimation and system identification using shorttime spectral analysis methods. The Bell System Technical Journal, 58(8):1743–1763, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-8-1743.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-8-1743.pdf.

Arnaud:1970:NOW

[Arn70] J. A. Arnaud. Nonorthogonal optical waveguides and resonators. The Bell System Technical Journal, 49(9):2311–2348, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2311.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol49-1970/
articles/bstj49-9-2311.pdf.

Arnaud:1974:PSM

[Arn74a] J. A. Arnaud. Pulse spreading in multimode, planar, optical fibers. The Bell System Technical Journal, 53(8):1599–1618, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1599.pdf.

Arnaud:1974:TSM

[Arn74b] J. A. Arnaud. Theory of the single-material, helicoidal fiber. The Bell System Technical Journal, 53(8):1643–1656, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1643.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1643.pdf.

Arnaud:1974:TCFa

[Arn74c] J. A. Arnaud. Transverse coupling in fiber optics Part I: Coupling between trapped modes. The Bell System Technical Journal, 53(2):217–224, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-217.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-2-217.pdf.

Arnaud:1974:TCFb

[Arn74d] J. A. Arnaud. Transverse coupling in fiber optics Part II: Coupling to mode sinks. The Bell System Technical Journal, 53(4):675–696, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-4-675.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-4-675.pdf.

Arnaud:1974:TCFc

[Arn74e] J. A. Arnaud. Transverse coupling in fiber optics Part III: Bending losses. The Bell System Technical Journal, 53(7):1379–1394, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-7-1379.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-7-1379.pdf.

Arnaud:1975:PBM

[Arn75a] J. A. Arnaud. Pulse broadening in multimode optical fibers. The Bell System Technical Journal, 54(7):1179–1205, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-7-1179.

pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1179.pdf.

Arnaud:1975:TCF

[Arn75b] J. A. Arnaud. Transverse coupling in fiber optics Part IV: Crosstalk. The Bell System Technical Journal, 54 (8):1431–1450, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-8-1431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-8-1431.pdf.

Atal:1970:APC

[AS70] B. S. Atal and M. R. Schroeder.

Adaptive predictive coding of speech signals. The Bell System Technical Journal, 49 (8):1973–1986, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1973.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1973.pdf.

Acampora:1972:WBE

[AS72] A. S. Acampora and P. T.

Sproul. Waveguide breakdown effects at high average power and long pulse length. The Bell System Technical Journal, 51(9):2065–2091, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-9-2065.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-9-2065.pdf.

Aughenbaugh:1978:LPM

[AS78] G. W. Aughenbaugh and H. T. Stump. Loop plant modeling: The facility analysis plan: New methodology for improving loop plant operations. The Bell System Technical Journal, 57(4):999–1024, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-999.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-999.pdf.

Amelio:1970:BJB

[ATS70] G. F. Amelio, M. F. Tompsett, and G. E. Smith. B.S.T.J. briefs: Experimental verification of the charge coupled device concept. The Bell System Technical Journal, 49(4):593–600, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-4-593.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-4-593.pdf.

Aung:1973:HTE

[Aun73] W. Aung. Heat transfer in electronic systems with emphasis on asymmetric heating. The Bell System Technical Journal, 52(6):907–925, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-907.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-907.pdf.

AviItzhak:1971:HTC

[Avi71] B. Avi Itzhak. Heavy traffic characteristics of a circular data network. The Bell System Technical Journal, 50(8):2521–2549, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2521.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2521.pdf.

Babler:1972:SFS

[Bab72]  G. M. Babler.  A study

of frequency selective fading for a microwave line-of-sight narrowband radio channel. The Bell System Technical Journal, 51(3):731–757, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-3-731.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-3-731.pdf.

Babler:1973:SFN

[Bab73]  G. M. Babler.  Selectively

faded nondiversity and space diversity narrowband microwave radio channels. The Bell System Technical Journal, 52 (2):239–261, February 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-2-239.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-2-239.pdf.

Baird:1971:PSF

[Bai71] Jack A. Baird. The Picturephone System: Foreword. The Bell System Technical Journal, 50(2):219–220, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-219.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-219.pdf.

Balaban:1976:SEE

[Bal76] P. Balaban. Statistical evaluation of the error rate of the fiberguide repeater using importance sampling. The Bell System Technical Journal, 55(6):745–766, July/August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-6-745.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol55-1976/
articles/bstj55-6-745.pdf.

Barnett:1970:MLS

[Bar70]  W. T. Barnett.  Microwave

line-of-sight propagation with and without frequency diversity. The Bell System Technical Journal, 49(8):1827–1871, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1827.pdf.

Barber:1971:STF

[Bar71] Raymond E. Barber. Short-term frequency stability of precision oscillators and frequency generators. The Bell System Technical Journal, 50(3):881–915, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-881.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-881.pdf.

Barnett:1972:MPG

[Bar72]  W. T. Barnett.  Multipath

propagation at 4, 6, and 11 GHz. The Bell System Technical Journal, 51(2):321–361, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-321.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-321.pdf.

Baumwolspiner:1972:SCN

[Bau72]  M. Baumwolspiner.  Stabil-

ity considerations in nonlinear feedback structures as applied to active networks. The Bell System Technical Journal, 51(9):2029–2063, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-9-2029.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-9-2029.pdf.

Bell:1978:LPM

[BB78] W. N. Bell and S. Blum. Loop plant modeling: A model of cable pairs added at the main frames for a large entity. The Bell System Technical Journal, 57(4):849–868, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-849.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-849.pdf.

Baumhauer:1979:EET

[BB79] J. C. Baumhauer and A. M. Brzezinski. The EL2 electret transmitter: Analytical modeling, optimization, and design. The Bell System Technical Journal, 58(7):1557–1578, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1557.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1557.pdf.

Bobeck:1979:CAM

[BBB^{+}79] A. H. Bobeck, S. L. Blank, A. D. Butherus, F. J. Ciak, and W. Strauss. Currentaccess magnetic bubble circuits. The Bell System Technical Journal, 58(6):1453–1540, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1453.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1453.pdf.

Benowitz:1975:DDS

[BBCC75] P. Benowitz, S. J. Butterfield, M. P. Cichetti, Jr., and T. G. Cross. Digital data system: Digital multiplexers. The Bell System Technical Journal, 54(5):893–918, May/June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-5-893.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-5-893.pdf.

Barnes:1977:WMW

[BBJH^{+}77] C. E. Barnes, P. Brostrup-Jensen, E. T. Harkless, R. W. Muise, and A. J. Nardi. WT4 millimeter waveguide system: Regenerative repeaters. The Bell System Technical Journal, 56 (10):2055–2075, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2055.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2055.pdf.

Bonomi:1977:WMW

[BBJM^{+}77] M. J. Bonomi, P. Brostrup-Jensen, D. R. Marcotte, R. W. Muise, and P. J. Tu. WT4 millimeter waveguide system: Protection switching, auxiliary communication, and maintenance. The Bell System Technical Journal, 56(10):2103–2117, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2103.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2103.pdf.

Brown:1977:WMW

[BBJST77] R. J. Brown, P. Brostrup-Jensen, J. J. Schottle, and P. J. Tu. WT4 millimeter waveguide system: Line and repeater equalization. The Bell System Technical Journal, 56(10):2077–2088, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2077.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2077.pdf.

Bobeck:1972:BJB

[BBL72] A. H. Bobeck, S. L. Blank, and H. J. Levinstein. B.S.T.J. briefs: Multilayer epitaxial garnet films for magnetic bubble devices — hard bubble suppression. The Bell System Technical Journal, 51(6):1431–1435, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1431.pdf.

Banks:1978:TNT

[BBW78] D. D. Banks, G. A. Bhat, and J. W. Wesner. Transaction network, telephones, and terminals: Physical design. The Bell System Technical Journal, 57(10): 3503–3515, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3503.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3503.pdf.

Baldinger:1970:TNP

[BC70] W. R. Baldinger and G. T. Clark.

TSPS no. 1: Physical design 2688 and 2689. The Bell System Technical Journal, 49(10): 2685–2709, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2685.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2685.pdf.

Boyle:1977:NET

[BCD^{+}77] J. F. Boyle, J. R. Colton, C. L. Dammann, B. J. Karafin, and H. Mann. No. 4 ESS: Transmission/switching interfaces and toll terminal equipment. The Bell System Technical Journal, 56(7):1057–1097, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1057.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1057.pdf.

Brouant:1978:SUC

[BCD^{+}78] M. Brouant, C. Chalhoub, P. Delage, D. N. Harper, H. Soulier, and R. L. Lynch. SG undersea cable system: Terminal transmission equipment. The Bell System Technical Journal, 57(7):2471–2496, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2471.
pdf;   http://www.alcatel-

lucent.com/bstj/vol57-1978/
articles/bstj57-7-2471.pdf.

Boyd:1977:WMW

[BCDT77] R. J. Boyd, Jr., W. E. Cohen, W. P. Doran, and R. D. Tuminaro. WT4 millimeter waveguide system: Waveguide design and fabrication. The Bell System Technical Journal, 56(10): 1873–1897, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-1873.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-1873.pdf.

Becker:1977:PTP

[BCE^{+}77] J. O. Becker, J. G. Chevalier, R. K. Eisenhart, J. H. Forster, A. W. Fulton, and W. L. Harrod. 1A processor: Technology and physical design. The Bell System Technical Journal, 56(2):207–236, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-207.pdf.

Baker:1970:TNS

[BCKM70] W. A. Baker, G. A. Culp, G. W.

Kinder, and F. H. Myers. TSPS no. 1: Stored program control no. 1A store. The Bell System Technical Journal, 49(10):

2509–2560, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2509.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2509.pdf.

Buck:1974:AIS

[BCKV74] I. D. Buck, G. D. Crudup, C. W. Kemp, and G. C. Vogel. Automatic intercept system: Development support systems. The Bell System Technical Journal, 53(1):155–170, January 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-1-155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-1-155.pdf.

Brendel:1974:AIS

[BCM^{+}74] P. J. Brendel, W. K. Comella, R. N. Markson, P. J. Moylan, and J. Orost. Automatic intercept system: Peripheral circuits. The Bell System Technical Journal, 53(1):71–106, January 1974. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol53/bstj53-1-
71.pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-1-71.pdf.

Balicki:1971:TMR

[BCS71] J. F. Balicki, P. S. Chun, and F. J. Salvo. TH-3 microwave radio system: Power system. The Bell System Technical Journal, 50(7):2235–2248, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2235.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2235.pdf.

Bowman:1977:PMS

[BDG^{+}77] P. W. Bowman, M. R. Dubman, F. M. Goetz, R. F. Kranzmann, E. H. Stredde, and R. J. Watters. 1A processor: Maintenance software. The Bell System Technical Journal, 56(2):255–287, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-255.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-255.pdf.

Bauman:1979:TSP

[BDJ79] S. M. Bauman, R. S. DiPietro, and R. J. Jaeger, Jr. Traffic service position system no. 1: Remote trunk arrangement: Overall description and operational characteristics. The Bell System Technical Journal, 58(6):1119–1135, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1119.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1119.pdf.

Baker:1978:TNT

[BDK78] W. E. Baker, R. M. Dudonis, and J. H. Kee. Transaction network, telephones, and terminals: Transaction stations. The Bell System Technical Journal, 57 (10):3487–3502, December 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol57/bstj57-10-3487.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol57-1978/articles/
                                        bstj57-10-3487.pdf.

Budlong:1977:PCS

[BDN^{+}77] A. H. Budlong, B. G. DeLugish, S. M. Neville, J. S. Nowak, J. L. Quinn, and F. W. Wendland. 1A processor: Control system. The Bell System Technical Journal, 56(2):135–179, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-2-135.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-2-135.pdf.

Berger:1979:TSP

[BDRP79] M. Berger, J. C. Dalby, Jr., V. L. Ransom, and E. M. Prell. Traffic service position system no. 1: Automated coin toll service: Overall description and operational characteristics. The Bell System Technical Journal, 58(6):1207–1223, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1207.pdf.

Buzzard:1978:TNT

[BDS78] C. A. Buzzard, J. A. Drager, and B. R. Saltzberg. Transaction networks, telephones, and terminals: Communication network and equipment. The Bell System Technical Journal, 57 (10):3349–3369, December 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3349.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3349.pdf.

Bebbington:1972:SBT

[Beb72] G. H. Bebbington. Soil burial tests: Effect of soil burial exposure on the properties of rubber, crosslinked polyethylene, and vulcanized wire coatings. The Bell System Technical Journal, 51(1):87–121, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol51/bstj51-1-

                                        87.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-87.pdf.

Bell:1976:RTS

[Bel76] A. Graham Bell. Researches in telephony and sciences vol. XII. The Bell System Technical Journal, 55(3):278–288, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-3-278.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-3-278.pdf.

Bell:1978:SSR

[Bel78] W. N. Bell. Steady-state response of a well-balanced wire pair to distributed interference. The Bell System Technical Journal, 57(3):653–668, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-3-653.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-3-653.pdf.

Benes:1970:TCN

[Ben70] V. E. Beneš. Traffic in connecting networks when existing calls are rearranged. The Bell System Technical Journal, 49(7):1471–1482, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol49/bstj49-7-1471.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1471.pdf.

Benes:1971:BJB

[Ben71a] V. E. Beneš. B.S.T.J. briefs:

A telephone traffic model based on randomly closing crosspoints, and its relationships to other models. The Bell System Technical Journal, 50(3):1094–1098, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-1094.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-1094.pdf.

Benes:1971:SFF

[Ben71b] V. E. Beneš. Stability of frequency feedback receivers under steps in input frequency. The Bell System Technical Journal, 50(5):1503–1524, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1503.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1503.pdf.

Benes:1973:SCN

[Ben73] V. E. Beneš. Semilattice characterization of nonblocking networks. The Bell System Technical Journal, 52(5):697–706, May/ June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-5-697.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-5-697.pdf.

Benes:1975:AGT

[Ben75a] V. E. Beneš. Applications of group theory to connecting networks. The Bell System Technical Journal, 54 (2):407–420, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-2-407.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-2-407.pdf.

Benes:1975:PRC

[Ben75b] V. E. Beneš. Proving the rearrangeability of connecting networks by group calculations. The Bell System Technical Journal, 54(2):421–434, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-2-421.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-2-421.pdf.

Benes:1975:TGT

[Ben75c] V. E. Beneš. Toward a grouptheoretic proof of the rearrangeability theorem for Clos’ network. The Bell System Technical Journal, 54(4):797–805, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-4-797.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-4-797.pdf.

Benes:1976:KIC

[Ben76] V. E. Beneš. On Kailath’s innovations conjecture hold. The Bell System Technical Journal, 55(7):981–1001, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-981.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-981.pdf.

Benes:1978:RNS

[Ben78] V. E. Beneš. Reduction of network states under symmetries. The Bell System Technical Journal, 57(1):111–149, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-1-111.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-1-111.pdf.

Berlekamp:1970:SMP

[Ber70a]  E. R. Berlekamp.  Some

mathematical properties of a scheme for reducing the bandwidth of motion pictures by Hadamard smearing. The Bell System Technical Journal, 49(6):969–986, July/August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-969.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-969.pdf.

Berrang:1970:TTH

[Ber70b]  J. E. Berrang.  Television

transmission of holograms using a narrow-band video signal. The Bell System Technical Journal, 49(5):879–887, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-879.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-879.pdf.

Berkowitz:1979:BJB

[Ber79]  H. W. Berkowitz.  B.S.T.J.

briefs: A counterexample to a conjecture on the blocking probabilities of linear graphs. The Bell System Technical Journal, 58(5):1107–1108, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-5-1107.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-5-1107.pdf.

Brewer:1978:SUC

[BEST78] S. T. Brewer, R. L. Easton, H. Soulier, and S. A. Taylor. SG undersea cable system: Requirements and performance. The Bell System Technical Journal, 57(7):2319–2354, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2319.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2319.pdf.

Brophy:1973:CED

[BFG73] F. J. Brophy, G. J. Foschini, and R. D. Gitlin. A compromise equalizer design incorporating performance invariance. The Bell System Technical Journal, 52(7):1077–1095, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1077.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1077.pdf.

Brown:1975:SDP

[BFR75] N. H. Brown, M. P. Fabisch, and C. J. Rifenberg. SAFE-GUARD data-processing system: Introduction and overview. The Bell System Technical Journal, 54(10):S9–S25, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol54/bstj54-10-S9.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-10-S9.pdf.

Batson:1972:DCB

[BG72a] J. E. D. Batson, Jr. and J. W. Gorman. D2 channal bank: Manufacturing and testing. The Bell System Technical Journal, 51(8):1743–1765, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-8-1743.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-8-1743.pdf.

Brandenburg:1972:BJB

[BG72b] L. H. Brandenburg and B. Gopinath. B.S.T.J. briefs: A table look up approach to loop switching. The Bell System Technical Journal, 51(9):2095–2099, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-9-2095.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-9-2095.pdf.

Bulfer:1979:TSP

[BGH79] A. F. Bulfer, W. E. Gibbons, and J. A. Hackett. Traffic service position system no. 1: Remote trunk arrangement: Hardware and software implementation. The Bell System Technical Journal, 58(6):1167–1205, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1167.pdf.

Barney:1970:NEA

[BGK70] D. R. Barney, P. K. Giloth, and H. G. Kienzle. No. 1 ESS ADF: System testing and early field operation experience. The Bell System Technical Journal, 49 (10):2975–3004, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2975.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2975.pdf.

Brandenburg:1972:APL

[BGK72] L. H. Brandenburg, B. Gopinath, and R. P. Kurshan. On the addressing problem of loop switching. The Bell System Technical Journal, 51(7): 1445–1469, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1445.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1445.pdf.

Bricker:1971:STT

[BGM^{+}71] P. D. Bricker, R. Gnanadesikan, M. V. Mathews, Miss S. Pruzansky, P. A. Tukey, K. W. Wachter, and J. L. Warner. Statistical techniques for talker identification. The Bell System Technical Journal, 50(4):1427–1454, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1427.pdf.

Bender:1971:PSW

[BH71] E. C. Bender and R. D. Howson. The Picturephone System: Wideband data service. The Bell System Technical Journal, 50(2):667–681, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-667.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-667.pdf.

Butler:1974:LAS

[BHKJ74] T. T. Butler, T. G. Hallin, J. J. Kulzer, and K. W. Johnson. LAMP: Application to switching-system development. The Bell System Technical Journal, 53(8):1535–1555, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1535.
pdf;   http://www.alcatel-

lucent.com/bstj/vol53-1974/
articles/bstj53-8-1535.pdf.

Bunin:1971:PSC

[BHO71] B. J. Bunin, R. B. Hirsch, and R. E. Olsen. The Picturephone System: Crosstalk considerations in the transmission of analog signals on paired cable. The Bell System Technical Journal, 50(2):427–457, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-427.pdf.

Brown:1971:PSN

[BHP71] D. W. Brown, J. R. Horvath, and T. S. Paxton. The Picturephone System: No. 101 ESS. The Bell System Technical Journal, 50(2):605–620, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-605.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-605.pdf.

Bisbee:1971:MLD

[Bis71a] D. L. Bisbee. Measurements of loss due to offsets and end separations of optical fibers. The Bell System Technical Journal, 50 (10):3159–3168, December 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3159.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3159.pdf.

Bisbee:1971:OFJ

[Bis71b] D. L. Bisbee. Optical fiber joining technique. The Bell System Technical Journal, 50(10): 3153–3158, December 1971. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3153.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3153.pdf.

Bedell:1971:TMR

[BJL71] H. R. Bedell, R. W. Judkins, and R. L. Lahlum. TH-3 microwave radio system: Microwave generator. The Bell System Technical Journal, 50(7):2205–2221, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2205.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2205.pdf.

Bodner:1978:TNT

[BJO78] H. A. Bodner, D. R. Johnson, and W. E. Omohundro. Transaction network, telephones, and terminals: Customer service center interface. The Bell System Technical Journal, 57(10): 3455–3474, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3455.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3455.pdf.

Balkovic:1971:CSH

[BKKM71] M. D. Balkovic, H. W. Klancer, S. W. Klare, and W. G. Mc-Gruther. 1969–70 connection survey: High-speed voiceband data transmission performance on the switched telecommunications network. The Bell System Technical Journal, 50(4):1349–1384, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1349.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1349.pdf.

Bender:1975:DDS

[BKL75] E. C. Bender, J. G. Kneuer, and W. J. Lawless. Digital data system: Local distribution system. The Bell System Technical Journal, 54(5):919–942, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-5-919.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-5-919.pdf.

Balaban:1971:SCD

[BKS71] P. Balaban, B. J. Karafin, and Mrs. D. B. Snyder. Statistical circuit design: A Monte Carlo tolerance analysis of the integrated, single-substrate, RC, touch-tone oscillator. The Bell System Technical Journal, 50(4):1263–1291, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1263.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1263.pdf.

Bohacek:1978:CCI

[BKT78] P. K. Bohacek, A. L. Kalro, and L. A. Tomko. Common channel interoffice signaling: Implementation planning. The Bell System Technical Journal, 57(2):251–261, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-2-251.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-2-251.pdf.

Buus:1970:SSR

[BKY70] R. G. Buus, J. J. Kassig, and P. A. Yeisley. SF system: Repeater and equalizer design. The Bell System Technical Journal, 49(5):631–661, May/ June 1970. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-631.pdf.

Blecher:1970:DPF

[Ble70] Franklin H. Blecher. Device photolithography: Forward. The Bell System Technical Journal, 49(9):1995–1996, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-1995.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-1995.pdf.

Blue:1972:TDS

[Blu72] J. L. Blue. Three-dimensional small-signal analysis of bipolar transistors. The Bell System Technical Journal, 51(4):889– 902, April 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol51/bstj51-4-
889.pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-889.pdf.

Blue:1977:ANQ

[Blu77]  James L. Blue.  Automatic

numerical quadrature. The Bell System Technical Journal, 56(9):1651–1678, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-9-1651.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-9-1651.pdf;
                                         https://archive.org/details/
                                        bstj56-9-1651/mode/2up.

Blue:1978:BIS

[Blu78] J. L. Blue. Boundary integral solutions of Laplace’s equation. The Bell System Technical Journal, 57(8):2797–2822, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-8-2797.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-8-2797.pdf.

Butherus:1970:LAB

[BLV70] A. D. Butherus, W. S. Lindenberger, and F. J. Vaccaro. Lead-acid battery: Electrochemical compatibility of plastics. The Bell System Technical Journal, 49(7):1377–1392, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1377.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1377.pdf.

Brooks:1971:SRB

[BM71] R. D. Brooks and H. G. Mattes. Spreading resistance between constant potential surfaces. The Bell System Technical Journal, 50(3):775–784, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-775.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-775.pdf.

Buckler:1978:AFSb

[BM78] M. J. Buckler and C. M. Miller.

Atlanta Fiber System Experiment: Optical crosstalk evaluation for two end-to-end lightguide system installations. The Bell System Technical Journal, 57(6):1759–1769, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1759.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1759.pdf.

Buchner:1971:ILB

[BN71] M. M. Buchner, Jr. and S. R.

Neal. Inherent load-balancing in step-by-step switching systems. The Bell System Technical Journal, 50(1):135–165, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-135.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-135.pdf.

Bonyhard:1973:DDR

[BN73] P. I. Bonyhard and T. J. Nelson. Dynamic data reallocation in bubble memories. The Bell System Technical Journal, 52(3):307–317, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-3-307.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-3-307.pdf.

Barry:1974:LSJ

[BNO74] J. F. Barry, S. Narayanan, and J. F. Oberst. L5 system: Jumbogroup frequency supply. The Bell System Technical Journal, 53(10):2109–2127, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-2109.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-2109.pdf.

Bobilin:1973:PST

[Bob73]  R. T. Bobilin.  Prefilters,

sampling, and transmission rates for intraframe codecs for picturephone service. The Bell System Technical Journal, 52(4):497–525, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-4-497.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol52-1973/
articles/bstj52-4-497.pdf.

Bobilin:1974:BJB

[Bob74] R. T. Bobilin. B.S.T.J. briefs:

Interframe picturephone coding using unconditional vertical and temporal subsampling techniques. The Bell System Technical Journal, 53 (2):395–400, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-395.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-395.pdf.

Bosik:1972:SDC

[Bos72]  B. S. Bosik.  The spectral

density of a coded digital signal. The Bell System Technical Journal, 51(4):921–932, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-921.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-921.pdf.

Bourne:1978:UTS

[Bou78] S. R. Bourne. UNIX timesharing system: The UNIX shell. The Bell System Technical Journal, 57(6):1971–1990, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol57/bstj57-6-1971.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1971.pdf;
                                         https://archive.org/details/
                                        bstj57-6-1971.

Boyce:1976:SRA

[Boy76] W. M. Boyce. Step response of an adaptive delta modulator. The Bell System Technical Journal, 55(4):373–393, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-4-373.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-4-373.pdf.

Brinsfield:1970:DPM

[BP70] Mrs. J. G. Brinsfield and S. Pardee. Device photolithography: The mask shop information system. The Bell System Technical Journal, 49 (9):2203–2220, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2203.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2203.pdf.

Boggs:1979:RAI

[BPM79] L. M. Boggs, H. M. Presby, and D. Marcuse. Rapid automatic index profiling of wholefiber samples: Part 1. The Bell System Technical Journal, 58(4):867–882, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-4-867.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-4-867.pdf.

Brown:1975:DDS

[BPS75] P. F. Brown, G. W. Phipps, and L. A. Spindel. Digital data system: Network planning. The Bell System Technical Journal, 54(5):861–878, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-5-861.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-5-861.pdf.

Brune:1979:TSP

[BPW79] W. L. Brune, R. J. Piereth, and A. G. Weygand. Traffic service position system no. 1: Remote trunk arrangement and position subsystem no. 2: Transmission and signaling considerations. The Bell System Technical Journal, 58(6):1137–1165, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1137.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1137.pdf.

Bodtmann:1974:RAS

[BR74] W. F. Bodtmann and C. L. Ruthroff. Rain attenuation on short radio paths: Theory, experiment, and design. The Bell System Technical Journal, 53(7):1329–1349, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-7-1329.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-7-1329.pdf.

Brackett:1970:CSH

[Bra70] C. A. Brackett. Characterization of second-harmonic effects in IMPATT diodes. The Bell System Technical Journal, 49(8):1777–1810, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1777.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1777.pdf.

Brady:1971:ETD

[Bra71] Paul T. Brady. Effects of transmission delay on conversational behavior on echo-free telephone circuits. The Bell System Technical Journal, 50(1):115–134, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-1-115.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol50-1971/
articles/bstj50-1-115.pdf.

Brady:1972:OMP

[Bra72]  Paul T. Brady.  Objective

measures of peak clipping and threshold crossings in continuous speech. The Bell System Technical Journal, 51(4):933–945, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-933.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-933.pdf.

Brackett:1973:ETI

[Bra73] C. A. Brackett. The elimination of tuning-induced burnout and bias-circuit oscillations in IMPATT oscillators. The Bell System Technical Journal, 52(3):271–306, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-271.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-3-271.pdf.

Brewer:1970:SSC

[Bre70]  S. T. Brewer.  SF subma-

rine cable system — forward. The Bell System Technical Journal, 49(5):601–604, May/June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol49/bstj49-5-601.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-5-601.pdf.

Breen:1971:PSC

[Bre71]  Charles Breen.  The Pic-

turephone System: Customer switching systems. The Bell System Technical Journal, 50 (2):553–565, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-553.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-553.pdf.

Bricker:1971:LES

[Bri71]  P. D. Bricker.  Listener

evaluation of stimulated telephone calling signals. The Bell System Technical Journal, 50(5):1559–1578, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-5-1559.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-5-1559.pdf.

Brilliant:1978:OEE

[Bri78]  M. B. Brilliant.  Obser-

vations of errors and error rates on T1 digital repeatered lines. The Bell System Technical Journal, 57(3):711–746, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-711.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-711.pdf.

Broyde:1970:DPE

[Bro70] Barret Broyde. Device photolithography: Electron-sensitive materials. The Bell System Technical Journal, 49(9): 2095–2104, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2095.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2095.pdf.

Broderick:1971:PSD

[Bro71a] C. W. Broderick. The Picturephone System: A digital transmission system for TD-2 radio. The Bell System Technical Journal, 50(2):481–499, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-481.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-481.pdf.

Brown:1971:PST

[Bro71b] H. E. Brown. The Picturephone System: Transmission plan. The Bell System Technical Journal, 50(2):351–394, February 1971. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-351.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-351.pdf.

Brown:1971:PSB

[Bro71c]  J. M. Brown.  The Pic-

turephone System: Baseband video transmission on loops and short-haul trunks. The Bell System Technical Journal, 50(2):395–425, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-395.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-395.pdf.

Beadle:1970:PST

[BS70a] W. E. Beadle and A. J. Schorr. Picturephone silicon target signal analysis. The Bell System Technical Journal, 49(6):921–952, July/August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-921.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-921.pdf.

Boyle:1970:BJB

[BS70b] W. S. Boyle and G. E. Smith. B.S.T.J. briefs: Charge coupled semiconductor devices. The Bell System Technical Journal, 49(4):587–593, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-4-587.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-4-587.pdf.

Burgess:1971:PSC

[BS71] P. N. Burgess and J. E.

Stickel. The Picturephone System: Central office switching. The Bell System Technical Journal, 50(2):533–552, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-533.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-533.pdf.

Berglund:1972:FPC

[BS72a] C. N. Berglund and R. J.

Strain. Fabrication and performance considerations of chargetransfer dynamic shift registers. The Bell System Technical Journal, 51(3):655–703, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-3-655.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-3-655.pdf.

Bloom:1972:DCB

[BS72b] S. D. Bloom and G. F. Swanson. D2 channel bank: Power conversion. The Bell System Technical Journal, 51(8):1713–1724, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-8-1713.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-8-1713.pdf.

Bisbee:1975:AGO

[BS75a] D. L. Bisbee and P. W. Smith. All-glass optical-fiber tapes. The Bell System Technical Journal, 54(3):479–484, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-3-479.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-3-479.pdf.

Brophy:1975:TDT

[BS75b] F. J. Brophy and A. C. Salazar. Two design techniques for digital phase networks. The Bell System Technical Journal, 54(4):767–781, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-4-767.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-4-767.pdf.

Bracker:1976:TFM

[BS76] W. E. Bracker and E. R.

Sears. Test facility for a message-switching system. The Bell System Technical Journal, 55(7):857–873, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-857.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-857.pdf.

Beaumont:1978:TNT

[BS78] L. R. Beaumont and K. W. Sussman. Transaction network, telephones, and terminals: Maintenance and administration. The Bell System Technical Journal, 57(10):3409–3425, December 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3409.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3409.pdf.

Buckler:1978:AFSa

[BSS78] M. J. Buckler, M. R. Santana, and M. J. Saunders. Atlanta Fiber System Experiment: Lightguide cable manufacture and performance. The Bell System Technical Journal, 57(6):1745–1757, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol57/bstj57-6-1745.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1745.pdf.

Berglund:1973:FCI

[BT73] C. N. Berglund and K. K. Thornber. A fundamental comparison of incomplete charge transfer in charge transfer devices. The Bell System Technical Journal, 52(2):147–182, February 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-2-147.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-2-147.pdf.

Buchner:1970:AES

[Buc70] M. M. Buchner, Jr. An asymmetric encoding scheme for word stuffing. The Bell System Technical Journal, 49(3):379–398, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-3-379.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-3-379.pdf.

Budrikis:1973:MAV

[Bud73] Z. L. Budrikis. Model approximations to visual spatiotemporal sine-wave threshold data. The Bell System Technical Journal, 52(9):1643–1667, November 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-9-
1643.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-9-1643.pdf.

Bullington:1971:PAV

[Bul71]  K. Bullington.  Phase and

amplitude variations in multipath fading of microwave signals. The Bell System Technical Journal, 50(6):2039–2053, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-2039.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-2039.pdf.

Bullington:1975:RSI

[Bul75]  K. Bullington.  Rain-scatter

interference in terrestrial microwave systems. The Bell System Technical Journal, 54 (1):177–187, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-1-177.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-177.pdf.

Burckhardt:1970:LSP

[Bur70] Christoph B. Burckhardt. Laser speckle pattern — A narrowband noise model. The Bell System Technical Journal, 49(2):309–316, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-2-309.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-2-309.pdf.

Burke:1971:ODC

[Bur71] P. J. Burke. The overflow distribution for constant holding time. The Bell System Technical Journal, 50(10):3195–3210, December 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3195.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3195.pdf.

Butler:1971:SCD

[But71] E. M. Butler. Statistical circuit design: Large change sensitivities for statistical designs. The Bell System Technical Journal, 50(4):1209–1224, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1209.pdf.

Biagetti:1970:LAB

[BW70a] R. V. Biagetti and M. C. Weeks. Lead-acid battery: Tetrabasic lead sulfate as a paste material for positive plates. The Bell System Technical Journal, 49(7):1305–1319, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1305.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1305.pdf.

Bridwell:1970:BDM

[BW70b] J. D. Bridwell and J. K.

Wolf. Burst distance and multiple-burst correction. The Bell System Technical Journal, 49(5):889–909, May/June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-889.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-889.pdf.

Bradford:1971:TMR

[BW71] C. E. Bradford and C. J. Waldron. TH-3 microwave radio system: The travelingwave tube amplifier. The Bell System Technical Journal, 50(7):2223–2234, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2223.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2223.pdf.

Brandenburg:1974:CGC

[BW74] L. H. Brandenburg and A. D. Wyner. Capacity of the Gaussian channel with memory: The multivariate case. The Bell System Technical Journal, 53(5):745–778, May/June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-745.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-745.pdf.

Byrne:1974:AIS

[BWW74] C. J. Byrne, W. A. Winckelmann, and R. M. Wolfe. Automatic intercept system: Organization and objectives. The Bell System Technical Journal, 53(1):1–18, January 1974. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-1-1.pdf;   http:
                                        //www.alcatel-lucent.com/bstj/
                                        vol53-1974/articles/bstj53-
                                        1-1.pdf.

Brockway:1978:ESS

[BY78] G. S. Brockway and G. M. Yanizeski. Elastic state of stress in a stalpeth cable jacket subjected to pure bending. The Bell System Technical Journal, 57(1):151–169, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol57/bstj57-1-151.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-1-151.pdf.

Bates:1974:LSS

[BZ74] R. K. Bates and D. J. Zorn. L5 system: Signal administration and interconnection. The Bell System Technical Journal, 53 (10):2129–2145, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol53/bstj53-10-2129.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-2129.pdf.

Cohen:1976:WDF

[CAK76] L. G. Cohen, H. W. Astle, and I. P. Kaminow. Wavelength dependence of frequency-response measurements in multimode optical fibers. The Bell System Technical Journal, 55(10):1509– 1523, December 1976. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol55/bstj55-10-1509.pdf;
http://www.alcatel-lucent.
com/bstj/vol55-1976/articles/
bstj55-10-1509.pdf.

Candy:1974:LPE

[Can74] J. C. Candy. Limiting the propagation of errors in onebit differential CODECs. The Bell System Technical Journal, 53(8):1667–1676, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-8-1667.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-8-1667.pdf.

Cappellari:1972:WMA

[Cap72] J. O. Cappellari, Jr. Where on the Moon? an Apollo systems engineering problem. The Bell System Technical Journal, 51(5):955–1126, May/June 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-5-955.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-5-955.pdf.

Carlin:1971:RLC

[Car71]  J. W. Carlin.  A relation

for the loss characteristics of circular electric and magnetic modes in dielectric lined waveguide. The Bell System Technical Journal, 50(5):1639–1643, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-5-1639.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-5-1639.pdf.

Candy:1972:MDD

[CB72] J. C. Candy and R. H. Bosworth. Methods for designing differential quantizers based on subjective evaluations of edge busyness. The Bell System Technical Journal, 51(7):1495–1516, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1495.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1495.pdf.

Chin:1979:CSH

[CBK79] A. K. Chin, G. W. Berkstresser, and V. G. Keramidas. Comparison of single heterostructure and double heterostructure GaAs — GaAlAs LEDs for optical data links. The Bell System Technical Journal, 58(7):1579–1591, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1579.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1579.pdf.

Campbell:1976:APD

[CCR76] M. R. Campbell, R. E. Crochiere, and L. R. Rabiner. An algorithmic procedure for designing hybrid FIR/IIR digital filters. The Bell System Technical Journal, 55(1):89–108, January 1976. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol55/bstj55-1-

                                        89.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-1-89.pdf.

Cieslak:1977:NES

[CCR^{+}77] T. J. Cieslak, L. M. Croxall, J. B. Roberts, M. W. Saad, and J. M. Scanlon. No. 4 ESS: Software organization and basic call handling. The Bell System Technical Journal, 56(7):1113–1138, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-7-1113.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-7-1113.pdf.

Cohen:1974:AIS

[CCWW74] H. Cohen, D. E. Confalone, B. D. Wagner, and W. W. Wood. Automatic intercept system: Operational programs. The Bell System Technical Journal, 53(1):19– 69, January 1974. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol53/bstj53-1-
                                        19.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-1-19.pdf.

Carlin:1971:LLM

[CD71] J. W. Carlin and P. D’Agostino. Low-loss modes in dielectric lined waveguide. The Bell System Technical Journal, 50 (5):1631–1638, May/June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1631.pdf.

Carlin:1973:NMO

[CD73] J. W. Carlin and P. D’Agostino.

Normal modes in overmoded dielectric-lined circular waveguide. The Bell System Technical Journal, 52(4):453–486, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-4-453.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-4-453.pdf.

Caspers:1979:ITD

[CD79] B. E. Caspers and P. B. Denes.

An interactive terminal for the design of advertisements. The Bell System Technical Journal, 58(10):2189–2216, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol58/bstj58-10-2189.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-10-2189.pdf.

Cosier:1978:SUC

[CDD^{+}78] J. E. H. Cosier, A. P. Davies, S. W. Dawson, Jr., R. F. Gleason, F. E. Kirkland, and T. A. McKenzie. SG undersea cable system: Installation and maintenance of the undersea system. The Bell System Technical Journal, 57(7):2523–2546, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2523.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2523.pdf.

Cohen:1978:PDP

[CDF^{+}78] L. G. Cohen, F. V. DiMarcello, J. W. Fleming, W. G. French, J. R. Simpson, and E. Weiszmann. Pulse dispersion properties of fibers with various material constituents. The Bell System Technical Journal, 57(5):1653–1662, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1653.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1653.pdf.

Comella:1970:TNP

[CDH70] W. K. Comella, C. M. Day, Jr., and J. A. Hackett. TSPS no. 1: Peripheral circuits. The Bell System Technical Journal, 49 (10):2561–2623, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2561.pdf;
http://www.alcatel-lucent.

com/bstj/vol49-1970/articles/
bstj49-10-2561.pdf.

Chappell:1974:LLC

[CES74] S. G. Chappell, C. H. Elmendorf, and L. D. Schmidt. LAMP: Logic-circuit simulators. The Bell System Technical Journal, 53(8):1451–1476, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1451.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1451.pdf.

Cohen:1979:ETT

[CF79] L. G. Cohen and J. W.

Fleming. Effect of temperature on transmission in lightguides. The Bell System Technical Journal, 58(4):945–951, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-4-945.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-4-945.pdf.

Cannone:1970:LAB

[CFB70] A. G. Cannone, D. O. Feder, and R. V. Biagetti. Leadacid battery: Positive grid design principles. The Bell System Technical Journal, 49(7): 1279–1303, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1279.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1279.pdf.

See erratum [Ano71k].

Clement:1977:PCA

[CFG^{+}77] G. F. Clement, P. S. Fuss, R. J. Griffith, R. C. Lee, and R. D. Royer. 1A processor: Control, administrative, and utility software. The Bell System Technical Journal, 56(2):237–254, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-2-237.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-2-237.pdf.

Candy:1971:TTC

[CFHM71] J. C. Candy, Mrs. M. A. Franke, B. G. Haskell, and F. W. Mounts. Transmitting television as clusters of frame-to-frame differences. The Bell System Technical Journal, 50(6):1889–1917, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-6-1889.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-6-1889.pdf.

Crawford:1978:CCI

[CFMS78] K. E. Crawford, C. J. Funk, P. R. Miller, and J. D. Sipes. Common channel interoffice signaling: 4A toll crossbar application. The Bell System Technical Journal, 57(2):283–323, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-2-283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-2-283.pdf.

Carney:1970:NEA

[CFP70] A. C. Carney, S. M. Fitch, and G. Parker. No. 1 ESS ADF: Teletypewriter stations and transmission facilities. The Bell System Technical Journal, 49(10): 2941–2973, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2941.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2941.pdf.

Calkin:1970:SSP

[CG70] E. T. Calkin and I. Golioto.

SF system: Power conversion. The Bell System Technical Journal, 49(5):749–765, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-749.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-749.pdf.

Chinnock:1975:POF

[CGBS75] E. L. Chinnock, D. Gloge, D. L. Bisbee, and P. W. Smith. Preparation of opticalfiber ends for low-loss tape splices. The Bell System Technical Journal, 54(3):471–477, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-3-471.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-3-471.pdf.

Carran:1974:AIS

[CGHW74] J. H. Carran, K. E. Greisen, W. G. Hall, and D. J. Wells. Automatic intercept system: Administering the intercept data base. The Bell System Technical Journal, 53(1):133–153, January 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-1-133.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-1-133.pdf.

Cooney:1971:TMH

[CGL71] R. T. Cooney, H. D. Griffiths, and F. H. Lanigan. TH-3 medium-haul application: Protection switching. The Bell System Technical Journal, 50(7):2315–2343, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol50/bstj50-7-2315.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2315.pdf.

Chu:1975:QOP

[CGL75] T. S. Chu, M. J. Gans, and W. E. Legg. Quasi-optical polarization diplexing of microwaves. The Bell System Technical Journal, 54(10):1665–1680, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol54/bstj54-10-1665.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-1665.pdf.

Crochiere:1977:TCW

[CGRS77] R. E. Crochiere, D. J. Goodman, L. R. Rabiner, and M. R. Sambur. Tandem connections of wideband and narrowband speech communications systems: Part 1 — narrowband-to-wideband link. The Bell System Technical Journal, 56(9):1701–1722, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1701.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1701.pdf.

Calkin:1978:SUC

[CGS^{+}78] E. T. Calkin, I. Golioto, W. J.

Schatz, R. E. Schroeder, and D. S. Shull. SG undersea cable system: Undersea system power. The Bell System Technical Journal, 57(7):2497–2522, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2497.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2497.pdf.

Compton:1977:NES

[CGWN77] H. B. Compton, P. K. Giloth, M. B. Waverly, and B. G. Niedfeldt. No. 4 ESS: System integration and early office experience. The Bell System Technical Journal, 56(7):1279–1295, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1279.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1279.pdf.

Chang:1972:FSD

[CH72] R. W. Chang and E. Y. Ho. On fast start-up data communication systems using pseudorandom training sequences. The Bell System Technical Journal, 51(9):2013–2027, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-9-2013.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-9-2013.pdf.

Chang:1974:LCO

[CH74] H. Y. Chang and G. W.

Heimbigner. LAMP: Controllability, observability, and maintenance engineering technique (COMET). The Bell System Technical Journal, 53 (8):1505–1534, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1505.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1505.pdf.

Chung:1977:BPS

[CH77] F. R. K. Chung and F. K.

Hwang. On blocking probabilities for switching networks. The Bell System Technical Journal, 56(8):1431–1446, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1431.pdf.

Chung:1978:GTT

[CH78a] F. R. K. Chung and F. K.

Hwang. A generalization of Takagi’s theorem on optimal channel graphs. The Bell System Technical Journal, 57(1):171–178, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-1-171.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-171.pdf.

Chung:1978:BPC

[CH78b] F. R. K. Chung and F. K. Hwang. On blocking probabilities for a class of linear graphs. The Bell System Technical Journal, 57(8):2915–2925, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-8-2915.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-8-2915.pdf.

See erratum [CH79].

Chung:1979:EFR

[CH79] F. R. K. Chung and F. K. Hwang. Erratum: F. R. K. Chung and F. K. Hwang, On Blocking Probabilities for a Class of Linear Graphs, BSTJ 57(8) 2915–2925 (1978). The Bell System Technical Journal, 58(6):1549, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1549.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1549.pdf.

See [CH78b].

Chang:1970:JOA

[Cha70a] Robert W. Chang. Joint optimization of automatic equalization and carrier acquisition for digital communication. The Bell System Technical Journal, 49(6):1069–1104, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1069.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-1069.pdf.

Chapman:1970:SSM

[Cha70b] Alan T. Chapman. SF system:

Manufacture of submarine cable repeaters and ocean block equalizers. The Bell System Technical Journal, 49(5):663–681, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-663.pdf.

Chang:1971:NES

[Cha71] Robert W. Chang. A new equalizer structure for fast start-up digital communications. The Bell System Technical Journal, 50(6):1969–2014, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1969.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1969.pdf.

Chang:1972:ADM

[Cha72] R. W. Chang. Analysis of a dual mode digital synchronization system employing digital rate-locked loops. The Bell System Technical Journal, 51(8):1881–1911, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1881.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1881.pdf.

Chao:1973:SPG

[Cha73]  Min Te Chao.  Statistical

properties of Gilbert’s burst noise model. The Bell System Technical Journal, 52 (8):1303–1324, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1303.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1303.pdf.

Chappell:1974:LAT

[Cha74] S. G. Chappell. LAMP: Automatic test generation for asynchronous digital circuits. The Bell System Technical Journal, 53(8):1477–1503, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1477.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1477.pdf.

Chen:1971:EOL

[Che71] W. Y. S. Chen. Estimated outage in long-haul radio relay systems with protection switching. The Bell System Technical Journal, 50(4):1455–1485, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1455.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1455.pdf.

Chen:1976:BJB

[Che76] W. Y. S. Chen. B.S.T.J. briefs:

A simple method for estimating five-minute point rain-rate distributions based on available climatological data. The Bell System Technical Journal, 55(1):129–134, January 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-1-129.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-1-129.pdf.

Chien:1970:UBE

[Chi70] Ta-Mu Chien. Upper bound on the efficiency of dc-constrained codes. The Bell System Technical Journal, 49(9):2267– 2287, November 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-9-2267.pdf;

                                        http://www.alcatel-lucent.
                                        com/bstj/vol49-1970/articles/
                                        bstj49-9-2267.pdf.

Cowan:1970:DPP

[CHJZ70] M. J. Cowan, D. R. Herriott, A. M. Johnson, and A. Zacharias. Device photolithography: The primary pattern generator — Part I — optical design. The Bell System Technical Journal, 49(9):2033–2041, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2033.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2033.pdf.

Chen:1974:LSR

[CHK^{+}74] R. M. M. Chen, C. F. Hempstead, Y. L. Kuo, M. L. Liou, R. P. Snicer, and E. D. Walsh. L5 system: Role of computing and precision measurements. The Bell System Technical Journal, 53(10):2249–2267, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-2249.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-2249.pdf.

Connor:1973:FFP

[CHM73] D. J. Connor, B. G. Haskell, and F. W. Mounts. A frame-to-frame picturephone coder for signals containing differential quantizing noise. The Bell System Technical Journal, 52(1):35–51, January 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-1-
35.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-35.pdf.

Cheng:1977:WMW

[CHM^{+}77] S. S. Cheng, E. T. Harkless, R. W. Muise, W. E. Studdiford, and D. N. Zuckerman. WT4 millimeter waveguide system: Field evaluation test system performance. The Bell System Technical Journal, 56(10): 2147–2156, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2147.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2147.pdf.

Cho:1972:OEW

[Cho72] Y. S. Cho. Optimal equalization of wideband coaxial cable channels using ‘bump’ equalizers. The Bell System Technical Journal, 51(6):1327–1345, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1327.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1327.pdf.

Cho:1974:MSE

[Cho74]  Y. S. Cho.  Mean-squared-

error equalization using manually adjusted equalizers. The Bell System Technical Journal, 53(5):847–865, May/June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-847.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-847.pdf.

Chadha:1979:AMP

[CHPT79] K. J. S. Chadha, C. F. Hunnicutt, S. R. Peck, and J. Tebes, Jr. Advanced mobile phone service: Mobile telephone switching office. The Bell System Technical Journal, 58(1):71–95, January 1979. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol58/bstj58-1-
                                        71.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-71.pdf.

Cavanaugh:1976:MSE

[CHS76] J. R. Cavanaugh, R. W. Hatch, and J. L. Sullivan. Models for the subjective effects of loss, noise, and talker echo on telephone connections. The Bell System Technical Journal, 55(9):1319–1371, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol55/bstj55-9-1319.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-9-1319.pdf.

Chu:1971:MPT

[Chu71a] T. S. Chu. Maximum power transmission between two reflector antennas in the Fresnel zone. The Bell System Technical Journal, 50(4):1407–1420, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1407.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1407.pdf.

Chu:1971:ROT

[Chu71b] T. S. Chu. Restoring the orthogonality of two polarizations in radio communication systems, I. The Bell System Technical Journal, 50(9):3063–3069, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-9-3063.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-9-3063.pdf.

Chu:1973:ROT

[Chu73] T. S. Chu. Restoring the orthogonality of two polarizations in radio communication systems, II. The Bell System Technical Journal, 52(3):319–327, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-3-319.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-3-319.pdf.

Chu:1974:RIC

[Chu74] T. S. Chu. Rain-induced crosspolarization at centimeter and millimeter wavelengths. The Bell System Technical Journal, 53(8):1557–1579, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-8-1557.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-8-1557.pdf.

Chung:1975:ORG

[Chu75] F. R. K. Chung. Optimal rearrangeable graphs. The Bell System Technical Journal, 54(9):1647–1661, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-9-1647.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-9-1647.pdf.

Chu:1977:CPR

[Chu77] T. S. Chu. Cancellation of polarization rotation in an offset paraboloid by a polarization grid. The Bell System Technical Journal, 56(6):977–986, July/ August 1977. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-6-977.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-6-977.pdf.

Chung:1978:ZBN

[Chu78]  F. R. K. Chung.  Zone-

balanced networks and block designs. The Bell System Technical Journal, 57(8):2957–2971, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2957.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2957.pdf.

Chu:1979:BJB

[Chu79a] T. S. Chu. B.S.T.J. briefs:

Effects of sandstorms on microwave propagation. The Bell System Technical Journal, 58(2):549–555, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-2-549.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-2-549.pdf.

Chung:1979:CSG

[Chu79b] F. R. K. Chung. On concentrators, superconcentrators, generalizers, and nonblocking networks. The Bell System Technical Journal, 58 (8):1765–1777, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-8-1765.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-8-1765.pdf.

Cummiskey:1973:AQD

[CJF73] P. Cummiskey, N. S. Jayant, and J. L. Flanagan. Adaptive quantization in differential PCM coding of speech. The Bell System Technical Journal, 52(7):1105–1118, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-7-1105.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-7-1105.pdf.

Cermak:1971:SCD

[CK71] I. A. Cermak and Mrs. D. B. Kirby. Statistical circuit design: Nonlinear circuits and statistical design. The Bell System Technical Journal, 50(4):1173–1195, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1173.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1173.pdf.

Cohen:1978:UTS

[CK78] H. Cohen and J. C. Kaufeld, Jr.

UNIX time-sharing system: The network operations center system. The Bell System Technical Journal, 57(6):2289–2304, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2289.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2289.pdf.

Clemetson:1971:EMW

[CKK^{+}71] W. J. Clemetson, N. D. Kenyon, K. Kurokawa, B. Owen, and W. O. Schlosser. An experimental MM-wave path length modulator. The Bell System Technical Journal, 50(9): 2917–2945, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-9-2917.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-9-2917.pdf.

Chen:1979:LTM

[CKS79] F. S. Chen, M. A. Karr, and P. W. Shumate. Laser transmitters for 70-MHz entrance links. The Bell System Technical Journal, 58(7):1617–1629, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1617.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-7-1617.pdf.

Clemens:1975:AOS

[CLB75] J. T. Clemens, E. F. Labuda, and C. N. Berglund. Aluminum oxide/silicon dioxide, double-insulator, MOS structure. The Bell System Technical Journal, 54(4):687–719, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-4-687.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-4-687.pdf.

Corbin:1970:NEA

[CLT70] J. E. Corbin, H. R. Lehman, and R. C. Townley. No. 1 ESS ADF: Autonomous data scanner and distributor. The Bell System Technical Journal, 49(10): 2857–2885, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol49/bstj49-10-2857.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol49-1970/articles/
                                        bstj49-10-2857.pdf.

Carlin:1973:EVL

[CM73] J. W. Carlin and A. Maione. Experimental verification of lowloss TM modes in dielectric-lined waveguide. The Bell System Technical Journal, 52(4):487– 496, April 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-4-
487.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-4-487.pdf.

Cherin:1975:AEL

[CM75a] A. H. Cherin and E. J. Murphy.

An analysis of the effect of lossy coatings on the transmission energy in a multimode optical fiber. The Bell System Technical Journal, 54(9):1531–1546, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-9-1531.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-9-1531.pdf.

Cherin:1975:QRA

[CM75b] A. H. Cherin and E. J. Murphy. Quasi-ray analysis of crosstalk between multimode optical fibers. The Bell System Technical Journal, 54(1):17–45, January 1975. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol54/bstj54-1-
17.pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-17.pdf.

Carlin:1977:WMW

[CM77] J. W. Carlin and S. C. Moorthy. WT4 millimeter waveguide: TE_{01} transmission in waveguide with axial curvature. The Bell System Technical Journal, 56 (10):1849–1872, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-1849.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-1849.pdf.

Chu:1978:MLD

[CM78] T. C. Chu and A. R. McCormick. Measurement of loss due to offset, end separation, and angular misalignment in graded index fibers excited by an incoherent source. The Bell System Technical Journal, 57(3):595–602, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-595.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-595.pdf.

Cook:1973:EMC

[CMG73] J. S. Cook, W. L. Mammel, and R. J. Grow. Effect of misalignments on coupling efficiency of single-mode optical fiber butt joints. The Bell System Technical Journal, 52(8):1439–1448, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1439.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1439.pdf.

Colson:1978:CCI

[CMR78] J. S. Colson, J. E. Massery, and G. A. Raack. Common channel interoffice signaling: Development tools. The Bell System Technical Journal, 57(2):429–447, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-2-429.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-2-429.pdf.

Cohen:1971:MAD

[Coh71] L. G. Cohen. Measured attenuation and depolarization of light transmitted along glass fibers. The Bell System Technical Journal, 50(1):23–42, January 1971. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol50/bstj50-1-
23.pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-23.pdf.

Cohen:1972:PCG

[Coh72]  L. G. Cohen.  Power cou-

pling from GaAs injection lasers into optical fibers. The Bell System Technical Journal, 51(3):573–594, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-3-573.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-3-573.pdf.

Coker:1972:EIC

[Cok72] C. H. Coker. An experimental interconnection of computers through a loop transmission system. The Bell System Technical Journal, 51(6):1167–1175, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1167.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1167.pdf.

Coleman:1973:UGR

[Col73] R. D. Coleman. Use of a gate to reduce the variance of delays in queues with random service. The Bell System Technical Journal, 52(8):1403–1422, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-8-1403.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-8-1403.pdf.

Connolly:1972:SBTa

[Con72a] R. A. Connolly. Soil burial tests: Soil burial of materials and structures. The Bell System Technical Journal, 51(1):1– 21, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol51/bstj51-1-

1.pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-1.pdf.

Connolly:1972:SBTb

[Con72b] R. A. Connolly. Soil burial tests: Trends in material behavior after eight years of soil exposure. The Bell System Technical Journal, 51(1):151–163, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-1-151.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-151.pdf.

Conners:1975:SDP

[Con75] R. R. Conners. SAFEGUARD data-processing system: Support software and support computers: An overview. The Bell System Technical Journal, 54(10): S149–S160, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S149.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S149.pdf.

Cooper:1970:QSC

[Coo70] R. B. Cooper. Queues served in cyclic order: Waiting times. The Bell System Technical Journal, 49(3):399–413, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-3-399.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-3-399.pdf.

Cook:1977:MIR

[Coo77] J. S. Cook. Minimum impulse response in graded-index fibers. The Bell System Technical Journal, 56(5):719–728, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-5-719.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-5-719.pdf.

Cox:1975:SEM

[Cox75] D. C. Cox. Some effects of measurement errors on rain depolarization experiments. The Bell System Technical Journal, 54(2):435–450, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-2-435.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-2-435.pdf.

Cox:1978:CEO

[Cox78] D. C. Cox. COMSTAR experiment: An overview of the Bell Laboratories 19- and 28-GHz COMSTAR beacon propagation experiments. The Bell System Technical Journal, 57(5):1231–1255, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1231.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1231.pdf.

Connor:1971:TCU

[CPS71] D. J. Connor, R. F. W.

Pease, and W. G. Scholes. Television coding using twodimensional spatial prediction. The Bell System Technical Journal, 50(3):1049–1061, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-1049.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-1049.pdf.

Cox:1971:DCA

[CR71] D. C. Cox and D. O. Reudink.

Dynamic channel assignment in high-capacity mobile communications systems. The Bell System Technical Journal, 50(6):1833–1857, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1833.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1833.pdf.

Cox:1972:DCA

[CR72] D. C. Cox and D. O. Reudink.

Dynamic channel assignment in two-dimensional large-scale mobile radio systems. The Bell System Technical Journal, 51(7):1611–1629, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1611.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1611.pdf.

Chan:1973:AMR

[CR73a] D. S. K. Chan and L. R. Rabiner. An algorithm for minimizing roundoff noise in cascade realizations of finite impulse response digital filters. The Bell System Technical Journal, 52(3):347–385, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-347.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-3-347.pdf.

Chan:1973:TRN

[CR73b] D. S. K. Chan and L. R. Rabiner. Theory of roundoff noise in cascade realizations of finite impulse response digital filters. The Bell System Technical Journal, 52(3):329–345, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-329.

pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-3-329.pdf.

Cherin:1976:IMP

[CR76] A. H. Cherin and P. J. Rich.

An injection-molded plastic connector for splicing optical cables. The Bell System Technical Journal, 55(8):1057–1067, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1057.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1057.pdf.

Cohen:1978:TNT

[CR78] K. L. Cohen and R. F. Ricca.

Transaction network, telephones, and terminals: Dial access interface. The Bell System Technical Journal, 57(10):3441– 3453, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3441.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3441.pdf.

Crater:1971:PSS

[Cra71] T. V. Crater. The Picturephone System: Service standards. The Bell System Technical Journal, 50(2):235–269, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-235.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-235.pdf.

Cherin:1979:VAP

[CRAC79] A. H. Cherin, P. J. Rich, C. J. Aloisio, and R. R. Cammons. A vacuum-assisted plastic repair splice for joining optical fiber ribbons. The Bell System Technical Journal, 58(8):1825–1838, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-8-1825.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-8-1825.pdf.

Crowley:1975:SDP

[Cro75] Thomas H. Crowley. SAFE-GUARD data-processing system: Foreword. The Bell System Technical Journal, 54(10):S3–S5, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-10-S3.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-10-S3.pdf.

Crochiere:1977:DSB

[Cro77] R. E. Crochiere. On the design of sub-band coders for low-bit-rate speech communication. The Bell System Technical Journal, 56(5):747–770, May/ June 1977. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-747.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-747.pdf.

Crochiere:1978:AKS

[Cro78a] R. Crochiere. An analysis of 16 kb/s sub-band coder performance: Dynamic range, tandem connections, and channel errors. The Bell System Technical Journal, 57(8):2927–2952, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2927.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2927.pdf.

Crochiere:1978:MRM

[Cro78b] R. E. Crochiere. A mid-range/ mid-tread quantizer switch for improved idle-channel performance in adaptive coders. The Bell System Technical Journal, 57(8):2953–2955, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2953.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2953.pdf.

Crowley:1978:UTS

[Cro78c] T. H. Crowley. UNIX timesharing system: Preface. The Bell System Technical Journal, 57(6):1897–1898, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1897.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1897.pdf.

Crochiere:1975:NID

[CRS75] R. E. Crochiere, L. R. Rabiner, and R. R. Shively. A novel implementation of digital phase shifters. The Bell System Technical Journal, 54(8):1497–1502, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-8-1497.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-8-1497.pdf.

Christopher:1977:ETS

[CRSB77] D. K. Christopher, L. R. Rabiner, P. Schweitzer, and D. E. Bock. An evaluation of two simple methods for detecting tones over telephone lines. The Bell System Technical Journal, 56(8):1513–1529, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1513.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1513.pdf.

Coleman:1971:BJB

[CS71] D. J. Coleman, Jr. and S. M.

Sze. B.S.T.J. briefs: A low-noise metal-semiconductor-metal (MSM) microwave oscillator. The Bell System Technical Journal, 50(5):1695–1699, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1695.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1695.pdf.

Crochiere:1977:VBC

[CS77a] R. E. Crochiere and M. R. Sambur. A variable-band coding scheme or speech encoding at 4.8 kb/s. The Bell System Technical Journal, 56(5):771–779, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-771.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-771.pdf.

Cross:1977:CSA

[CS77b] P. S. Cross and R. V. Schmidt.

Coupled surface-acoustic-wave resonators. The Bell System Technical Journal, 56 (8):1447–1482, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1447.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-8-1447.pdf.

See errata [Ano78n].

Croxall:1978:CCI

[CS78] L. M. Croxall and R. E. Stone. Common channel interoffice signaling: No. 4 ESS application. The Bell System Technical Journal, 57(2):361–377, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-2-361.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-2-361.pdf.

Clark:1979:TSP

[CSL79] G. T. Clark, K. Streisand, and D. H. Larson. Traffic service position system no. 1: Station signaling and announcement subsystem: Hardware for automated coin toll service. The Bell System Technical Journal, 58(6):1225–1249, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1225.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1225.pdf.

Cagle:1971:PSS

[CSW71] W. B. Cagle, R. R. Strokes, and B. A. Wright. The Picturephone System: The station: 2C video telephone station set. The Bell System Technical Journal, 50(2):271–312, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-271.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-271.pdf.

Chang:1974:LSD

[CSW74] H. Y. Chang, G. W. Smith, Jr., and R. B. Walford. LAMP: System description. The Bell System Technical Journal, 53(8):1431–1449, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1431.pdf.

Cirillo:1972:DCB

[CT72] A. J. Cirillo and D. K. Thovson.

D2 channel bank: Digital functions. The Bell System Technical Journal, 51(8):1701–1712, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1701.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1701.pdf.

Chow:1979:BJB

[CT79] W. F. Chow and W. W. Troutman. B.S.T.J. briefs: MAC-4: A single-chip microcomputer. The Bell System Technical Journal, 58(4):959–962, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-4-959.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-4-959.pdf.

Cummiskey:1975:SIA

[Cum75] P. Cummiskey. Single-integration, adaptive delta modulation. The Bell System Technical Journal, 54(8):1463–1474, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-8-1463.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-8-1463.pdf.

Cushman:1970:LAB

[Cus70] R. H. Cushman. Lead-acid battery: Techniques for bonding the positive plates. The Bell System Technical Journal, 49(7):1419–1446, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1419.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1419.pdf.

Chai:1973:IIM

[CW73] D. T. Chai and J. M. Wier. Interactive information management systems. The Bell System Technical Journal, 52(10): 1681–1689, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1681.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1681.pdf.

Chu:1978:CEC

[CWE^{+}78] T. S. Chu, R. W. Wilson, R. W. England, D. A. Gray, and W. E. Legg. COMSTAR experiment: The Crawford Hill 7-meter millimeter wave antenna. The Bell System Technical Journal, 57(5):1257–1288, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1257.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1257.pdf.

Crochiere:1976:DCS

[CWF76] R. E. Crochiere, S. A. Webber, and J. L. Flanagan. Digital coding of speech in sub-bands. The Bell System Technical Journal, 55(8):1069–1085, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1069.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1069.pdf.

Dabby:1975:PMS

[Dab75]  F. W. Dabby.  Permanent

multiple splices of fused-silica fibers. The Bell System Technical Journal, 54(2):451–455, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-2-451.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-2-451.pdf.

Dahringer:1972:SBT

[Dah72] D. W. Dahringer. Soil burial tests: Effect of soil burial exposure on the properties of adhesives and pressure-sensitive tapes. The Bell System Technical Journal, 51(1):123–149, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-1-123.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-123.pdf.

Danielson:1975:BLF

[Dan75] W. E. Danielson. Bell Laboratories fiftieth anniversary. The Bell System Technical Journal, 54(1):1–2, January 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-1-1.pdf;   http:
                                        //www.alcatel-lucent.com/bstj/
                                        vol54-1975/articles/bstj54-
                                        1-1.pdf.

Darlington:1970:AAA

[Dar70] Sidney Darlington. Analytical approximations to approximations in the Chebyshev sense. The Bell System Technical Journal, 49(1):1–32, January 1970. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol49/bstj49-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-1-1.pdf.

Daskalakis:1971:PSK

[Das71] A. Daskalakis. The Picturephone System: Key telephone systems. The Bell System Technical Journal, 50(2):567–584, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-567.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-567.pdf.

Dawson:1972:BJB

[Daw72]  R. W. Dawson.  B.S.T.J.

briefs: Effect of ambient temperature on infrared transmission through a glass fiber. The Bell System Technical Journal, 51(2):569–571, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-569.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-569.pdf.

Dickieson:1971:SCD

[DC71] A. C. Dickieson and J. Chernak. Statistical circuit design: History and introduction. The Bell System Technical Journal, 50(4):1099–1103, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1099.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1099.pdf.

Daumer:1978:SCS

[DC78] W. R. Daumer and J. R. Cavanaugh. A subjective comparison of selected digital codecs for speech. The Bell System Technical Journal, 57(9):3119–3165, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-9-3119.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-9-3119.pdf.

Dubnowski:1979:VRC

[DC79] J. J. Dubnowski and R. E. Crochiere. Variable rate coding of speech. The Bell System Technical Journal, 58(3):577–600, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol58/bstj58-3-577.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-577.pdf.

Dowd:1970:DPP

[DCRZ70] P. G. Dowd, M. J. Cowan, P. E.

Rosenfeld, and A. Zacharias. Device photolithography: The primary pattern generator: Part III: The control system. The Bell System Technical Journal, 49(9):2061–2067, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2061.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2061.pdf.

DeLange:1975:EPG

[DDH75] O. E. DeLange, A. F. Dietrich, and D. C. Hogg. An experiment on propagation of 60-GHz waves through rain. The Bell System Technical Journal, 54(1):165–176, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-1-165.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-165.pdf.

DeBonte:1972:SSH

[DeB72] W. J. DeBonte. The static stability of half-bubbles. The Bell System Technical Journal, 51(9):1933–1955, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-9-1933.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-9-1933.pdf.

DeCoste:1972:SBT

[DeC72] J. B. DeCoste. Soil burial tests: Effect of soil burial exposure on the properties of plastics for wire and cable. The Bell System Technical Journal, 51(1):63–86, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol51/bstj51-1-
                                        63.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-63.pdf.

Denenberg:1976:SME

[Den76] J. N. Denenberg. Spectral moment estimators: A new approach to tone detection. The Bell System Technical Journal, 55(2):143–155, February 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-2-143.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-2-143.pdf.

Descloux:1973:TMB

[Des73]  A. Descloux.  Traffic mea-

surement biases induced by partial sampling. The Bell System Technical Journal, 52 (8):1375–1402, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1375.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1375.pdf.

Descloux:1975:VLM

[Des75]  A. Descloux.  Variance of

load measurements in Markovian service systems. The Bell System Technical Journal, 54(7):1277–1300, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1277.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1277.pdf.

Descloux:1976:SPV

[Des76] A. Descloux. Some properties of the variance of the switchcount load. The Bell System Technical Journal, 55(1):59–88, January 1976. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol55/bstj55-1-
59.pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-1-59.pdf.

Doyle:1975:SDP

[DG75] W. S. Doyle and J. R. Gibbons. SAFEGUARD dataprocessing system: Process design in the structure of real-time software systems. The Bell System Technical Journal, 54(10): S101–S109, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S101.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S101.pdf.

Dragone:1979:IRA

[DG79] C. Dragone and M. J. Gans. Imaging reflector arrangements to form a scanning beam using a small array. The Bell System Technical Journal, 58(2):501–515, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-2-501.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-2-501.pdf.

Dolotta:1978:UTS

[DHM78] T. A. Dolotta, R. C. Haight, and J. R. Mashey. UNIX time-sharing system: The Programmer’s Workbench. The Bell System Technical Journal, 57(6):2177–2200, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2177.
pdf;   http://www.alcatel-

lucent.com/bstj/vol57-1978/
articles/bstj57-6-2177.pdf.

Dickman:1975:SDP

[Dic75] B. N. Dickman. SAFEGUARD data-processing system: CEN-TRAN — A case history in extendible language design. The Bell System Technical Journal, 54(10):S161–S172, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S161.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S161.pdf.

Dietrich:1971:TMR

[Die71]  N. R. Dietrich.  TH-3 mi-

crowave radio system: Microwave integrated circuits. The Bell System Technical Journal, 50(7):2175–2194, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2175.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2175.pdf.

Dixon:1976:DML

[Dix76]  R. W. Dixon.  Derivative

measurements of light-current-voltage characteristics of (Al,Ga)As double-heterostructure lasers. The Bell System Technical Journal, 55(7):973–980, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-7-973.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-7-973.pdf.

Davis:1975:SDP

[DJ75] E. J. Davis and H. M. Jackson II. SAFEGUARD dataprocessing system: Management overview. The Bell System Technical Journal, 54(10):S223– S229, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-10-S223.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol54-1975/articles/
                                        bstj54-10-S223.pdf.

DAltroy:1974:LSU

[DJNP74] F. A. D’Altroy, R. M. Jacobs, J. M. Nacci, and E. J. Panner. L5 system: Ultralinear transistors. The Bell System Technical Journal, 53(10):2195– 2202, December 1974. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-2195.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-2195.pdf.

Durney:1970:TNS

[DKP^{+}70] G. R. Durney, H. W. Kettler, E. M. Prell, G. Riddell, and W. B. Rohn. TSPS no. 1: Stored program control no. 1A. The Bell System Technical Journal, 49 (10):2445–2507, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2445.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2445.pdf.

Danielsen:1976:RSN

[DL76] D. Danielsen and W. A. Liss.

Remreed switching networks for no. 1 and no. 1A ESS: Remreed network electronic control. The Bell System Technical Journal, 55(5):565–595, May/ June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-565.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-565.pdf.

Dunn:1978:LPM

[DL78] D. M. Dunn and J. M. Landwehr.

Loop plant modeling: Statistical analyses of costs in loop plant operations. The Bell System Technical Journal, 57(4):965–998, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-965.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-965.pdf.

Donohue:1975:SDP

[DM75] B. P. Donohue III and J. F. Mc-Donald. SAFEGUARD dataprocessing system: Processsystem testing and the system exerciser. The Bell System Technical Journal, 54(10): S111–S122, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-10-S111.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol54-1975/articles/
                                        bstj54-10-S111.pdf.

Duffy:1978:SNP

[DM78] F. P. Duffy and R. A. Mercer. A study of network performance and customer behavior during direct-distance-dialing call attempts in the U.S.A. The Bell System Technical Journal, 57(1):1–33, January 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol57/bstj57-1-
                                        1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-1.pdf.

DeLoach:1978:SAP

[DMK78] B. C. DeLoach, Jr., R. C. Miller, and S. Kaufman. Sound alerter powered over an optical fiber. The Bell System Technical Journal, 57(9):3309–3316, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3309.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3309.pdf.

Dammann:1972:DCB

[DMM72] C. L. Dammann, L. D. McDaniel, and C. L. Maddox. D2 channel bank: Multiplexing and coding. The Bell System Technical Journal, 51(8):1675–1699, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1675.pdf.

Duffy:1975:EPT

[DMNT75] F. P. Duffy, G. K. Mc-Nees, I. Nasell, and T. W. Thatcher, Jr. Echo performance of toll telephone connections in the United States. The Bell System Technical Journal, 54(2):209–243, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-209.pdf.

Duff:1979:TSE

[DO79] D. G. Duff and H. M. Olson.

Transistor surface effects responsible for anomalous third-order intermodulation distortion in undersea cable telephone system. The Bell System Technical Journal, 58(9):2025–2069, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-9-2025.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-9-2025.pdf.

Doane:1978:MED

[Doa78]  J. L. Doane.  Measurement

of echoes due to spurious TEon modes in a long-distance 60-mm waveguide communication system. The Bell System Technical Journal, 57(9): 3253–3266, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3253.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3253.pdf.

Dorros:1971:PSN

[Dor71] Irwin Dorros. The Picturephone System: The network. The Bell System Technical Journal, 50(2):221–233, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-221.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-221.pdf.

Dunn:1977:WMW

[DPR77] C. N. Dunn, O. G. Petersen, and D. C. Redline. WT4 millimeter waveguide system: Semiconductor devices for the WT4 repeater. The Bell System Technical Journal, 56(10):2119–2134, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2119.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2119.pdf.

Dougherty:1971:PSM

[DPS71] H. J. Dougherty, E. B. Peterson, and M. G. Schachtman. The Picturephone System: Maintenance plan. The Bell System Technical Journal, 50(2):621–644, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-621.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-621.pdf.

DiPiazza:1979:AMP

[DPZ79] G. C. DiPiazza, A. Plitkins, and G. I. Zysman. Advanced mobile phone service: The cellular test bed. The Bell System Technical Journal, 58(1):215–248, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-1-215.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-215.pdf.

Dahlbom:1978:CCI

[DR78] C. A. Dahlbom and J. S. Ryan. Common channel interoffice signaling: History and description of a new signaling system. The Bell System Technical Journal, 57(2):225–250, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-2-225.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-2-225.pdf.

Dragone:1972:CHG

[Dra72a] C. Dragone. Conditions of high gain in mixers and their relation to the jump phenomenon. The Bell System Technical Journal, 51(10):2139–2167, December 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol51/bstj51-10-2139.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol51-1972/articles/
                                        bstj51-10-2139.pdf.

Dragone:1972:PSS

[Dra72b] C. Dragone. Performance and stability of Schottky barrier mixers. The Bell System Technical Journal, 51(10):2169– 2196, December 1972. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol51/bstj51-10-2169.pdf;
http://www.alcatel-lucent.
com/bstj/vol51-1972/articles/
bstj51-10-2169.pdf.

Dragone:1974:IAM

[Dra74] C. Dragone. An improved antenna for microwave radio systems consisting of two cylindrical reflectors and a corrugated horn. The Bell System Technical Journal, 53(7):1351–1377, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-7-1351.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-7-1351.pdf.

Dragone:1977:CBM

[Dra77a] C. Dragone. Characteristics of a broadband microwave corrugated feed: A comparison between theory and experiment. The Bell System Technical Journal, 56(6):869–888, July/ August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-6-869.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-6-869.pdf.

Dragone:1977:RTM

[Dra77b] C. Dragone. Reflection, transmission, and mode conversion in a corrugated feed. The Bell System Technical Journal, 56(6):835–867, July/August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-6-835.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-6-835.pdf.

Dragone:1978:OMA

[Dra78] C. Dragone. Offset multireflector antennas with perfect pattern symmetry and polarization discrimination. The Bell System Technical Journal, 57(7):2663–2684, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-7-2663.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-7-2663.pdf.

Dahringer:1970:LAB

[DS70] D. W. Dahringer and J. R. Shroff. Lead-acid battery: Jarcover seals. The Bell System Technical Journal, 49(7): 1393–1403, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1393.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1393.pdf.

Derosier:1973:LLS

[DS73] R. M. Derosier and J. Stone.

Low-loss splices in optical fibers. The Bell System Technical Journal, 52(7):1229–1236, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1229.pdf.

Dudderar:1978:CPA

[DSS78] T. D. Dudderar, J. B. Seery, and P. G. Simpkins. Caustic patterns associated with melt zones in solidified glass samples — Part I: Symmetric cases. The Bell System Technical Journal, 57(9):3209–3225, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3209.pdf.

Drazy:1971:TMR

[DSW71] E. J. Drazy, R. E. Sheehey, and H. C. Wang. TH-3 microwave radio system: Networks. The Bell System Technical Journal, 50(7):2137–2153, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2137.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2137.pdf.

Duffy:1971:CSA

[DT71] F. P. Duffy and T. W. Thatcher, Jr. 1969–70 connection survey: Analog transmission performance on the switched telecommunications network. The Bell System Technical Journal, 50(4):1311–1347, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1311.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1311.pdf.

Duttweiler:1972:WTJ

[Dut72] D. L. Duttweiler. Waiting time jitter. The Bell System Technical Journal, 51(1):165–207, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-1-165.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-165.pdf.

Duttweiler:1976:JPP

[Dut76] D. L. Duttweiler. The jitter performance of phase-locked loops extracting timing from baseband data waveforms. The Bell System Technical Journal, 55(1):37– 58, January 1976. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol55/bstj55-1-
                                        37.pdf;   http://www.alcatel-

lucent.com/bstj/vol55-1976/
articles/bstj55-1-37.pdf.

Delatore:1979:TSP

[DVW79] J. P. Delatore, D. VanHaften, and L. A. Weber. Traffic service position system no. 1: System verification and evaluation procedures. The Bell System Technical Journal, 58(6):1335–1346, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1335.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1335.pdf.

DiMarcello:1978:AFS

[DW78] F. V. DiMarcello and J. C.

Williams. Atlanta Fiber System Experiment: Reproducibility of optical fibers prepared by a modified chemical vapor deposition process. The Bell System Technical Journal, 57(6):1723–1734, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1723.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1723.pdf.

Dinn:1978:CEC

[DZ78] N. F. Dinn and G. A. Zimmerman. COMSTAR experiment: COMSTAR beacon receiver diversity experiment. The Bell System Technical Journal, 57(5):1341–1367, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1341.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1341.pdf.

Ebert:1970:CGC

[Ebe70] P. M. Ebert. The capacity of the Gaussian channel with feedback. The Bell System Technical Journal, 49(8):1705–1712, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1705.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1705.pdf.

Eckler:1978:SAB

[Eck78] A. R. Eckler. A statistical analysis of Bell System building fires, 1971–1977. The Bell System Technical Journal, 57(8):2769–2796, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-8-2769.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-8-2769.pdf.

Ehrbar:1978:SUC

[EFG78] R. D. Ehrbar, A. E. Ford, and G. Gerbier. SG undersea cable system: Introduction and development plan. The Bell System Technical Journal, 57(7):2313–2317, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2313.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2313.pdf.

Ehrlich:1979:AMP

[EFW79] N. Ehrlich, R. E. Fisher, and T. K. Wingard. Advanced mobile phone service: Cell-site hardware. The Bell System Technical Journal, 58(1):153–199, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-1-153.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-1-153.pdf.

Ewin:1970:NEA

[EG70] J. C. Ewin and P. K. Giloth. No.

1 ESS ADF: System organization and objectives. The Bell System Technical Journal, 49 (10):2733–2752, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2733.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2733.pdf.

Eisenberg:1972:ACD

[Eis72] M. Eisenberg. Almost-coherent detection of phase-shift-keyed signals using an injectionlocked oscillator. The Bell System Technical Journal, 51 (8):1867–1880, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-8-1867.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-8-1867.pdf.

Eisenberg:1977:ETN

[Eis77] M. Eisenberg. Engineering traffic networks for more then one busy hour. The Bell System Technical Journal, 56(1):1–20, January 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol56/bstj56-1-
                                        1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-1-1.pdf.

Evans:1976:ANA

[EKN76] J. G. Evans, F. W. Kerfoot, and R. L. Nichols. Automated network analyzers for the 0.9- to 12.4-GHz range. The Bell System Technical Journal, 55(6):691–721, July/August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-6-691.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol55-1976/
articles/bstj55-6-691.pdf.

Eleftherion:1977:WMW

[Ele77] M. P. Eleftherion. WT4 millimeter waveguide system: Waveguide medium manufacturing process research and development. The Bell System Technical Journal, 56(10):2179–2196, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2179.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2179.pdf.

Elsner:1977:DAM

[Els77] W. B. Elsner. A descent algorithm for the multihour sizing of traffic networks. The Bell System Technical Journal, 56(8):1405–1430, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1405.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1405.pdf.

Eichenbaum:1977:ALS

[ES77] B. R. Eichenbaum and M. R.

Santana. Analysis of longitudinal stress imparted to fibers in twisting an optical communication cable unit. The Bell System Technical Journal, 56(8):1503–1512, October 1977. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-8-1503.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-8-1503.pdf.

Falconer:1976:AGA

[Fal76a] D. D. Falconer. Analysis of a gradient algorithm for simultaneous passband equalization and carrier phase recovery. The Bell System Technical Journal, 55(4):409–428, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-4-409.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-4-409.pdf.

Falconer:1976:JAE

[Fal76b] D. D. Falconer. Jointly adaptive equalization and carrier recovery in two-dimensional digital communication systems. The Bell System Technical Journal, 55(3):317–334, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-3-317.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-3-317.pdf.

Falconer:1978:AEC

[Fal78] D. D. Falconer. Adaptive equalization of channel nonlinearities in QAM data transmission systems. The Bell System Technical Journal, 57(7): 2589–2611, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2589.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2589.pdf.

Fang:1976:ASV

[Fan76]  G. S. Fang.  Alarm statis-

tics of the violation monitor and remover. The Bell System Technical Journal, 55 (8):1197–1217, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1197.pdf.

Fang:1978:RMB

[Fan78]  G. S. Fang.  Reliability of

a microprocessor-based protection switching system. The Bell System Technical Journal, 57(7):2633–2661, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2633.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2633.pdf.

Friedenson:1975:RAF

[FDDM75] R. A. Friedenson, R. W. Daniels, R. J. Dow, and P. H. McDonald. RC active filters for the D3 channel bank. The Bell System Technical Journal, 54(3):507–529, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-3-507.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-3-507.pdf.

Fennick:1970:BJB

[Fen70]  J. H. Fennick.  B.S.T.J.

briefs: Intersymbol interference and the P/AR meter. The Bell System Technical Journal, 49(6):1245–1248, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-1245.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1245.pdf.

Fenderson:1971:TMR

[Fen71] G. L. Fenderson. TH-3 microwave radio system: The IF main amplifier. The Bell System Technical Journal, 50(7):2195–2204, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2195.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol50-1971/
articles/bstj50-7-2195.pdf.

Ferguson:1972:ORB

[Fer72] M. J. Ferguson. Optimal reception for binary partial response channels. The Bell System Technical Journal, 51 (2):493–505, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-493.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-2-493.pdf.

Falconer:1973:TMM

[FF73] D. D. Falconer and G. J. Foschini. Theory of minimum mean-square-error QAM systems employing decision feedback equalization. The Bell System Technical Journal, 52(10): 1821–1849, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1821.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1821.pdf.

Favin:1971:PSL

[FG71] D. L. Favin and J. F. Gilmore.

The Picturephone System: Line and trunk maintenance arrangements. The Bell System Technical Journal, 50 (2):645–665, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-645.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-645.pdf.

Falconer:1978:ORD

[FG78] D. D. Falconer and R. D. Gitlin. Optimum reception of digital data signals in the presence of timing-phase hits. The Bell System Technical Journal, 57(9):3181–3208, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-9-3181.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-9-3181.pdf.

Fraser:1978:BST

[FGM78] A. G. Fraser, B. Gopinath, and J. A. Morrison. Buffering of slow terminals. The Bell System Technical Journal, 57(8):2865–2885, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-8-2865.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-8-2865.pdf.

Foschini:1975:ODD

[FGS75] G. J. Foschini, R. D. Gitlin, and J. Salz. Optimum direct detection for digital fiberoptic communication systems. The Bell System Technical Journal, 54(8):1389–1430, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-8-1389.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-8-1389.pdf.

Foschini:1973:STD

[FGW73] G. J. Foschini, R. D. Gitlin, and S. B. Weinstein. On the selection of a two-dimensional signal constellation in the presence of phase jitter and Gaussian noise. The Bell System Technical Journal, 52(6):927–965, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-927.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-927.pdf.

Fleming:1971:CSL

[FH71] H. C. Fleming and R. M.

Hutchinson, Jr. 1969–70 connection survey: Low-speed data transmission performance on the switched telecommunications network. The Bell System Technical Journal, 50(4):1385–1405, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1385.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1385.pdf.

Franks:1979:MRM

[FHH^{+}79] R. L. Franks, H. Heffes, J. M. Holtzman, S. Horing, and E. J. Messerli. A model relating measurement and forecast errors to the provisioning of direct final trunk groups. The Bell System Technical Journal, 58(2):351–377, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-351.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-351.pdf.

Fox:1977:WMW

[FHT77] P. E. Fox, S. Harris, and D. J. Thomson. WT4 millimeter waveguide system: Mechanical gauging techniques. The Bell System Technical Journal, 56 (10):2007–2023, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-2007.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-2007.pdf.

Flanagan:1975:SSD

[FIS75] J. L. Flanagan, K. Ishizaka, and K. L. Shipley. Synthesis of speech from a dynamic model of vocal cords and vocal tract. The Bell System Technical Journal, 54(3):485–506, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-3-485.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-3-485.pdf.

Fisher:1979:AMP

[Fis79] R. E. Fisher. Advanced mobile phone service: A subscriber set for the equipment test. The Bell System Technical Journal, 58(1):123–143, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-1-123.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-1-123.pdf.

Fleischer:1979:FAS

[FL79] P. E. Fleischer and K. R. Laker.

A family of active switched capacitor biquad building blocks. The Bell System Technical Journal, 58(10):2235–2269, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol58/bstj58-10-2235.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-10-2235.pdf.

Flanagan:1979:AFA

[Fla79] J. L. Flanagan. Acoustic filters to aid digital voice. The Bell System Technical Journal, 58(4):903–944, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-4-903.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-4-903.pdf.

Friedman:1971:MMC

[FM71] A. D. Friedman and P. R. Menon. Mathematical models of computation using magnetic bubble interactions. The Bell System Technical Journal, 50(6):1701–1719, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1701.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1701.pdf.

Falconer:1973:ACM

[FM73] D. D. Falconer and F. R. Magee, Jr. Adaptive channel memory truncation for maximum likelihood sequence estimation. The Bell System Technical Journal, 52(9):1541–1562, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-9-1541.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-9-1541.pdf.

Forys:1975:ATG

[FM75] L. J. Forys and E. J. Messerli.

Analysis of trunk groups containing short-holding-time trunks. The Bell System Technical Journal, 54(6):1127–1153, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-6-1127.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-6-1127.pdf.

Falconer:1979:AEC

[FM79] D. D. Falconer and K. H.

Mueller. Adaptive echo cancellation/AGC structures for twowire, full-duplex data transmission. The Bell System Technical Journal, 58(7):1593– 1616, September 1979. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol58/bstj58-7-1593.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-7-1593.pdf.

French:1974:BJB

[FMOT74] W. G. French, J. B. MacChesney, P. B. O’Connor, and G. W. Tasker. B.S.T.J. briefs: Optical waveguides with very low losses. The Bell System Technical Journal, 53(5):951–954, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol53/bstj53-5-951.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-951.pdf.

Froehlich:1978:TNT

[FMSZ78] F. E. Froehlich, C. B. McDowell, F. P. Sansone, and G. D. Zally. Transaction network, telephones, and terminals: The switched network transaction telephone system. The Bell System Technical Journal, 57(10): 3475–3486, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol57/bstj57-10-3475.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol57-1978/articles/
                                        bstj57-10-3475.pdf.

Foschini:1971:COP

[Fos71] G. J. Foschini. Crosstalk in outside plant cable systems. The Bell System Technical Journal, 50(7):2421–2448, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2421.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2421.pdf.

Flint:1974:LSP

[FP74] A. F. Flint and H. S. Pustarfi. L5 system: 39A precision oscillator. The Bell System Technical Journal, 53(10): 2097–2107, December 1974. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol53/bstj53-10-2097.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-2097.pdf.

Fluhr:1979:AMP

[FP79] Z. C. Fluhr and P. T. Porter.

Advanced mobile phone service: Control architecture. The Bell System Technical Journal, 58(1):43–69, January 1979. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol58/bstj58-1-
43.pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-1-43.pdf.

Franks:1973:ONC

[FR73a] R. L. Franks and R. W.

Rishel. Optimum network call-carrying capacity. The Bell System Technical Journal, 52(7):1195–1214, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1195.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1195.pdf.

Franks:1973:OMT

[FR73b] R. L. Franks and R. W.

Rishel. Overload model of telephone network operation. The Bell System Technical Journal, 52(9):1589–1615, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-9-1589.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-9-1589.pdf.

Fitch:1975:DDS

[FR75] S. M. Fitch and D. L. Rechtenbaugh. Digital data system: Testing and maintenance. The Bell System Technical Journal, 54(5):845–860, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-5-845.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-5-845.pdf.

Fredericks:1979:ASS

[FR79] A. A. Fredericks and G. A. Reisner. Approximations to stochastic service systems, with an application to a retrial model. The Bell System Technical Journal, 58(3):557–576, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-557.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-557.pdf.

Fraser:1978:UTS

[Fra78] A. G. Fraser. UNIX timesharing system: Circuit design aids. The Bell System Technical Journal, 57(6):2233–2249, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2233.pdf.

Freeman:1975:SDP

[Fre75] R. D. Freeman. SAFEGUARD data-processing system: An experiment in software development. The Bell System Technical Journal, 54(10):S199– S210, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S199.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S199.pdf.

Freedman:1978:LPM

[Fre78] H. T. Freedman. Loop plant modeling: Optimal operating policies for serving areas using connect-through administration. The Bell System Technical Journal, 57(4):911–926, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-911.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-911.pdf.

Friedenfelds:1978:LPM

[Fri78] J. Friedenfelds. Loop plant modeling: A simple model for studying feeder capacity expansion. The Bell System Technical Journal, 57(4):807–823, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-807.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-807.pdf.

Flanagan:1972:WTA

[FRSD72] J. L. Flanagan, L. R. Rabiner, R. W. Schafer, and J. D. Denman. Wiring telephone apparatus from computergenerated speech. The Bell System Technical Journal, 51 (2):391–397, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-2-391.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-391.pdf.

Fukui:1979:DBD

[Fuk79] H. Fukui. Determination of the basic device parameters of a GaAs MESFET. The Bell System Technical Journal, 58(3):771–797, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-771.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-771.pdf.

Fitzwilliam:1978:TNT

[FW78] J. W. Fitzwilliam and R. L.

Wagner. Transaction network, telephones, and terminals: Overview. The Bell System Technical Journal, 57(10): 3325–3329, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3325.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3325.pdf.

Fox:1978:SUC

[FYMS78] W. M. Fox, W. H. Yocom, P. R. Munk, and E. F. Sartori. SG undersea cable system: Semiconductor devices and passive components. The Bell System Technical Journal, 57(7):2405–2434, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2405.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2405.pdf.

Gans:1976:CPR

[Gan76]  M. J. Gans.  Cross polar-

ization in reflector-type beam waveguides and antennas. The Bell System Technical Journal, 55(3):289–316, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-3-289.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-3-289.pdf.

Garrett:1970:LAB

[Gar70]  C. G. B. Garrett.  Lead-

acid battery: Foreword. The Bell System Technical Journal, 49(7):1249–1252, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1249.pdf.

Garey:1972:OTP

[Gar72] M. R. Garey. Optimal test point selection for sequential manufacturing processes. The Bell System Technical Journal, 51(1):291–300, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-1-291.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-291.pdf.

Gardner:1975:MLO

[Gar75] W. B. Gardner. Microbending loss in optical fibers. The Bell System Technical Journal, 54(2):457–465, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-457.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-457.pdf.

Gawron:1975:SDP

[Gaw75] L. J. Gawron. SAFEGUARD data-processing system: System error control. The Bell System Technical Journal, 54(10): S123–S131, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S123.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S123.pdf.

Gashler:1976:RSN

[GAW^{+}76] R. J. Gashler, W. E. Archer, N. Wasserman, D. H. Yano, and R. C. Zolnoski. Remreed switching networks for no. 1 and no. 1A ESS: Remreed switches. The Bell System Technical Journal, 55(5):537–564, May/June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-537.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-537.pdf.

Grau:1976:RSN

[GBN^{+}76] T. G. Grau, E. Barzilai, R. I. Narney, R. N. Rath, D. E. Tompsett, and R. G. Whitfield. Remreed switching networks for no. 1 and no. 1A ESS: Remreed contact and switching network evaluation. The Bell System Technical Journal, 55(5):663–680, May/June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-5-663.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-5-663.pdf.

Geary:1979:MED

[Gea79] J. M. Geary. Monolithic electronic devises based on domain wall motion in a ferroeletric crystal. The Bell System Technical Journal, 58(2):467–489, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-467.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-467.pdf.

Gersho:1972:SSD

[Ger72] Allen Gersho. Stochastic stability of delta modulation. The Bell System Technical Journal, 51(4):821–841, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol51/bstj51-4-821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-821.pdf.

Geyling:1976:BFD

[Gey76] F. T. Geyling. Basic fluiddynamic considerations in the drawing of optical fibers. The Bell System Technical Journal, 55(8):1011–1056, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1011.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1011.pdf.

Gersho:1970:PFR

[GG70] Allen Gersho and David J.

Goodman. Projecting filters for recursive prediction of discrete-time processes. The Bell System Technical Journal, 49(9):2377–2403, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2377.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2377.pdf.

Goodman:1973:QND

[GG73] David J. Goodman and Larry J.

Greenstein. Quantizing noise of DELTA M/PCM encoders. The Bell System Technical Journal, 52(2):183–204, February 1973. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-2-183.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-2-183.pdf.

Gretter:1977:WMW

[GGL^{+}77] R. W. Gretter, R. P. Guenther, M. Lutchansky, D. Olasin, and A. B. Watrous. WT4 millimeter waveguide system: Mechanical design of sheathed waveguide medium. The Bell System Technical Journal, 56(10): 1899–1922, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-1899.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-1899.pdf.

Gersho:1979:CIT

[GGO79] A. Gersho, B. Gopinath, and A. M. Odlyzko. Coefficient inaccuracy in transversal filtering. The Bell System Technical Journal, 58(10):2301–2316, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol58/bstj58-10-2301.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol58-1979/articles/
                                        bstj58-10-2301.pdf.

Gusler:1970:SCC

[GH70] L. T. Gusler and D. C. Hogg.

Some calculations on coupling between satellite communications and terrestrial radio-relay systems due to scattering by rain. The Bell System Technical Journal, 49(7):1491–1511, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1491.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1491.pdf.

Gitlin:1975:TRS

[GH75] R. D. Gitlin and J. F. Hayes.

Timing recovery and scramblers in data transmission. The Bell System Technical Journal, 54(3):569–593, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-3-569.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-3-569.pdf.

Guenther:1977:WMW

[GH77] R. P. Guenther and W. M.

Hauser. WT4 millimeter waveguide system: Reliability and maintenance of the WT4 transmission medium. The Bell System Technical Journal, 56(10): 2035–2053, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2035.pdf;

                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-2035.pdf.

Greene:1977:NEN

[GHHJ77] T. V. Greene, D. G. Haenschke, B. H. Hornbach, and C. E. Johnson. No. 4 ESS: Network management and traffic administration. The Bell System Technical Journal, 56(7):1169–1202, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-7-1169.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-7-1169.pdf.

Goldstein:1972:MMI

[GHL72] A. J. Goldstein, I. D. Harmon, and A. B. Lesk. Manmachine interaction in humanface identification. The Bell System Technical Journal, 51 (2):399–427, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-2-399.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-399.pdf.

Gitlin:1973:PED

[GHM73] R. D. Gitlin, E. Y. Ho, and J. E. Mazo. Passband equalization of differentially phasemodulated data signals. The Bell System Technical Journal, 52(2):219–238, February 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-2-219.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-2-219.pdf.

Guinta:1977:NED

[GHRS77] J. A. Guinta, S. F. Heath III, J. T. Raleigh, and M. T. Smith, Jr. No. 4 ESS: Data/trunk administration and maintenance. The Bell System Technical Journal, 56(7):1203–1237, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1203.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1203.pdf.

Gerdine:1977:WMW

[GHWY77] M. A. Gerdine, L. W. Hinderks, S. D. Williams, and D. T. Young. WT4 millimeter waveguide system: Electrical transmission measurement system. The Bell System Technical Journal, 56 (10):2025–2034, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2025.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2025.pdf.

Gibson:1978:LPM

[Gib78] A. E. Gibson. Loop plant modeling: Cost models for loop plant work operations using semi-Markov processes. The Bell System Technical Journal, 57(4):927–940, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-927.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-927.pdf.

Gilbert:1975:LSP

[Gil75]  E. N. Gilbert.  Line-of-

sight paths over random terrain. The Bell System Technical Journal, 54(4):735–765, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-4-735.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-4-735.pdf.

Gilbert:1979:PPT

[Gil79] E. N. Gilbert. The packing problem for twisted pairs. The Bell System Technical Journal, 58 (10):2143–2162, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol58/bstj58-10-2143.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol58-1979/articles/
                                        bstj58-10-2143.pdf.

Giust:1971:TMR

[Giu71] O. Giust. TH-3 microwave radio system: Modulators. The Bell System Technical Journal, 50(7):2155–2173, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2155.pdf.

Gopinath:1976:TTR

[GK76] B. Gopinath and R. P. Kurshan. A touch-tone receivergenerator with digital channel filters. The Bell System Technical Journal, 55(4):455–467, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-4-455.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-4-455.pdf.

Glance:1971:PSM

[Gla71a]  B. Glance.  Power spectra

of multilevel digital phasemodulated signals. The Bell System Technical Journal, 50 (9):2857–2878, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-9-2857.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-9-2857.pdf.

Glance:1971:DPD

[Gla71b]  Bernard Glance.  Digital

phase demodulator. The Bell System Technical Journal, 50(3):933–949, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-933.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-933.pdf.

Gloge:1970:CMB

[Glo70]  Detlef Gloge.  Crosstalk in

multiple-beam waveguides. The Bell System Technical Journal, 49(1):55–71, January 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol49/bstj49-1-55.pdf;   http:
                                        //www.alcatel-lucent.com/bstj/
                                        vol49-1970/articles/bstj49-
                                        1-55.pdf.

Gloge:1972:OPF

[Glo72] D. Gloge. Optical power flow in multimode fibers. The Bell System Technical Journal, 51(8):1767–1783, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-8-1767.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-8-1767.pdf.

Gloge:1973:IRC

[Glo73] D. Gloge. Impulse response of clad optical multimode fibers. The Bell System Technical Journal, 52(6):801–816, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-801.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-801.pdf.

Gloge:1975:OFP

[Glo75] D. Gloge. Optical-fiber packaging and its influence on fiber straightness and loss. The Bell System Technical Journal, 54(2):245–262, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-245.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-245.pdf.

Gloge:1976:OTL

[Glo76] D. Gloge. Offset and tilt loss in optical fiber splices. The Bell System Technical Journal, 55(7):905–916, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-905.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-905.pdf.

Gershwin:1974:PLT

[GLW74] S. B. Gershwin, R. V. Laue, and Eric Wolman. Peak-load traffic administration of a rural multiplexer with concentration. The Bell System Technical Journal, 53(2):261–281, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-261.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-261.pdf.

Gopinath:1971:WTP

[GM71] B. Gopinath and D. Mitra. When are transistors passive? The Bell System Technical Journal, 50(8):2835–2847, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2835.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2835.pdf.

Gloge:1973:IRF

[GM73] D. Gloge and E. A. J. Marcatili. Impulse response of fibers with ring-shaped parabolic index distribution. The Bell System Technical Journal, 52 (7):1161–1168, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1161.
pdf;   http://www.alcatel-

lucent.com/bstj/vol52-1973/
articles/bstj52-7-1161.pdf.

Gopinath:1977:DTS

[GM77] B. Gopinath and J. A. Morrison. Discrete-time single server queues with correlated inputs. The Bell System Technical Journal, 56(9):1743–1768, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1743.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1743.pdf.

Goodman:1976:SEP

[GMN76] D. J. Goodman, B. J. Mc-Dermott, and L. H. Nakatani. Subjective evaluation of PCM coded speech. The Bell System Technical Journal, 55 (8):1087–1109, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1087.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1087.pdf.

Gopinath:1973:FQB

[GMS73] B. Gopinath, Debasis Mitra, and M. M. Sondhi. Formulas on queues in burst processes– I. The Bell System Technical Journal, 52(1):9–33, January 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol52/bstj52-1-
                                        9.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-1-9.pdf.

Gilbert:1974:CWD

[GMS74] E. N. Gilbert, F. J. MacWilliams, and N. J. A. Sloane. Codes which detect deception. The Bell System Technical Journal, 53(3):405–424, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-3-405.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-3-405.pdf.

Goell:1974:IAO

[Goe74a] J. E. Goell. Input amplifiers for optical PCM receivers. The Bell System Technical Journal, 53(9):1771–1793, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-9-1771.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-9-1771.pdf.

Goell:1974:ORH

[Goe74b] J. E. Goell. An optical repeater with high-impedance input amplifier. The Bell System Technical Journal, 53(4):629–643, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-4-629.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-629.pdf.

Goldman:1973:VSD

[Gol73] Joel Goldman. A Volterra series description of crosstalk interference in communications systems. The Bell System Technical Journal, 52(5):649–668, May/ June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-5-649.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-5-649.pdf.

Goldman:1974:SPS

[Gol74] Joel Goldman. Statistical properties of a sum of sinusoids and Gaussian noise and its generalization to higher dimensions. The Bell System Technical Journal, 53(4):557–580, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-4-557.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-557.pdf.

Goldsmith:1977:QFS

[Gol77]  P. F. Goldsmith.  A qua-

sioptical feed system for radioastronomical observations at millimeter wavelengths. The Bell System Technical Journal, 56(8):1483–1501, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-8-1483.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-8-1483.pdf.

Goodman:1971:DAA

[Goo71] David J. Goodman. A digital approach to adaptive delta modulation. The Bell System Technical Journal, 50(4):1421–1426, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1421.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1421.pdf.

Goodrich:1974:MTI

[Goo74]  L. M. Goodrich.  A map

technique for identifying variables of symmetry. The Bell System Technical Journal, 53(5):801–826, May/June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-801.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-801.pdf.

Gopinath:1971:CLM

[Gop71] B. Gopinath. On the control of linear multiple input-output systems. The Bell System Technical Journal, 50(3):1063–1081, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-1063.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-1063.pdf.

Garrison:1974:LST

[GOS74] J. L. Garrison, A. Olsen, Jr., and T. H. Simmonds, Jr. L5 system: Transmission networks and magnetic components. The Bell System Technical Journal, 53 (10):2203–2248, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol53/bstj53-10-2203.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-2203.pdf.

Gummel:1970:ICC

[GP70] H. K. Gummel and H. C.

Poon. An integral charge control model of bipolar transistors. The Bell System Technical Journal, 49(5):827–852, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-827.pdf.

Graham:1971:APL

[GP71] R. L. Graham and H. O. Pollak. On the addressing problem for loop switching. The Bell System Technical Journal, 50(8):2495–2519, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2495.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2495.pdf.

Gordon:1978:TNT

[GR78] T. H. Gordon and R. E. Reid. Transaction network, telephones, and terminals: Polled access interface. The Bell System Technical Journal, 57(10):3427– 3439, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3427.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3427.pdf.

Graham:1970:MSM

[Gra70a] R. L. Graham. A mathematical study of a model of magnetic domain interactions. The Bell System Technical Journal, 49(8):1627–1644, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1627.pdf.

Gray:1970:TTV

[Gra70b] D. A. Gray. Transit-time variations in line-of-sight tropospheric propagation paths. The Bell System Technical Journal, 49(6):1059–1068, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1059.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-1059.pdf.

Graham:1972:PDS

[Gra72] N. Y. Graham. Perspective drawing of surfaces with hidden line elimination. The Bell System Technical Journal, 51(4):843–861, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-843.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-843.pdf.

Greenstein:1973:SON

[Gre73] L. J. Greenstein. Slope overload noise in linear delta modulators with Gaussian inputs. The Bell System Technical Journal, 52(3):387–421, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-387.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-3-387.pdf.

Greenstein:1974:SBS

[Gre74] L. J. Greenstein. Spectrum of a binary signal block coded for DC suppression. The Bell System Technical Journal, 53(6):1103–1126, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-6-1103.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-6-1103.pdf.

Gross:1970:DPC

[GRW70] A. G. Gross, J. Raamot, and Mrs. S. B. Watkins. Device photolithography: Computer systems for pattern generator control. The Bell System Technical Journal, 49(9): 2011–2029, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2011.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2011.pdf.

Gunn:1971:PSM

[GRW71] J. F. Gunn, J. S. Ronne, and D. C. Weller. The Picturephone System: Mastergroup digital transmission on modern coaxial systems. The Bell System Technical Journal, 50(2):501–520, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-501.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-501.pdf.

Gopinath:1970:DSH

[GS70] B. Gopinath and M. M. Sondhi.

Determination of the shape of the human vocal tract from acoustical measurements. The Bell System Technical Journal, 49(6):1195–1214, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1195.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-1195.pdf.

Gitlin:1971:TRP

[GS71a] R. D. Gitlin and J. Salz.

Timing recovery in PAM systems. The Bell System Technical Journal, 50(5):1645–1669, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1645.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1645.pdf.

Gordon:1971:PSS

[GS71b] A. M. Gordon and J. B. Singleton. The Picturephone System: Station set components. The Bell System Technical Journal, 50(2):313–349, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-313.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-313.pdf.

Gibson:1973:IMS

[GS73a] T. A. Gibson and P. F. Stockhausen. Information management system: MASTER LINKS — A hierarchical data system. The Bell System Technical Journal, 52(10):1691–1724, December 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1691.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1691.pdf.

Goldstein:1973:PMT

[GS73b] A. J. Goldstein and D. G. Schweikert. A proper model for testing the planarity of electrical circuits. The Bell System Technical Journal, 52(1):135–142, January 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-1-135.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-135.pdf.

Glance:1974:LNI

[GS74a] B. Glance and W. W. Snell, Jr. Low-noise, integrated, millimeter-wave receiver. The Bell System Technical Journal, 53(7):1321–1328, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-7-1321.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-7-1321.pdf.

Green:1974:LSL

[GS74b] J. H. Green and R. W. Sanders.

L5 system: Line-protection switching. The Bell System Technical Journal, 53(10):2011– 2034, December 1974. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol53/bstj53-10-2011.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-2011.pdf.

Gans:1975:SFF

[GS75] M. J. Gans and R. A. Semplak. Some far-field studies of an offset launcher. The Bell System Technical Journal, 54(7):1319–1340, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1319.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-7-1319.pdf.

Gloge:1973:OFE

[GSBC73] D. Gloge, P. W. Smith, D. L. Bisbee, and E. L. Chinnock. Optical fiber end preparation for low-loss splices. The Bell System Technical Journal, 52(9):1579–1588, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-9-1579.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-9-1579.pdf.

Goodman:1979:OSP

[GSC^{+}79] D. J. Goodman, C. Scagliola, R. E. Cochiere, L. R. Rabiner, and J. Goodman. Objective and subjective performance of tandem connections of waveform coders with an LPC vocoder. The Bell System Technical Journal, 58(3):601–629, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-3-601.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-3-601.pdf.

Guess:1975:EFL

[Gue75] H. A. Guess. The effect of frame load and balance on dial-tone delay in no. 5 crossbar. The Bell System Technical Journal, 54 (10):1755–1793, December 1975.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-1755.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-1755.pdf.

Gummel:1970:CCR

[Gum70] H. K. Gummel. A charge control relation for bipolar transistors. The Bell System Technical Journal, 49(1):115–120, January 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-1-115.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-1-115.pdf.

Gray:1974:SCS

[GW74] R. M. Gray and A. D. Wyner.

Source coding for a simple network. The Bell System Technical Journal, 53(9): 1681–1721, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-9-1681.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-9-1681.pdf.

Gitlin:1979:RTW

[GW79] R. D. Gitlin and S. B. Weinstein.

On the required tap-weight precision for digitally implemented, adaptive, mean-squared equalizers. The Bell System Technical Journal, 58(2):301–321, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-301.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-301.pdf.

Haggerty:1975:SDP

[Hag75] J. P. Haggerty. SAFEGUARD data-processing system: Central logic and control operating system. The Bell System Technical Journal, 54(10):S89–S99, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-10-S89.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-10-S89.pdf.

Halfin:1970:OMC

[Hal70] Shlomo Halfin. An optimization method for cascaded filters. The Bell System Technical Journal, 49(2):185–190, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-2-185.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-2-185.pdf.

Halfin:1975:AMC

[Hal75]  S. Halfin.  An approximate

method for calculation delays for a family of cyclic-type queues. The Bell System Technical Journal, 54(10):1733–1754, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol54/bstj54-10-1733.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-1733.pdf.

Hamming:1970:DN

[Ham70] R. W. Hamming. On the distribution of numbers. The Bell System Technical Journal, 49(8):1609–1625, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1609.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1609.pdf.

Haskell:1972:BCS

[Has72] B. G. Haskell. Buffer and channel sharing by several interframe picturephone coders. The Bell System Technical Journal, 51(1):261–289, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-1-261.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-261.pdf.

Haskell:1975:EMN

[Has75] B. G. Haskell. Entropy measurements for nonadaptive and adaptive, frame-to-frame, linearpredictive coding of video telephone signals. The Bell System Technical Journal, 54(6): 1155–1174, July/August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-6-1155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-6-1155.pdf.

Hasegawa:1978:FEL

[Has78] A. Hasegawa. Free electron laser. The Bell System Technical Journal, 57(8):3069–3089, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-3069.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-3069.pdf.

Hayes:1974:PME

[Hay74] J. F. Hayes. Performance models of an experimental computer communication network. The Bell System Technical Journal, 53(2):225–259, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-225.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-225.pdf.

Heffes:1973:AFC

[Hef73] H. Heffes. Analysis of firstcome first-served queuing systems with peaked inputs. The Bell System Technical Journal, 52(7):1215–1228, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1215.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1215.pdf.

Helfand:1979:NIS

[Hel79] E. Helfand. Numerical integration of stochastic differential equations. The Bell System Technical Journal, 58(10): 2289–2299, December 1979. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol58/bstj58-10-2289.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-10-2289.pdf.

Henderson:1973:DEF

[Hen73a] D. M. Henderson. Dispersion and equalization in fiber optic communication systems. The Bell System Technical Journal, 52 (10):1867–1876, December 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1867.pdf;
http://www.alcatel-lucent.

                                        com/bstj/vol52-1973/articles/
                                        bstj52-10-1867.pdf.

Henry:1973:BJB

[Hen73b] Paul S. Henry. B.S.T.J. briefs: Attenuation through the clear atmosphere at 30, 19, and 13 GHz for low elevation angles. The Bell System Technical Journal, 52(6):1031–1035, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-6-1031.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-6-1031.pdf.

Herriott:1970:DPL

[Her70] Donald R. Herriott. Device photolithography: Lenses for the photolithographic system. The Bell System Technical Journal, 49(9):2105–2116, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2105.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2105.pdf.

Heyman:1975:DMA

[Hey75] D. P. Heyman. A diffusion model approximation for the GI/G/1 queue in heavy traffic. The Bell System Technical Journal, 54(9):1637–1646, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-9-1637.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-9-1637.pdf.

Hogg:1977:IRD

[HGLM77] D. C. Hogg, A. J. Giger, A. C.

Longton, and E. E. Muller. The influence of rain on design of 11-GHz terrestrial radio relay. The Bell System Technical Journal, 56(9):1575–1580, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1575.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1575.pdf.

Heffes:1973:PTC

[HH73] H. Heffes and J. M. Holtzman. Peakedness of traffic carried by a finite trunk group with renewal input. The Bell System Technical Journal, 52(9):1617–1642, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-9-1617.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-9-1617.pdf.

Heffes:1971:DAC

[HHJ71] H. Heffes, S. Horing, and D. L.

Jagerman. On the design and analysis of a class of PCM systems. The Bell System Technical Journal, 50(3):917–932, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-917.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-917.pdf.

Hopkins:1974:AIS

[HHM^{+}74] J. W. Hopkins, P. D. Hunter, R. E. Machol, J. J. DiSalvo, and R. J. Piereth. Automatic intercept system: File subsystem. The Bell System Technical Journal, 53(1):107–132, January 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-1-107.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-1-107.pdf.

Hill:1971:CEP

[Hil71]  F. S. Hill, Jr.  The com-

putation of error probability for digital transmission. The Bell System Technical Journal, 50(6):2055–2077, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-6-2055.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-6-2055.pdf.

Hill:1977:ETL

[Hil77] D. W. Hill. The effects of traffic load variation on measurement accuracy. The Bell System Technical Journal, 56(4):561–573, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-4-561.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-4-561.pdf.

Hamori:1971:TMR

[HJ71] A. Hamori and R. M. Jensen.

TH-3 microwave radio system: Microwave transmitter and receiver. The Bell System Technical Journal, 50(7): 2117–2135, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2117.pdf.

Hoover:1975:SDP

[HJ75] E. S. Hoover and R. A. Jacoby.

SAFEGUARD data-processing system: SAFEGUARD data reduction system. The Bell System Technical Journal, 54(10): S181–S189, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S181.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S181.pdf.

Huttenhoff:1977:NEP

[HJJ^{+}77] J. H. Huttenhoff, J. Janik, Jr., G. D. Johnson, W. R. Schleicher, M. F. Slana, and F. H. Tendick, Jr. No. 4 ESS: Peripheral system. The Bell System Technical Journal, 56(7):1029–1055, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-7-1029.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-7-1029.pdf.

Hogg:1973:BJB

[HL73] D. C. Hogg and W. E. Legg. B.S.T.J. briefs: A finline radiator. The Bell System Technical Journal, 52(7): 1249–1253, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-7-1249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-7-1249.pdf.

Hwang:1977:CGB

[HL77] F. K. Hwang and T. C. Liang. Construction for group-balanced connecting networks. The Bell System Technical Journal, 56(9):1643–1649, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-9-1643.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol56-1977/
articles/bstj56-9-1643.pdf.

Hwang:1978:CSZ

[HL78] F. K. Hwang and T. C. Liang.

The construction for symmetrical zone-balanced networks. The Bell System Technical Journal, 57(8):2973–2981, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2973.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2973.pdf.

Harrod:1979:BME

[HL79] W. L. Harrod and A. G. Lubowe.

The BELLPAC modular electronic packaging system. The Bell System Technical Journal, 58(10):2271–2288, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol58/bstj58-10-2271.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-10-2271.pdf.

Harper:1978:SUC

[HLL78] D. N. Harper, B. O. Larson, and M. Laurette. SG undersea cable system: Commissioning: Final system alignment and evaluation. The Bell System Technical Journal, 57(7):2547–2564, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-7-2547.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-7-2547.pdf.

Hamilton:1974:LSL

[HM74] B. H. Hamilton and S. Mottel. L5 system: Line-power feed. The Bell System Technical Journal, 53(10):1987–2009, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-1987.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-1987.pdf.

Hubbard:1970:MMT

[HMG70] W. M. Hubbard, G. D. Manderville, and J. E. Goell. Multilevel modulation techniques for millimeter guided waves. The Bell System Technical Journal, 49(1):33–54, January 1970. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol49/bstj49-1-
                                        33.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-1-33.pdf.

Halfin:1972:MGJ

[HMS72] S. Halfin, C. J. McCallum, Jr., and M. Segal. Modeling the growth of jumpers on the main distributing frame. The Bell System Technical Journal, 51(7):1517–1534, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1517.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1517.pdf.

Hilsinger:1977:PTI

[HMSV77] H. A. Hilsinger, K. D. Mozingo, C. F. Starnes, and G. A. Van-Dine. 1A processor: Testing and integration. The Bell System Technical Journal, 56(2):289–312, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-289.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-289.pdf.

Hill:1976:TCP

[HN76] D. W. Hill and S. R. Neal. Traffic capacity of a probabilityengineered truck group. The Bell System Technical Journal, 55(7):831–842, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-831.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-831.pdf.

Hunt:1978:CAM

[HN78] R. M. Hunt and J. W. Nippert. Computer-aided magnetic circuit design for a bell ringer. The Bell System Technical Journal, 57(1):179–203, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-1-179.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-1-179.pdf.

Harkless:1977:WMW

[HNW77] E. T. Harkless, A. J. Nardi, and H. C. Wang. WT4 millimeter waveguide system: Channelization. The Bell System Technical Journal, 56(10):2089– 2101, December 1977. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol56/bstj56-10-2089.pdf;
http://www.alcatel-lucent.
com/bstj/vol56-1977/articles/
bstj56-10-2089.pdf.

Hirt:1970:SST

[HO70] W. B. Hirt and D. O. Oldfather. SF system: Transmission tests, computations and equalization during installation. The Bell System Technical Journal, 49(5):783–798, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-783.
pdf;   http://www.alcatel-

lucent.com/bstj/vol49-1970/
articles/bstj49-5-783.pdf.

Ho:1971:OEE

[Ho71] E. Y. Ho. Optimum equalization and the effect of timing and carrier phase on synchronous data systems. The Bell System Technical Journal, 50(5):1671–1689, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1671.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1671.pdf.

Hwang:1977:PIA

[HO77] F. K. Hwang and A. M. Odlyzko.

A probability inequality and its application to switching networks. The Bell System Technical Journal, 56(5):821–826, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-821.pdf.

Holtzman:1971:ADE

[Hol71]  J. M. Holtzman.  Analysis

of dependence effects in telephone trunking networks. The Bell System Technical Journal, 50(8):2647–2662, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-8-2647.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-8-2647.pdf.

Holtzman:1973:BJB

[Hol73] J. M. Holtzman. B.S.T.J. briefs: The accuracy of the equivalent random method with renewal inputs. The Bell System Technical Journal, 52(9): 1673–1679, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-9-1673.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-9-1673.pdf.

Holden:1975:OFP

[Hol75] W. S. Holden. An opticalfrequency pulse-position-modulation experiment. The Bell System Technical Journal, 54 (2):285–296, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-2-285.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-2-285.pdf.

Howland:1970:DPO

[HP70] F. L. Howland and K. M. Poole. Device photolithography: An overview of the new mask-making system. The Bell System Technical Journal, 49(9):1997–2009, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-1997.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-1997.pdf.

Henning:1972:DCB

[HP72] H. H. Henning and J. W. Pan.

D2 channel bank: System aspects. The Bell System Technical Journal, 51(8):1641–1657, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1641.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1641.pdf.

Heindel:1973:IMS

[HR73] L. E. Heindel and J. T. Roberto.

Information management system: The-off-the-shelf system— A packaged information management system. The Bell System Technical Journal, 52(10): 1743–1763, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1743.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1743.pdf.

Herrmann:1973:PDR

[HRC73] O. Herrmann, L. R. Rabiner, and D. S. K. Chan. Practical design rules for optimum finite impulse response low-pass digital filters. The Bell System Technical Journal, 52(6):769–799, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-769.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-769.pdf.

Huseby:1970:LAB

[HRH70] T. W. Huseby, J. T. Ryan, and P. Hubbauer. Lead-acid battery: Polyvinyl chloride battery jars and covers. The Bell System Technical Journal, 49(7):1359–1376, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1359.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1359.pdf.

Hayes:1971:TAR

[HS71] J. F. Hayes and D. N. Sherman. Traffic analysis of a ring switched data transmission system. The Bell System Technical Journal, 50(9):2947–2978, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-9-2947.
pdf;   http://www.alcatel-

lucent.com/bstj/vol50-1971/
articles/bstj50-9-2947.pdf.

Hayes:1972:SDM

[HS72a] J. F. Hayes and D. N. Sherman. A study of data multiplexing techniques and delay performance. The Bell System Technical Journal, 51 (9):1983–2011, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-9-1983.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-9-1983.pdf.

Ho:1972:BJB

[HS72b] E. Y. Ho and D. A. Spaulding. B.S.T.J. briefs: Data transmission performance in the presence of carrier phase jitter and Gaussian noise. The Bell System Technical Journal, 51(8):1927–1931, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1927.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1927.pdf.

Hahn:1975:SDP

[HS75a] J. R. Hahn, Jr. and F. E. Slojkowski. SAFEGUARD dataprocessing system: Maintenance and diagnostic subsystem. The Bell System Technical Journal, 54(10):S63–S72, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-10-S63.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-10-S63.pdf.

Haskell:1975:LBR

[HS75b] B. G. Haskell and R. L. Schmidt. A low-bit-rate interframe coder for videotelephone. The Bell System Technical Journal, 54(8):1475–1495, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-8-1475.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-8-1475.pdf.

Hawley:1978:LPE

[HS78a] G. T. Hawley and K. E. Stiefel. Loop plant electronics: Voice frequency electronics for loop applications. The Bell System Technical Journal, 57(4):1071–1107, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-1071.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-1071.pdf.

Heffron:1978:TNT

[HS78b] W. G. Heffron, Jr. and N. E. Snow. Transaction network, telephones, and terminals: Transaction network service. The Bell System Technical Journal, 57 (10):3331–3347, December 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3331.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3331.pdf.

Heller:1970:TNS

[HST70] K. A. Heller, D. A. Schmitt, and R. M. Taylor. TSPS no. 1: System testing. The Bell System Technical Journal, 49(10): 2711–2728, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2711.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2711.pdf.

Hubbard:1971:MMD

[Hub71] W. M. Hubbard. A method for measurement of the duration of picosecond pulses by beat frequency detection of laser output. The Bell System Technical Journal, 50(1):1–21, January 1971. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol50/bstj50-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-1.pdf.

Hubbard:1973:UOF

[Hub73]  W. M. Hubbard.  Utiliza-

tion of optical-frequency carriers for low- and moderatebandwidth channels. The Bell System Technical Journal, 52(5):731–765, May/June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-5-731.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-5-731.pdf.

Huff:1979:AMP

[Huf79] D. L. Huff. Advanced mobile phone service: The developmental system. The Bell System Technical Journal, 58(1):249–269, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-1-249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-249.pdf.

Haugk:1976:RSN

[HW76] G. Haugk and E. G. Walsh. Remreed switching networks for no. 1 and no. 1A ESS: System overview. The Bell System Technical Journal, 55 (5):503–509, May/June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-5-503.

pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-503.pdf.

Henn:1978:LPE

[HW78] R. W. Henn and D. H.

Williamson. Loop plant electronics: Physical design considerations. The Bell System Technical Journal, 57(4):1185–1223, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-1185.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-1185.pdf.

Hwang:1971:BRN

[Hwa71] F. K. Hwang. On Beneš rearrangeable networks. The Bell System Technical Journal, 50(1):201–207, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-201.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-201.pdf.

Hwang:1976:RSM

[Hwa76] F. K. Hwang. On the rearrangeability of some multistage connecting networks. The Bell System Technical Journal, 55(9):1411–1422, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol55/bstj55-9-1411.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-9-1411.pdf.

Hwang:1979:TSM

[Hwa79] F. K. Hwang. Three-stage multiconnection networks which are nonblocking in the wide sense. The Bell System Technical Journal, 58(10):2183–2187, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol58/bstj58-10-2183.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol58-1979/articles/
                                        bstj58-10-2183.pdf.

Ho:1970:NAE

[HY70] E. Y. Ho and Y. S. Yeh. A new approach for evaluating the error probability in the presence of intersymbol interference and additive Gaussian noise. The Bell System Technical Journal, 49(9):2249–2265, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2249.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2249.pdf.

Haynie:1971:SCD

[HY71a] G. D. Haynie and S. Yang. Statistical circuit design: Confirmation of design using computercontrolled test sets. The Bell System Technical Journal, 50(4):1197–1208, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1197.pdf.

Ho:1971:EPM

[HY71b] E. Y. Ho and Y. S. Yeh. Error probability of a multilevel digital system with intersymbol interference and Gaussian noise. The Bell System Technical Journal, 50(3):1017–1023, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-1017.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-1017.pdf.

Ishizaka:1972:SVS

[IF72] K. Ishizaka and J. L. Flanagan. Synthesis of voiced sounds from a two-mass model of the vocal chords. The Bell System Technical Journal, 51(6):1233–1268, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1233.pdf.

Ishizaka:1977:APL

[IF77] K. Ishizaka and J. L. Flanagan. Acoustic properties of longitudinal displacement in vocal cord vibration. The Bell System Technical Journal, 56(6):889–918, July/ August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-6-889.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-6-889.pdf.

Irvin:1978:FMR

[IL78] J. C. Irvin and A. Loya. Failure mechanisms and reliability of low-noise GaAs FETs. The Bell System Technical Journal, 57(8):2823–2846, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-8-2823.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-8-2823.pdf.

Illium:1975:DDS

[ILR75] H. C. Illium, W. B. Lueft, and D. W. Rice. Digital data system: Physical design. The Bell System Technical Journal, 54(5):943–964, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-5-943.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol54-1975/
articles/bstj54-5-943.pdf.

Isaacs:1973:CUC

[IS73] J. C. Isaacs, Jr. and N. A.

Strakhov. Crosstalk in uniformly coupled lossy transmission lines. The Bell System Technical Journal, 52(1):101–115, January 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-1-101.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-101.pdf.

Ierley:1970:SPM

[IZ70] W. H. Ierley and H. Zucker.

A stationary phase method for the computation of the far field of open Cassegrain antennas. The Bell System Technical Journal, 49(3):431–454, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-3-431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-3-431.pdf.

Jackson:1970:IRN

[Jac70] Leland B. Jackson. On the interaction of roundoff noise and dynamic range in digital filters. The Bell System Technical Journal, 49(2):159–184, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-2-159.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-2-159.pdf.

Jacobs:1978:AFS

[Jac78] Ira Jacobs. Atlanta Fiber System Experiment: Overview. The Bell System Technical Journal, 57(6):1717–1721, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1717.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1717.pdf.

Jagerman:1970:ITA

[Jag70] David Jagerman. Information theory and approximation of bandlimited functions. The Bell System Technical Journal, 49(8):1911–1941, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1911.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1911.pdf.

Jagerman:1974:SPE

[Jag74] D. L. Jagerman. Some properties of the Erlang Loss Function. The Bell System Technical Journal, 53(3):525–551, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-3-525.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-3-525.pdf.

Jagerman:1975:NBT

[Jag75] D. L. Jagerman. Nonstationary blocking in telephone traffic. The Bell System Technical Journal, 54(3):625–661, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-3-625.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-3-625.pdf.

Jagerman:1978:ITL

[Jag78] D. L. Jagerman. An inversion technique for the Laplace transform with application to approximation. The Bell System Technical Journal, 57(3):669–710, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-669.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-669.pdf.

Jayant:1970:ADM

[Jay70] N. S. Jayant. Adaptive delta modulation with a one-bit memory. The Bell System Technical Journal, 49(3):321–342, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-3-321.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-3-321.pdf.

Jayant:1973:AQO

[Jay73] N. S. Jayant. Adaptive quantization with a one-word memory. The Bell System Technical Journal, 52(7):1119–1144, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-7-1119.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-7-1119.pdf.

Jayant:1974:CBB

[Jay74] N. S. Jayant. On the correlation between bit sequences in consecutive delta modulations of a speech signal. The Bell System Technical Journal, 53(5):937–946, May/June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-937.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-937.pdf.

Jayant:1975:ACT

[Jay75a] N. S. Jayant. An autocorrelation criterion for the timediversity reception of speech over burst-error channels. The Bell System Technical Journal, 54(9):1583–1595, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-9-1583.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-9-1583.pdf.

Jayant:1975:SST

[Jay75b]  N. S. Jayant.  Step-size

transmitting differential coders for mobile telephony. The Bell System Technical Journal, 54(9):1557–1581, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-9-1557.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-9-1557.pdf.

Jayant:1977:PAD

[Jay77] N. S. Jayant. Pitch-adaptive DPCM coding of speech with two-bit quantization and fixed spectrum prediction. The Bell System Technical Journal, 56(3):439–454, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-3-439.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-3-439.pdf.

Joyce:1971:LDR

[JB71] W. B. Joyce and W. J.

Bertram. Linearized dispersion relation and Green’s function for discrete-charge-transfer devices with incomplete transfer. The Bell System Technical Journal, 50(6):1741–1759, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1741.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1741.pdf.

Jayant:1979:AAC

[JC79] N. S. Jayant and S. W. Christensen. Adaptive aperture coding for speech waveforms-I. The Bell System Technical Journal, 58(7):1631–1645, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1631.pdf.

Jaeger:1970:TNS

[JJ70] R. J. Jaeger, Jr. and A. E. Joel, Jr. TSPS no. 1: System organization and objectives. The Bell System Technical Journal, 49 (10):2417–2443, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2417.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2417.pdf.

Johnson:1978:UTSb

[JL78] S. C. Johnson and M. E.

Lesk. UNIX time-sharing system: Language development tools. The Bell System Technical Journal, 57(6):2155–2175, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2155.pdf.

Jaeger:1978:POW

[JMM78] R. E. Jaeger, J. B. MacChesney, and T. J. Miller. The preparation of optical waveguide preforms by plasma deposition. The Bell System Technical Journal, 57(1):205–210, January 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-1-205.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-1-205.pdf.

Jansen:1971:TMR

[JP71] R. M. Jansen and R. C. Prime.

TH-3 microwave radio system: System considerations. The Bell System Technical Journal, 50(7):2085–2116, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2085.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2085.pdf.

Jordan:1978:RN

[JPS^{+}78] A. S. Jordan, R. H. Peaker, R. H. Saul, H. J. Braun, and H. H. Wade. The reliability of 302A numerics. The Bell System Technical Journal, 57(8):2983–2999, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-8-2983.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-8-2983.pdf.

Jayant:1971:PSO

[JR71] N. S. Jayant and A. E. Rosenberg. The preference of slope overload to granularity in the delta modulation of speech. The Bell System Technical Journal, 50(10):3117–3125, December 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3117.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3117.pdf.

Jayant:1972:ADQ

[JR72] N. S. Jayant and L. R. Rabiner. The application of dither to the quantization of speech signals. The Bell System Technical Journal, 51(6):1293–1304, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1293.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1293.pdf.

Johnson:1978:UTSa

[JR78] S. C. Johnson and D. M.

Ritchie. UNIX time-sharing system: Portability of C programs and the UNIX system. The Bell System Technical Journal, 57(6):2021–2048, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2021.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2021.pdf.

Johnson:1970:DPP

[JZ70] A. M. Johnson and A. Zacharias.

Device photolithography: The primary pattern generator: Part IV: Alignment and performance evaluation. The Bell System Technical Journal, 49(9): 2069–2075, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2069.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2069.pdf.

Kaiser:1974:LLS

[KA74] P. Kaiser and H. W. Astle. Low-loss single-material fibers made from pure fused silica. The Bell System Technical Journal, 53(6):1021–1039, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-6-1021.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-6-1021.pdf.

Kaiser:1970:ITG

[Kai70] P. Kaiser. An improved thermal gas lens for optical beam waveguides. The Bell System Technical Journal, 49(1):137–153, January 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-1-137.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-1-137.pdf.

Kailath:1971:SEI

[Kai71] Thomas Kailath. Some extensions of the innovations theorem. The Bell System Technical Journal, 50(4):1487–1494, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1487.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1487.pdf.

Kaneko:1970:UFS

[Kan70] H. Kaneko. A unified formulation of segment companding laws and synthesis of codecs and digital compandors. The Bell System Technical Journal, 49(7):1555–1588, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1555.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1555.pdf.

Karafin:1971:SCD

[Kar71] B. J. Karafin. Statistical circuit design: The optimum assignment of component tolerances for electrical networks. The Bell System Technical Journal, 50(4):1225–1242, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1225.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1225.pdf.

Kaufman:1974:FEM

[Kau74]  S. Kaufman.  Finite-element

method for the determination of thermal stresses in anisotropic solids of revolution. The Bell System Technical Journal, 53(5):827–845, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-5-827.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-827.pdf.

Kaufman:1977:FTD

[Kau77]  J. S. Kaufman.  Faulty-

trunk detection algorithms using EADAS/ICUR traffic data. The Bell System Technical Journal, 56(6):919–976, July/August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-6-919.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-6-919.pdf.

Kaskey:1978:CCI

[KCM^{+}78] B. Kaskey, J. S. Colson, R. F. Mills, F. H. Myers, J. T. Raleigh, A. F. Schweizer, and R. A. Tauson. Common channel interoffice signaling: Technology and hardware. The Bell System Technical Journal, 57(2):379–428, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-2-379.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-2-379.pdf.

Kenyon:1970:LCS

[Ken70] N. D. Kenyon. A lumped-circuit study of basic oscillator behavior. The Bell System Technical Journal, 49(2):255–272, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-255.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-255.pdf.

Kent:1973:CDB

[Ken73] W. H. Kent. Charge distribution in buried-channel chargecoupled devices. The Bell System Technical Journal, 52(6): 1009–1024, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-1009.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-1009.pdf.

Kerwin:1970:DPT

[Ker70] R. E. Kerwin. Device photolithography: Thin photosensitive materials. The Bell System Technical Journal, 49 (9):2179–2192, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2179.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2179.pdf.

Kessler:1971:TPB

[Kes71] J. E. Kessler. The transmission performance of Bell System toll connecting trunks. The Bell System Technical Journal, 50(8):2741–2777, October 1971. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-8-2741.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-8-2741.pdf.

Koontz:1970:LAB

[KFBL70] D. E. Koontz, D. O. Feder, L. D. Babusci, and H. J. Luer. Leadacid battery: Reserve batteries for Bell System use: Design of the new cell. The Bell System Technical Journal, 49(7):1253–1278, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1253.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1253.pdf.

Kurshan:1976:DST

[KG76] R. P. Kurshan and B. Gopinath. Digital single-tone generatordetectors. The Bell System Technical Journal, 55(4):469–496, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-4-469.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-4-469.pdf.

Kelcourse:1974:LSO

[KH74] F. C. Kelcourse and F. J. Herr. L5 system: Overall description and system design. The Bell System Technical Journal, 53 (10):1901–1933, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol53/bstj53-10-1901.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-1901.pdf.

Kak:1977:SEU

[KJ77] S. C. Kak and N. S. Jayant.

On speech encryption using waveform scrambling. The Bell System Technical Journal, 56(5):781–808, May/June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-781.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-781.pdf.

Kernighan:1970:EHP

[KL70] B. W. Kernighan and S. Lin.

An efficient heuristic procedure for partitioning graphs. The Bell System Technical Journal, 49(2):291–307, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-291.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-291.pdf.

Kernighan:1973:HSS

[KL73] B. W. Kernighan and S. Lin.

Heuristic solution of a signal design optimization. The Bell System Technical Journal, 52(7):1145–1159, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1145.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1145.pdf.

Klein:1972:SBT

[Kle72] T. H. Klein. Soil burial tests: Effect of soil burial exposure on the properties of structural grade reinforced plastic laminates. The Bell System Technical Journal, 51(1):51–62, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol51/bstj51-1-
51.pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-51.pdf.

Klein:1976:RSN

[Kle76] H. J. Klein. Remreed switching networks for no. 1 and no. 1A ESS: Physical design of remreed switching networks. The Bell System Technical Journal, 55(5):607–636, May/June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-607.
pdf;   http://www.alcatel-

lucent.com/bstj/vol55-1976/
articles/bstj55-5-607.pdf.

Kernighan:1978:UTS

[KLO78] B. W. Kernighan, M. E. Lesk, and J. F. Ossanna, Jr. UNIX time-sharing system: Document preparation. The Bell System Technical Journal, 57(6): 2115–2135, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2115.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2115.pdf;
https://www.tuhs.org/Archive/
Documentation/Papers/BSTJ/
bstj57-6-2115.pdf.

Kossyk:1970:DPP

[KLRS70] G. J. W. Kossyk, J. P. Laico, L. Rongved, and J. W. Stafford. Device photolithography: The primary pattern generator: Part II: Mechanical design. The Bell System Technical Journal, 49(9):2043–2059, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2043.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2043.pdf.

Kennedy:1976:RSN

[KLS76] J. C. Kennedy, W. A. Liss, and J. R. Smith. Remreed switching networks for no. 1 and no.

1A ESS: Remreed line scanner. The Bell System Technical Journal, 55(5):597–606, May/ June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-597.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-5-597.pdf.

Kaiser:1973:BJB

[KMM73] P. Kaiser, E. A. J. Marcatili, and S. E. Miller. B.S.T.J. briefs: A new optical fiber. The Bell System Technical Journal, 52(2):265–269, February 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-2-265.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-2-265.pdf.

Kuczura:1972:ACC

[KN72] A. Kuczura and S. R. Neal. The accuracy of call-congestion measurements for loss systems with renewal input. The Bell System Technical Journal, 51(10): 2197–2208, December 1972. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol51/bstj51-10-2197.pdf;
http://www.alcatel-lucent.
com/bstj/vol51-1972/articles/
bstj51-10-2197.pdf.

Kleinman:1978:NAP

[KN78] D. A. Kleinman and D. F.

Nelson. Nonlinear analysis of a photovoltaic optical telephone receiver. The Bell System Technical Journal, 57(5):1569–1596, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1569.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1569.pdf.

Knowlton:1977:CDO

[Kno77]  K. C. Knowlton.  Com-

puter displays optically superimposed on input devices. The Bell System Technical Journal, 56(3):367–383, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-3-367.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-3-367.pdf.

Kienzle:1970:NEA

[KNS^{+}70] H. G. Kienzle, K. L. Nicodemus, M. T. Smith, Jr., E. W. Weber, and H. M. Zydney. No. 1 ESS ADF: Message processing program organization. The Bell System Technical Journal, 49 (10):2753–2829, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/

                                        Vol49/bstj49-10-2753.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol49-1970/articles/
                                        bstj49-10-2753.pdf.

Kogelnik:1976:FRN

[Kog76] H. Kogelnik. Filter response of nonuniform almost-periodic structures. The Bell System Technical Journal, 55 (1):109–126, January 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-1-109.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-1-109.pdf.

Koontz:1978:LPMb

[Koo78a] W. L. G. Koontz. Loop plant modeling: An approach to modeling operating costs in the loop network. The Bell System Technical Journal, 57(4):891–909, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-891.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-891.pdf.

Koontz:1978:LPMa

[Koo78b] W. L. G. Koontz. Loop plant modeling: Economic evaluation of subscriber pair gain system applications. The Bell System Technical Journal, 57(4):825–848, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-825.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-825.pdf.

Kettley:1970:TNO

[KPS70] A. W. Kettley, E. J. Pasternak, and M. F. Sikorsky. TSPS no. 1: Operational programs. The Bell System Technical Journal, 49 (10):2625–2683, December 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol49/bstj49-10-2625.pdf;
http://www.alcatel-lucent.
com/bstj/vol49-1970/articles/
bstj49-10-2625.pdf.

Krambeck:1971:ZLT

[Kra71] R. H. Krambeck. Zero loss transfer across gaps in a CCD. The Bell System Technical Journal, 50(10):3169–3175, December 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol50/bstj50-10-3169.pdf;
http://www.alcatel-lucent.
com/bstj/vol50-1971/articles/
bstj50-10-3169.pdf.

Kropfl:1972:EDB

[Kro72] W. J. Kropfl. An experimental data block switching system. The Bell System Technical Journal, 51(6):1147–1165, July/ August 1972. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1147.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1147.pdf.

Krupp:1976:ATS

[Kru76]  R. S. Krupp.  Analysis of

toll switching networks. The Bell System Technical Journal, 55(7):843–856, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-7-843.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-7-843.pdf.

Krupp:1979:PKP

[Kru79] R. S. Krupp. Properties of Kruithof’s projection method. The Bell System Technical Journal, 58(2):517–538, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-517.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-517.pdf.

Kogelnik:1970:BJB

[KS70] H. Kogelnik and T. P. Sosnowski. B.S.T.J. briefs: Holographic thin film couplers. The Bell System Technical Journal, 49(7):1602–1608, September 1970. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1602.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1602.pdf.

Knapp:1971:TMH

[KS71] J. W. Knapp and K. L.

Seastrand. TH-3 mediumhaul application: Frequencydiplexed auxiliary channel. The Bell System Technical Journal, 50(7):2287–2313, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2287.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2287.pdf.

Knerr:1978:LNG

[KS78] R. H. Knerr and C. B. Swan.

A low-noise gallium arsenide field effect transistor amplifier for 4-GHz radio. The Bell System Technical Journal, 57(3):479–490, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-479.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-479.pdf.

Kahng:1974:IDD

[KSBL74] D. Kahng, W. J. Sundburg, D. M. Boulin, and J. R. Ligenza.

Interfacial dopants for dualdielectric, charge-storage cells. The Bell System Technical Journal, 53(9):1723–1739, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-9-1723.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-9-1723.pdf.

Krupp:1973:SNP

[KT73] R. S. Krupp and L. A. Tomko. Switching networks of planar shifting arrays. The Bell System Technical Journal, 52(6): 991–1007, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-991.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-991.pdf.

Klosterman:1976:RSN

[KU76] C. H. Klosterman and J. E. Unrue, Jr. Remreed switching networks for no. 1 and no. 1A ESS: Transmission design and environmental protection of remreed networks. The Bell System Technical Journal, 55(5):637–661, May/June 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-5-637.
pdf;   http://www.alcatel-

lucent.com/bstj/vol55-1976/
articles/bstj55-5-637.pdf.

Kuczura:1972:QMR

[Kuc72]  A. Kuczura.  Queues with

mixed renewal and Poisson inputs. The Bell System Technical Journal, 51(6):1305–1326, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1305.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1305.pdf.

Kuczura:1973:BIM

[Kuc73a] A. Kuczura. Batch input to a multiserver queue with constant service times. The Bell System Technical Journal, 52(1):83–99, January 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-1-
83.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-83.pdf.

Kuczura:1973:IPO

[Kuc73b] Anatol Kuczura. The interrupted Poisson as an overflow process. The Bell System Technical Journal, 52(3):437–448, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-437.
pdf;   http://www.alcatel-

lucent.com/bstj/vol52-1973/
articles/bstj52-3-437.pdf.

Kuehn:1979:MSN

[Kue79] P. J. Kuehn. Multiqueue systems with nonexhaustive cyclic service. The Bell System Technical Journal, 58(3):671–698, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-671.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-671.pdf.

Kuoni:1975:SDP

[Kuo75] J. P. Kuoni. SAFEGUARD dataprocessing system: A means to effective computer resource utilization. The Bell System Technical Journal, 54(10):S191– S196, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S191.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S191.pdf.

Kurshan:1970:BJB

[Kur70]  R. P. Kurshan.  B.S.T.J.

briefs: All terminal bubbles programs yield the elementary symmetric polynomials. The Bell System Technical Journal, 49(8):1991–1994, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1991.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1991.pdf.

Kerdock:1978:AFS

[KW78] R. S. Kerdock and D. H.

Wolaver. Atlanta Fiber System Experiment: Results of the Atlanta Experiment. The Bell System Technical Journal, 57(6):1857–1879, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1857.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1857.pdf.

Kwei:1972:SBT

[Kwe72] T. K. Kwei. Soil burial tests:

Effect of soil burial exposure on the properties of electrical grade reinforced plastic laminates. The Bell System Technical Journal, 51(1):47–49, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol51/bstj51-1-
47.pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-47.pdf.

Krambeck:1972:DST

[KWP72] R. H. Krambeck, R. H. Walden, and K. A. Pickar. A doped surface two-phase CCD. The Bell System Technical Journal, 51(8):1849–1866, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-8-1849.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-8-1849.pdf.

Laane:1970:MQN

[Laa70] R. R. Laane. Measured quantizing noise spectrum for singleintegration delta-modulation coders. The Bell System Technical Journal, 49(2):191–195, February 1970. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol49/bstj49-2-
                                        191.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-2-191.pdf.

Laane:1971:BJB

[Laa71] R. R. Laane. B.S.T.J. briefs: HILO — an improved transmission scheme for semiconductor switching networks. The Bell System Technical Journal, 50(3):1089–1093, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-1089.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-1089.pdf.

Laker:1979:ECA

[Lak79] K. R. Laker. Equivalent circuits for the analysis and synthesis of switched capacitor networks. The Bell System Technical Journal, 58(3):729–769, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-729.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-729.pdf.

Lapsa:1976:CMC

[Lap76] P. M. Lapsa. Calculation of multidisturber crosstalk probabilities — application to subscriberloop gain. The Bell System Technical Journal, 55 (7):875–903, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-875.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-875.pdf.

Lastovka:1976:OAV

[Las76] J. B. Lastovka. An optical apparatus for very-small-angle light scattering — design, analysis, and performance. The Bell System Technical Journal, 55(9):1225–1293, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-9-1225.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-9-1225.pdf.

Lycklama:1978:UTSa

[LB78] H. Lycklama and D. L. Bayer. UNIX time-sharing system: The MERT operating system. The Bell System Technical Journal, 57(6):2049–2086, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2049.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-2049.pdf.

Liu:1970:LBR

[LC70] S. C. Liu and D. K. Cohoon. Limiting behaviors of randomly excited hyperbolic tangent systems. The Bell System Technical Journal, 49(4):543–560, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-4-543.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-4-543.pdf.

Lycklama:1978:UTSc

[LC78] H. Lycklama and C. Christensen. UNIX time-sharing system: A minicomputer satellite processor system. The Bell System Technical Journal, 57(6):2103–2113, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2103.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2103.pdf.

Lee:1969:ECF

[Lee69] W. C. Y. Lee. An extended correlation function of two random variables applied to mobile radio transmission. The Bell System Technical Journal, 48(10): 3423–3440, December 1969. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol48/bstj48-10-3423.pdf;
http://www.alcatel-lucent.
com/bstj/vol48-1969/articles/
bstj48-10-3423.pdf. See erra-

tum [Ano70l].

Lee:1971:ASR

[Lee71]  W. C. Y. Lee.  Antenna

spacing requirement for a mobile radio base-station diversity. The Bell System Technical Journal, 50(6):1859–1876, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1859.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1859.pdf.

Leeper:1973:UDD

[Lee73] David G. Leeper. A universe digital data scrambler. The Bell System Technical Journal, 52 (10):1851–1865, December 1973.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol52/bstj52-10-1851.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol52-1973/articles/
                                        bstj52-10-1851.pdf.

Lee:1975:EJC

[Lee75] T. P. Lee. Effect of junction capacitance on the rise time of LED’s and on the turn-on delay of injection lasers. The Bell System Technical Journal, 54(1):53– 68, January 1975. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol54/bstj54-1-
                                        53.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-1-53.pdf.

Lenahan:1972:DCF

[Len72] T. A. Lenahan. Dimensionality of crosstalk functions. The Bell System Technical Journal, 51(6):1347–1362, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1347.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1347.pdf.

Lenahan:1977:ETP

[Len77a] T. A. Lenahan. Experimental test of propagation-parameter calculations for shielded balanced pair cables. The Bell System Technical Journal, 56(4):627–636, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-4-627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-4-627.pdf.

Lenahan:1977:TUCa

[Len77b] T. A. Lenahan. The theory of uniform cables — Part I: Calculation of propagation parameters. The Bell System Technical Journal, 56(4):597–610, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-4-597.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-4-597.pdf.

Lenahan:1977:TUCb

[Len77c] T. A. Lenahan. The theory of uniform cables — Part II: Calculation of charge components. The Bell System Technical Journal, 56(4):611–625, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-4-611.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-4-611.pdf.

Levinson:1978:ESA

[Lev78] S. E. Levinson. The effects of syntactic analysis on word recognition accuracy. The Bell System Technical Journal, 57(5):1627–1644, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1627.pdf.

Lewis:1970:FPP

[Lew70] J. A. Lewis. The flat plate problem for a semiconductor. The Bell System Technical Journal, 49(7):1483–1490, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1483.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1483.pdf.

Lewis:1976:BLC

[Lew76] J. A. Lewis. The boundary layer in a concentrated, multicomponent electrolyte. The Bell System Technical Journal, 55(6):803–824, July/August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-6-803.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-6-803.pdf.

Liu:1972:EEP

[LF72] S. C. Liu and L. W. Fagel. Earthquake environment for physical design: A statistical analysis. The Bell System Technical Journal, 51(9):1957–1982, November 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-9-1957.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-9-1957.pdf.

Limb:1972:BDG

[Lim72]  J. O. Limb.  Buffering of

data generated by the coding of moving images. The Bell System Technical Journal, 51(1):239–259, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-1-239.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-239.pdf.

Limb:1973:PCU

[Lim73]  J. O. Limb.  Picture cod-

ing: The use of a viewer model in source encoding. The Bell System Technical Journal, 52(8):1271–1302, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1271.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-8-1271.pdf.

Lin:1971:SBF

[Lin71] S. H. Lin. Statistical behavior of a fading signal. The Bell System Technical Journal, 50(10): 3211–3270, December 1971. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3211.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3211.pdf.

Lin:1973:SBR

[Lin73] S. H. Lin. Statistical behavior of rain attenuation. The Bell System Technical Journal, 52(4):557–581, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-4-557.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-4-557.pdf.

Lin:1975:MCR

[Lin75] S. H. Lin. A method for calculating rain attenuation distributions on microwave paths. The Bell System Technical Journal, 54(6):1051–1086, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-6-1051.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol54-1975/
articles/bstj54-6-1051.pdf.

Lin:1976:DRR

[Lin76a]  S. H. Lin.  Dependence of

rain-rate distribution on raingauge integration time. The Bell System Technical Journal, 55(1):135–142, January 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-1-135.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-1-135.pdf.

Lin:1976:RRD

[Lin76b] S. H. Lin. Rain-rate distributions and extreme-value statistics. The Bell System Technical Journal, 55(8):1111–1124, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1111.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1111.pdf.

Lin:1977:GRN

[Lin77a]  S. H. Lin.  11-GHz ra-

dio: Nationwide long-term rain rate statistics and empirical calculation of 11-GHz microwave rain attenuation. The Bell System Technical Journal, 56(9):1581–1604, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-9-1581.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-9-1581.pdf.

Lin:1977:IMD

[Lin77b] S. H. Lin. Impact of microwave depolarization during multipath fading on digital radio performance. The Bell System Technical Journal, 56(5):645–674, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-5-645.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-5-645.pdf.

Lin:1978:MRR

[Lin78]  S. H. Lin.  More on rain

rate distributions and extreme value statistics. The Bell System Technical Journal, 57 (5):1545–1568, May/June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1545.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1545.pdf.

Little:1978:CCI

[Lit78] R. S. Little. Common channel interoffice signaling: Field implementation. The Bell System Technical Journal, 57 (2):449–465, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-2-449.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-2-449.pdf.

Liu:1970:SSR

[Liu70] S. C. Liu. Synthesis of stochastic representations of ground motions. The Bell System Technical Journal, 49(4):521–541, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-4-521.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-4-521.pdf.

Liu:1971:TVS

[Liu71] S. C. Liu. Time-varying spectra and linear transformation. The Bell System Technical Journal, 50(7):2365–2374, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2365.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2365.pdf.

Liu:1974:ARB

[Liu74]  S. C. Liu.  Application of

response-bound method in shock and vibration analysis of telephone structures. The Bell System Technical Journal, 53 (7):1403–1426, September 1974.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-7-1403.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-7-1403.pdf.

Lloyd:1977:RVF

[Llo77]  S. P. Lloyd.  Rate vs. fi-

delity for the binary source. The Bell System Technical Journal, 56(3):427–437, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-3-427.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-3-427.pdf.

Laane:1970:DMC

[LM70] R. R. Laane and B. T. Murphy. Delta modulation codec for telephone transmission and switching applications. The Bell System Technical Journal, 49(6):1013–1031, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-1013.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1013.pdf.

Lewis:1976:FLC

[LM76] J. A. Lewis and J. McKenna. The field of a line charge near the tip of a dielectric wedge. The Bell System Technical Journal, 55(3):335–342, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-3-335.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-3-335.pdf.

Lawrence:1979:NIC

[LM79] V. B. Lawrence and K. V. Mina.

A new and interesting class of limit cycles in recursive digital filters. The Bell System Technical Journal, 58(2):379–408, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-2-379.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-2-379.pdf.

Luderer:1978:UTS

[LMT78] G. W. R. Luderer, J. F.

Maranzano, and B. A. Tague. UNIX time-sharing system: The UNIX operating system as a base for applications. The Bell System Technical Journal, 57(6):2201–2207, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2201.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2201.pdf.

LaCava:1976:LTU

[LMY76] B. R. LaCava, W. D. Miller, and B. Yaged. Last-trunk-usage measurements in step-by-step switching systems. The Bell System Technical Journal, 55(10): 1553–1572, December 1976. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol55/bstj55-10-1553.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol55-1976/articles/
                                        bstj55-10-1553.pdf.

Liu:1972:COM

[LN72] S. C. Liu and F. Neghabat. A cost optimization model for seismic design of structures. The Bell System Technical Journal, 51(10):2209–2225, December 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol51/bstj51-10-2209.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol51-1972/articles/
                                        bstj51-10-2209.pdf.

Luer:1977:NES

[LO77] H. J. Luer and R. Ostapiak. No. 4 ESS: System power. The Bell System Technical Journal, 56(7):1099–1111, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-7-1099.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol56-1977/
articles/bstj56-7-1099.pdf.

Logan:1971:SCD

[Log71] John Logan. Statistical circuit design: Characterization and modeling for statistical design. The Bell System Technical Journal, 50(4):1105–1147, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1105.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1105.pdf.

Logan:1977:IZC

[Log77] B. F. Logan, Jr. Information in the zero crossings of bandpass signals. The Bell System Technical Journal, 56(4):487–510, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-4-487.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-4-487.pdf.

Logan:1978:TAM

[Log78]  B. F. Logan, Jr.  Theory

of analytic modulation systems. The Bell System Technical Journal, 57(3):491–576, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-491.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-3-491.pdf.

Long:1978:LPM

[Lon78]  N. G. Long.  Loop plant

modeling: Overview. The Bell System Technical Journal, 57(4):797–806, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-797.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-797.pdf.

Liss:1977:WMW

[LOOS77] W. J. Liss, D. Olasin, J. W. Osmun, and S. Shapiro. The WT4 millimeter waveguide system: The WT4 repeater station. The Bell System Technical Journal, 56(10):2135–2146, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-2135.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-2135.pdf.

Limb:1971:SIC

[LP71] J. O. Limb and R. F. W. Pease. A sample interframe coder for video telephony. The Bell System Technical Journal, 50(6):1877–1888, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol50/bstj50-6-1877.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1877.pdf.

Limb:1974:CIF

[LPW74] J. O. Limb, R. F. W. Pease, and K. A. Walsh. Combining intraframe and frame-to-frame coding for television. The Bell System Technical Journal, 53(6):1137–1173, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-6-1137.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-6-1137.pdf.

Lundgren:1979:DRO

[LR79] C. W. Lundgren and W. D.

Rummler. Digital radio outage due to selective fading — observation vs prediction from laboratory simulation. The Bell System Technical Journal, 58(5):1073–1100, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-5-1073.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-5-1073.pdf.

Levinson:1978:EWR

[LRF78] S. E. Levinson, A. E. Rosenberg, and J. L. Flanagan. Evaluation of a word recognition system using syntax analysis. The Bell System Technical Journal, 57(5):1619–1626, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1619.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1619.pdf.

Lebert:1970:SSO

[LS70] A. W. Lebert and G. J. Schaible. SF system: Ocean cable and couplings. The Bell System Technical Journal, 49(5):699–719, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-5-699.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-5-699.pdf.

Landau:1971:SCE

[LS71] H. J. Landau and D. Slepian. Some computer experiments in picture processing for bandwidth reduction. The Bell System Technical Journal, 50 (5):1525–1540, May/June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-5-1525.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-5-1525.pdf.

Lynch:1970:SSS

[LTV70] R. L. Lynch, J. L. Thomas, and C. A. Von Roesgen. SF system: Shore terminal facilities and fault location. The Bell System Technical Journal, 49(5):721–748, May/June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-721.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-721.pdf.

Lueder:1970:DTF

[Lue70a] Ernst Lueder. A decomposition of a transfer function minimizing distortion and inband losses. The Bell System Technical Journal, 49(3):455–469, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-3-455.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-3-455.pdf.

Luer:1970:LAB

[Lue70b] H. J. Luer. Lead-acid battery: Incorporating the new battery into the telephone plant. The Bell System Technical Journal, 49(7):1447–1470, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1447.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1447.pdf.

Lueder:1972:GSO

[Lue72] E. Lueder. The general secondorder twin-T and its application to frequency-emphasizing networks. The Bell System Technical Journal, 51(1):301–316, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-1-301.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-301.pdf.

Lundgren:1970:SSA

[Lun70] C. W. Lundgren. A satellite system for avoiding serial Sun-transit outages and eclipses. The Bell System Technical Journal, 49(8):1943–1972, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1943.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1943.pdf.

Lundquist:1971:CSN

[Lun71] Leif Lundquist. Channel spacing and necessary bandwidth in FDM-FM systems. The Bell System Technical Journal, 50(3):869–880, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-869.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol50-1971/
articles/bstj50-3-869.pdf.

Lewis:1970:FSE

[LW70] J. A. Lewis and E. Wasserstrom. The field singularity at the edge of an electrode on a semiconductor surface. The Bell System Technical Journal, 49(6):1183–1194, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1183.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-1183.pdf.

Liu:1974:SDS

[LW74] C. K. Liu and T. L. Wang.

Sampled-data system approach to model time-division switches. The Bell System Technical Journal, 53(4):611–627, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-4-611.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-611.pdf.

Lycklama:1978:UTSb

[Lyc78] H. Lycklama. UNIX time-sharing system: UNIX on a microprocessor. The Bell System Technical Journal, 57(6):2087–2101, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol57/bstj57-6-2087.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-2087.pdf.

MacWilliams:1970:BCW

[Mac70] Mrs. F. J. MacWilliams. Binary codes which are ideals in the group algebra of an Abelian group. The Bell System Technical Journal, 49(6): 987–1011, July/August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-987.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-987.pdf.

MacDonald:1979:AMP

[Mac79] V. H. MacDonald. Advanced mobile phone service: The cellular concept. The Bell System Technical Journal, 58(1):15–41, January 1979. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol58/bstj58-1-
                                        15.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-15.pdf.

Morse:1978:SUC

[MAGS78] G. E. Morse, S. Ayers, R. F. Gleason, and J. R. Stauffer. SG undersea cable system: Cable and coupling design. The Bell System Technical Journal, 57(7):2435–2469, September 1978. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2435.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2435.pdf.

Manley:1974:UNI

[Man74] J. M. Manley. The use of negative-impedance units inserted uniformly into a transmission line reduce attenuation. The Bell System Technical Journal, 53(9):1845–1892, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-9-1845.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-9-1845.pdf.

Manhire:1978:PTC

[Man78]  L. M. Manhire.  Physical

and transmission characteristics of customer loop plant. The Bell System Technical Journal, 57(1):35–59, January 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol57/bstj57-1-
35.pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-1-35.pdf.

Marcatili:1970:DGC

[Mar70a] E. A. J. Marcatili. Dielectric guide with curved axis and truncated parabolic index. The Bell System Technical Journal, 49(8):1645–1663, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1645.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1645.pdf.

Marcuse:1970:EDM

[Mar70b] Dietrich Marcuse. Excitation of the dominant mode of a round fiber by a Gaussian beam. The Bell System Technical Journal, 49(8):1695–1703, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1695.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1695.pdf.

Marcuse:1970:RLT

[Mar70c] Dietrich Marcuse. Radiation losses of tapered dielectric slab waveguides. The Bell System Technical Journal, 49(2):273–290, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-273.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-273.pdf.

See erratum [Ano70k].

Marcuse:1970:RLD

[Mar70d] Dietrich Marcuse. Radiation losses of the dominant mode in round dielectric waveguides. The Bell System Technical Journal, 49(8):1665–1693, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1665.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1665.pdf.

Marcuse:1971:AUC

[Mar71a] D. Marcuse. Attenuation of unwanted cladding modes. The Bell System Technical Journal, 50(8):2565–2583, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2565.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2565.pdf.

Marcuse:1971:BLA

[Mar71b] D. Marcuse. Bending losses of the asymmetric slab waveguide. The Bell System Technical Journal, 50(8):2551–2563, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2551.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2551.pdf.

Marcuse:1971:CDM

[Mar71c] D. Marcuse. The coupling of degenerate modes in two parallel dielectric waveguides. The Bell System Technical Journal, 50(6):1791–1816, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-6-1791.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-6-1791.pdf.

Marcuse:1971:CCS

[Mar71d] D. Marcuse. Crosstalk caused by scattering in slab waveguides. The Bell System Technical Journal, 50(6):1817–1831, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-6-1817.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-6-1817.pdf.

Marcuse:1972:DCP

[Mar72a] D. Marcuse. Derivation of coupled power equations. The Bell System Technical Journal, 51(1):229–237, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-1-229.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-229.pdf.

Marcuse:1972:FPC

[Mar72b] D. Marcuse. Fluctuations of the power of coupled modes. The Bell System Technical Journal, 51(8):1793–1800, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1793.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1793.pdf.

Marcuse:1972:HOL

[Mar72c] D. Marcuse. Higher-order loss processes and the loss penalty of multimode operation. The Bell System Technical Journal, 51(8):1819–1836, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1819.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1819.pdf.

Marcuse:1972:HOS

[Mar72d] D. Marcuse. Higher-order scattering losses in dielectric waveguides. The Bell System Technical Journal, 51(8):1801–1817, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1801.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1801.pdf.

Marcuse:1972:PDR

[Mar72e] D. Marcuse. Power distribution and radiation losses in multimode dielectric slab waveguides. The Bell System Technical Journal, 51(2):429–454, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-2-429.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-429.pdf.

Marcuse:1972:PPT

[Mar72f] D. Marcuse. Pulse propagation in a two-mode waveguide. The Bell System Technical Journal, 51(8):1785–1791, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-8-1785.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-8-1785.pdf.

Marcuse:1972:PPM

[Mar72g] D. Marcuse. Pulse propagation in multimode dielectric waveguides. The Bell System Technical Journal, 51(6):1199–1232, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1199.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1199.pdf.

Marcuse:1973:CMT

[Mar73a] D. Marcuse. Coupled mode theory of round optical fibers. The Bell System Technical Journal, 52(6):817–842, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-817.pdf.

Marcuse:1973:CCI

[Mar73b] D. Marcuse. Coupling coefficients for imperfect asymmetric slab waveguides. The Bell System Technical Journal, 52(1):63– 82, January 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-1-
63.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-63.pdf.

Marcuse:1973:IRO

[Mar73c] D. Marcuse. The impulse response of an optical fiber with parabolic index profile. The Bell System Technical Journal, 52(7):1169–1174, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1169.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1169.pdf.

Marcuse:1973:LIR

[Mar73d] D. Marcuse. Losses and impulse response of a parabolic index fiber with random bends. The Bell System Technical Journal, 52(8):1423–1437, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1423.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1423.pdf.

Marcuse:1973:SLC

[Mar73e] D. Marcuse. Scattering losses caused by the support structure of an uncladded fiber. The Bell System Technical Journal, 52(2):205–217, February 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-2-205.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-2-205.pdf.

Marcatili:1974:SCW

[Mar74a] E. A. J. Marcatili. Slabcoupled waveguides. The Bell System Technical Journal, 53(4):645–674, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-4-645.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-645.pdf.

Marcuse:1974:BOW

[Mar74b] D. Marcuse. Bent optical waveguide with lossy jacket. The Bell System Technical Journal, 53(6):1079–1101, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-6-1079.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-6-1079.pdf.

Marcuse:1974:LIR

[Mar74c] D. Marcuse. Losses and impulse response in parabolic index fibers with square cross section. The Bell System Technical Journal, 53(2):195–215, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-195.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-195.pdf.

Marcuse:1974:RSI

[Mar74d] D. Marcuse. Rayleigh scattering and the impulse response of optical fibers. The Bell System Technical Journal, 53(4):705–715, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-4-705.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-705.pdf.

Marcuse:1974:RMP

[Mar74e] D. Marcuse. Reduction of multimode pulse dispersion by intentional mode coupling. The Bell System Technical Journal, 53(9):1795–1815, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-9-1795.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-9-1795.pdf.

Marcuse:1974:TSM

[Mar74f] D. Marcuse. Theory of the single-material fiber. The Bell System Technical Journal, 53(8):1619–1641, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-8-1619.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-8-1619.pdf.

Marcuse:1975:CMT

[Mar75a] D. Marcuse. Coupled-mode theory for anisotropic optical waveguides. The Bell System Technical Journal, 54(6):985–995, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-6-985.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-6-985.pdf.

Marcuse:1975:EPI

[Mar75b] D. Marcuse. Excitation of parabolic-index fibers with incoherent sources. The Bell System Technical Journal, 54 (9):1507–1530, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-9-1507.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-9-1507.pdf.

Marcuse:1976:ETT

[Mar76a] D. Marcuse. Exact theory of TE-wave-scattering from blazed dielectric gratings. The Bell System Technical Journal, 55(9):1295–1317, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-9-1295.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-9-1295.pdf.

Marcuse:1976:MLS

[Mar76b] D. Marcuse. Microbending losses of single-mode, step-index and multimode, parabolic-index fibers. The Bell System Technical Journal, 55(7):937–955, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-937.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-7-937.pdf.

Marcuse:1976:MMR

[Mar76c] D. Marcuse. Mode mixing with reduced losses in parabolic-index fibers. The Bell System Technical Journal, 55(6):777–802, July/ August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-6-777.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-6-777.pdf.

Marcuse:1976:SAL

[Mar76d] D. Marcuse. Scattering and absorption losses of multimode optical fibers and fiber lasers. The Bell System Technical Journal, 55(10):1463–1488, December 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol55/bstj55-10-1463.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol55-1976/articles/
                                        bstj55-10-1463.pdf.

Marcuse:1976:SSL

[Mar76e] D. Marcuse. Steady-state losses of optical fibers and fiber resonators. The Bell System Technical Journal, 55(10):1445– 1462, December 1976. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol55/bstj55-10-1445.pdf;

http://www.alcatel-lucent.
com/bstj/vol55-1976/articles/
bstj55-10-1445.pdf.

Marcatili:1977:MDO

[Mar77a] E. A. J. Marcatili. Modal dispersion in optical fibers with arbitrary numerical aperture and profile dispersion. The Bell System Technical Journal, 56(1):49– 63, January 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol56/bstj56-1-
49.pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-1-49.pdf.

Marcuse:1977:LAS

[Mar77b] D. Marcuse. Loss analysis of single-mode fiber splices. The Bell System Technical Journal, 56(5):703–718, May/June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-703.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-703.pdf.

Marcuse:1977:TDG

[Mar77c] D. Marcuse. Thick dielectric grating on asymmetric slab waveguide. The Bell System Technical Journal, 56(3):329– 353, March 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol56/bstj56-3-

                                        329.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-3-329.pdf.

Marsh:1978:LPM

[Mar78] B. L. Marsh. Loop plant modeling: The feeder allocation process. The Bell System Technical Journal, 57(4):869–890, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-869.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-869.pdf.

Mattes:1970:TEP

[Mat70] H. G. Mattes. Thermal and electrical properties of coated conductive substrates for integrated circuit chip mounting. The Bell System Technical Journal, 49(6):1151–1182, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-1151.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1151.pdf.

Maurer:1974:LSJ

[Mau74] R. E. Maurer. L5 system: Jumbogroup multiplex terminal. The Bell System Technical Journal, 53(10):2065–2096, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/

Vol53/bstj53-10-2065.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-2065.pdf.

Maydan:1971:MIR

[May71] D. Maydan. Micromachining and image recording on thin films by laser beams. The Bell System Technical Journal, 50(6):1761–1789, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1761.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1761.pdf.

Mazo:1971:TSA

[Maz71]  J. E. Mazo.  Theory for

some asynchronous time-division switches. The Bell System Technical Journal, 50(3):983–1016, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-983.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-983.pdf.

Mazo:1975:FTN

[Maz75a] J. E. Mazo. Faster-than-Nyquist signaling. The Bell System Technical Journal, 54 (8):1451–1462, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol54/bstj54-8-1451.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-8-1451.pdf.

Mazo:1975:GDF

[Maz75b] J. E. Mazo. A geometric derivation of Forney’s upper bound. The Bell System Technical Journal, 54(6):1087–1094, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-6-1087.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-6-1087.pdf.

Mazo:1975:OTP

[Maz75c] J. E. Mazo. Optimum timing phase for an infinite equalizer. The Bell System Technical Journal, 54(1):189–201, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-1-189.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-1-189.pdf.

Mazo:1977:ABT

[Maz77]  J. E. Mazo.  On the an-

gle between two Fourier subspaces. The Bell System Technical Journal, 56(3):411–426, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-3-411.

pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-3-411.pdf.

Mazo:1978:JCT

[Maz78a] J. E. Mazo. Jitter comparison of tones generated by squaring and by fourth-power circuits. The Bell System Technical Journal, 57(5):1489–1498, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1489.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1489.pdf.

Mazo:1978:SHO

[Maz78b] J. E. Mazo. On the stability of higher order digital filters which use saturation arithmetic. The Bell System Technical Journal, 57(3):747–763, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-3-747.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-3-747.pdf.

Mazo:1979:ITE

[Maz79a] J. E. Mazo. On the independence theory of equalizer convergence. The Bell System Technical Journal, 58(5):963–993, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol58/bstj58-5-963.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-5-963.pdf.

Mazo:1979:STO

[Maz79b] J. E. Mazo. Some theoretical observations on spreadspectrum communications. The Bell System Technical Journal, 58(9):2013–2023, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-9-2013.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-9-2013.pdf.

Mitchell:1971:FDE

[MB71] O. M. Mracek Mitchell and David A. Berkley. A fullduplex echo suppressor using center-clipping. The Bell System Technical Journal, 50(5): 1619–1630, May/June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-5-1619.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-5-1619.pdf.

Mounts:1979:OSF

[MBN79] F. W. Mounts, E. G. Bowen, and A. N. Netravali. An ordering scheme for facsimile coding. The Bell System Technical Journal, 58(9):2113–2128, November 1979. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-9-2113.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-9-2113.pdf.

Morrison:1973:BJBb

[MC73] J. A. Morrison and T. S. Chu.

B.S.T.J. briefs: Perturbation calculations of rain-induced differential attenuation and differential phase shift at microwave frequencies. The Bell System Technical Journal, 52(10):1907– 1913, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1907.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1907.pdf.

Morrison:1974:SPE

[MC74] J. A. Morrison and M. J. Cross.

Scattering of a plane electromagnetic wave by axisymmetric raindrops. The Bell System Technical Journal, 53(6):955–1019, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-6-955.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-6-955.pdf.

Morrison:1973:BJBa

[MCC73] J. A. Morrison, M. J. Cross, and T. S. Chu. B.S.T.J. briefs: Rain-induced differential attenuation and differential phase shift at microwave frequencies. The Bell System Technical Journal, 52(4):599–604, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-4-599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-4-599.pdf.

Marr:1972:FBI

[MCKP72] G. Marr, G. T. Cheney, E. F. King, and E. G. Parks. A fast bipolar-IGFET buffer-driver. The Bell System Technical Journal, 51(2):363–370, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-363.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-2-363.pdf.

McMahon:1978:UTS

[MCM78] L. E. McMahon, L. L. Cherry, and R. Morris. UNIX timesharing system: Statistical text processing. The Bell System Technical Journal, 57(6): 2137–2154, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2137.
pdf;   http://www.alcatel-

lucent.com/bstj/vol57-1978/
articles/bstj57-6-2137.pdf.

Means:1975:TCT

[Mea75]  D. R. Means.  TI car-

bon transmitter model for use in computer-sided analysis of telephone set transmission characteristics. The Bell System Technical Journal, 54(7):1301–1318, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1301.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1301.pdf.

Messerli:1972:BJB

[Mes72] E. J. Messerli. B.S.T.J. briefs:

Proof of a convexity property of the Erlang B formula. The Bell System Technical Journal, 51(4):951–953, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-951.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-951.pdf.

Messerschmitt:1973:GTIa

[Mes73a] D. G. Messerschmitt. A geometric theory of intersymbol interference, Part I: Zero-forcing and decision-feedback equalization. The Bell System Technical Journal, 52(9):1483–1519, November 1973. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-9-1483.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-9-1483.pdf.

Messerschmitt:1973:GTIb

[Mes73b] D. G. Messerschmitt. A geometric theory of intersymbol interference, Part II: Performance of the maximum likelihood detector. The Bell System Technical Journal, 52(9):1521–1539, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-9-1521.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-9-1521.pdf.

Meseke:1975:SDP

[Mes75] D. W. Meseke. SAFEGUARD data-processing system: The data-processing system performance requirements in retrospect. The Bell System Technical Journal, 54(10):S29–S37, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-10-S29.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-10-S29.pdf.

Messerli:1978:AVU

[Mes78] E. J. Messerli. An approximation for the variance of the UPCO offered load estimate. The Bell System Technical Journal, 57(7):2575–2587, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2575.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2575.pdf.

Mettler:1979:GCS

[Met79] S. C. Mettler. A general characterization of splice loss for multimode optical fibers. The Bell System Technical Journal, 58 (10):2163–2182, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol58/bstj58-10-2163.pdf;
http://www.alcatel-lucent.
com/bstj/vol58-1979/articles/
bstj58-10-2163.pdf.

Mueller:1970:CLC

[MF70] R. K. Mueller and G. J. Foschini. The capacity of linear channels with additive Gaussian noise. The Bell System Technical Journal, 49(1):81–94, January 1970. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol49/bstj49-1-
81.pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-1-81.pdf.

Marcatili:1973:MTG

[MG73] E. A. J. Marcatili and D. Gloge. Multimode theory of gradedcore fibers. The Bell System Technical Journal, 52(9): 1563–1578, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-9-1563.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-9-1563.pdf.

Morris:1975:WGI

[MG75] J. E. Morris and J. W. Gewartowski. A 1-watt, 6-Gigahertz IMPATT amplifier for short-haul radio applications. The Bell System Technical Journal, 54(4):721–733, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-4-721.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-4-721.pdf.

Mitra:1978:APS

[MG78] Debasis Mitra and B. Gotz. An adaptive PCM system designed for noisy channels and digital implementations. The Bell System Technical Journal, 57(7):2727–2763, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-7-2727.

pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2727.pdf.

Melchior:1978:AFS

[MHSS78] H. Melchior, A. R. Hartman, D. P. Schinke, and T. E. Seidel. Atlanta Fiber System Experiment: Planar epitaxial silicon avalanche photodiode. The Bell System Technical Journal, 57(6):1791–1807, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1791.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1791.pdf.

Milner:1970:LAB

[Mil70] Paul C. Milner. Lead-acid battery: Float behavior of the lead-acid battery system. The Bell System Technical Journal, 49(7):1321–1334, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1321.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1321.pdf.

Miller:1972:CSB

[Mil72] C. M. Miller. Capacitances of a shielded balanced-pair transmission line. The Bell System Technical Journal, 51(3):759–776, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-3-759.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-3-759.pdf.

Miller:1974:DDG

[Mil74] S. E. Miller. Delay distortion in generalized lens-like media. The Bell System Technical Journal, 53(2):177–193, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-2-177.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-2-177.pdf.

See erratum [Ano75k].

Miller:1975:FOC

[Mil75a]  C. M. Miller.  A fiber-

optic-cable connector. The Bell System Technical Journal, 54(9):1547–1555, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-9-1547.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-9-1547.pdf.

Miller:1975:LTS

[Mil75b]  C. M. Miller.  Loose tube

splices for optical fibers. The Bell System Technical Journal, 54(7):1215–1225, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1215.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1215.pdf.

Miller:1975:ELI

[Mil75c] G. Miller. The effects of longitudinal imbalance on crosstalk. The Bell System Technical Journal, 54(7):1227–1251, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1227.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1227.pdf.

Miller:1976:LFR

[Mil76a]  C. M. Miller.  Laminated

fiber ribbon for optical communication cables. The Bell System Technical Journal, 55 (7):929–935, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-929.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-929.pdf.

Miller:1976:TVT

[Mil76b] C. M. Miller. Transmission vs transverse offset for parabolicprofile fiber splices with unequal core diameters. The Bell System Technical Journal, 55(7):917–927, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-7-917.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-7-917.pdf.

Miller:1978:FOA

[Mil78] C. M. Miller. Fiber-optic array splicing with etched silicon chips. The Bell System Technical Journal, 57(1):75–90, January 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol57/bstj57-1-
                                        75.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-75.pdf.

Miner:1972:SBT

[Min72] R. J. Miner. Soil burial tests: Effect of soil burial exposure on the properties of molded plastics. The Bell System Technical Journal, 51(1):23–42, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol51/bstj51-1-
                                        23.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-1-23.pdf.

Mitra:1974:MAA

[Mit74] Debasis Mitra. Mathematical analysis of an adaptive quantizer. The Bell System Technical Journal, 53(5):867–898, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-5-867.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-5-867.pdf.

Mitra:1975:NRM

[Mit75] Debasis Mitra. New results from a mathematical study of an adaptive quantizer. The Bell System Technical Journal, 54(2):335–368, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-335.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-335.pdf.

Mitra:1977:CDI

[Mit77] Debasis Mitra. Criteria for determining if a high-order digital filter using saturation arithmetic is free of overflow oscillations. The Bell System Technical Journal, 56(9):1679–1699, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1679.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1679.pdf.

Miller:1979:BJB

[ML79] R. C. Miller and R. B. Lawry.

B.S.T.J. briefs: Optically powered speech communication over a fiber lightguide. The Bell System Technical Journal, 58(7):1735–1741, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1735.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1735.pdf.

Marcuse:1971:EWI

[MM71a] D. Marcuse and E. A. J. Marcatili. Excitation of waveguides for integrated optics with laser beams. The Bell System Technical Journal, 50(1):43–57, January 1971. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol50/bstj50-1-
43.pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-43.pdf.

Millard:1971:PSD

[MM71b] J. B. Millard and H. I. Maunsell. The Picturephone System: Digital encoding of the video signal. The Bell System Technical Journal, 50(2):459–479, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-2-459.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-2-459.pdf.

McKenna:1972:ESS

[MM72a] J. McKenna and J. A. Morrison. Exact solutions to some deterministic and random transmission line problems. The Bell System Technical Journal, 51(6):1269–1292, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1269.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1269.pdf.

Morrison:1972:CLE

[MM72b] J. A. Morrison and J. McKenna.

Coupled line equations with random coupling. The Bell System Technical Journal, 51 (1):209–228, January 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-1-209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-209.pdf.

Marcuse:1973:TWO

[MM73] D. Marcuse and W. L. Mammel.

Tube waveguide for optical transmission. The Bell System Technical Journal, 52(3):423–435, March 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-3-423.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-3-423.pdf.

Marr:1975:DLP

[MM75] G. Marr and G. L. Mowery. A depletion-load, pchannel, bipolar-IGFET technology. The Bell System Technical Journal, 54(1):69–79, January 1975. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol54/bstj54-1-
                                        69.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-1-69.pdf.

Massey:1978:CSS

[MM78a] W. A. Massey and J. A. Morrison. Calculation of steadystate probabilities for content of buffer with correlated inputs. The Bell System Technical Journal, 57(9):3097–3117, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-9-3097.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-9-3097.pdf.

Miller:1978:LMP

[MM78b] C. M. Miller and S. C. Mettler. A loss model for parabolicprofile fiber splices. The Bell System Technical Journal, 57(9):3167–3180, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-9-3167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-9-3167.pdf.

Maldonado:1979:FST

[MM79a] J. R. Maldonado and D. Maydan. Fast simultaneous thickness measurements of gold and nickel layers on copper substrates. The Bell System Technical Journal, 58(8):1851–1868, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1851.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-8-1851.pdf.

Moran:1979:HRS

[MM79b] J. M. Moran and D. Maydan. High resolution, steep profile, resist patterns. The Bell System Technical Journal, 58(5):1027–1036, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-5-1027.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-5-1027.pdf.

Meitzler:1970:ISD

[MMF70] A. H. Meitzler, J. R. Maldonado, and D. B. Fraser. Image storage and display devices using fine-grain, ferroelectric ceramics. The Bell System Technical Journal, 49(6):953–967, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-953.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-953.pdf.

Mahoney:1975:DDS

[MMM75] J. J. Mahoney, Jr., J. J. Mansell, and R. C. Matlack. Digital data system: User’s view of the network. The Bell System Technical Journal, 54(5):833–844, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-5-833.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-5-833.pdf.

Mounts:1977:DQR

[MNP77] F. W. Mounts, A. N. Netravali, and B. Prasada. Design of quantizers for real-time Hadamard-transform coding of pictures. The Bell System Technical Journal, 56(1):21–48, January 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol56/bstj56-1-
                                        21.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-1-21.pdf.

Manfred:1978:LPE

[MNS78] M. T. Manfred, G. A. Nelson, and C. H. Sharpless. Loop plant electronics: Digital loop carrier systems. The Bell System Technical Journal, 57(4):1129–1156, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-1129.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-1129.pdf.

Mounts:1978:SEO

[MNW78] F. W. Mounts, A. N. Netravali, and K. A. Walsh. Some extensions of the ordering techniques for compression of twolevel facsimile pictures. The Bell System Technical Journal, 57(8):3057–3067, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-3057.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-3057.pdf.

Moorthy:1970:ATC

[Moo70] S. C. Moorthy. Analysis of a thin circular loop antenna over a homogeneous Earth. The Bell System Technical Journal, 49(6):1215–1233, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1215.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1215.pdf.

Morrison:1970:AAN

[Mor70] J. A. Morrison. Asymptotic analysis of a nonlinear autonomous vibratory system. The Bell System Technical Journal, 49(1):73– 80, January 1970. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol49/bstj49-1-
                                        73.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-1-73.pdf.

Moreland:1971:PSM

[Mor71]  J. P. Moreland.  Perfor-

mance of a system of mutually synchronized clocks. The Bell System Technical Journal, 50(7):2449–2464, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2449.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2449.pdf.

McCabe:1977:NEP

[MOR^{+}77] P. S. McCabe, J. B. Otto, S. Roy, G. A. Sellers, Jr., and K. W. Zweifel. No. 4 ESS: Program administration, test, and evaluation. The Bell System Technical Journal, 56(7):1239–1277, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1239.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1239.pdf.

Moreland:1978:EPP

[Mor78a] J. P. Moreland. Estimation of point-to-point telephone traffic. The Bell System Technical Journal, 57(8):2847–2863, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2847.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2847.pdf.

Morgen:1978:LPE

[Mor78b] D. H. Morgen. Loop plant electronics: Maintenance and administration of loop electronics. The Bell System Technical Journal, 57(4):1055–1070, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-1055.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-1055.pdf.

Morrison:1978:CLA

[Mor78c] J. A. Morrison. A combinatorial lemma and its application to concentrating trees of discrete-time queues. The Bell System Technical Journal, 57(5):1645–1652, May/ June 1978. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1645.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1645.pdf.

Morrison:1979:IEE

[Mor79] J. A. Morrison. Integral equations for electromagnetic scattering by perfect conductors with two-dimensional geometry. The Bell System Technical Journal, 58(2):409–425, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-409.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-409.pdf.

Moschytz:1970:GAT

[Mos70] G. S. Moschytz. A general approach to twin-T design and its application to hybrid integrated linear active networks. The Bell System Technical Journal, 49(6):1105–1149, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-1105.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1105.pdf.

Mottl:1977:DPC

[Mot77] T. O. Mottl. Dual-polarized channel outages during multipath fading. The Bell System Technical Journal, 56 (5):675–701, May/June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-675.pdf.

May:1971:MCI

[MP71] A. S. May and M. J. Pagones.

Model for computation of interference to radio-relay systems from geostationary satellites. The Bell System Technical Journal, 50(1):81–102, January 1971. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol50/bstj50-1-
81.pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-81.pdf.

Marcuse:1975:MCO

[MP75] D. Marcuse and H. M. Presby.

Mode coupling in an optical fiber with core distortions. The Bell System Technical Journal, 54(1):3–15, January 1975. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol54/bstj54-1-
3.pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-3.pdf.

Myers:1978:AFS

[MP78] D. L. Myers and F. P. Partus. Atlanta Fiber System Experiment: Preform fabrication and fiber drawing by western electric product engineering control center. The Bell System Technical Journal, 57(6):1735–1744, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1735.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1735.pdf.

Moseley:1977:PNL

[MPS77] R. H. Moseley, S. B. Pirkau, and R. B. Swerdlow. Pseudorandom noise loading for noise power ratio and law of addition. The Bell System Technical Journal, 56(4):511–533, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-4-511.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-4-511.pdf.

McIlroy:1978:UTS

[MPT78] M. D. McIlroy, E. N. Pinson, and B. A. Tague. UNIX time-sharing system: Forward. The Bell System Technical Journal, 57(6):1899–1904, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol57/bstj57-6-1899.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1899.pdf.

McGonegal:1978:SVH

[MRM78] C. A. McGonegal, L. R. Rabiner, and B. J. McDermott. Speaker verification by human listeners over several speech transmission systems. The Bell System Technical Journal, 57(8):2887–2900, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2887.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2887.pdf.

McGonegal:1979:EST

[MRR79] C. A. McGonegal, A. E. Rosenberg, and L. R. Rabiner. The effects of several transmission systems on an automatic speaker verification system. The Bell System Technical Journal, 58(9):2071–2087, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-9-2071.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-9-2071.pdf.

Meyers:1977:NEM

[MRY77] M. N. Meyers, W. A. Routt, and K. W. Yoder. No. 4 ESS: Maintenance software. The Bell System Technical Journal, 56(7):1139–1167, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1139.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1139.pdf.

McKenna:1972:ADL

[MS72] J. McKenna and N. L. Schryer. On the accuracy of the depletion layer approximation for charge coupled devices. The Bell System Technical Journal, 51(7):1471–1485, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-7-1471.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-7-1471.pdf.

McKenna:1973:PCCb

[MS73a] J. McKenna and N. L. Schryer. The potential in a chargecoupled device with no mobile minority carriers. The Bell System Technical Journal, 52(10): 1765–1793, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1765.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1765.pdf.

McKenna:1973:PCCa

[MS73b] J. McKenna and N. L. Schryer.

The potential in a charge coupled device with no mobile minority carriers and zero plate separation. The Bell System Technical Journal, 52(5):669–696, May/ June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-5-669.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-5-669.pdf.

McKenna:1975:AFA

[MS75a] J. McKenna and N. L. Schryer.

Analysis of field-aided, chargecoupled device transfer. The Bell System Technical Journal, 54(4):667–685, April 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-4-667.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-4-667.pdf.

Mueller:1975:CEN

[MS75b] K. H. Mueller and D. A. Spaulding. Cyclic equalization–A new rapidly converging equalization technique for synchronous data communication. The Bell System Technical Journal, 54(2):369–406, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol54/bstj54-2-369.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-2-369.pdf.

Mazo:1976:ODC

[MS76] J. E. Mazo and J. Salz. On optical data communication via direct detection of light pulses. The Bell System Technical Journal, 55(3):347–369, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-3-347.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-3-347.pdf.

Maxemchuk:1979:AID

[MS79a] N. F. Maxemchuk and J. A. Stuller. An adaptive intraframe DPCM codec based upon nonstationary image model. The Bell System Technical Journal, 58(6):1395–1412, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1395.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1395.pdf.

Maxemchuk:1979:RTE

[MS79b] N. F. Maxemchuk and J. A. Stuller. Reduction of transmission error propagation in adaptively predicted, DPCM encoded pictures. The Bell System Technical Journal, 58(6):

1413–1423, July/August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1413.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1413.pdf.

Miller:1979:SSC

[MS79c] G. Miller and T. C. Spang. Spurious signal criteria for voiceband telephone equipment. The Bell System Technical Journal, 58(4):809–837, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-4-809.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-4-809.pdf.

MacWilliams:1972:MIN

[MSG72] Mrs. F. J. MacWilliams, N. J. A.

Sloane, and J. M. Goethals. The MacWilliams identities for nonlinear codes. The Bell System Technical Journal, 51(4):803–819, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-803.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-803.pdf.

McDermott:1978:POE

[MSG78] B. McDermott, C. Scagliola, and D. Goodman. Perceptual and objective evaluation of speech processed by adaptive differential PCM. The Bell System Technical Journal, 57(5):1597–1618, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1597.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1597.pdf.

McKenna:1974:DCT

[MSW74] J. McKenna, N. L. Schryer, and R. H. Walden. Design considerations for a two-phase, buriedchannel, charge-coupled device. The Bell System Technical Journal, 53(8):1581–1597, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-8-1581.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-8-1581.pdf.

Morrison:1977:PHF

[MSW77] J. A. Morrison, J. B. Seery, and L. O. Wilson. Propagation of high-frequency elastic surface waves along cylinders with various cross-sectional shapes. The Bell System Technical Journal, 56(1):77–114, January 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol56/bstj56-1-

77.pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-1-77.pdf.

Maione:1978:AFS

[MSW78] T. L. Maione, D. D. Sell, and D. H. Wolaver. Atlanta fiber system experiment: Practical 45-Mb/s regenerator for lightwave transmission. The Bell System Technical Journal, 57(6):1837–1856, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1837.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1837.pdf.

Maddox:1972:DCB

[MT72] C. L. Maddox and D. K.

Thovson. D2 channel bank: Per-channel equipment. The Bell System Technical Journal, 51(8):1659–1673, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1659.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1659.pdf.

Mueller:1973:NAO

[Mue73] K. H. Mueller. A new approach to optimum pulse shaping in sampled systems using time-domain filtering. The Bell System Technical Journal, 52(5):723–729, May/June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-5-723.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-5-723.pdf.

Mueller:1975:NFC

[Mue75] K. H. Mueller. A new, fastconverging mean-square algorithm for adaptive equalizers with partial-response signaling. The Bell System Technical Journal, 54(1):143–153, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-1-143.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-1-143.pdf.

Mueller:1979:CEC

[Mue79]  K. H. Mueller.  Combining

echo cancellation and decision feedback equalization. The Bell System Technical Journal, 58(2):491–500, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-2-491.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-491.pdf.

Muller:1978:CEN

[Mul78] E. E. Muller. COMSTAR experiment: Notes on the COM-STAR beacon experiment. The Bell System Technical Journal, 57(5):1369–1370, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1369.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1369.pdf.

Muska:1977:EOF

[Mus77] W. M. Muska. An experimental optical-fiber link for low-bit-rate applications. The Bell System Technical Journal, 56(1):65– 75, January 1977. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol56/bstj56-1-
65.pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-1-65.pdf.

Musa:1975:SDP

[MW75] J. D. Musa and F. N. Woomer, Jr. SAFEGUARD dataprocessing system: Software project management. The Bell System Technical Journal, 54 (10):S245–S259, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S245.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S245.pdf.

Mattes:1978:TNT

[MW78a] H. G. Mattes and B. A. Wright. Transaction network, telephones, and terminals: Transaction printer. The Bell System Technical Journal, 57(10):3517– 3529, December 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol57/bstj57-10-3517.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol57-1978/articles/
                                        bstj57-10-3517.pdf.

Miller:1978:CCI

[MW78b] P. R. Miller and R. E. Wallace. Common channel interoffice signaling: Signaling network. The Bell System Technical Journal, 57(2):263–282, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-2-263.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-2-263.pdf.

Miller:1979:GAL

[MWW^{+}79] R. C. Miller, R. H. Willens, H. A. Watson, L. A. D’Asaro, and M. Feldman. A galliumarsenide laser facsimile printer. The Bell System Technical Journal, 58(9):1909–1998, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-9-1909.

pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-9-1909.pdf.

Nagel:1971:SAI

[Nag71] C. M. Nagel, Jr. Some analytical investigations of phase contrast imaging. The Bell System Technical Journal, 50(6): 1943–1967, July/August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1943.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1943.pdf.

Neal:1971:CCS

[Nea71a] Scotty Neal. Combining correlated streams of nonrandom traffic. The Bell System Technical Journal, 50(6):2015–2037, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-2015.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-2015.pdf.

Neal:1971:SDN

[Nea71b] Scotty Neal. Spectral density of a nonlinear function of a Gaussian process. The Bell System Technical Journal, 50(3):1025–1047, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-1025.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-1025.pdf.

Neal:1972:EGM

[Nea72] Scotty Neal. The equivalent group method for estimating the capacity of partial-access service systems which carry overflow traffic. The Bell System Technical Journal, 51(3):777–783, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-3-777.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-3-777.pdf.

Netravali:1973:NOA

[Net73] Arun Netravali. A note on optimal approximating manifolds of a function class. The Bell System Technical Journal, 52(7):1237–1242, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-7-1237.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-7-1237.pdf.

Netravali:1977:ODF

[Net77]  A. N. Netravali.  Optimum

digital filters for interpolative A/D converters. The Bell System Technical Journal, 56(9):1629–1641, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1629.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1629.pdf.

Neumann:1971:SSS

[Neu71]  Peter G. Neumann.  Self-

synchronizing sequential coding with low redundancy. The Bell System Technical Journal, 50(3):951–981, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-951.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-951.pdf.

Nichols:1975:SDP

[Nic75] B. C. Nichols. SAFEGUARD data-processing system: Structured programming and program production librarians. The Bell System Technical Journal, 54 (10):S211–S219, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S211.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S211.pdf.

Neal:1973:ERN

[NK73a] S. R. Neal and A. Kuczura.

Erratum: S. R. Neal and A. Kuczura, A Theory of Traffic-Measurement Errors for Loss Systems With Renewal Input, BSTJ 52(6) 967–990 (1973). The Bell System Technical Journal, 52(10):1915, December 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1915.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1915.pdf.  See

[NK73b].

Neal:1973:TTM

[NK73b] S. R. Neal and A. Kuczura. A theory of traffic-measurement errors for loss systems with renewal input. The Bell System Technical Journal, 52(6):967–990, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-967.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-967.pdf.

See erratum [NK73a].

Netravali:1976:OTC

[NMB76] A. N. Netravali, F. W. Mounts, and E. G. Bowen. Ordering techniques for coding of two-tone facsimile pictures. The Bell System Technical Journal, 55(10): 1539–1552, December 1976. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol55/bstj55-10-1539.pdf;
http://www.alcatel-lucent.

com/bstj/vol55-1976/articles/
bstj55-10-1539.pdf.

Netravali:1977:TCD

[NMB77] A. N. Netravali, F. W. Mounts, and J. D. Beyer. Techniques for coding dithered two-level pictures. The Bell System Technical Journal, 56(5):809–819, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-5-809.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-5-809.pdf.

Netravali:1979:AOT

[NMW79] A. N. Netravali, F. W. Mounts, and K. A. Walsh. Adaptation of ordering techniques for facsimile pictures with no single element runs. The Bell System Technical Journal, 58(4):857–865, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-4-857.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-4-857.pdf.

See errata [Ano79l].

Noll:1975:CSV

[Nol75a] P. Noll. A comparative study of various quantization schemes for speech encoding. The Bell System Technical Journal, 54(9):1597–1614, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-9-1597.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-9-1597.pdf.

Noll:1975:ECE

[Nol75b]  P. Noll.  Effects of chan-

nel errors on the signal-to-noise performance of speechencoding systems. The Bell System Technical Journal, 54 (9):1615–1636, November 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-9-1615.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-9-1615.pdf.

Noll:1978:PQS

[Nol78] P. Noll. On predictive quantizing schemes. The Bell System Technical Journal, 57(5): 1499–1532, May/June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1499.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1499.pdf.

Nagelberg:1978:UTS

[NP78] E. R. Nagelberg and M. A. Pilla. UNIX time-sharing system: RBCS/RCMAS — converting to the MERT operating system. The Bell System Technical Journal, 57(6):2275–2287, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2275.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2275.pdf.

Netravali:1977:SEA

[NPM77] A. N. Netravali, B. Presada, and F. W. Mounts. Some experiments in adaptive and predictive Hadamard transform coding of pictures. The Bell System Technical Journal, 56(8):1531–1547, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1531.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1531.pdf.

Netravali:1979:MCTa

[NR79] A. N. Netravali and J. D. Robbins. Motion-compensated television coding: Part 1. The Bell System Technical Journal, 58(3):631–670, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-631.pdf.

Nickell:1972:CRC

[NS72] R. E. Nickell and D. C. Stickler. Controlled response of a ceramic microphone. The Bell System Technical Journal, 51(2):543–562, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-2-543.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-543.pdf.

Netravali:1976:OQD

[NS76] A. N. Netravali and R. Saigal. Optimum quantizer design using a fixed-point algorithm. The Bell System Technical Journal, 55(9):1423–1435, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-9-1423.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-9-1423.pdf.

Netravali:1979:MCTb

[NS79] A. N. Netravali and J. A. Stuller. Motion-compensated transform coding. The Bell System Technical Journal, 58 (7):1703–1718, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-7-1703.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol58-1979/
articles/bstj58-7-1703.pdf.

Oberst:1970:PRP

[Obe70] James F. Oberst. Pull-in range of a phase-locked loop with a binary phase comparator. The Bell System Technical Journal, 49(9):2289–2302, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2289.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2289.pdf.

OSullivan:1970:LAB

[OBW70] T. D. O’Sullivan, R. V. Biagetti, and M. C. Weeks. Lead-acid battery: Electrochemical characterization of the Bell System battery: Field trials of the battery. The Bell System Technical Journal, 49(7):1335–1358, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1335.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1335.pdf.

Ogawa:1977:STM

[Oga77] K. Ogawa. Simplified theory of the multimode fiber coupler. The Bell System Technical Journal, 56(5):729–745, May/ June 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-5-729.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-5-729.pdf.

Ohm:1974:PMB

[Ohm74] E. A. Ohm. A proposed multiplebeam microwave antenna for Earth stations and satellites. The Bell System Technical Journal, 53(8):1657–1665, October 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-8-1657.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-8-1657.pdf.

Olsen:1971:SCD

[Ols71] R. G. Olsen. Statistical circuit design: The application of Monte Carlo techniques to the study of impairments in the waveguide transmission system. The Bell System Technical Journal, 50(4):1293–1310, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-4-1293.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-4-1293.pdf.

Olson:1975:SDP

[Ols75] J. W. Olson. SAFEGUARD data-processing system: Architecture of the central logic and control. The Bell System Technical Journal, 54(10):S41–S61, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-10-S41.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-10-S41.pdf.

ONeill:1971:SCD

[O’N71] L. A. O’Neill. Statistical circuit design: A case study of the use of computer aids in circuit design — pulse equalizers for the T2 digital transmission line. The Bell System Technical Journal, 50(4):1243–1262, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1243.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1243.pdf.

Ostermayer:1974:ORI

[OP74] F. W. Ostermayer, Jr. and D. A. Pinnow. Optimum refractive-index difference for graded-index fibers resulting from concentration-fluctuation scattering. The Bell System Technical Journal, 53(7): 1395–1402, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-7-1395.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-7-1395.pdf.

Osborne:1971:ROP

[Osb71] T. L. Osborne. Rain outage performance of tendem and path diversity 18-GHz short hop radio systems. The Bell System Technical Journal, 50(1):59–79, January 1971. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol50/bstj50-1-
                                        59.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-1-59.pdf.

Osborne:1977:GRA

[Osb77] T. L. Osborne. 11-GHz radio: Application of rain attenuation data to 11-GHz radio path engineering. The Bell System Technical Journal, 56(9):1605–1627, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-9-1605.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-9-1605.pdf.

Opferman:1971:CRSa

[OTW71a] D. C. Opferman and N. T. Tsao-Wu. On a class of rearrangeable switching networks — Part I: Control algorithm. The Bell System Technical Journal, 50(5):1579–1600, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1579.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1579.pdf.

Opferman:1971:CRSb

[OTW71b] D. C. Opferman and N. T.

Tsao-Wu. On a class of rearrangeable switching networks — Part II: Enumeration studies and fault diagnosis. The Bell System Technical Journal, 50(5):1601–1618, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1601.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1601.pdf.

Owen:1972:IMR

[Owe72] B. Owen. The identification of modal resonances in ferrite loaded waveguide Y-junctions and their adjustment for circulation. The Bell System Technical Journal, 51(3):595–627, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-3-595.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-3-595.pdf.

Park:1978:ICP

[Par78a] K. I. Park. Intelligible crosstalk performance of voice-frequency customer loops. The Bell System Technical Journal, 57 (8):3001–3029, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-3001.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-3001.pdf.

Parker:1978:AFL

[Par78b] J. C. Parker, Jr. Analytical foundation for low-frequency powertelephone interference. The Bell System Technical Journal, 57(5):1663–1697, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1663.pdf.

Pinnel:1973:MRE

[PB73] M. R. Pinnel and J. E. Bennett. The metallurgy of remendur: Effects of processing variations. The Bell System Technical Journal, 52(8):1325–1340, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1325.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1325.pdf.

Persky:1974:CCF

[PB74] G. Persky and D. J. Bartelink.

Controlled current filaments in PNIPN structures with application to magnetic field detection. The Bell System Technical Journal, 53(3):467–502, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-3-467.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-3-467.pdf.

Pinnel:1975:IGM

[PB75] M. R. Pinnel and J. E. Bennett.

Influences of glass-to-metal sealing on the structure and magnetic properties of an Fe/Co/V alloy. The Bell System Technical Journal, 54(6):997–1009, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-6-997.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-6-997.pdf.

Pease:1972:CVS

[Pea72]  R. F. W. Pease.  Con-

ditional vertical subsampling — A technique to assist in the coding of television signals. The Bell System Technical Journal, 51(4):787–802, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol51/bstj51-4-787.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-4-787.pdf.

Pehlert:1970:ABT

[Peh70]  W. K. Pehlert, Jr.  Anal-

ysis of a burst-trapping error correction procedure. The Bell System Technical Journal, 49(4):493–519, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-4-493.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-4-493.pdf.

Pekarich:1978:UTS

[Pek78] S. P. Pekarich. UNIX timesharing system: No. 4 ESS diagnostic environment. The Bell System Technical Journal, 57(6):2265–2274, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2265.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-2265.pdf.

Penzias:1970:BJB

[Pen70]  A. A. Penzias.  B.S.T.J.

briefs: First result from 15.3-GHz Earth–Space propagation study. The Bell System Technical Journal, 49(6):1242–1245, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1242.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-1242.pdf.

Personick:1971:BJB

[Per71a]  S. D. Personick.  B.S.T.J.

briefs: An image band interpretation of optical heterodyne noise. The Bell System Technical Journal, 50(1):213–216, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-213.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-213.pdf.

Personick:1971:NRA

[Per71b]  S. D. Personick.  New re-

sults on avalanche multiplication statistics with applications to optical detection. The Bell System Technical Journal, 50(1):167–189, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-167.pdf.

Personick:1971:SGC

[Per71c] S. D. Personick. Statistics of a general class of avalanche detectors with applications to optical communication. The Bell System Technical Journal, 50 (10):3075–3095, December 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3075.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3075.pdf.

Personick:1971:TDD

[Per71d] S. D. Personick. Time dispersion in dielectric waveguides. The Bell System Technical Journal, 50(3):843–859, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-843.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-843.pdf.

Persky:1972:HCE

[Per72] G. Persky. Hot carrier effects in the integral chargecontrol model for bipolar transistors. The Bell System Technical Journal, 51(2):455–460, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-2-455.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-2-455.pdf.

Personick:1973:AQA

[Per73a] S. D. Personick. Applications for quantum amplifiers in simple digital optical communication systems. The Bell System Technical Journal, 52(1):117–133, January 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-117.pdf.

Personick:1973:BLE

[Per73b] S. D. Personick. Baseband linearity and equalization in fiber optic digital communication systems. The Bell System Technical Journal, 52(7):1175–1194, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-7-1175.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1175.pdf.

Personick:1973:RDDa

[Per73c] S. D. Personick. Receiver design for digital fiber optic communication systems, I. The Bell System Technical Journal, 52(6):843–874, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-843.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-843.pdf.

Personick:1973:RDDb

[Per73d] S. D. Personick. Receiver design for digital fiber optic communication systems, II. The Bell System Technical Journal, 52(6):875–886, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-6-875.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-6-875.pdf.

Personick:1974:OTM

[Per74] S. D. Personick. Optimal tradeoff of mode-mixing optical filtering and index difference in digital fiber optic communication systems. The Bell System Technical Journal, 53(5):785–800, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-785.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-785.pdf.

Personick:1975:TDT

[Per75] S. D. Personick. Two derivations of the time-dependent coupledpower equations. The Bell System Technical Journal, 54(1):47– 52, January 1975. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol54/bstj54-1-
                                        47.pdf;   http://www.alcatel-

lucent.com/bstj/vol54-1975/
articles/bstj54-1-47.pdf.

Personick:1976:CEN

[Per76] S. D. Personick. Comparison of equalizing and nonequalizing repeaters for optical fiber systems. The Bell System Technical Journal, 55(7):957–971, September 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-7-957.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-7-957.pdf.

Personick:1977:PTO

[Per77]  S. D. Personick.  Photon

tube — an optical-fiber timedomain reflectometer. The Bell System Technical Journal, 56(3):355–366, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-3-355.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-3-355.pdf.

Pferd:1979:ESF

[Pfe79] W. Pferd. The evolution and special features of Bell System telephone equipment buildings. The Bell System Technical Journal, 58(2):427–466, February 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol58/bstj58-2-427.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-2-427.pdf.

Phillips:1975:SDP

[Phi75] A. K. Phillips. SAFEGUARD data-processing system: Debugging a real-time multiprocessor system. The Bell System Technical Journal, 54(10):S133– S145, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-10-S133.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol54-1975/articles/
                                        bstj54-10-S133.pdf.

Pierce:1972:NBS

[Pie72]  J. R. Pierce.  Network for

block switching of data. The Bell System Technical Journal, 51(6):1133–1145, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1133.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1133.pdf.

Pirz:1979:DWC

[Pir79] F. Pirz. Design of a wideband, constant beamwidth, array microphone for use in the near field. The Bell System Technical Journal, 58(8):1839–1850, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1839.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-8-1839.pdf.

Purvis:1970:RCS

[PK70] M. B. Purvis and R. W. Kordos. Reed-contact switch series for the I. F. band. The Bell System Technical Journal, 49(2):229–254, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-229.pdf.

Pease:1971:EST

[PL71] R. F. W. Pease and J. O. Limb.

Exchange of spatial and temporal resolution in television coding. The Bell System Technical Journal, 50(1):191–200, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-191.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-191.pdf.

Presby:1979:RAI

[PMAB79] H. M. Presby, D. Marcuse, H. W. Astle, and L. M. Boggs. Rapid automatic index profiling of whole-fiber samples: Part 2. The Bell System Technical Journal, 58(4):883–902, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-4-883.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-4-883.pdf.

Prasada:1978:LRT

[PMN78] B. Prasada, F. W. Mounts, and A. N. Netravali. Level reassignment: A technique for bitrate reduction. The Bell System Technical Journal, 57(1):61–73, January 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol57/bstj57-1-
                                        61.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-61.pdf.

Poole:1970:DPP

[Poo70] K. M. Poole. Device photolithography: The primary pattern generator: Introduction. The Bell System Technical Journal, 49(9):2031, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2031.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2031.pdf.

Puerling:1973:IMS

[PR73] B. W. Puerling and J. T. Roberto. Information management system: The natural dialogue system. The Bell System Technical Journal, 52(10): 1725–1741, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1725.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1725.pdf.

Prabhu:1974:SDP

[PR74] V. K. Prabhu and H. E.

Rowe. Spectra of digital phase modulation by matrix methods. The Bell System Technical Journal, 53(5):899–935, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-5-899.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-5-899.pdf.

Prabhu:1970:PDM

[Pra70] V. K. Prabhu. On the performance of digital modulation systems that expand bandwidth. The Bell System Technical Journal, 49(6):1033–1057, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-6-1033.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-6-1033.pdf.

Prabhu:1971:SCE

[Pra71] V. K. Prabhu. Some considerations of error bounds in digital systems. The Bell System Technical Journal, 50(10): 3127–3151, December 1971. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3127.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3127.pdf.

Prabhu:1976:SOD

[Pra76]  V. K. Prabhu.  Spectral

occupancy of digital anglemodulation signals. The Bell System Technical Journal, 55(4):429–453, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-4-429.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-4-429.pdf.

Presby:1976:GUP

[Pre76] H. M. Presby. Geometrical uniformity of plastic coatings on optical fibers. The Bell System Technical Journal, 55(10): 1525–1537, December 1976. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol55/bstj55-10-1525.pdf;
                                        http://www.alcatel-lucent.

com/bstj/vol55-1976/articles/
bstj55-10-1525.pdf.

Poulsen:1970:DPR

[PS70] M. E. Poulsen and J. W.

Stafford. Device photolithography: Reduction cameras: Mechanical design of the 3.5X and 1.4X reduction cameras. The Bell System Technical Journal, 49(9):2129–2143, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2129.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2129.pdf.

Presby:1975:MSG

[PSMO75] H. M. Presby, R. D. Standley, J. B. MacChesney, and P. B. O’Connor. Material structure of germanium-doped optical fibers and preforms. The Bell System Technical Journal, 54(10): 1681–1692, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-10-1681.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-1681.pdf.

Potter:1970:NEA

[PSW70] J. L. Potter, Mrs. F. B.

Strebendt, and J. R. Williams. No. 1 ESS ADF: Magnetic tape subsystem. The Bell System Technical Journal, 49(10):2915– 2940, December 1970. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol49/bstj49-10-2915.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol49-1970/articles/
                                        bstj49-10-2915.pdf.

Rabiner:1972:DFI

[Rab72] L. R. Rabiner. The design of finite impulse response digital filters using linear programming techniques. The Bell System Technical Journal, 51(6):1177–1198, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1177.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1177.pdf.

Rainal:1973:CDA

[Rai73] A. J. Rainal. Computing distortion in analog FM communication systems. The Bell System Technical Journal, 52(5):627–648, May/June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-5-627.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-5-627.pdf.

Rainal:1976:TRC

[Rai76]  A. J. Rainal.  Temperature

rise at a constriction in a current-carrying printed conductor. The Bell System Technical Journal, 55(2):233–269, February 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-2-233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-2-233.pdf.

Rainal:1979:TPV

[Rai79] A. J. Rainal. Transmission properties of various styles of printed wiring boards. The Bell System Technical Journal, 58(5):995–1025, May/June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-5-995.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-5-995.pdf.

Ramaswamy:1974:SLF

[Ram74] V. Ramaswamy. Strip-loaded film waveguide. The Bell System Technical Journal, 53(4):697– 704, April 1974. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol53/bstj53-4-
697.pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-697.pdf.

Rao:1978:LPE

[Rao78] T. N. Rao. Loop plant electronics: Analog loop carrier systems. The Bell System Technical Journal, 57(4):1109–1128, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-4-1109.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-4-1109.pdf.

Rawson:1970:DPR

[Raw70] Eric G. Rawson. Device photolithography: Reduction cameras: Optical design and adjustment. The Bell System Technical Journal, 49(9): 2117–2128, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2117.pdf.

Ruthroff:1970:LPM

[RB70] C. L. Ruthroff and W. F. Bodtmann. A linear phase modulator for large baseband bandwidths. The Bell System Technical Journal, 49(8):1893–1903, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1893.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1893.pdf.

Ruthroff:1974:ACC

[RB74] C. L. Ruthroff and W. F.

Bodtmann. Adaptive coding for coherent detection of digital phase modulation. The Bell System Technical Journal, 53(3):449–466, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-3-449.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-3-449.pdf.

Rife:1976:MTP

[RB76] D. C. Rife and R. R. Boorstyn.

Multiple tone parameter estimation from discrete-time observations. The Bell System Technical Journal, 55(9): 1389–1410, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-9-1389.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-9-1389.pdf.

Rutledge:1977:WMW

[RBH77] D. R. Rutledge, H. A. Baxter, and W. M. Hauser. WT4 millimeter waveguide system: Waveguide installation. The Bell System Technical Journal, 56 (10):1923–1951, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/

                                        Vol56/bstj56-10-1923.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-1923.pdf.

Rabiner:1976:DAP

[RC76] L. R. Rabiner and R. E. Crochiere. On the design of all-pass signals with peak amplitude constraints. The Bell System Technical Journal, 55(4):395–407, April 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-4-395.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-4-395.pdf.

Runge:1978:AFS

[RC78] P. K. Runge and S. S. Cheng. Atlanta Fiber System Experiment: Demountable single-fiber optic connectors and their measurement on location. The Bell System Technical Journal, 57(6):1771–1790, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1771.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1771.pdf.

Reis:1972:SDS

[Rei72] G. C. Reis. Stability of distributed systems with feedback via Michailov’s criterion. The Bell System Technical Journal, 51(4):903–919, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-903.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-903.pdf.

Rice:1970:RPV

[Ric70] S. O. Rice. Response of periodically varying systems to shot noise — application to switched RC circuits. The Bell System Technical Journal, 49(9):2221–2248, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2221.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2221.pdf.

Rice:1973:DPB

[Ric73a] S. O. Rice. Distortion produced by band limitation of an FM wave. The Bell System Technical Journal, 52(5):605–626, May/ June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-5-605.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-5-605.pdf.

Rice:1973:DRE

[Ric73b] S. O. Rice. Distribution of ∑ a_{n}/n, a_{n} randomly equal to ±1. The Bell System Technical Journal, 52(7):1097– 1103, September 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-7-1097.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-7-1097.pdf.

Rice:1973:EEI

[Ric73c] S. O. Rice. Efficient evaluation of integrals of analytic functions by the trapezoidal rule. The Bell System Technical Journal, 52(5):707–722, May/ June 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-5-707.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-5-707.pdf.

Rice:1973:VSM

[Ric73d]  S. O. Rice.  Volterra sys-

tems with more than one input port-distortion in a frequency converter. The Bell System Technical Journal, 52 (8):1255–1270, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1255.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1255.pdf.

Rice:1975:NEI

[Ric75] S. O. Rice. Numerical evaluation of integrals with infinite limits and oscillating integrands. The Bell System Technical Journal, 54(1):155–164, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-1-155.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-1-155.pdf.

Rifenberg:1975:SDP

[Rif75]  C. J. Rifenberg.  SAFE-

GUARD data-processing system: The dictionary approach to digital maintenance. The Bell System Technical Journal, 54(10):S73–S85, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-10-S73.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-10-S73.pdf.

Rife:1979:PSF

[Rif79]  D. C. Rife.  Periodic se-

quences that facilitate data set spectrum measurements. The Bell System Technical Journal, 58(9):2089–2096, November 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-9-2089.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-9-2089.pdf.

Ritchie:1978:UTSb

[Rit78] D. M. Ritchie. UNIX timesharing system: A retrospective. The Bell System Technical Journal, 57(6):1947–1969, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1947.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1947.pdf.

Rabiner:1972:PEE

[RJ72] L. R. Rabiner and J. A. Johnson. Perceptual evaluation of the effects of dither on low bit rate PCM systems. The Bell System Technical Journal, 51(7):1487–1494, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-7-1487.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-7-1487.pdf.

Ritchie:1978:UTSc

[RJLK78] D. M. Ritchie, S. C. Johnson, M. E. Lesk, and B. W. Kernighan. UNIX time-sharing system: The C programming language. The Bell System Technical Journal, 57(6): 1991–2019, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1991.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1991.pdf.

Rabiner:1974:SCB

[RKHD74] L. R. Rabiner, J. F. Kaiser, O. Herrmann, and M. T. Dolan. Some comparisons between FIR and IIR digital fibers. The Bell System Technical Journal, 53(2):305–331, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-305.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-305.pdf.

Rubin:1972:COL

[RM72] H. Rubin and H. E. Meadows.

Controllability and observability in linear time-variable networks with arbitrary symmetry groups. The Bell System Technical Journal, 51(2):507–542, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-507.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-2-507.pdf.

Ritchie:1978:CCI

[RM78] A. E. Ritchie and J. Z. Menard.

Common channel interoffice signaling: An overview. The Bell System Technical Journal, 57(2):221–224, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-2-221.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-2-221.pdf.

Robertson:1971:FAA

[Rob71]  G. H. Robertson.  A fast

amplitude approximation for quadrature pairs. The Bell System Technical Journal, 50 (8):2849–2852, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2849.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2849.pdf.

Rodriguez:1978:TNT

[Rod78] E. J. Rodriguez. Transaction network, telephones, and terminals: Transaction network operational programs. The Bell System Technical Journal, 57 (10):3371–3407, December 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-10-3371.pdf;
http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-10-3371.pdf.

Rosencwaig:1972:BJB

[Ros72] A. Rosencwaig. B.S.T.J. briefs:

The effect of a second magnetic layer on hard bubbles. The Bell System Technical Journal, 51(6):1440–1444, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1440.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1440.pdf.

Rosenberg:1976:EAS

[Ros76]  A. E. Rosenberg.  Evalua-

tion of an automatic speakerverification system over telephone lines. The Bell System Technical Journal, 55(6): 723–744, July/August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-6-723.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-6-723.pdf.

Rovegno:1978:UTS

[Rov78] H. D. Rovegno. UNIX timesharing system: A support environment for MAC-8 systems. The Bell System Technical Journal, 57(6):2251–2263, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2251.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-2251.pdf.

Rowe:1975:PSD

[RP75] H. E. Rowe and V. K. Prabhu. Power spectrum of a digital, frequency-modulation signal. The Bell System Technical Journal, 54(6):1095–1125, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-6-1095.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-6-1095.pdf.

Rader:1970:FMG

[RRS70] C. M. Rader, L. R. Rabiner, and R. W. Schafer. A fast method of generating digital random numbers. The Bell System Technical Journal, 49(9):2303–2310, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-9-2303.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-9-2303.pdf.

Rabiner:1974:BMF

[RS74] L. R. Rabiner and R. W. Schafer. On the behavior of minimax FIR digital Hilbert transformers. The Bell System Technical Journal, 53(2):363–390, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-363.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-363.pdf.

Rabiner:1975:ADE

[RS75] L. R. Rabiner and M. R. Sambur.

An algorithm for determining the endpoints of isolated utterances. The Bell System Technical Journal, 54(2):297–315, February 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-2-297.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-2-297.pdf.

Ramaswamy:1976:POC

[RS76] V. Ramaswamy and R. D.

Standley. A phased, optical, coupler-pair switch. The Bell System Technical Journal, 55(6):767–775, July/August 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-6-767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-6-767.pdf.

Ramaswamy:1978:RPP

[RS78] V. Ramaswamy and R. D.

Standley. Radiation patterns from parallel, optical waveguide directional couplers — parameter measurements. The Bell System Technical Journal, 57(7):2685–2693, September 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-7-2685.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-7-2685.pdf.

Rosenberg:1979:ARS

[RS79] A. E. Rosenberg and C. E. Schmidt. Automatic recognition of spoken spelled names for obtaining directory listings. The Bell System Technical Journal, 58(8):1797–1823, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1797.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-8-1797.pdf.

Rabiner:1977:ESA

[RSA77] L. R. Rabiner, C. E. Schmidt, and B. S. Atal. Evaluation of a statistical approach to voiced-unvoiced-silence analysis for telephone-quality speech. The Bell System Technical Journal, 56(3):455–482, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-3-455.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-3-455.pdf.

Rabiner:1977:TCW

[RSCG77] L. R. Rabiner, M. R. Sambur, R. E. Crochiere, and D. J. Goodman. Tandem connections of wideband and narrowband speech communication systems: Part 2 — wideband-to-narrowband link. The Bell System Technical Journal, 56(9):1723–1741, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1723.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1723.pdf.

Rabiner:1971:CSS

[RSF71] L. R. Rabiner, R. W. Shafer, and J. L. Flanagan. Computer synthesis of speech by concatenation of formant-coded words. The Bell System Technical Journal, 50(5):1541–1558, May/ June 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-5-1541.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-5-1541.pdf.

Rosenthal:1974:ALB

[RSR74] L. H. Rosenthal, R. W. Schafer, and L. R. Rabiner. An algorithm for locating the beginning and end of an utterance using ADPCM coded speech. The Bell System Technical Journal, 53(6):1127–1135, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-6-1127.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-6-1127.pdf.

Riddell:1979:TSP

[RSS79] G. Riddell, C. R. Swanson, and R. T. Steinbrenner. Traffic service position system no. 1: Operator training facilities. The Bell System Technical Journal, 58(6):1347–1357, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1347.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1347.pdf.

Ramaswamy:1978:PES

[RSSF78] V. Ramaswamy, R. D. Standley, D. Sze, and W. G. French. Polarization effects in short length, single mode fibers. The Bell System Technical Journal, 57(3):635–651, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-3-635.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-3-635.pdf.

Rosenberger:1971:PAE

[RT71] J. R. Rosenberger and E. J.

Thomas. Performance of an adaptive echo canceller operating in a noisy, linear, timeinvariant environment. The Bell System Technical Journal, 50(3):785–813, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-785.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-785.pdf.

Ritchie:1977:NES

[RT77] A. E. Ritchie and L. S.

Tuomenoksa. No. 4 ESS: System objectives and organization. The Bell System Technical Journal, 56(7):1017–1027, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-7-1017.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-7-1017.pdf.

Ritchie:1978:UTSa

[RT78] D. M. Ritchie and K. Thompson. The UNIX time-sharing system. The Bell System Technical Journal, 57(6):1905–1929, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1905.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1905.pdf.

Rummler:1979:NSF

[Rum79] W. D. Rummler. A new selective fading model: Application to propagation data. The Bell System Technical Journal, 58(5):1037–1071, May/ June 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-5-1037.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-5-1037.pdf.

Rustako:1978:ESP

[Rus78] A. J. Rustako, Jr. An Earth– Space propagation measurement at Crawford Hill using the 12-GHz CTS satellite beacon. The Bell System Technical Journal, 57(5):1431–1448, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1431.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1431.pdf.

Ruthroff:1970:RAR

[Rut70]  C. L. Ruthroff.  Rain at-

tenuation and radio path design. The Bell System Technical Journal, 49(1):121–135, January 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-1-121.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol49-1970/
articles/bstj49-1-121.pdf.

Ruthroff:1971:MPF

[Rut71] C. L. Ruthroff. Multiple-path fading on line-of-sight microwave radio systems as a function of path length and frequency. The Bell System Technical Journal, 50(7):2375–2398, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2375.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2375.pdf.

Rife:1970:UDF

[RV70] D. C. Rife and G. A. Vincent. Use of the Discrete Fourier Transform in the measurement of frequencies and levels of tones. The Bell System Technical Journal, 49(2):197–228, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-197.pdf.

Reinhart:1974:CND

[RW74] F. K. Reinhart and L. O.

Wilson. Coupling of nearly degenerate modes in parallel asymmetric dielectric waveguides. The Bell System Technical Journal, 53(4):717–739, April 1974. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-4-717.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-4-717.pdf.

Rabiner:1979:ACT

[RW79] L. R. Rabiner and J. G. Wilpon. Application of clustering techniques to speaker-trained isolated word recognition. The Bell System Technical Journal, 58 (10):2217–2233, December 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol58/bstj58-10-2217.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol58-1979/articles/
                                        bstj58-10-2217.pdf.

Reudink:1977:SSB

[RY77] D. O. Reudink and Y. S. Yeh. A scanning spot-beam satellite system. The Bell System Technical Journal, 56(8):1549–1560, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-8-1549.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-8-1549.pdf.

Salz:1973:OMS

[Sal73] J. Salz. Optimum mean-square decision feedback equalization. The Bell System Technical Journal, 52(8):1341–1373, October 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-8-1341.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-8-1341.pdf.

Salazar:1974:DTR

[Sal74] Andres C. Salazar. Design of transmitter and receiver filters for decision feedback equalization. The Bell System Technical Journal, 53(3):503–523, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-3-503.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-3-503.pdf.

Saleh:1975:PIM

[Sal75] A. A. M. Saleh. Polarizationindependent, multilayer dielectrics at oblique incidence. The Bell System Technical Journal, 54(6):1027–1049, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-6-1027.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-6-1027.pdf.

Sambur:1975:ELP

[Sam75] M. R. Sambur. An efficient linear-prediction vocoder. The Bell System Technical Journal, 54(10):1693–1723, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-10-1693.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol54-1975/articles/
                                        bstj54-10-1693.pdf.

Sandberg:1970:TAN

[San70a] I. W. Sandberg. Theorems on the analysis of nonlinear transistor networks. The Bell System Technical Journal, 49(1):95–114, January 1970. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol49/bstj49-1-
                                        95.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-1-95.pdf.

Sandberg:1970:TCT

[San70b] I. W. Sandberg. Theorems on the computation of the transient response of nonlinear networks containing transistors and diodes. The Bell System Technical Journal, 49(8):1739–1776, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1739.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1739.pdf.

Sandberg:1978:NCO

[San78a] I. W. Sandberg. A note concerning optical-waveguide modulation transfer functions. The Bell System Technical Journal, 57(8):3047–3056, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-3047.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-3047.pdf.

Sandberg:1978:SIS

[San78b] I. W. Sandberg. On the stability of interconnected systems. The Bell System Technical Journal, 57(8):3031–3046, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-3031.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-3031.pdf.

Sandberg:1979:CIP

[San79]  I. W. Sandberg.  A char-

acterization of the invariance of positivity for functional differential equations. The Bell System Technical Journal, 58(8):1885–1894, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1885.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-8-1885.pdf.

Smith:1975:MPT

[SBGC75] P. W. Smith, D. L. Bisbee, D. Gloge, and E. L. Chinnock. A molded-plastic technique for connecting and splicing optical-fiber tapes and cables. The Bell System Technical Journal, 54(6):971–984, July/ August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-6-971.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-6-971.pdf.

Siller:1978:DEO

[SBR78a] C. A. Siller, Jr., P. E. Butzien, and J. E. Richard. Design and experimental optimization of a canister antenna for 18-GHz operation. The Bell System Technical Journal, 57(3):603–633, March 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-3-603.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-3-603.pdf.

Smith:1978:AFS

[SBR78b] R. G. Smith, C. A. Brackett, and H. W. Reinbold. Atlanta Fiber System Experiment: Optical detector package. The Bell System Technical Journal, 57(6):1809–1822, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1809.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1809.pdf.

Scagliola:1979:EAS

[Sca79] C. Scagliola. Evaluation of adaptive speech coders under noisy channel conditions. The Bell System Technical Journal, 58(6):1369–1394, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1369.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1369.pdf.

Shumate:1978:AFS

[SCD78] P. W. Shumate, Jr., F. S.

Chen, and P. W. Dorman. Atlanta Fiber System Experiment: GaAlAs laser transmitter for lightwave transmission systems. The Bell System Technical Journal, 57(6):1823–1836, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1823.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1823.pdf.

Scharfetter:1970:PFC

[Sch70]  D. L. Scharfetter.  Power-

frequency characteristics of the TRAPATT diode mode of high efficiency power generation in germanium and silicon avalanche diodes. The Bell System Technical Journal, 49(5):799–826, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-799.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-799.pdf.

Schumer:1971:CAM

[Sch71] M. A. Schumer. A combinatorial analysis of the main distributing frame: Spare requirements for conversion to preferential assignment from random assignment. The Bell System Technical Journal, 50(7):2465–2484, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-7-2465.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-7-2465.pdf.

Schlosser:1972:DDW

[Sch72] W. O. Schlosser. Delay distortion in weakly guiding optical fibers due to elliptic deformation of the boundary. The Bell System Technical Journal, 51(2):487–492, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-2-487.
pdf;   http://www.alcatel-

lucent.com/bstj/vol51-1972/
articles/bstj51-2-487.pdf.

Schlosser:1973:GIM

[Sch73] W. O. Schlosser. Gain-induced modes in planar structures. The Bell System Technical Journal, 52(6):887–905, July/August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-887.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-887.pdf.

Schroeder:1974:SPO

[Sch74] W. E. Schroeder. Spurious parametric oscillations in IM-PATT diode circuits. The Bell System Technical Journal, 53(7):1187–1210, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-7-1187.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-7-1187.pdf.

Schroeder:1978:ASS

[Sch78]  C. M. Schroeder.  Accu-

rate silicon spacer chips for an optical-fiber cable connector. The Bell System Technical Journal, 57(1):91–97, January 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol57/bstj57-1-

                                        91.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-1-91.pdf.

Simpkins:1977:OCD

[SDMS77] P. G. Simpkins, T. D. Dudderar, J. McKenna, and J. B. Seery. On the occurrence of caustics in the drawdown zone of silica fibers. The Bell System Technical Journal, 56(4):535–560, April 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-4-535.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-4-535.pdf.

Seery:1978:CPA

[SDS78] J. B. Seery, T. D. Dudderar, and P. G. Simpkins. Caustic patterns associated with melt zones in solidified glass samples — Part II: Asymmetric cases. The Bell System Technical Journal, 57(9):3227–3251, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-9-3227.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-9-3227.pdf.

Seidel:1971:MFF

[Sei71] H. Seidel. A microwave feedforward experiment. The Bell System Technical Journal, 50(9):2879–2916, November 1971. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-9-2879.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-9-2879.pdf.

Semplak:1971:DFM

[Sem71]  R. A. Semplak.  Dual fre-

quency measurements of rain induced microwave attenuation on a 2.6-kilometer propagation path. The Bell System Technical Journal, 50(8):2599–2606, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2599.pdf.

Semplak:1973:BJB

[Sem73] R. A. Semplak. B.S.T.J. briefs:

The effect of rain on circular polarization at 18 GHz. The Bell System Technical Journal, 52(6):1029–1031, July/ August 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-6-1029.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-6-1029.pdf.

Semplak:1974:SMD

[Sem74] R. A. Semplak. Simultaneous measurements of depolarization by rain using linear and circular polarizations at 18 GHz. The Bell System Technical Journal, 53(2):400–404, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-400.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-400.pdf.

Semplak:1977:GMM

[Sem77]  R. A. Semplak.  100-GHz

measurements on a multiplebeam offset antenna. The Bell System Technical Journal, 56(3):385–398, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-3-385.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-3-385.pdf.

Semplak:1979:BJB

[Sem79] R. A. Semplak. B.S.T.J. briefs: A 30-GHz scale model, pyramidal, horn-reflector antenna. The Bell System Technical Journal, 58(6):1551–1556, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1551.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1551.pdf.

Sen:1970:MCS

[Sen70]  Tapas K. Sen.  Masking

of crosstalk by speech and noise. The Bell System Technical Journal, 49(4):561–584, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-4-561.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-4-561.pdf.

Sequin:1972:BJB

[Seq72] C. H. Sequin. B.S.T.J. briefs:

Blooming suppression in charge coupled area imaging devices. The Bell System Technical Journal, 51(8):1923–1926, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1923.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1923.pdf.

Sequin:1974:EIL

[Seq74] C. H. Sequin. Experimental investigation of a linear 500element 3-phase charge-coupled device. The Bell System Technical Journal, 53(4):581–610, April 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-4-581.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-4-581.pdf.

Setzer:1970:CTT

[Set70] David E. Setzer. Computed transmission through rain at microwave and visible frequencies. The Bell System Technical Journal, 49(8):1873–1892, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1873.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1873.pdf.

Setzer:1971:ASD

[Set71] David E. Setzer. Anisotropic scattering due to rain at radio-relay frequencies. The Bell System Technical Journal, 50(3):861–868, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-861.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-861.pdf.

Steele:1977:DSS

[SG77] R. Steele and D. J. Goodman. Detection and selective smoothing of transmission errors in linear PCM. The Bell System Technical Journal, 56(3):399–409, March 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-3-399.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol56-1977/
articles/bstj56-3-399.pdf.

Sondhi:1974:FQB

[SGM74] M. M. Sondhi, B. Gopinath, and Debasis Mitra. Formulas on queues in burst processes — II. The Bell System Technical Journal, 53(3):425–448, March 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-3-425.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-3-425.pdf.

Steele:1979:DFU

[SGM79] R. Steele, D. J. Goodman, and C. A. McGonegal. DPCM with forced updating and partial correction of transmission errors. The Bell System Technical Journal, 58(3):721–728, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-721.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-721.pdf.

Standley:1974:ERS

[SH74] R. D. Standley and W. S.

Holden. Experimental results on a single-material optical fiber. The Bell System Technical Journal, 53(5):779–784, May/ June 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-5-779.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-5-779.pdf.

Staehler:1979:TSP

[SH79] R. E. Staehler and W. S. Hayward, Jr. Traffic service position system no. 1: Recent developments, an overview. The Bell System Technical Journal, 58(6):1109–1118, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-6-1109.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-6-1109.pdf.

Shanholt:1970:ESL

[Sha70] Gerald A. Shanholt. Eventual stability for Lipschitz functional differential systems. The Bell System Technical Journal, 49(8):1905–1909, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-8-1905.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-8-1905.pdf.

Shah:1972:LPM

[Sha72] S. R. Shah. A linear phase modulator for a short-hop microwave radio system. The Bell System Technical Journal, 51(8):1837–1848, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1837.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1837.pdf.

Shah:1975:EEN

[Sha75] J. Shah. Effects of environmental nuclear radiation on optical fibers. The Bell System Technical Journal, 54(7): 1207–1213, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1207.pdf.

Shnidman:1970:OPE

[Shn70] David A. Shnidman. Optimum phase equalization of filters for digital signals. The Bell System Technical Journal, 49(7):1531–1554, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1531.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1531.pdf.

Sambur:1976:SEM

[SJ76] M. R. Sambur and N. S. Jayant.

Speech encryption by manipulations of LPC parameters.

The Bell System Technical Journal, 55(9):1373–1388, November 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-9-1373.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-9-1373.pdf.

Steele:1979:SBP

[SJS79] R. Steele, N. S. Jayant, and C. E. Schmidt. Statistical block protection coding for DPCM-encoded speech. The Bell System Technical Journal, 58(7):1647–1657, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-7-1647.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-7-1647.pdf.

Stuck:1974:SAT

[SK74] B. W. Stuck and B. Kleiner. A statistical analysis of telephone noise. The Bell System Technical Journal, 53(7): 1263–1320, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-7-1263.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-7-1263.pdf.

Snow:1975:DDS

[SK75] N. E. Snow and N. Knapp, Jr. Digital data system: System overview. The Bell System Technical Journal, 54 (5):811–832, May/June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-5-811.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-5-811.pdf.

Sandberg:1978:PMT

[SKCM78] I. W. Sandberg, I. P. Kaminow, L. G. Cohen, and W. L. Mammel. On the phase of the modulation transfer function of a multimode optical-fiber guide. The Bell System Technical Journal, 57(1):99–109, January 1978. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol57/bstj57-1-
99.pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-1-99.pdf.

Slepian:1972:MDS

[Sle72a]  D. Slepian.  On maxen-

tropic discrete stationary processes. The Bell System Technical Journal, 51(3):629–653, March 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-3-629.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-3-629.pdf.

Slepian:1972:DM

[Sle72b] David Slepian. On delta modulation. The Bell System Technical Journal, 51(10):2101– 2137, December 1972. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol51/bstj51-10-2101.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol51-1972/articles/
                                        bstj51-10-2101.pdf.

Slepian:1978:PSW

[Sle78] D. Slepian. Prolate spheroidal wave functions, Fourier analysis, and uncertainty — V: The discrete case. The Bell System Technical Journal, 57 (5):1371–1430, May/June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1371.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1371.pdf.

Sloane:1972:FPT

[Slo72] N. J. A. Sloane. On finding the paths through a network. The Bell System Technical Journal, 51(2):371–390, February 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-2-371.

pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-2-371.pdf.

Sloane:1976:SDE

[Slo76] N. J. A. Sloane. A simple description of an error-correcting code for high-density magnetic tape. The Bell System Technical Journal, 55(2):157–165, February 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-2-157.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-2-157.pdf.

Smithgall:1979:AOT

[Smi79]  D. H. Smithgall.  Applica-

tion of optimization theory to the control of the optical fiber drawing process. The Bell System Technical Journal, 58(6):1425–1435, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1425.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1425.pdf.

Stuller:1979:TDM

[SN79] J. A. Stuller and A. N.

Netravali. Transform domain motion estimation. The Bell System Technical Journal, 58(7):1673–1702, September 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-7-1673.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-7-1673.pdf.

Snare:1978:CCI

[Sna78] R. C. Snare. Common channel interoffice signaling: Peripheral maintenance and administration support system. The Bell System Technical Journal, 57(2):325–360, February 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-2-325.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-2-325.pdf.

Snell:1971:LLM

[Sne71]  W. W. Snell, Jr.  Low-

loss microstrip filters developed by frequency scaling. The Bell System Technical Journal, 50(6):1919–1931, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-6-1919.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-6-1919.pdf.

Someda:1973:SLL

[Som73] C. G. Someda. Simple, low-loss joints between single-mode optical fibers. The Bell System Technical Journal, 52(4):583–596, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-4-583.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-4-583.pdf.

Subramanian:1973:PDC

[SOP73] M. Subramanian, K. C. O’Brien, and P. J. Puglis. Phase dispersion characteristics during fade in a microwave line-of-sight radio channel. The Bell System Technical Journal, 52(10): 1877–1902, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1877.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1877.pdf.

Saunders:1977:BJB

[SP77] M. J. Saunders and W. L.

Parham. B.S.T.J. briefs: Adhesive sandwich optical fiber ribbons. The Bell System Technical Journal, 56(6):1013–1014, July/ August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-6-1013.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-6-1013.pdf.

Spang:1976:LNE

[Spa76] T. C. Spang. Loss-noise-echo study of the direct distance dialing network. The Bell System Technical Journal, 55(1):1– 36, January 1976. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol55/bstj55-1-
                                        1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-1-1.pdf.

Spencer:1977:NEP

[Spe77] A. E. Spencer, Jr. No. 4 ESS: Prologue. The Bell System Technical Journal, 56(7): 1015–1016, September 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-7-1015.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-7-1015.pdf.

Schafer:1971:DDF

[SR71] R. W. Schafer and L. R. Rabiner. Design of digital filter banks for speech analysis. The Bell System Technical Journal, 50(10): 3097–3115, December 1971. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol50/bstj50-10-3097.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol50-1971/articles/
                                        bstj50-10-3097.pdf.

Schafer:1974:BMR

[SR74] R. W. Schafer and L. R. Rabiner.

On the behavior of minimax relative error FIR digital differentiators. The Bell System Technical Journal, 53(2):333–361, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-333.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-333.pdf.

Sambur:1975:SID

[SR75] M. R. Sambur and L. R. Rabiner. A speaker-independent digit-recognition system. The Bell System Technical Journal, 54(1):81–102, January 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol54/bstj54-1-81.pdf;   http:
//www.alcatel-lucent.com/bstj/
vol54-1975/articles/bstj54-
1-81.pdf.

Schmidt:1979:DST

[SRB79] C. E. Schmidt, L. R. Rabiner, and D. A. Berkley. A digital simulation of the telephone system. The Bell System Technical Journal, 58(4):839–855, April 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-4-839.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-4-839.pdf.

Schafer:1975:FDF

[SRH75] R. W. Schafer, L. R. Rabiner, and O. Herrmann. FIR digital filter banks for speech analysis. The Bell System Technical Journal, 54(3):531–544, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-3-531.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-3-531.pdf.

Schwartz:1978:AFS

[SRMC78] M. I. Schwartz, W. A. Reenstra, J. H. Mullins, and J. S. Cook. Atlanta Fiber System Experiment: The Chicago Lightwave Communications Project. The Bell System Technical Journal, 57(6):1881–1888, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1881.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1881.pdf.

Samaroo:1970:DPE

[SRPR70] W. Samaroo, J. Raamot, P. Parry, and G. Robertson. Device photolithography: The electron beam pattern generator. The Bell System Technical Journal, 49(9):2077–2094, November 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol49/bstj49-9-2077.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2077.pdf.

Schneider:1971:SHI

[SS71a] M. V. Schneider and W. W.

Snell, Jr. A scaled hybrid integrated multiplier from 10 to 30 GHz. The Bell System Technical Journal, 50(6):1933–1942, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1933.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1933.pdf.

Strain:1971:NDA

[SS71b] R. J. Strain and N. L. Schryer.

A nonlinear diffusion analysis of charge-coupled-device transfer. The Bell System Technical Journal, 50(6):1721–1740, July/ August 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-6-1721.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-6-1721.pdf.

Schneider:1974:BJB

[SS74] M. V. Schneider and W. W.

Shell, Jr. B.S.T.J. briefs: Stripline downconverter with subharmonic pump. The Bell System Technical Journal, 53(6):1179–1183, July/ August 1974. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-6-1179.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-6-1179.pdf.

Schatz:1978:LPE

[SS78] J. G. Schatz and M. A. Schwartz. Loop plant electronics: Planning for loop electronic systems. The Bell System Technical Journal, 57(4):1035–1054, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-1035.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-1035.pdf.

Salz:1971:DTC

[SSP71] J. Salz, J. R. Sheehan, and D. J. Paris. Data transmission by combined AM and PM. The Bell System Technical Journal, 50(7):2399–2419, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2399.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2399.pdf.

Sharpe:1970:LAB

[SSV70] L. H. Sharpe, J. R. Shroff, and F. J. Vaccaro. Leadacid battery: Post seals for the new Bell System battery.

The Bell System Technical Journal, 49(7):1405–1417, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-7-1405.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-7-1405.pdf.

Standley:1974:BJB

[Sta74] R. D. Standley. B.S.T.J. briefs:

Fiber ribbon optical transmission lines. The Bell System Technical Journal, 53(6):1183–1185, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-6-1183.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-6-1183.pdf.

Staehler:1977:POO

[Sta77] R. E. Staehler. 1A processor: Organization and objectives. The Bell System Technical Journal, 56(2):119–134, February 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-2-119.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-2-119.pdf.

Stiles:1978:LPM

[Sti78] J. A. Stiles. Loop plant modeling: Economic design of distribution cable networks. The Bell System Technical Journal, 57(4):941–963, April 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-4-941.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-4-941.pdf.

Storey:1976:DMC

[Sto76] T. F. Storey. Design of a microprogram control for a processor in an electronic switching system. The Bell System Technical Journal, 55(2):183–232, February 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-2-183.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-2-183.pdf.

Stuck:1976:DSPa

[Stu76a] B. W. Stuck. Distinguishing stable probability measures — Part I: Discrete time. The Bell System Technical Journal, 55(8):1125–1182, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol55/bstj55-8-1125.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol55-1976/
                                        articles/bstj55-8-1125.pdf.

Stuck:1976:DSPb

[Stu76b]  B. W. Stuck.  Distinguish-

ing stable probability measures — Part II: Continuous time. The Bell System Technical Journal, 55(8):1183–1196, October 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-8-1183.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-8-1183.pdf.

Sullivan:1971:LSM

[Sul71]  J. L. Sullivan.  A labora-

tory system for measuring loudness loss of telephone connections. The Bell System Technical Journal, 50(8):2663–2739, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2663.pdf.

Stanaway:1979:TSP

[SVW79] J. J. Stanaway, Jr., J. J. Victor, and R. J. Welsch. Traffic service position system no. 1: Software development tools. The Bell System Technical Journal, 58(6):1307–1333, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1307.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1307.pdf.

Seastrand:1971:TMH

[SW71a] K. L. Seastrand and D. S. Williams. TH-3 medium-haul application: System considerations. The Bell System Technical Journal, 50(7):2271–2285, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2271.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2271.pdf.

Swift:1971:TMH

[SW71b] R. A. Swift and J. A. Word. TH-3 medium-haul application: Equipment and building considerations. The Bell System Technical Journal, 50(7): 2345–2353, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2345.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2345.pdf.

Slepian:1973:CTM

[SW73] D. Slepian and J. K. Wolf. A coding theorem for multiple access channels with correlated sources. The Bell System Technical Journal, 52(7):1037–1076, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol52/bstj52-7-1037.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol52-1973/
articles/bstj52-7-1037.pdf.

Semmelman:1971:SCD

[SWD71] C. L. Semmelman, E. D. Walsh, and G. T. Daryanani. Statistical circuit design: Linear circuits and statistical design. The Bell System Technical Journal, 50(4):1149–1171, April 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-4-1149.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-4-1149.pdf.

Saltzberg:1975:DDS

[SZ75] B. R. Saltzberg and H. M.

Zydney. Digital data system: Network synchronization. The Bell System Technical Journal, 54(5):879–892, May/ June 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-5-879.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-5-879.pdf.

Szentirmai:1973:SMF

[Sze73] G. Szentirmai. Synthesis of multiple-feedback active filters. The Bell System Technical Journal, 52(4):527–555, April 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol52/bstj52-4-527.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol52-1973/
                                        articles/bstj52-4-527.pdf.

Thomas:1974:LSC

[TAB74] J. L. Thomas, R. E. Anderson, and P. J. Baun. L5 system: Centralized transmission surveillance. The Bell System Technical Journal, 53(10):2035– 2064, December 1974. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol53/bstj53-10-2035.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol53-1974/articles/
                                        bstj53-10-2035.pdf.

Tabor:1972:BJB

[TBCR72] W. J. Tabor, A. H. Bobeck, G. P. Vella Coleiro, and A. Rosencwaig. B.S.T.J. briefs: A new type of cylindrical magnetic domain (bubble isomers). The Bell System Technical Journal, 51(6):1427–1431, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-6-1427.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-6-1427.pdf.

Thiele:1971:EGT

[TBTG71] A. A. Thiele, A. H. Bobeck, E. Della Torre, and U. F. Gianola. The energy and general translation force of cylindrical magnetic domains. The Bell System Technical Journal, 50(3):711–724, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-711.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-711.pdf.

Tatsuguchi:1976:WGG

[TG76] I. Tatsuguchi and J. W. Gewartowski.

A 10-W, 6-GHz, GaAs IM-PATT amplifier for microwave radio systems. The Bell System Technical Journal, 55 (2):167–182, February 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-2-167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-2-167.pdf.

Thiele:1971:DIT

[Thi71]  A. A. Thiele.  Device im-

plications of the theory of cylindrical magnetic domains. The Bell System Technical Journal, 50(3):725–773, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-3-725.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-3-725.pdf.

Thomas:1971:ADE

[Tho71a] David T. Thomas. Analysis and design of elementary blinders for large horn reflector antennas. The Bell System Technical Journal, 50(9):2979–2995, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-9-2979.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-9-2979.pdf.

Thomas:1971:AEC

[Tho71b] E. J. Thomas. An adaptive echo canceller in a nonideal environment (nonlinear or time variant). The Bell System Technical Journal, 50(8):2779–2795, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-8-2779.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-8-2779.pdf.

Thomas:1971:SCA

[Tho71c] E. J. Thomas. Some considerations on the application of the Volterra representation or nonlinear networks to adaptive echo cancellers. The Bell System Technical Journal, 50(8):2797–2805, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-8-2797.

pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2797.pdf.

Thornber:1973:ERD

[Tho73a] K. K. Thornber. Error rates of digital signals in charge transfer devices. The Bell System Technical Journal, 52(10):1795– 1809, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1795.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1795.pdf.

Thornber:1973:OLC

[Tho73b] K. K. Thornber. Operational limitations of charge transfer devices. The Bell System Technical Journal, 52(9): 1453–1482, November 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol52/bstj52-9-1453.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-9-1453.pdf.

Thornber:1974:TNC

[Tho74a] K. K. Thornber. Theory of noise in charge-transfer devices. The Bell System Technical Journal, 53(7):1211–1262, September 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol53/bstj53-7-1211.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-7-1211.pdf.

Thornber:1974:TMF

[Tho74b] K. K. Thornber. Treatment of microscopic fluctuations in noise theory. The Bell System Technical Journal, 53(6):1041–1078, July/ August 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol53/bstj53-6-1041.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol53-1974/
                                        articles/bstj53-6-1041.pdf.

Thomson:1977:WMW

[Tho77a] D. J. Thomson. WT4 millimeter waveguide system: Spectrum estimation techniques for characterization and development of WT4 waveguide — II. The Bell System Technical Journal, 56 (10):1983–2005, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-1983.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-1983.pdf.

Thomson:1977:SET

[Tho77b] David J. Thomson. Spectrum estimation techniques for characterization and development of WT4 waveguide — 1. The Bell System Technical Journal, 56(9):1769–1815, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1769.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1769.pdf.

Thompson:1978:UTS

[Tho78] K. Thompson. UNIX timesharing system: UNIX implementation. The Bell System Technical Journal, 57(6): 1931–1946, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1931.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1931.pdf.

Tompsett:1973:MTI

[TKK73] M. F. Tompsett, B. B. Kosicki, and D. Kahng. Measurements of transfer inefficiency of 250element undercut-isolated charge coupled devices. The Bell System Technical Journal, 52(1):1– 7, January 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-1.pdf.

Thornber:1974:BTS

[TKN74] K. K. Thornber, D. Kahng, and C. T. Neppell. Bias-temperature-stess studies of change retention in dual-dielectric, chargestorage cells. The Bell System Technical Journal, 53(9): 1741–1770, November 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-9-1741.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-9-1741.pdf.

Tribolet:1979:CTP

[TNMC79] J. M. Tribolet, P. Noll, B. J. Mc-Dermott, and R. E. Crochiere. A comparison of the performance of four low-bit-rate speech waveform coders. The Bell System Technical Journal, 58(3):699–712, March 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-3-699.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-3-699.pdf.

Tong:1970:PBT

[Ton70]  S. Y. Tong.  Performance

of bust-trapping codes. The Bell System Technical Journal, 49(4):477–491, April 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-4-477.
pdf;   http://www.alcatel-

lucent.com/bstj/vol49-1970/
articles/bstj49-4-477.pdf.

Tow:1975:DES

[Tow75] J. Tow. Design and evaluation of shifted-companion-form active filters. The Bell System Technical Journal, 54(3):545–568, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-3-545.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-3-545.pdf.

Tsiang:1979:AMP

[Tsi79] S. H. Tsiang. Advanced mobile phone service: Development support systems. The Bell System Technical Journal, 58(1):201–213, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-1-201.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-1-201.pdf.

Turrin:1975:MSR

[Tur75] R. H. Turrin. A multibeam, spherical-reflector satellite antenna for the 20- and 30-GHz bands. The Bell System Technical Journal, 54(6): 1011–1026, July/August 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol54/bstj54-6-1011.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-6-1011.pdf.

Urich:1971:PSS

[Uri71]  J. F. Urich.  The Picture-

phone System: Switching plan. The Bell System Technical Journal, 50(2):521–531, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-521.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-521.pdf.

VanHaften:1975:SDP

[Van75a] D. Van Haften. SAFEGUARD data-processing system: Software change control. The Bell System Technical Journal, 54 (10):S231–S236, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol54/bstj54-10-S231.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol54-1975/articles/
                                        bstj54-10-S231.pdf.

VanSciver:1975:SDP

[Van75b] P. A. Van Sciver. SAFEGUARD data-processing system: Systems programming in PL/1. The Bell System Technical Journal, 54 (10):S173–S180, December 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol54/bstj54-10-S173.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S173.pdf.

Ventrice:1972:SBT

[Ven72] F. X. Ventrice. Soil burial tests: Effect of soil burial exposure on the properties of casting resins. The Bell System Technical Journal, 51(1):43–46, January 1972. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol51/bstj51-1-
43.pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-1-43.pdf.

VanSlooten:1972:DCB

[VG72] D. J. VanSlooten and K. A.

Gluckow. D2 channel bank: Physical design and introductory program. The Bell System Technical Journal, 51 (8):1725–1742, October 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-8-1725.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-8-1725.pdf.

Vigants:1970:NFS

[Vig70]  Arvids Vigants.  The num-

ber of fades in space-diversity reception. The Bell System Technical Journal, 49(7): 1513–1530, September 1970.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1513.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1513.pdf.

Vigants:1971:NDF

[Vig71] Arvids Vigants. Number and duration of fades at 6 and 4 GHz. The Bell System Technical Journal, 50(3):815–841, March 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-3-815.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-3-815.pdf.

Vigants:1975:SDE

[Vig75] A. Vigants. Space-diversity engineering. The Bell System Technical Journal, 54(1):103–142, January 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-1-103.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-1-103.pdf.

Vigants:1979:TUF

[VP79] A. Vigants and M. V. Pursley. Transmission unavailability of frequency-diversity protected microwave FM radio systems caused by multipath fading. The Bell System Technical Journal, 58(8):1779–1796, October 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-8-1779.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-8-1779.pdf.

Wirtz:1974:LSP

[WA74] R. J. Wirtz and W. G. Albert. L5 system: Physical design. The Bell System Technical Journal, 53(10):2147–2194, December 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol53/bstj53-10-2147.pdf;
http://www.alcatel-lucent.
com/bstj/vol53-1974/articles/
bstj53-10-2147.pdf.

Wadsack:1978:FCD

[Wad78a] R. L. Wadsack. Fault coverage in digital integrated circuits. The Bell System Technical Journal, 57(5):1475–1488, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-5-1475.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-5-1475.pdf.

Wadsack:1978:FML

[Wad78b] R. L. Wadsack. Fault modeling and logic simulation of CMOS and MOS integrated circuits. The Bell System Technical Journal, 57(5):1449–1474, May/ June 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-5-1449.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-5-1449.pdf.

Waggener:1970:BBE

[Wag70]  H. A. Waggener.  BSTJ

briefs: Electrochemically controlled thinning of silicon. The Bell System Technical Journal, 49(3):473–475, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-3-473.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-3-473.pdf.

Wagar:1971:TET

[Wag71]  H. N. Wagar.  Transient

effect in telephone switching circuits when relay windings are disconnected. The Bell System Technical Journal, 50(9):2997–3062, November 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-9-2997.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-9-2997.pdf.

Wahl:1977:TYP

[Wah77] A. J. Wahl. Ten years of power aging of the same group of submarine cable semiconductor devices. The Bell System Technical Journal, 56(6):987–1005, July/ August 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-6-987.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-6-987.pdf.

Walvick:1970:CEC

[Wal70] E. A. Walvick. On the capacity of an ensemble of channels with differing parameters. The Bell System Technical Journal, 49(3):415–429, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-3-415.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-3-415.pdf.

Waldhauer:1977:AAF

[Wal77] F. D. Waldhauer. Anticausal analysis of feedback amplifiers. The Bell System Technical Journal, 56(8):1337–1386, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1337.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1337.pdf.

Walker:1979:AMP

[Wal79] J. T. Walker. Advanced mobile phone service: A service test mobile telephone control unit. The Bell System Technical Journal, 58(1):145–152, January 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol58/bstj58-1-145.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-145.pdf.

Wannier:1970:AMG

[Wan70]  Gregory H. Wannier.  On

an anomaly in the mobility of gaseous ions. The Bell System Technical Journal, 49(3):343–354, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-3-343.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-3-343.pdf.

Warters:1977:WMW

[War77] W. D. Warters. WT4 millimeter waveguide system: Introduction. The Bell System Technical Journal, 56(10):1825–1827, December 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol56/bstj56-10-1825.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol56-1977/articles/
                                        bstj56-10-1825.pdf.

White:1973:DGA

[WC73] Gerard White and Gen M.

Chin. A DC-to-2.3-GHz amplifier using an ‘embedding’ scheme. The Bell System Technical Journal, 52(1):53–62, January 1973. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol52/bstj52-1-
53.pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-1-53.pdf.

Webber:1977:BJB

[WHF77] S. A. Webber, C. J. Harris, and J. L. Flanagan. B.S.T.J. briefs: Use of variable-quality coding and time-interval modification in packet transmission of speech. The Bell System Technical Journal, 56(8):1569–1573, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-8-1569.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-8-1569.pdf.

White:1971:OMH

[Whi71] Gerard White. Optical modulation at high information rates. The Bell System Technical Journal, 50(8):2607–2645, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2607.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-8-2607.pdf.

Whitaker:1975:AOD

[Whi75]  B. A. Whitaker.  Analysis

and optimal design of a multiserver, multiqueue system with finite waiting space in each queue. The Bell System Technical Journal, 54(3):595–623, March 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol54/bstj54-3-595.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol54-1975/
                                        articles/bstj54-3-595.pdf.

Wilson:1970:BJB

[Wil70a]  R. W. Wilson.  B.S.T.J.

briefs: A three-radiometer pathdiversity experiment. The Bell System Technical Journal, 49(6):1239–1242, July/ August 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-6-1239.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-6-1239.pdf.

Wilson:1970:NTE

[Wil70b] Alan N. Wilson, Jr. New theorems on the equations of nonlinear DC transistor networks. The Bell System Technical Journal, 49(8):1713–1738, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1713.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1713.pdf.

Wilkinson:1971:SCL

[Wil71] R. I. Wilkinson. Some comparisons of load and loss data with current teletraffic theory. The Bell System Technical Journal, 50(8):2807–2834, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-8-2807.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-8-2807.pdf.

Willson:1972:SEQ

[Wil72] A. N. Willson, Jr. Some effects of quantization and adder overflow on the forced response of digital filters. The Bell System Technical Journal, 51(4):863–887, April 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-4-863.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-4-863.pdf.

Wilson:1977:VLN

[Wil77] Lynn O. Wilson. Vibrations of a lithium niobate fiber. The Bell System Technical Journal, 56(8):1387–1404, October 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol56/bstj56-8-1387.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol56-1977/
                                        articles/bstj56-8-1387.pdf.

Witsenhausen:1979:SRT

[Wit79] H. S. Witsenhausen. On the structure of real-time source coders. The Bell System Technical Journal, 58(6):1437– 1451, July/August 1979. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol58/bstj58-6-1437.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol58-1979/articles/
                                        bstj58-6-1437.pdf.

Walden:1972:BJB

[WKS^{+}72] R. H. Walden, R. H. Krambeck, R. J. Strain, J. McKenna, N. L. Schryer, and G. E. Smith. B.S.T.J. briefs: The buried channel charge coupled device. The Bell System Technical Journal, 51(7):1635–1640, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-7-1635.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-7-1635.pdf.

Wasserstrom:1970:PDC

[WM70] E. Wasserstrom and J. McKenna. The potential due to a charged metallic strip on a semiconductor surface. The Bell System Technical Journal, 49(5):853–877, May/ June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-853.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-853.pdf.

Wahl:1970:SST

[WMLT70] A. J. Wahl, W. McMahon, N. G. Lesh, and W. J. Thompson. SF system: Transistors, diodes and components. The Bell System Technical Journal, 49(5):683–698, May/June 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-5-683.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-5-683.pdf.

Wolfe:1972:BJB

[WN72] R. Wolfe and J. C. North.

B.S.T.J. briefs: Suppression of hard bubbles in magnetic garnet films by ion implantation. The Bell System Technical Journal, 51(6):1436–1440, July/ August 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol51/bstj51-6-1436.
pdf;   http://www.alcatel-
lucent.com/bstj/vol51-1972/
articles/bstj51-6-1436.pdf.

Wolman:1972:CPM

[Wol72]  Eric Wolman.  The camp-

on problem for multiple-address traffic. The Bell System Technical Journal, 51(6):1363– 1422, July/August 1972. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol51/bstj51-6-1363.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol51-1972/articles/
                                        bstj51-6-1363.pdf.

Warwick:1971:PSC

[WP71] P. S. Warwick and G. W. Phipps. The Picturephone System: Computer access. The Bell System Technical Journal, 50(2):683–700, February 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-2-683.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-2-683.pdf.

Williams:1975:SDP

[WP75] W. H. Mac Williams and J. E. Petersen. SAFEGUARD dataprocessing system: A costplus-award-fee contract for a large software development program. The Bell System Technical Journal, 54(10):S237– S244, December 1975. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/

Vol54/bstj54-10-S237.pdf;
http://www.alcatel-lucent.
com/bstj/vol54-1975/articles/
bstj54-10-S237.pdf.

Walk:1973:SGT

[WR73] R. Walk and J. Rootenburg.

Stability of a general type of pulse-width-modulated feedback system. The Bell System Technical Journal, 52(10): 1811–1819, December 1973. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol52/bstj52-10-1811.pdf;
http://www.alcatel-lucent.
com/bstj/vol52-1973/articles/
bstj52-10-1811.pdf.

Wrixon:1971:MAA

[Wri71] G. T. Wrixon. Measurements of atmospheric attenuation on an Earth–Space path at 90 GHz using a Sun tracker. The Bell System Technical Journal, 50(1):103–114, January 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol50/bstj50-1-103.
pdf;   http://www.alcatel-
lucent.com/bstj/vol50-1971/
articles/bstj50-1-103.pdf.

Wiley:1970:ENC

[WS70] J. D. Wiley and J. A. Seman.

The enumeration of neighbors on cubic and hexagonal-based lattices. The Bell System Technical Journal, 49(3):355–378, March 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-3-355.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-3-355.pdf.

Wonsiewicz:1978:UTS

[WSS78] B. C. Wonsiewicz, A. R. Storm, and J. D. Sieber. UNIX timesharing system: Microcomputer control of apparatus, machinery, and experiments. The Bell System Technical Journal, 57(6):2209–2232, July/ August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-2209.pdf.

Wyner:1971:IIP

[Wyn71] A. D. Wyner. On the intersymbol interference problem for the Gaussian channel. The Bell System Technical Journal, 50(7):2355–2363, September 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-7-2355.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-7-2355.pdf.

Wyner:1975:UBE

[Wyn75a] A. D. Wyner. Upper bound on error probability for detection with unbounded intersymbol interference. The Bell System Technical Journal, 54(7):1341–1351, September 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-7-1341.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-7-1341.pdf.

Wyner:1975:WTC

[Wyn75b] A. D. Wyner. The wire-tap channel. The Bell System Technical Journal, 54(8):1355–1387, October 1975. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol54/bstj54-8-1355.
pdf;   http://www.alcatel-
lucent.com/bstj/vol54-1975/
articles/bstj54-8-1355.pdf.

Wyner:1976:NCB

[Wyn76] A. D. Wyner. A note on the capacity of the bandlimited Gaussian channel. The Bell System Technical Journal, 55(3):343–346, March 1976. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol55/bstj55-3-343.
pdf;   http://www.alcatel-
lucent.com/bstj/vol55-1976/
articles/bstj55-3-343.pdf.

Wyner:1978:SDP

[Wyn78] A. D. Wyner. Signal design for PAM data transmission to minimize excess bandwidth. The Bell System Technical Journal, 57(9):3277–3307, November 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-9-3277.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-9-3277.pdf.

Yamin:1972:DAF

[Yam72a] Michael Yamin. Derivation of all figures formed by the intersection of generalized polygons. The Bell System Technical Journal, 51(7):1595–1610, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-7-1595.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol51-1972/
                                        articles/bstj51-7-1595.pdf.

Yamin:1972:XCP

[Yam72b] Michael Yamin. XYTOLR — A computer program for integrated circuit mask design checkout. The Bell System Technical Journal, 51(7):1581–1593, September 1972. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol51/bstj51-7-1581.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol51-1972/
articles/bstj51-7-1581.pdf.

Youngs:1979:TSP

[YBBW79] E. A. Youngs, W. J. Bushnell, and A. Barone-Wing. Traffic service position system no. 1: Automated coin toll service: Human factors studies. The Bell System Technical Journal, 58(6):1291–1305, July/ August 1979. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol58/bstj58-6-1291.
pdf;   http://www.alcatel-
lucent.com/bstj/vol58-1979/
articles/bstj58-6-1291.pdf.

Yeh:1970:AAR

[Yeh70] Y. S. Yeh. An analysis of adaptive retransmission arrays in a fading environment. The Bell System Technical Journal, 49(8):1811–1825, October 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-8-1811.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-8-1811.pdf.

Yeh:1971:III

[YH71] Y. S. Yeh and E. Y. Ho. Improved intersymbol interference error bounds in digital systems. The Bell System Technical Journal, 50(8):2585–2598, October 1971. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol50/bstj50-8-2585.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol50-1971/
                                        articles/bstj50-8-2585.pdf.

Young:1979:AMP

[You79] W. R. Young. Advanced mobile phone service: Introduction background and objectives. The Bell System Technical Journal, 58(1):1–14, January 1979. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol58/bstj58-1-
                                        1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol58-1979/
                                        articles/bstj58-1-1.pdf.

Zucker:1970:BJB

[Zuc70a] H. Zucker. B.S.T.J. briefs: On the computation of the far field of open Cassegrain antennas. The Bell System Technical Journal, 49(7):1595–1602, September 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol49/bstj49-7-1595.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol49-1970/
                                        articles/bstj49-7-1595.pdf.

Zucker:1970:ORV

[Zuc70b] H. Zucker. Optical resonators with variable reflectivity mirrors. The Bell System Technical Journal, 49(9):2349–2376, November 1970. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-9-2349.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-9-2349.pdf.

Zucker:1974:TDA

[Zuc74] H. Zucker. Time domain analysis and synthesis of notch filters. The Bell System Technical Journal, 53(2):283–304, February 1974. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol53/bstj53-2-283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol53-1974/
articles/bstj53-2-283.pdf.